### v1

# Bayesian Optimization in Language space: An Eval-Efficient AI Self-Improvement Framework

Abstract
Large Language Models (LLMs) have recently enabled self-improving AI, i.e., AI that iteratively generates, evaluates, and refine its own outcomes. Recent studies have shown that self-improving AI focusing on prompt optimization (i.e., context engineering) can outperform state-of-the-art reinforcement-learning fine-tuned LLMs. Here, their ‘performance’ is typically measured by query efficiency—the number of LLM-generated solution samples required to reach a specified performance threshold. However, in many societal applications, the primary limitation is not generating new solutions but evaluating them. For instance, evaluating an ad’s effectiveness requires significant human feedback, which is far more costly and time-consuming than generating a candidate ad. To optimize for the evaluation efficiency objective, a natural approach is to extend Bayesian Optimization (BO), a framework proven optimal for evaluation efficiency, to the language domain. However, the difficulty of directly estimating suitable acquisition functions in LLMs’ minds makes this extension challenging. This paper overcomes this challenge by proving that the combination of the simple and widely used Best-of-N selection strategy and simple textual gradients (i.e., textual edits from a critic model) statistically emulates the behavior of the gradients on the canonical Upper Confidence Bound (UCB) acquisition function, which induces optimal exploration in terms of evaluation efficiency. Based on this result, we propose TextGrad-Best-of-N Bayesian Optimization (T-BoN BO), a simple and eval-efficient language-space Bayesian optimization framework for AI self-improvement. We also empirically validate T-BoN BO by applying it to automated ad alignment tasks for persona distribution, demonstrating its superior performance compared to popular state-of-the-art baselines. In sum, T-BoN BO is a simple but principled framework for AI self-improving system design that bridges prompt optimization with classical Bayesian optimization.
Keywords: Self-improving AI, Evaluation efficiency, Bayesian optimization, Context Engineering
1 Introduction
1.1 Self-improving AI
Data-centric decision making has long relied on a human-driven cycle of historical data analysis, solution proposal, and experimentation. This human-driven iterative approach has supported a wide range of business decisions, ranging from website design (hauser2009website) to product recommendations (nandy2021b). We briefly describe this paradigm using the example of ad-optimization; see the left panel of Figure 1. The process starts with a manager generating a set of ad creatives based on a creative brief. This task is often done by ad agencies or marketers, who rely on their understanding of the product, theories of consumer behavior, market conditions, and feedback from prior ad campaigns. This ad generation process typically takes weeks, and serves as the main bottleneck in the iterative ad-optimization process. Once we have a set of candidate ads, the manager moves to evaluation, i.e., feeds these ads into advertising platforms or consumer surveys to identify the best performing one, and this step usually takes a few days. Finally, to complete this feedback loop, the manager analyzes the performance of the ads from this campaign and uses the insights learned to generate a new slate of candidate ads; and the iterative process continues.
Although effective in many settings, this human-driven workflow suffers from drawbacks. First, the speed with which firms can iterate over this process depends critically on how quickly humans can learn from the performance of prior solutions and propose new candidate solutions. Historically, generating new candidate solutions (or ads in the case of our example) has been time-consuming, and this step formed a key bottleneck in speeding up the process. Further, there exists no formal framework for how human agents can learn from the performance of prior solutions and leverage them to propose better candidates, and this step is often left to managerial intuition. Indeed, much of the research and managerial focus on data-driven business solutions has largely centered on the evaluation step, i.e., given a set of candidate solutions/options, how can we identify the most effective one – and solutions such as A/B tests and Multi-Armed Bandits (MAB) experiments have been proposed and tested (kohavi2020trustworthy; fiez_etal_2024); but how to leverage insights from prior rounds of evaluation systematically and improve the set of candidate solutions remains an unsolved problem.
Ad generation (Step 1)(Weeks)Analysis (Step 3)(days)Ad evaluation (Step 2)(days)HumanBottleneckAd generation (Step 1)(seconds)Analysis (Step 3)(minutes–hours)Ad evaluation (Step 2)(days)AIBottleneck
Figure 1: Comparison of human-driven and AI-driven ad self-improvement cycles in digital advertising.
Recent advances in generative AI have begun to transform this human-driven paradigm by providing potential solutions to the issues discussed above. First, candidate generation is fast and scalable with generative AI models. For example, in the case of digital ads, generative models can produce high-quality creative outputs, including ad copy, images, and complete campaign concepts that adhere to brand guidelines in a matter of minutes (jansen2024automated; hartmann2025power) (see right panel of Figure 1). [^1] Indeed, a substantial share of creatives is now generated by AI, reflecting improvements in both quality and speed (coffee2025aiadbuying; Majic2025VibeMarketing). Second, Large language Models (LLMs) and related generative systems can now conduct sophisticated data analysis, summarizing performance patterns and identifying potential drivers of success (ferber2024context; guo2024ds; ghosh2024exploring; wiedemer2025videomodelszeroshotlearners), providing a pathway for analyzing historical data and improving candidate solutions. As such, with the new capabilities enabled by AI, the traditional bottlenecks of analysis and generation of candidate solutions are significantly reduced. Given these developments, there is immense interest in the development and deployment of self-improving AI (silver2025welcome), which automates the complete analysis–generation–evaluation loop at scale, allowing systems to update themselves continuously in response to emerging data (MantiaChatterjeeLee2025; chen2025xbench).
1.2 Research Agenda and Challenges
In this paper, we study a prompt-optimization-based self-improving AI framework for societal and business problems, motivated by recent successes of prompt-optimization-based self-improving AI algorithms that outperform state-of-the-art reinforcement learning fine-tuning methods (agrawal2025gepa). While the use of modern generative AI systems has reduced the overhead associated with content generation and analysis (Steps 1 and 3), there are still three challenges that we need to overcome to design an efficient and effective self-improving AI that can automate decision-making in societal and business problems.
  • $\bullet$
    Challenge 1 – Evaluation-efficiency
    First, in many real-world societal and business problems, the primary bottleneck is no longer generating new solutions efficiently (query efficiency) but evaluating them efficiently (evaluation efficiency). For instance, in the case of digital advertising, generating ad creatives using LLM queries is super fast, but evaluating an ad’s effectiveness requires significant human feedback or large-scale experimentation (johnson2023inferno). As such, any AI system that we design needs to be evaluation-efficient, i.e., reach a high-performing solution using as few costly evaluations as possible, rather than trying to optimize for a few inexpensive generations as possible. While there is a nascent but growing literature in computer science on self-improving AI systems, this literature mainly focuses on generation efficiency as the main bottleneck (agrawal2025gepa), and has largely overlooked evaluation efficiency.
  • $\bullet$
    Challenge 2 – Lack of gradient functions
    Second, notice that we are essentially solving an optimization problem in language space, i.e., we want to take an input prompt, evaluate its performance, and improve it to reach a higher performance. [^2] Self-improving AI can thus be viewed as principled prompt optimization or context engineering since it involves improving textual inputs/prompts based on historic performance data (zhou2022large; khattab2023dspy; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro). In standard numerical settings, the natural way to solve such problems is to use gradient functions. However, in language spaces, the notion of a gradient is ill-defined because they lack the inner-product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction). Thus, we first need to define a meaningful notion of a “gradient” in the domain of text.
  • $\bullet$
    Challenge 3 – Balancing exploration and exploitation
    Third, while it is tempting to move fully along the textual gradient that worked well previously, i.e., exploit the knowledge gained from prior prompts, this can be sub-optimal due to uncertainty about that knowledge. Thus, we also need to bake some exploration in the choice of gradient direction. Again, while there exist bandit-based approaches that balance exploration and exploitation in the case where we have a fixed candidate set and the exploration space is numerical (jamieson2014best; russo2016simple; russo2018tutorial; mason2020finding; simchi2023multi; fiez_etal_2024), these do not automatically translate to our setting since our goal is to explore in the language space and our exploration focuses on changing the set of candidate solutions rather than deciding how much to explore each option within a fixed candidate set. Thus, a key challenge is to account for uncertainty to guide the search over the language space. Canonical uncertainty measures, such as the Upper Confidence Bound (UCB) or the posterior distribution, are not readily available from off-the-shelf LLMs, making their direct application to language space non-trivial. Further, we note that self-consistency variance or embedding distances are poor proxies, and training a Gaussian Process over high-dimensional text embeddings is brittle, difficult to implement, expensive, and data-hungry.
In this paper, our goal is to address the above challenges and build an evaluation-efficient self-improving AI that effectively balances exploration and exploitation in the language space.
1.3 Our Framework and Algorithm
We start with the observation that prompt optimization for self-improving AI follows a familiar pattern – we iteratively choose a candidate prompt that generates a candidate solution, observe a noisy performance score associated with this solution, and then decide where to query next while minimizing evaluations. Throughout, we have to be cognizant of three concerns: (1) we need to balance exploration and exploitation when choosing the next candidate solution, (2) we have no closed-form model linking prompts to scores, and (3) the main scarce resource is the number of evaluations rather than candidate generation. Our key insight is that these features of our problem are similar to the regime in which classic Bayesian Optimization (BO) operates successfully (frazier2018tutorial; srinivas2012information; papenmeier2025understanding). BO is a principled framework for tackling black-box optimization problems where evaluation efficiency is an important consideration. (See Web Appendix E for a detailed theoretical discussion of BO.) Thus, we seek to extend BO from numerical inputs to the space of natural language prompts to design an evaluation-efficient self-improving AI.
At its core, BO assumes a surrogate model (e.g., Gaussian Process or Reproducing Hilbert Kernel Space) to approximate the objective function and provide uncertainty estimates of that approximation. From this surrogate, BO constructs a function often called the acquisition function and selects as the next candidate the point with the highest acquisition value. One of the most widely used BO methods is Bayesian Optimization with Upper Confidence Bound (UCB) acquisition function (srinivas2012information). In practice, since the global maximizer of the UCB acquisition function is often not available in a closed form, we typically use gradient-based BO with UCB acquisition function (wilson2018maximizing). Gradient-based BO methods have shown surprising effectiveness for high-dimensional problems when the gradient-based BO is adapted to avoid vanishing gradients, for example, through random multi-start (where each initialization yields a local maximum of UCB) and careful gradient sampling strategies (papenmeier2025understanding). However, the standard gradient-based UCB BO is not directly applicable to the language space, as the notion of a gradient is ill-defined because they lack the inner-product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction). Therefore, we first need to define a notion of gradients in the language space and devise an appropriate acquisition function in the text space.
We now describe how we modify and adapt the BO framework to address the challenges discussed earlier. We propose a simple, evaluation-efficient iterative prompt-optimization framework that we call TextGrad-Best-of-N Bayesian Optimization (T-BoN BO) that addresses all three challenges discussed earlier. First, the BO framework naturally addresses the first challenge of the evaluation efficiency. It is known that BO with UCB acquisition function (srinivas2012information) is a theoretically evaluation-efficient algorithm, i.e., optimally balances exploration and exploitation, minimizing regret (whitehouse2023sublinear; wang2023regret). Specifically, for both expected cumulative regret and simple regret, the regret upper bound of BO with UCB acquisition function tightly matches the regret lower bound of scarlett2017lower.
To solve the second challenge, T-BoN BO builds on the recently proposed TextGrad (textgrad) framework that adapt gradient descent to discrete text by asking an external auxiliary LLM critic to propose targeted edits that improve a candidate; applying these edits induces a local, directional update, i.e., a “textual gradient”, in language space. This makes gradient-style search feasible without continuous relaxations or mapping the LLM prompts to the embedding space. textgrad show that textual gradients based on their approach enable consistent improvements across diverse scientific domains such as code-completion benchmarks, improving accuracy on graduate-level QA datasets, optimizing drug-like molecules, and refining radiotherapy planning for clinical objectives. These results establish textual gradients as an effective surrogate for numerical gradients when optimizing systems that operate in language space.
To address the third challenge, T-BoN BO chooses the next prompt to evaluate via a series of textual gradients which are selected under the Best-of-N principle (snell2024scaling). Moving in the textual gradient direction suggested by the critic exploits the expected improvement implied by past results but ignores uncertainty in those expectations. T-BoN bakes in exploration into the choice of gradient direction by adopting an optimism-based rule in the spirit of the optimism-in-the-face-of-uncertainty principle: it samples multiple local textual edits and moves along the direction the LLM judges most promising, i.e., the Best-of-$N$ candidate. We show that Best-of-$N$ over locally sampled textual edits statistically emulates gradient ascent on a UCB acquisition function, so that Best-of-$N$ effectively acts as an implicit optimistic acquisition maximizer in language space.
We now briefly describe our T-BoN BO algorithm. At each optimization iteration, in Step 1 (meta-reflection phase), we use the critic model to analyze the history of previously evaluated solutions and their scores, and update a meta-reflection that summarizes what has tended to work well or poorly. In Step 2 (Best-of-$N$ textual-gradient phase), we perform a sequence of textual gradient updates: for each gradient step, (i) we propose $N$ textual-gradient edits that transform the current prompt into $N$ candidate prompts, (ii) each candidate prompt is passed to a solution-generation model to produce a corresponding candidate solution, (iii) conditioned on the meta-reflections, the critic model predicts which candidate solution is most likely to perform best, and (iv) we set the prompt associated with this predicted-best solution as the new current prompt. After iterating Step 2, we obtain a refined prompt and its associated solution to evaluate. In Step 3 (evaluation phase), we evaluate this solution to obtain a scalar score, which is then used by T-BoN BO to update its history and meta-reflection for subsequent optimization steps.
1.4 Theoretical Guarantees and Empirical Evaluation
We provide theoretical guarantees on the performance of our approach. We formally show that, under mild regularity conditions, the Best-of-$N$ selection over locally sampled textual edits induces, in probability, ascent directions of the UCB acquisition function with an exploration parameter $\beta_{N}=\Theta(\sqrt{\ln N})$ (Theorem 2). In other words, T-BoN BO emulates gradient-based UCB Bayesian Optimization in the implicit embedding space induced by the LLM, even though it operates purely in language space and does not construct an explicit surrogate or uncertainty model. Since UCB-BO is known to achieve sublinear regret and optimal evaluation-efficiency guarantees, T-BoN BO inherits these theoretical properties up to constant factors and approximation error stemming from textual gradients. In sum, we formally establish that T-BoN BO is theoretically equivalent to the gradient-based implementation of Bayesian Optimization with the Upper Confidence Bound (UCB) acquisition function (srinivas2012information; wilson2018maximizing), which is proven to be evaluation-efficient (whitehouse2023sublinear; wang2023regret). This equivalence establishes T-BoN BO as a valid, eval-efficient, gradient-based Bayesian optimization procedure in the language space.
In addition to its theoretical efficiency, T-BoN BO is simple and easy to implement since does not require any explicit uncertainty modeling, embeddings, or Gaussian process surrogates. We now demonstrate the empirical performance of our approach using a series of numerical experiments. Concretely, we instantiate T-BoN BO in a digital advertising setting, where the goal is to iteratively refine image-generation prompts for ad creatives so as to maximize ad responsiveness for a population of customers. We test and measure how well and how fast a method optimizes for (i.e., aligns with) the preference distribution induced by LLM-based simulation of personas in the Digital Twins persona dataset called Twin-2k-500 (toubia2025twin). [^3] While LLM-based persona simulation can be biased and may not reflect real-world human behavior (li2025llm; peng2025mega), it still induces a valid preference distribution. As such, it can serve as a valid benchmark on which we can compare the performance of different AI-based optimization methods. We demonstrate that T-BoN BO significantly outperforms strong state-of-the-art baselines such as Best-of-N (snell2024scaling) and GEPA (agrawal2025gepa) in terms of evaluation efficiency, i.e., reaching fixed performance with a smaller number of steps. Specifically, we show that:
  • $\bullet$
    T-BoN BO, the proposed evaluation efficient framework, achieves nearly twice more improvements from Best-of-N baselines compared to GEPA (agrawal2025gepa), the state-of-the-art prompt optimization-based self-improving AI algorithm, which outperforms GRPO (liu2024deepseek), the state-of-the-art reinforcement learning fine-tuning algorithm.
  • $\bullet$
    T-BoN BO outperforms the state-of-the-art baseline even when we simulate preferences without any persona information, i.e., when the evaluator is a single LLM judge with no auxiliary context.
Together, these results show that T-BoN BO can reach good performance both in settings where there is significant heterogeneity in the preferences of the target population as well as in settings with homogeneous preferences.
In sum, our paper makes four main contributions to the literature on AI, optimization, and data-driven decision making. First, from a conceptual perspective, we identify evaluation efficiency as the key objective to optimize in self-improving AI design in societal and business systems. Second, we propose T-BoN BO, a purely text-based self-improving AI framework that combines textual-gradient search with Best-of-$N$ selection to automate and optimize business and marketing decisions (e.g., ad creatives, product descriptions). Third, we provide theory showing that Best-of-$N$ over locally sampled textual edits selects, in probability, ascent directions of the UCB acquisition with exploration weight $\beta=\Theta(\sqrt{\ln N})$; thus, we prove that T-BoN BO emulates gradient-based UCB and inherits evaluation-efficiency guarantees. Finally, we demonstrate empirical gains on automated ad-alignment across eight scenarios: T-BoN BO improves faster and attains higher final scores compared to strong baselines, including Best-of-64 and GEPA, and remains effective both in settings with and without contextual information.
The rest of the paper is organized as follows. Section 2 discusses the related literature. Section 3 describes the preliminaries related to the ideas discussed in this paper. Section 4 formalizes the AI system and its objective, along with self-improving algorithms. Section 5 introduces the main algorithm proposed in this paper, T-BoN BO, and its components. Section 6 proves that T-BoN BO theoretically emulates the gradient-based UCB Bayesian Optimization. Sections  7 and 8 empirically demonstrate T-BoN BO’s effectiveness on ad-alignment tasks versus strong baselines. Finally, Section 9 concludes with a discussion of implications and limitations.
2 Related Literature
Our paper relates and contributes to multiple streams of literature, including the literature on AI and LLMs in computer science, the operations literature on optimization, and the marketing literature on ad optimization. We now discuss each one below.
The literature on self-improving AI systems has often been motivated by an observation: the progress in AI based only on human data is hitting limits, so AI systems must create and learn from their own data. In other words, powerful AI systems “should have their own stream of experience that progresses, like humans, over a long time-scale” (silver2025welcome). That is, we need to design the self-improving AI system that iteratively 1) generate data of outcomes, evaluate outcomes with grounded signals, and 2) self-tune the three knobs –models, tools, and prompts– in the system to achieve long-term objectives under real-world feedback (wang2025maestro). Here, tuning models is about altering language models’ input-output correspondences (huang-etal-2023-large; lee2023rlaif; acikgoz2025self); tuning tools is about specifying the tools the AI system has access to, including memory tools (hou2025model; ouyang2025reasoningbank; qiu2025alita); tuning prompts is about engineering prompts in the model’s context window, including system messages, task instructions, constraints, schemas, and few-shot exemplars (zhou2022large; khattab2023dspy; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro).
In this paper, we study self-improving systems, focusing on tuning prompts, which have recently attracted attention after agrawal2025gepa showed that iterative prompt optimization can be more efficient than state-of-the-art model fine-tuning techniques such as GRPO (liu2024deepseek). Specifically, given any limited budget for the number of LLM generation query calls, their proposed iterative prompt optimization method, named GEPA (Genetic-Pareto), outperformed GRPO across their experiments. Those experiments included benchmarks such as complex questions across multiple Wikipedia articles, strict instruction-following under constraints, retrieval-augmented fact verification to support or refute claims, and maintaining answer quality while avoiding privacy leakage. In all those benchmarks, the number of LLM query calls for generating candidate answers was considered as the primary performance bottleneck. Likewise, prior literature on prompt-based self-improving AI (agrawal2025gepa; zhang2025agentic; wang2025maestro) has focused on measuring efficiency in terms of generation efficiency, i.e., minimizing the number of LLM-generated rollouts. On the other hand, our paper proposes and focuses on the evaluation efficiency objective. That is, instead of focusing on reducing the number of LLM outcome-generation rollouts, we aim to reduce the number of evaluations required to achieve an effective outcome.
Our paper also relates to the literature on relating LLMs to Bayesian Optimization (BO). A series of papers focuses on applying LLM to enhance BO for numerical optimization problems. liularge focuses on leveraging LLMs’ contextual understanding and few-shot learning capabilities to improve BO’s efficiency, enabling better warm-starting and surrogate modeling; agliettifunbo uses an LLM to discover new acquisition functions for Bayesian optimization. The LLM writes candidate acquisition function formulas as code, which are evaluated on various optimization problems. Conversely, singh2025mechanistic applies BO to enhance LLM’s internal behavior to boost the model’s zero-shot and few-shot performance. In contrast to these papers, our paper focuses on considering the prompt optimization problem that arises in LLM-based self-improving AI system design as a Bayesian Optimization problem. In line with this focus, schneider2025hyperbandbased and kong2025metapromptoptimizationllmbasedsequential formulate prompt optimization among the fixed set of candidate prompts as Bayesian optimization. While schneider2025hyperbandbased utilizes ideas from adversarial bandits, kong2025metapromptoptimizationllmbasedsequential feeds each prompt text into an embedding model, and the GP surrogate operates on those fixed embedding features to guide selection. In contrast to these approaches, our work focuses on self-improving AI systems that actively generate new candidate prompts as the process proceeds, rather than pre-specifying a fixed set of candidate prompts.
In terms of method evaluation benchmarks, our paper relates to the methodological literature on ad optimization in digital marketing. A series of papers focused on the efficient execution of ad evaluation, i.e., iteratively improving ad choice efficiently given a fixed set of candidate ads. schwartz2017customer adopts Thompson Sampling with a hierarchical general linear model for a real display advertising campaign for a financial-services firm, which purchased display advertising across 59 websites. geng2020online partitions JD.com’s platform users into disjoint sub-populations and applies a contextual Thompson sampling algorithm to maximize the expected advertiser payoff. aramayo2023multiarmed developed a contextual Thompson sampling algorithm to dynamically determine the list of house ads to display on the homepage of an electronic retailer to maximize the accumulated click-through rate of the ads. ba2022advertising used parametric Thompson Sampling, where a parametric structure links features (media, audience attributes) to conversion likelihood to optimally allocate ad-media and target-audience combinations in high-dimensional settings with low success rates. The main difference between bandit-based papers and ours is the existence of a pre-defined set of candidate arms. Self-improving AI systems focus on creating new solutions by analyzing previous results. On the other hand, bandit methods aim to balance exploration of a predefined set of candidate ads optimally. More broadly, there is also a difference in problem scope. Self-improving AI systems iterate among three steps: generation, evaluation, and analysis. In contrast, bandit methods focus on efficiently executing the evaluation step. Indeed, an AI self-improvement framework’s evaluation step may employ a bandit method to improve efficiency further.
3 Preliminaries
3.1 AI system.
An AI system $\Phi$ can be formally defined as a feed-forward modular workflow, which consists of one or more interconnected Large Language Model (LLM) components. In line with prior work (better_together; khattab2022demonstrate; miprov2; langprobe; agrawal2025gepa), we define an AI system $\Phi$ as a tuple $(M,C,\mathcal{X},\mathcal{Y})$. Here, $M=\langle M_{1},\ldots,M_{|M|}\rangle$ represents the set of LLM-powered components, $C$ is the control logic that orchestrates the workflow and sequences the tasks, and $\mathcal{X}$ and $\mathcal{Y}$ define the system input space and the system output space. Each component $M_{i}=(\pi_{i},\theta_{i},\mathcal{I}_{i},\mathcal{O}_{i})$ is an independent unit where $\pi_{i}$ is the prompt, $\theta_{i}$ represents the fixed weights of the underlying language model, and $\mathcal{I}_{i},\mathcal{O}_{i}$ specify the component’s input space and output space.
We denote the prompts in the whole system $\Pi=\left\langle\pi_{1},\ldots,\pi_{|M|}\right\rangle$. As the prompts $\Pi$ are what we system designers can control for the system $\Phi$, we denote $\Phi$ as the functional $\Phi(\cdot)$ when we would like to emphasize system $\Phi$’s dependency on prompts $\Pi$. Accordingly, we call $\Phi(\Pi)$ the ‘system outcome’ corresponding to prompt $\Pi$. Note that $\Phi(\Pi)$ is itself a function that maps $\mathcal{X}$ to $\mathcal{Y}$, i.e., $\Phi(\Pi):\mathcal{X}\mapsto\mathcal{Y}$.
Figure 2 illustrates an example AI system with five LLM-powered components. Here, its control logic $C=\{\mathcal{X}\rightarrow\left\{M_{1},M_{2}\right\},M_{1}\rightarrow M_{3},M_{2}\rightarrow M_{4},\left\{M_{3},M_{4}\right\}\rightarrow M_{5}\rightarrow\mathcal{Y}\}$ describes that the the system input $\mathcal{X}$ becomes the component input for $M_{1}$ and $M_{2}$, $M_{3}$’s input is $M_{1}$’s output, $M_{4}$’s input is $M_{2}$’s output, $M_{5}$’s input is $M_{3}$’s output and $M_{4}$’s output, and the final output $\mathcal{Y}$ is the output of $M_{5}$. The system $\Phi$’s input is the prompt $\Pi=\langle\pi_{1},\pi_{2},\pi_{3},\pi_{4},\pi_{5}\rangle$.
[…]
Figure 2: An example AI system with control logic $C=\{\mathcal{X}\rightarrow\left\{M_{1},M_{2}\right\},M_{1}\rightarrow M_{3},M_{2}\rightarrow M_{4},\left\{M_{3},M_{4}\right\}\rightarrow M_{5}\rightarrow\mathcal{Y}\}$ and prompt $\Pi=\langle\pi_{1},\pi_{2},\pi_{3},\pi_{4},\pi_{5}\rangle$.
3.2 TextGrad: Textual gradients in language space.
A central question for extending gradient-based Bayesian optimization to text is how to define a usable notion of “gradient” in language. A recent influential paper by (textgrad) addresses this problem by proposing a framework called the TextGrad; and it has has attracted broad interest across multiple research communities. Instead of computing numerical derivatives, TextGrad asks a critic LLM to propose targeted edits that would improve a candidate text. These edits act as local, directional signals, i.e., textual gradients, that guide search in a discrete language space. Importantly, this approach does not require embedding relaxations or external surrogate models, making it directly applicable to prompt optimization and other text-centric tasks.
For instance, consider an ad creative generation prompt for an image model: “A simple photo of a plant-based burger on a white plate.” A critic LLM might suggest: “Make the burger appear delicious, and show it being enjoyed at a barbecue.” Incorporating this edit yields: “A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends.” This change is analogous to following a gradient direction—small, constructive, and aligned with improving performance according to the evaluation signal.
Empirical studies provided in textgrad show that textual gradients enable consistent improvements across diverse scientific domains. In programming tasks, they raise success rates on challenging code-completion benchmarks. In knowledge-intensive reasoning, they improve accuracy on graduate-level QA datasets. In applied domains, they have been used to optimize drug-like molecules (via text encodings) and to refine radiotherapy planning prompts for clinical objectives. These results establish textual gradients as an effective surrogate for numerical gradients when optimizing systems that operate in and over language, justifying our use of gradient-style search in the discrete text space.
3.3 Bayesian Optimization.
Bayesian Optimization (BO) is a framework for maximizing an unknown and expensive-to-evaluate objective function under a limited evaluation budget. Formally, consider a black-box function $f:\mathcal{Z}\to\mathbb{R}$, where in our setting $f$ corresponds to the system-level objective $J(\Pi)$ over prompts. BO maintains a surrogate model, typically a Gaussian Process (GP) or another probabilistic regressor defined on $\mathcal{Z}$, that summarizes current beliefs about $f$. For any candidate $z\in\mathcal{Z}$, the surrogate yields a posterior mean $\mu(z)$ and an associated uncertainty measure $\sigma(z)$.
The next evaluation point is selected by maximizing an acquisition function $A(z)$ constructed from $\mu(z)$ and $\sigma(z)$, rather than by directly maximizing $\mu(z)$. Acquisition functions encode an explicit trade-off between exploitation (preferring points with high predicted value) and exploration (preferring points with high uncertainty). A widely used choice is the Upper Confidence Bound (UCB) acquisition function is:
$A_{\beta}(z)\;=\;\mu(z)\;+\;\beta\,\sigma(z),$ (1)
where the exploration parameter $\beta>0$ controls the strength of the exploration term (srinivas2012information; wilson2018maximizing). In many applications, the maximizer of $A_{\beta}$ is not available in closed form, and one instead employs gradient-based BO, which performs local ascent on $A_{\beta}$ (often from multiple random initializations) to obtain approximate maximizers (wilson2018maximizing; papenmeier2025understanding). In gradient-based BO, each evaluation is followed by multiple gradient steps that identify local optima of the acquisition function. Gradient-based BO methods have shown surprising effectiveness for high-dimensional problems under random multi-starts and careful choice of gradients (papenmeier2025understanding).
In our setting, we do not construct an explicit surrogate model; rather, we interpret $\mu(\cdot)$ and $\sigma(\cdot)$ as implicit beliefs realized within the LLM and design our method to emulate the behavior of gradient-based UCB in this implicit space.
4 Overview of Self-Improving AI Algorithms
4.1 The optimization objective.
Consider a system designer’s decision making problem, where she chooses the prompt set $\Pi$ for the LLM-based AI system $\Phi$. [^4] We call $\Pi$ in a plural form ‘prompts’ or ‘prompt set’ instead of a singular form ‘prompt’, as there are often multiple prompt components in one system. See AI system paragraph in Section 3.1 for a detailed description of an AI system. [^5] LLM-based AI systems can also be modified by tuning the LLM’s internal parameters using LLM fine-tuning methods such as GRPO (liu2024deepseek). In this paper, we focus on prompt optimization, which fixes those internal parameters throughout. We call $\Phi(\Pi)$ the ‘system outcome’ of $\Phi$ corresponding to a prompt set $\Pi$. In a digital marketing setting, $\Phi$ is an image generation AI, and $\Phi(\Pi)$ is the ad generated by the prompt set $\Pi$. As a system, $\Phi(\Pi)$ has inputs and corresponding outputs. Specifically, $\Phi(\Pi)$ maps the input space $\mathcal{X}$ to a distribution over an output space $\mathcal{Y}$. For each $x\in\mathcal{X}$, we denote by $y(\Phi(\Pi),x)$ the random $\mathcal{Y}$-valued vector output of the system at input $x$. In a digital marketing setting, input $x\in\mathcal{X}$ is a platform user, and a realization of $y(\Phi(\Pi),x)$ represents that user’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\Pi)$.
The distribution of inputs over $\mathcal{X}$ may depend on the prompts set $\Pi$. For example, in a digital marketing setting, a recommender system on the ad platform determines which users are shown the ad $\Phi(\Pi)$. To capture such dependences, we denote the input distribution over $\mathcal{X}$ by $\mathcal{D}_{\mathcal{X}}(\Pi)$.
To evaluate the output, we use a scoring rule: a function $r:\mathcal{Y}\to[0,1]$ that assigns a scalar effectiveness score to each $y\in\mathcal{Y}$. In the digital marketing example, $r$ is a marketer-defined function that maps observed user responses (clicks, conversions, spend) to a normalized scalar effectiveness score.
Our optimization objective is to discover the prompt $\Pi^{*}$ that maximizes the objective function
$J(\Pi):=\mathbb{E}[r(y(\Phi(\Pi),x))].$ (2)
Note that $J$ cannot be optimized by directly computing gradients, because the input distribution $\mathcal{D}_{\mathcal{X}}(\Pi)$ can shift with $\Pi$; even when $J(\Pi)$ is differentiable with respect to $\Pi$ and $y$’s randomness is not dependent on $\Pi$, we have:
$\nabla_{\Pi}J(\Pi)=\underbrace{\mathbb{E}\big[\nabla_{\Pi}r(y(\Phi(\Pi),x))\big]}_{\text{model term}}\;+\;\underbrace{\mathbb{E}\big[r(y(\Phi(\Pi),x))\,\nabla_{\Pi}\log d_{\Pi}(x)\big]}_{\text{distribution term}},$ (3)
where $d_{\Pi}$ is the density of $\mathcal{D}_{\mathcal{X}}(\Pi)$. Since $d_{\Pi}$ and $\nabla_{\Pi}\log d_{\Pi}(x)$ are unobservable, we cannot compute gradients from observed rewards. This motivates the use of an external LLM that can analyze prior data and propose directions for improvement without explicitly computing gradients, leading to self-improving AI algorithms we describe below that iteratively optimize for $J$.
4.2 Critic LLM and Self-improving AI algorithm
Given a scoring rule $r$, a self-improving AI algorithm $\mathcal{A}$ optimizes $J(\Pi)$ through iteratively analyzing the previous history, proposing new prompts, and observing their scores. This is often enabled by assuming an external critic LLM, denoted $M_{\text{critic}}$, that can analyze previous data and propose the next solution.
Critic LLM.
Specifically, a critic LLM is assumed to have the following capabilities:
  • 1)
    The critic LLM can look at $\Phi(\Pi_{1})$ and $\Phi(\Pi_{2})$ side by side and predict which one will be better.
  • 2)
    The critic LLM can read the history of past attempts and summarize the main lessons as a short list of rules or guidelines.
  • 3)
    The critic LLM can suggest a small set of prompt edits that are likely to improve the current prompt.
In the digital-marketing example, these correspond to an LLM’s ability to (1) compare two ad creatives and predict which will perform better (calderon2025alternative), (2) examine past performance data and derive guidelines for effective ads in context (jiang2024many; ferber2024context), and (3) propose prompt improvements based on those guidelines (textgrad; li2025test).
Self-improving AI algorithm
Given an initial prompt defined as $\Pi_{0}$, a self-improving AI algorithm proceeds in discrete optimization iterations, indexed by $t=1,2,\ldots$. At the beginning of iteration $t$, the algorithm has access to the history of past prompts and their observed scores and then selects a new prompt set $\Pi_{t}$, deploys the corresponding system $\Phi(\Pi_{t})$, and receives feedback about its performance. At each algorithm iteration $t$, we define the score $s_{t}$ of prompts $\Pi_{t}$ as the empirical average reward at $t$, i.e.,
$s_{t}:=\frac{1}{L}\sum_{l=1}^{L}r\left(y\left(\Phi\left(\Pi_{t}\right),x_{t}^{(l)}\right)\right),\quad x_{t}^{(l)}\sim\mathcal{D}_{\mathcal{X}}\left(\Pi_{t}\right),$ (4)
where $L$ denotes the number of input samples of $x_{t}^{(l)}$ we evaluate for each $t$. Note that we compare the performance of algorithms in terms of the progress of $s_{t}$ across iterations $t$.
Given access to a history of past evaluations $H_{t-1}=\{(\Pi_{\tau},s_{\tau})\}_{i=\tau}^{t-1}$ where $s_{\tau}$ is the empirically evaluated score for the prompts $\Pi_{\tau}$, a self-improving AI algorithm is defined as an algorithm $\mathcal{A}$ iterates the following steps at each $t$ to optimize for $J(\Pi)$:
  1. 1.
    Analysis: The critic LLM analyzes the previous history $H_{t-1}$.
  2. 2.
    Generation: The critic LLM proposes the next prompts $\Pi_{t}$ based on the analysis.
  3. 3.
    Evaluation: The system $\Phi(\Pi_{t})$ is deployed, and its performance is measured, yielding a score $s_{t}$. The new observation is added to the history: $H_{t}=H_{t-1}\cup\{(\Pi_{t},s_{t})\}$.
[…]
Figure 3: A self-improving AI algorithm implemented on the example AI system in Figure 2.
Figure 3 illustrates how the example AI system described in Figure 2 is positioned in the self-improving AI.
5 TextGrad-Best-of-N Bayesian Optimization (T-BoN BO)
We now describe the self-improving AI framework proposed in this paper, TextGrad-Best-of-N Bayesian Optimization (T-BoN BO) (Algorithm 1). Along with the system to optimize $\Phi$ and the scoring rule $r$, it requires a critic model $M_{\text{critic}}$.
Algorithm.
T-BoN BO starts from an initial prompt $\Pi_{0}$. Initially at iteration $t=0$, it gets the initial system $c_{0}=\Phi(\Pi_{0})$ and evaluates it on $\mathcal{D}_{\mathcal{X}}(\Pi_{0})$ to obtain initial score $s_{0}=\mathrm{Eval}(c_{0},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$. The history and the reflection is initialized as $H_{0}=\{(\Pi_{0},c_{0},s_{0})\}$ and $R_{0}=\emptyset$.
We first fix the hyperparameters: the number of candidates per gradient step ($N$), the number of gradient steps per iteration ($G$), and the number of iterations ($T$). Then at each iteration $t$, it first sets $\Pi_{t,0}$ as $\Pi_{t-1}$, the previous iteration’s prompt, and proceeds as follows:
  • 1.
    Best-of-N Textual Gradient steps. (Detailed in Section 5.1) Repeat steps 1) – 4) for $g=1,\ldots,G$:
    1. 1)
      Textual gradient generation: From the current inner prompt $\Pi_{t,g-1}$, query the textual-gradient generation operator $\nabla_{\text{text}}(\Pi_{t,g-1};R_{t-1},M_{\text{critic}})$ independently $N$ times to obtain textual gradients $\{\delta^{(i)}_{t,g}\}_{i=1}^{N}$.
    2. 2)
      Candidate prompt generation: Apply each textual gradient $\delta_{t,g}^{(i)}$ to $\Pi_{t,g-1}$ for all $i$ to get $\{\Pi_{t,g}^{(i)}\}_{i=1}^{N}$.
    3. 3)
      Candidate system outcome generation: Produce the system outcomes $\{c_{t,g}^{(i)}:=\Phi(\Pi_{t,g}^{(i)})\}_{i=1}^{N}$.
    4. 4)
      Best-of-N selection: Find the best of $\{(\Pi_{t,g}^{(i)},c_{t,g}^{(i)})\}_{i=1}^{N}$ and set the best one as $(\Pi_{t,g},c_{t,g})$.
  • 2.
    Evaluation. (Detailed in Section 5.2) Set $c_{t}=c_{t,G}$ and $\Pi_{t}=\Pi_{t,g}$. Evaluate $c_{t}$ to get the evaluation result $s_{t}$, i.e., $s_{t}=\mathrm{Eval}(c_{t},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$. Append $(\Pi_{t},c_{t},s_{t})$ to $H_{t-1}$ to obtain $H_{t}$. If $s_{t}>s_{t-1}$, keep $(\Pi_{t},c_{t})$. Otherwise, set $(\Pi_{t},c_{t})=(\Pi_{t-1},c_{t-1})$ and set $s_{t}=s_{t-1}$. [^6] This ‘gated’ acceptance/rollback was adopted from how textgrad implemented TextGrad.
  • 3.
    Meta-reflection. (Detailed in Section 5.3) Update the reflection $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$ for use at iteration t+1t{+}1.
After $T$ optimization steps, $\{c_{t}\}_{t=1}^{T}$ and $\{s_{t}\}_{t=1}^{T}$ are reported as the algorithm’s optimization outcome.
Note that T-BoN BO takes $N\times G$ generation of prompt/outcome candidates before one evaluation. This careful selection of a prompt/outcome candidate to evaluate is the key to evaluation efficiency in T-BoN BO: it deliberately trades inexpensive generation compute for fewer evaluations.
[…]
Figure 4: Summary of T-BoN BO: $T$ iterations, $G$ gradient steps, and Best-of-$N$ candidates at each gradient step. T-BoN BO deliberately trades inexpensive generation compute for fewer evaluations: each iteration generates $N\times G$ prompt/outcome candidates with no evaluation.
Input: System $\Phi$; initial prompt $\Pi_{0}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$ ; Output: Optimized $\Pi_{T},c_{T},s_{T}$ 1 $c_{0}\leftarrow\Phi(\Pi_{0})$; 2 $s_{0}\leftarrow\text{Eval}(c_{0},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$; 3 $H_{0}\leftarrow\{(\Pi_{0},c_{0},s_{0})\}$, $R_{0}\leftarrow\emptyset$; 4 for $t=1$ to $T$ do 5 $\Pi_{t,0}\leftarrow\Pi_{t-1}$; 6 for $g=1$ to $G$ do 7    for $i=1$ to $N$ do 8       $\delta^{(i)}_{t,g}\leftarrow\nabla_{\text{text}}(\Pi_{t,g-1};R_{t-1},M_{\text{critic}})$ ; 9       $\Pi_{t,g}^{(i)}\leftarrow\operatorname{Apply}(\Pi_{t,g-1},\delta_{t,g}^{(i)})$ ; 10       $c_{t,g}^{(i)}\leftarrow\Phi(\Pi_{t,g}^{(i)})$ ; 11        12      end for 13    $(\Pi_{t,g},c_{t,g})\leftarrow\textsc{Best-of-N}(\{(\Pi_{t,g}^{(i)},c_{t,g}^{(i)})\},R_{t-1},M_{\text{critic}})$ 14   end for 15 $c_{t}\leftarrow c_{t,G}$, $\Pi_{t}\leftarrow\Pi_{t,G}$; 16 $s_{t}\leftarrow\text{Eval}(c_{t},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$; 17 $H_{t}\leftarrow H_{t-1}\cup\{(\Pi_{t},c_{t},s_{t})\}$; 18 if $s_{t}<s_{t-1}$ then 19    $\Pi_{t}\leftarrow\Pi_{t-1}$, $c_{t}\leftarrow c_{t-1}$, $s_{t}\leftarrow s_{t-1}$; 20     21   end if 22 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$; 23  24 end for return $\{c_{t}\}_{t=1}^{T}$ Algorithm 1 T-BoN BO
Input: System $\Phi$; $\{\Pi_{0}^{j}\}_{j=1}^{J}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$; $J$ trajectories Output: Optimized triples $\{(\Pi^{j}_{T},c^{j}_{T},s^{j}_{T})\}_{j=1}^{J}$ 1 for $j=1$ to $J$ do 2 $c_{0}^{j}\leftarrow\Phi(\Pi_{0}^{j})$; 3 $s_{0}^{j}\leftarrow\text{Eval}(c_{0}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$; 4 $H_{0}\leftarrow\{(\Pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$, $R\leftarrow\emptyset$; 5  6 end for 7 8for $t=1$ to $T$ do 9 for $j=1$ to $J$ do 10    $\Pi^{j}_{t,0}\leftarrow\Pi^{j}_{t-1}$; 11    for $g=1$ to $G$ do 12       for $i=1$ to $N$ do 13          $\delta_{t,g}^{j,(i)}\leftarrow\nabla_{\text{text}}(\Pi_{t,g-1}^{j};R,M_{\text{critic}})$ ; 14          $\Pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}(\Pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)})$ ; 15          $c_{t,g}^{j,(i)}\leftarrow\Phi(\Pi_{t,g}^{j,(i)})$ ; 16           17         end for 18       $(\Pi_{t,g}^{j},c_{t,g}^{j})\leftarrow\textsc{Best-of-N}(\{(\Pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\},R_{t-1},M_{\text{critic}})$ 19      end for 20    $c_{t}^{j}\leftarrow c^{j}_{t,G}$, $\Pi_{t}^{j}\leftarrow\Pi_{t,G}^{j}$; 21    $s_{t}^{j}\leftarrow\text{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$; 22    $H_{t}\leftarrow H_{t-1}\cup\{(\Pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}$; 23    if $s_{t}^{j}<s_{t-1}^{j}$ then 24       $\Pi_{t}^{j}\leftarrow\Pi_{t-1}^{j}$, $c_{t}^{j}\leftarrow c_{t-1}^{j}$, $s_{t}^{j}\leftarrow s_{t-1}^{j}$; 25        26      end if 27     28   end for 29 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$; 30  31 end for return $\{c_{t}^{j_{t}^{*}}\}_{t=1}^{T}$, where $j_{t}^{*}:=\operatorname{argmax}_{j}s_{t}^{j}$ Algorithm 2 Parallel T-BoN BO
Figure 5: T-BoN algorithm (Left) and Parallel T-BoN algorithm (Right).
5.1 Best-of-N Textual Gradient steps.
At each gradient step $g=1,\ldots,G$, the algorithm expands the current prompt $\Pi_{t,g-1}$ into a neighborhood of $N$ variants, then contracts back to a single survivor. The process begins with textual gradient generation: the critic model $M_{\text{critic}}$ produces $N$ improvement directions,
$\delta^{(i)}_{t,g}\;\sim\;\nabla_{\text{text}}\!\big(\Pi_{t,g-1};\,R_{t-1},\,M_{\text{critic}}\big),\qquad i=1,\ldots,N,$ (5)
where randomness is introduced by a nonzero sampling temperature and nucleus filtering (top-$p$ sampling) [^7] In the experiments, we used Gemini Flash 2.5 with temperature =0.8=0.8 and top $p=0.9$. , which ensures that the gradients are not identical but represent diverse, locally valid edits.
Each textual gradient is then applied to form a candidate prompt
$\Pi^{(i)}_{t,g}=\operatorname{Apply}\!\big(\Pi_{t,g-1},\,\delta^{(i)}_{t,g}\big)$, (6)
with the rewrite operator configured for stability (low/zero temperature) so the edits remain coherent. The generator model then produces the system outcomes
$c^{(i)}_{t,g}:=\Phi\!\big(\Pi^{(i)}_{t,g}\big)$. (7)
The system outcome candidates are reduced to one winner by the Best-of-N choice operation. For the Best-of-N choice operation, following liu2025pairjudgermperformbestofn, we employ the pairwise tournament that uses critic model $M_{\text{critic}}$. First, the $N$ candidates are permuted to produce randomly chosen tournament matches. Each match compares two outcomes $(c^{(i)}_{t,g},c^{(j)}_{t,g})$ by querying $M_{\text{critic}}$ for a head-to-head judgment (For the prompt, please refer to Figure 12 in Appendix A). illustrates the prompt example for ad generation tasks. To mitigate noise, the comparison is repeated $K$ times (typically an even number to evenly swap orders to remove position bias (shi2025judgingjudgessystematicstudy), and break even by random; here we chose $K=4$), producing votes $\{V_{k}(i,j)\}_{k=1}^{K}$, where $V_{k}(i,j)=i$ if critic prefers $i$ over $j$, and vice versa. The winner of the pairwise match is then defined by the majority rule $\operatorname{argmax}_{u\in\{i,j\}}\sum_{k=1}^{K}\mathbf{1}_{\{V_{k}(i,j)=u\}}$. Winners advance round by round until only one candidate remains. The survivor $(\Pi_{t,g},c_{t,g})=(\Pi^{(i^{\star})}_{t,g},c^{(i^{\star})}_{t,g})$ is carried forward for the next gradient step $g+1$.
5.2 Evaluation
After completing $G$ gradient steps, we arrive at the $(\Pi_{t,G},c_{t,G})$, which we redefine as $(\Pi_{t},c_{t})$. We evaluate $(\Pi_{t},c_{t})$: recall from Section 4 that $(\Pi_{t},c_{t})$ is evaluated as
$s_{t}\;:=\;\frac{1}{L}\sum_{l=1}^{L}r\!\left(y\left(c_{t},x_{t}^{(l)}\right)\right),\qquad x_{t}^{(l)}\sim\mathcal{D}_{\mathcal{X}}(\Pi_{t}),$ (8)
where $r\!\left(y\left(c_{t},x_{t}^{(l)}\right)\right)$ is the evaluation of $c_{t}$ for the input $x_{t}^{(l)}$ and $L$ denotes the number of inputs evaluated. After evaluation, the triple $(\Pi_{t},c_{t},s_{t})$ is appended to the history $H_{t-1}$ to form $H_{t}$. Following textgrad, we adopt a simple acceptance rule that controls progress: if $s_{t}\geq s_{t-1}$, the new candidate is adopted; otherwise, the system rolls back to the previous iteration’s outcome, i.e., $(\Pi_{t-1},c_{t-1},s_{t-1})$.
5.3 Meta-reflection.
At the end of each iteration $t$, we compress the full history $H_{t}$ of prompts, generated system outcomes, and evaluation results into a natural language reflection $R_{t}$. $R_{t}$ consists of a small set of actionable rules. In the ad generation example, those rules may take the form “emphasize social context,” “avoid cluttered backgrounds,” or “prefer warm lighting.” This reflection serves as a global memory; in subsequent iterations, the critic LLM takes this reflection as a part of its context when generating textual gradients. (For the prompt used for meta-reflection, see Figure 13 in Appendix A).
There are many ways to construct a reflection from past results. In our experiments, following li2025test, we instruct the critic model $M_{\text{critic}}$ to compare the five lowest- and five highest-performing outcomes from previous iterations (or, when fewer than ten are available, all outcomes) [^8] Limiting the number of outcomes in the critic model’s context window helps mitigate context rotting (hong2025context)., identify recurring visual and textual patterns, and distill them into a concise set of actionable rules.
5.4 Parallel T-BoN BO.
Although a local maximum found by gradient steps from a single starting point is often effective in practice, it may under-explore the ample prompt space. Following the practices (wilson2018maximizing; papenmeier2025understanding), we also introduce a parallel variant of T-BoN BO that runs $J$ optimization trajectories simultaneously (Algorithm 2). Here, each of $J$ trajectories starts from a different initial prompt (denoted as $\{\Pi_{0}^{j}\}_{j=1}^{J}$). After independent gradient steps and evaluations at each iteration, evaluations from all trajectories are merged into the shared history, and a single meta-reflection is produced. Different trajectories diversify the search, while the shared reflection ensures that discoveries made by one trajectory are immediately shared with the others. Combined with the careful choice of gradients in T-BoN BO, this parallel process ensures the broader coverage of the search space and faster convergence to high-quality solutions in high-dimensional BO (papenmeier2025understanding) with a slight loss of optimality in the worst case (wilson2018maximizing).
At the end of $T$ optimization steps, the $J$-trajectory parallel T-BoN BO returns $\{(\Pi_{T}^{j},c_{T}^{j},s_{T}^{j})\}_{j=1}^{J}$. For the purpose of reporting, for each $t\leq T$, we select $j_{t}^{\ast}:=\operatorname{argmax}_{j}s_{t}^{j}$ and report its output $c_{t}^{j_{t}^{\ast}}$ as the step $t$’s outcome. That is, $\{c_{t}^{j_{t}^{\ast}}\}_{t=1}^{T}$ is reported as the final outcome.
6 Theory: T-BoN BO’s Asymptotic Equivalence to UCB Bayesian Optimization
In this section, we prove that the Best-of-$N$ over textual gradients selects, in probability, the ascent direction of the UCB acquisition with exploration weight $\beta$ scaling as $\Theta(\sqrt{\ln N})$ (Theorem 2). What allows us to do rigorous theoretical analysis is that LLM prompt embeddings are invertible and bi-Lipschitz: LLM’s prompt embeddings are known to be injective, and therefore invertible (nikolaou2025language); LLM embeddings are Lipschitz (tang2025understanding), implying they are bi-Lipschitz. Hence, small textual edits can be represented as small changes in the embedding space and vice versa. This enables us to map changes in prompts to changes in embeddings in Euclidean space, and later bring the theoretical guarantees back in the prompt space.
Intuitive explanation.
Denote $x$ to be an embedding of a prompt. Pick a small radius $\varepsilon>0$ around $x$ and sample $N$ nearby edits $x_{i}=x+\varepsilon u_{i}$ with $u_{i}$ on the unit sphere. Score each point
$Y_{i}=\mu(x_{i})+\xi_{i}\,\sigma(x_{i}),\quad\xi_{i}\text{ i.i.d. sub-Gaussian and independent of }\{u_{i}\}.$ (9)
Let $M_{N}=\max_{i}\xi_{i}$ and $q_{N}=F^{-1}(1-1/N)$. Since $M_{N}\simeq q_{N}$ and $q_{N}=\Theta(\sqrt{\ln N})$, Best-of-$N$ acts like picking the sampled $x_{i}$ with the largest
$\mu(x_{i})+q_{N}\,\sigma(x_{i}).$ (10)
On the small sphere around $x$ the function $A_{\beta}(x^{\prime})=\mu(x^{\prime})+\beta\,\sigma(x^{\prime})$ has a single best direction. Call it $v_{\beta}$. Draw a narrow cone around $v_{\beta}$. Then, with a large enough $N$, at least one $u_{i}$ lands in that cone with high probability. The winner comes from that cone. Set $\beta=q_{N}$. Then Best-of-$N$ behaves like UCB with weight $\beta$: it picks the sampled point most similar to the $A_{\beta}$ best direction.
Rigorous derivation.
Let $\mathcal{M}$ be the set of syntactically valid prompts. Suppose that there exists a text embedding $\mathsf{E}:\mathcal{M}\to\mathbb{R}^{d}$, where $x\;:=\;\mathsf{E}(\Pi)\in\mathbb{R}^{d}$ is the embedding of a prompt $\Pi\in\mathcal{M}$. We first state the assumptions motivated by how language model embeddings behave in practice, and then state the main theorem, Theorem 2.
At each gradient step of T-BoN BO, we sample edits that correspond to $u_{1},\ldots,u_{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\rho$ on the embedding space $\subset\mathbb{R}^{d}$, where $\rho$ is assumed to have a density bounded below by a positive constant with respect to surface measure. Assumption 1 implies that small textual edits induce near‑linear moves in $\mathbb{R}^{d}$, consistent with the recent findings that prompt embeddings preserve phrase similarity under small changes (reimers2019sentence; he2025position; tang2025understanding; nikolaou2025language).
Assumption 1 (Embedding map and edit realization).
For the current prompt $\Pi$ with $x=\mathsf{E}(\Pi)$, constants $C<\infty$ and $\varepsilon_{0}>0$ such that for every unit vector $u\in\mathbb{R}^{d}$ and every $\varepsilon\in(0,\varepsilon_{0}]$ the edit operator $\mathrm{Apply}$ realizes a first–order move:
$\mathsf{E}\!\big(\mathrm{Apply}(\Pi,\varepsilon u)\big)=x+\varepsilon u+r(\varepsilon,u),\qquad\sup_{\|u\|=1}\|r(\varepsilon,u)\|\leq C\,\varepsilon^{2}.$ (11)
Next, the assumption 2 is a standard regularity/technical condition: let’s call a function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $C^{1,1}$ if it is continuously differentiable and its gradient is Lipschitz: $\exists L<\infty$ such that $\|\nabla f(y)-\nabla f(z)\|\leq L\|y-z\|$ for all $y,z$.
Assumption 2 ($\mu,\sigma$ are of $C^{1,1}$).
Let $\mu(x)$ and $\sigma(x)$ denote the mean and standard deviation of the evaluation variable $Y$ at embedding $x$. We assume $\mu,\sigma$ are $C^{1,1}$ locally. We use $g=\nabla\mu(x)$ and $h=\nabla\sigma(x)$.
Assumption 3 specifies a heteroscedastic sub-Gaussian model for candidate scores (with i.i.d. mean-zero, unit-variance, unbounded-support noise independent of the sampled directions) and an ideal Best-of-$N$ choice oracle.
Assumption 3 (Noise model and choice oracle).
Let $\{\xi_{i}\}_{i=1}^{N}$ be i.i.d. mean-zero, unit-variance, Gaussian random variables with unbounded support, independent of $\{u_{i}\}$. For candidates $x_{i}:=x+\varepsilon u_{i}$ we observe
$Y_{i}\;=\;\mu(x_{i})\;+\;\sigma(x_{i})\,\xi_{i},$ (12)
and the Best-of-$N$ oracle returns $i^{\star}\in\arg\max_{i\in[N]}Y_{i}$.
Assumption 4 (Small-step regime and limit order).
We take limits with $N\to\infty$ first, then $\varepsilon\downarrow 0$.
Lemma 1 (Gaussian maxima and spacing (vershynin2018high)).
Let $\xi_{1},\dots,\xi_{N}$ be i.i.d. mean-zero, unit-variance Gaussian, $M_{N}=\max_{i\leq N}\xi_{i}$, $S_{N}$ the second largest, and $q_{N}:=F^{-1}(1-1/N)$ the (11/N)(1-1/N)-quantile of $\xi_{1}$. Then $q_{N}=\Theta(\sqrt{\ln N})$, $M_{N}-q_{N}\to 0$ in probability, and $M_{N}-S_{N}=O_{p}(1/q_{N})$.
Theorem 2 (Best-of-$N$ asymptotically induces a UCB gradient direction).
Under Assumptions 1-4, define $\beta_{N}:=q_{N}$ from Lemma 1. Let $i^{\star}\in\arg\max_{i}Y_{i}$ and set $\widehat{u}_{N}:=u_{i^{\star}}$. Define the step
$\Delta x^{(N)}\;:=\;\mathsf{E}\!\big(\mathrm{Apply}(\Pi,\varepsilon\widehat{u}_{N})\big)-x\;=\;\varepsilon\widehat{u}_{N}+r(\varepsilon,\widehat{u}_{N}).$ (13)
Then, with the limit order $N\to\infty$ first and $\varepsilon\downarrow 0$ second,
$\frac{\Delta x^{(N)}}{\|\Delta x^{(N)}\|}\;=\;\widehat{u}_{N}\;\xrightarrow{p}\;\frac{\nabla A_{\beta_{N}}(x)}{\|\nabla A_{\beta_{N}}(x)\|}\;=\;\frac{g+\beta_{N}h}{\|g+\beta_{N}h\|}.$ (14)
Moreover,
$A_{\beta_{N}}(x+\Delta x^{(N)})\;\geq\;A_{\beta_{N}}(x)\;+\;\varepsilon\,\|\nabla A_{\beta_{N}}(x)\|\;-\;o_{p}(\varepsilon).$ (15)
7 Case Study: Ad Optimization in Digital Marketing
[…]
Figure 6: An ad generation AI system $\Phi$ and prompt $\Pi=\langle\pi\rangle$.
[…]
Figure 7: A self-improving AI implemented on the ad generation AI system in Figure 6.
In this section, we present a case study on how ad platforms can use T-BoN BO for ad optimization and automate the ad generation, analysis, and evaluation cycle. Given an a priori fine-tuned image generation AI model that is aligned with its client’s brand, the objective is to efficiently refine the prompts for ad generation to better align with the users’ preference distribution.
7.1 Problem definition
Formally, define the brand fine-tuned ad-generation AI system as $\Phi$, whose behavior is controlled by a prompt set $\Pi$ (Figure 6). For a given prompt set $\Pi$, we write $\Phi(\Pi)$ for the corresponding ad creative shown on the ad platform. As a system, $\Phi(\Pi)$ maps the input space $\mathcal{X}$ (potential ad viewers on the platform) to a distribution over the output space $\mathcal{Y}$ (observable ad responses). For each $x\in\mathcal{X}$, we denote by $y(\Phi(\Pi),x)$ the random $\mathcal{Y}$-valued output of the system at input $x$. In a digital marketing setting, $x$ is a platform user, and a realization of $y(\Phi(\Pi),x)$ represents that user’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\Pi)$. When a user $x\in\mathcal{X}$ is shown the ad creative $\Phi(\Pi)$, we observe an outcome $y\in\mathcal{Y}$, which is a realization of $y(\Phi(\Pi),x)$.
As discussed earlier, the distribution of exposed users depends on the ad $\Phi(\Pi)$ and therefore on the prompt set $\Pi$. We denote by $\mathcal{D}_{\mathcal{X}}(\Pi)$ the unknown distribution of users $x$ who are shown the ad generated from $\Pi$ when the ad $\Phi(\Pi)$ is run on the platform.
To evaluate system performance, a marketer pre-defines a scoring rule $r:\mathcal{Y}\to[0,1]$ that maps observed user responses (clicks, conversions, spend) to a normalized effectiveness score. The ad optimization objective of the agency for a given client and campaign can then be written as
$J(\Pi)\;:=\;\mathbb{E}\big[\,r\big(y(\Phi(\Pi),x)\big)\big],$ (16)
where the expectation is taken over $x\sim\mathcal{D}_{\mathcal{X}}(\Pi)$ and any additional randomness in $y(\Phi(\Pi),x)$.
The self–improving ad system built by the agency can therefore be viewed as an iterative algorithm that searches over prompts $\pi$ in order to maximize $J(\pi)$ while minimizing the number of costly ad evaluations on the platform (Figure 7). At iteration $t$, a candidate prompt $\pi_{t}$ is selected, the corresponding ad creative $\Phi(\pi_{t})$ is generated, and the campaign is (conceptually) exposed to a batch of users $\{x_{t}^{(l)}\}_{l=1}^{L}\sim\mathcal{D}_{\mathcal{X}}(\pi_{t})$. The empirical evaluation
$s_{t}\;:=\;\frac{1}{L}\sum_{l=1}^{L}r\big(y(\Phi(\pi_{t}),x_{t}^{(l)})\big)$ (17)
serves as a noisy estimate of $J(\pi_{t})$ and is fed back into the optimization algorithm (T-BoN BO) to update the history $H_{t}$ and the meta–reflection $R_{t}$.
7.2 Application of T-BoN BO
We now describe how the general T-BoN BO framework in Section 5 is instantiated for ad agency’s ad optimization problem. Specifically, we focus on the parallel version of T-BoN BO (Algorithm 2), where $J$ optimization trajectories start from different initial prompts and run simultaneously.
Initialization.
The decision-maker/platform starts from initial prompts $\{\pi_{0}^{j}\}_{j=1}^{J}$, i.e., prompts already in use or high-quality baselines provided by humans. (e.g., “A simple, photorealistic image of a plant-based burger patty. [^9] Note that in the experiments in Section 8, we use Worst5-of-N prompts as initial starting points, which are higher-quality, much more detailed initial prompts; See Appendix A.) For each initial prompt $\pi_{0}^{j}$, the system generates an ad $\Phi(\pi_{0}^{j})$ and obtains an initial score $s_{0}^{j}$ by deploying the ad and observing outcomes. These tuples $\{(\pi_{0}^{j},\Phi(\pi_{0}^{j}),s_{0}^{j})\}_{j=1}^{J}$ form the initial history $H_{0}$ for the parallel T-BoN BO procedure.
Best-of-NN textual-gradient steps.
Within each optimization iteration $t$, T-BoN BO refines each trajectory’s prompt $\pi_{t-1}^{j}$ through $G$ Best-of-N textual-gradient steps (Figure 4). For a given trajectory $j$ and inner step $g$, the critic model $M_{\text{critic}}$ (a large language model such as Gemini 2.5 Flash) receives as context: (i) the current prompt $\pi_{t,g-1}^{j}$, (ii) the current meta-reflection $R_{t-1}$, and (iii) a short description of the brand and campaign scenario. Conditioned on this context, it proposes $N$ small natural-language edits
$\delta_{t,g}^{j,(1)},\ldots,\delta_{t,g}^{j,(N)}\;\sim\;\nabla_{\text{text}}(\pi_{t,g-1}^{j};R_{t-1},M_{\text{critic}}),$ (18)
which form the textual gradients. Each edit is then applied to the current prompt to obtain $N$ candidate prompts
$\pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}\big(\pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)}\big),\qquad i=1,\dots,N.$ (19)
These textual gradients correspond to small, directed moves in language space that, according to the critic’s judgment, are likely to increase ad effectiveness. For example, conditioned on the reflection $R_{t-1}$ (e.g., “social occasions and visible grill marks tend to perform better than plain pack shots”), the critic $M_{\text{critic}}$ may propose the following textual-gradient candidates:
  1. 1.
    Emphasize social context and enjoyment: “Show the burger being enjoyed at an outdoor barbecue with friends, not alone on a plate.”
  2. 2.
    Highlight sensory appeal and indulgence: “Make the patty look extra juicy with clear grill marks, a toasted bun, and melty toppings.”
  3. 3.
    Make sustainability salient but secondary: “Subtly include eco-friendly cues (like a small ‘plant-based’ tag or greenery in the background) without overpowering the food.”
Each of these textual gradients $\delta^{(i)}$ is then applied to the original prompt to produce candidate prompts. For example, applying the first gradient yields the prompt:
‘‘A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends, everyone smiling and reaching for food.’’
Applying the second gradient yields the prompt:
‘‘A close-up, photorealistic shot of a plant-based burger with deep grill marks, a toasted brioche bun, melty vegan cheese, fresh toppings, and steam rising to suggest warmth and juiciness.’’
To select the next prompt that will go forward to the next gradient step among $N$ candidate prompts $\{\pi_{t,g}^{j,(i)}\}_{i=1}^{N}$, T-BoN BO applies the Best-of-$N$ choice rule over the $N$ corresponding candidate ads $\{c_{t,g}^{j,(i)}\}_{i=1}^{N}$. Following liu2025pairjudgermperformbestofn, we implement Best-of-$N$ via a pairwise tournament: in each match, the critic model $M_{\text{critic}}$ compares two creatives side by side and predicts which will perform better, given the campaign goals and the current meta-reflection $R_{t-1}$ (see Figure 12 in Appendix A). After multiple randomized pairwise comparisons, the winner $c_{t,g}^{j}$ and its corresponding prompt $\pi_{t,g}^{j}$ are carried forward to the next gradient step. Repeating this expansion-and-selection process for $g=1,\dots,G$ yields the refined prompt $\pi_{t}^{j}$ and ad $\Phi(\pi_{t}^{j})$ for trajectory $j$ at iteration $t$.
Evaluation as the costly step.
Once all $G$ gradient steps are completed, T-BoN BO performs a single expensive evaluation for each trajectory by deploying the ad $\Phi(\pi_{t}^{j})$ and computing the score $s_{t}^{j}$. In the real ad-agency use case, this evaluation step would correspond to running the creative in an A/B test, multi-armed bandit experiment, or survey-based study, and mapping the resulting performance metrics to a scalar score via the scoring rule $r$. The new triple $(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})$ is appended to the history $H_{t}$, and a simple acceptance rule is applied: if $s_{t}^{j}$ fails to improve upon $s_{t-1}^{j}$, the trajectory rolls back to $(\pi_{t-1}^{j},\Phi(\pi_{t-1}^{j}),s_{t-1}^{j})$.
Meta-reflection shared across trajectories.
After all $J$ trajectories have been evaluated at iteration $t$, the algorithm updates the global meta-reflection $R_{t}=\textsc{MetaReflect}(H_{t})$, as in Section 5.3. Concretely, $M_{\text{critic}}$ is prompted to review a subset of the best- and worst-performing ads and prompts in $H_{t}$ and to summarize the key patterns and rules that distinguish successful creatives from unsuccessful ones (see Figure 13 in Appendix A). For example, the reflection may emphasize that scenes with social eating and visible grill marks perform better than isolated product shots, or that emphasizing “satisfying” and “juicy” yields higher effectiveness scores. This reflection is then injected into the context of subsequent textual-gradient generation and Best-of-$N$ comparisons, allowing all trajectories to share accumulated knowledge even though they explore different regions of the prompt space.
Reported outcome.
After $T$ iterations, the system returns the sequence of best-performing ads $\{c_{t}^{j_{t}^{\ast}}\}_{t=1}^{T}$, where $j_{t}^{\ast}=\arg\max_{j}s_{t}^{j}$ is the trajectory that achieves the highest score at iteration $t$.
8 Experiments
We validate the empirical effectiveness of the proposed T-BoN BO framework for a real-world digital marketing problem described in 7. Specifically, we examine if and how much T-BoN BO improves outcomes under a fixed evaluation budget compared to state-of-the-art baselines.
8.1 Evaluation Module using Persona data and LLM-based Simulation
To evaluate how well an ad performs on the target population and improve prompts, we need a target population and a way to measure the preferences of that population for a given ad. There are many ways to do this, such as field experiments, consumer surveys, or simulated responses from a pre-defined target population. In our experiments, we use LLM-generated preferences over the Twin-2k-500 persona distribution provided by toubia2025twin. An important advantage of using this persona distribution is that it allows us to test the performance of different algorithms fast without the need for large-scale field experiments or expensive surveys. [^10] Note that naive field deployment cannot be used to test the relative performance of algorithms since an algorithm’s measured advantage would be a mix of its own effect and distributional effects induced by the ad platform’s ad targeting systems.
This dataset includes the data of 2,058 U.S. adults answering 500 survey questions spanning demographics, personality, cognitive ability, economic preferences, classic heuristics-and-biases tasks, and a pricing study. These survey questions of each persona are passed to a multi-modal LLM (Gemini 2.5 Flash) as context so that the LLM can simulate the ad effectiveness for the persona. [^11] While LLM-based persona simulation is often biased and does not reflect real-world human behavior (li2025llm; peng2025mega), it still induces a valid preference distribution for which, as a benchmark, we would like to see how well a method can optimize. We divided the Twin-2k-500 personas into 80% (1647 personas) for the training set and 20% (411 personas) for the test set. Every time we evaluate an ad during the algorithm (i.e., training stage), we randomly sample 200 personas from the training set and simulate the ad’s effectiveness for them. Every time we want to test an ad, such as a final ad output from the algorithm, we simulate its effectiveness for all the test set personas.
8.2 Experiment Setup
We consider eight synthetic ad campaign scenarios that cover a diverse range of products across distinct categories, defined by the creative brief that outlines the strategic and creative direction for each scenario. (For the creative brief for each scenario, see Appendix B.)
  • Scenario 1: “GreenBite,” a new plant-based burger patty.
  • Scenario 2: “AuraSonics X1,” high-end, noise-canceling wireless earbuds.
  • Scenario 3: “Odyssey E-SUV,” a new all-electric family SUV.
  • Scenario 4: “Oasis Eco-Lodge,” a secluded, luxury resort with beautiful natural surroundings.
  • Scenario 5: “Momentum,” a mobile-first banking app for freelancers and the gig economy.
  • Scenario 6: “MindGarden," a subscription-based meditation and mindfulness app.
  • Scenario 7: “Aeterno,” a classic, automatic Swiss-made wristwatch with a heritage design.
  • Scenario 8: “SyncFlow,” a project management and collaboration software platform for remote teams.
For each scenario, from its creative brief, we generated 64 high-quality initial prompts that were used as input to the state-of-the-art image generation AI model named Imagen 4  [^12] See Vertex AI’s description (https://console.cloud.google.com/vertex-ai/publishers/google/model-garden/imagen-4.0-generate-preview-06-06) for details. to generate the initial 64 high-quality ads. (For prompts, see Figure 14 and 15 in Appendix A.) To maintain a rigorous baseline, we made the 64 initial prompts detailed enough to yield strong ads. This ensures that the algorithm’s subsequent improvements are not explained as trivial quality gains from having initial under-specified prompts. That is, the measured improvements are attributed to the algorithm’s ability to align ads with the underlying persona distribution better.
When simulating an ad’s effectiveness for a persona, we ask the LLM (with the prompt in Figure 8) what the ad’s “effectiveness” [^13] The vagueness of the term “effectiveness” constitutes an evaluation distribution induced by LLM-persona based on what LLM ‘thinks’ about the term, which we target to align the algorithms with. That is, as long as we consistently query the same LLM to predict “effectiveness” using the same prompt, such vagueness is justified. for the persona will be, with the score of [1, 2, 3, 4, 5]. Here, the score of one is the least effective, and the score of five is the most effective. As the LLM’s internal assignment of probability for the scores is available as ‘logprobs’ [^14] See Vertex AI’s logprobs description (https://cloud.google.com/vertex-ai/generative-ai/docs/model-reference/inference#logprobs) for details., we use them to calculate the mean score, and consider it as the evaluation score of the ad for the persona (which is a fraction number like 3.41), instead of considering the actual response score from the LLM (which is an integer number among 1-5). We then take the average of the evaluation scores for the personas to get the final evaluation score of the ad.
Persona prompt for simulating ad effectiveness. PERSONA DATA: $\{$
Which part of the United States do you currently live in? [Answer]
What is the highest level of schooling or degree that you have completed? [Answer]
$\ldots$ (Many other survey questions and answers) $\ldots$
[…]
$\ldots$ (Many other survey questions and answers) $\ldots$
AD IMAGE: [image] TASK: Return only one item from ["1","2","3","4","5"] for ad effectiveness.
Effective Score Scale Definition:
1: Extremely Unlikely. The persona would actively ignore or be annoyed by this ad.
2: Unlikely. The persona would likely scroll past without a second thought.
3: Mediocre. It is hard to decide whether the personal would click or don’t click.
4: Likely. The persona is intrigued and has a good chance of clicking to learn more.
5: Extremely Likely. The persona is the ideal target; a click is almost certain.
No explanation. Just the score.
Figure 8: Persona simulation prompt for the digital marketing problem.
Among the initial ads, we identified the best initial ad, which we call Best-of-64, and the worst five initial ads, which we call Worst5-of-64, by running a best-arm identification bandit algorithm (russo2016simple) by alternating the best-arm identification objective and worst 5-arm identification objective for evaluations of 5,000 sequential random samples from the training persona set.
8.3 Baselines
As baselines, we consider the performance of two popular state-of-the-art algorithms. One is the Best-of-64 (snell2024scaling) baseline, which is one of the most popular and well-performing test-time alignment methods that has been reported to outperform reinforcement learning fine-tuning methods under large enough $N$ (gao2023scaling; mudgal2023controlled; eisenstein2023helping; beirami2024theoretical). Specifically, the test persona set-based evaluation result for the Best-of-64 ad (identified above in Section 8.2 using training persona set) was used as the Best-of-64 baseline. We also evaluated the ads in Worst5-of-64 in the same way and took the mean of their scores to obtain the Worst5-of-64 baseline.
Another baseline is GEPA (Genetic-Pareto) (agrawal2025gepa). As discussed earlier, this reflective and evolutionary prompt evolution algorithm has recently demonstrated superior performance compared to GRPO (liu2024deepseek), a state-of-the-art reinforcement fine-tuning technique. GEPA keeps a set of candidate prompts, and samples a prompt to mutate based on its evaluation result for a problem (in our case, a persona) set they call the Pareto set. A Pareto set stores ‘important’ problems. Since we randomly sample personas from the training persona set, we cannot store specific personas and retrieve them later. Therefore, we instead use the version of GEPA that 1) randomly samples a persona and 2) mutates all candidate prompts equally and throws away the old prompts to avoid unbounded growth of the candidate set.
For the experiment, we run five-trajectory ($J=5$) parallel T-BoN BO and GEPA starting from Worst5-of-64, i.e., the worst five initial prompts we identified in Section 8.2. This starting point allows us to attribute observed gains to the algorithms rather than to favorable starting prompts. We test if T-BoN BO outperforms the Best-of-N baseline and GEPA baseline throughout the optimization process. For both T-BoN BO and GEPA, we ran 10 optimization steps, evaluating 200 randomly sampled personas per step (2,000 total persona evaluations). We used a multi-modal LLM (Gemini-2.5-Flash) to serve as the critic model for both T-BoN BO and GEPA.
8.4 Results and Analysis
Trend analysis
Figure 9: Progress comparison of T-BoN BO and GEPA with Best-of-64 baseline. T-BoN BO implements parallel T-BoN BO with $J=5$. The performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the testing persona set. Since it is the mean of means, we don’t plot the standard errors.
Figure 9 plots how much T-BoN BO, GEPA, and Best-of-N improved their performance from the initial Worst5-of-N score across eight scenarios. (For the detailed data used to generate this plot, refer to Table 2 and 3 in Appendix C.) At each step (the $x$-axis), the performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the testing persona set. For example, the step-2 score is the average of the eight step-2 mean evaluation scores. As the figure illustrates, both T-BoN BO and GEPA improve consistently across steps, with T-BoN BO demonstrating faster early gains and achieving a significantly better final performance than GEPA. Notably, by the third step on average, which is equivalent to 600 persona evaluations, T-BoN BO and GEPA both outperform the Best-of-64 baseline.
Statistical analysis
Model
$\text{score}=\alpha+\beta_{1}\mathbf{1}\{t=10\}+\beta_{2}\mathbf{1}\{\mathrm{TBoN}\}\,\mathbf{1}\{t=10\}+\text{persona FE}+\text{scenario FE}+\varepsilon$
Effect Estimate Std. err T-stat
$\beta_{1}$: GEPA gain $(t=10)$ from Worst5-of-64 0.3168*** 0.0325 9.752
$\beta_{2}$: T-BoN BO gain - GEPA gain $(t=10)$ from Worst5-of-64 0.0743*** 0.0248 2.991
$\beta_{1}+\beta_{2}$: T-BoN BO gain $(t=10)$ from Worst5-of-64 0.3911*** 0.0355 11.011
Summary statistics
Observations 13,152
Personas 411
Scenarios 8
  • Notes: Two-way cluster-robust standard errors.
  • * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
Table 1: Two-way fixed effect (TWFE) analysis result with persona and scenario fixed effects with two-way clustered SEs (cameron2011robust). Standard error of $\beta_{1}+\beta_{2}$ was computed using Delta method.
In addition to the aggregate trend analysis provided in Figure 9, we now provide statistical evidence that T-BoN BO indeed outperforms the GEPA. For this, we conducted a Two-way Fixed Effect (TWFE) analysis using T-BoN BO and GEPA’s final ($t=10$) scores to compare the effectiveness of each method rigorously. Specifically, we consider the following TWFE model:
$\text{score}=\alpha+\beta_{1}\mathbf{1}\{t=10\}+\beta_{2}\mathbf{1}\{\mathrm{TBoN}\}\,\mathbf{1}\{t=10\}+\text{persona FE}+\text{scenario FE}+\varepsilon.$ (20)
where $\mathbf{1}\{t=10\}$ is the indicator variable of whether it is $t=0$ (GEPA and T-BoN BO’s initial starting point, i.e., Worst5-of-64 score) or $t=10$ (GEPA and T-BoN BO’s final score), $\beta_{1}$ is the GEPA’s gain from Worst5-of-64 after $t=10$, $\beta_{2}$ is the T-BoN BO’s advantage over GEPA after $t=10$, and $\beta_{1}+\beta_{2}$ is the T-BoN BO’s gain from Worst5-of-64 after $t=10$. The unobserved residual $\varepsilon$ is assumed to follow the normal distribution. To remove potential correlation among unobserved residuals within the same persona or scenario, we conducted a cluster-robust Two-way Fixed Effect (TWFE) analysis with two-way clustering (cameron2011robust; cameron2015practitioner). The result of TWFE analysis is shown in Table 1. As we can see, both of T-BoN BO’s gain and GEPA ’s gain from the starting point (Worst5-of-64) after $T=10$ are statistically significant. In addition, we conclude that T-BoN BO has a significant 23.5% more gain in score compared to GEPA’s gain.
Qualitative analysis
The significant gains of GEPA and T-BoN BO can be interpreted as demonstrating their ability to align with the persona distribution. However, such significance can also be easily gained by starting from a poor starting point, i.e., a poor Worst5-of-N prompt and corresponding ad image. One way to check whether this happens is to compare the quality of the prompts; we provide them in Appendix A to demonstrate that this is not the case. Another qualitative approach is to compare the generated ad images and verify whether the differences among ads are in quality or in the message and representation. Figure 10 compares the ad images of the baseline methods. As we can see, there is no clear visual quality difference among those ad images; the differences are mainly in the ad’s message and how it is represented. This supports the claim that T-BoN BO and GEPA improvements from Worst5-of-N demonstrated in Figure 9 are attributed to alignment, not trivial quality gains from correcting poor or under-specified prompts.
GreenBite (Plant-based patty) MindGarden (Meditation app) AuraSonics X1 (Noise-canceling earbuds)
Worst5-of-64 (starting point) Refer to caption Score: 2.672 Refer to caption Score: 2.670 Refer to caption Score: 2.765
Best-of-64 Refer to caption Score: 2.898 Refer to caption Score: 2.930 Refer to caption Score: 2.934
GEPA (after T=10) Refer to caption Score: 3.046 Refer to caption Score: 3.068 Refer to caption Score: 3.028
T-BoN BO (after T=10) Refer to caption Score: 3.168 Refer to caption Score: 3.132 Refer to caption Score: 3.104
Figure 10: T-BoN BO and baseline’s generated Ad images for three fictional brands: GreenBite burger patty (Scenario 1), MindGarden meditation app (Scenario 6), and AuraSonics noise-canceling earbuds (Scenario 2). For the Ad images of the other five brand scenarios, see Figure 18 in Appendix C.
Figure 11: An ablation study experimenting T-BoN BO with $J=5$ and Best-of-64 for Gemini 2.5 Flash instead of persona distribution with temperature 0. (GEPA was excluded, as there is no problem-dependent heterogeneity.) The performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the testing persona set. Again, since it is the mean of means, we don’t plot the standard deviations.
8.5 Ablation study
The main experiment in Section 8.1-8.4 evaluates the candidate ad at each iteration as the mean of scores across 411 personas (20% of the persona data in Twin-2k-500 (toubia2025twin)) that were injected into the LLM’s context. Each persona consists of a large amount of information, with 500 survey questions accounting for around 50,00050{,}000 tokens. Would T-BoN BO still outperform strong state-of-the-art baselines when the LLM receives less persona context information?
To test this, we ran a no-persona ablation experiment in which the evaluator is a single LLM (Gemini 2.5 Flash) with no auxiliary persona context. Concretely, at each iteration, we present only the ad image candidate to LLM judge (Gemini 2.5 Flash) together with the same 1–5 effectiveness rubric, set temperature to 0, and compute the same scalar score as the log-probability–weighted expectation over $\{1,\dots,5\}$ as before. [^15] With temperature =0=0, the judge’s output is deterministic for a fixed image and rubric; the logprob vector is still available, so the expected score is stable across repeated calls. This collapses the evaluation to a deterministic mapping and removes heterogeneity from personas and sampling, allowing us to test whether T-BoN BO’s performance depends on specific persona information-related setup. To ensure there is no leakage of preference, we used GPT-5 instead of Gemini 2.5 Flash for generating meta-reflection, generating textual gradients, and making Best-of-N choices.
We again compare T-BoN BO (J=5J{=}5 trajectories) against the Best-of-64 initial baseline; we exclude GEPA, which requires multiple personas (multiple problems, in GEPA’s language.)
Figure 11 plots T-BoN BO across 10 optimization steps starting from the Worst5-of-64 prompts under the ablation setup. As in Figure 9, the main experiment’s results, the value shown in Figure 11 is the mean of that step’s scores over eight scenarios. As can see in Figure 11, T-BoN BO again outperforms Best-of-64 baseline from optimization step 3. Except for the fact that in general the evaluation scores without persona information are more optimistic than scores with persona information for both T-BoN BO and Best-of-64, the ablation experiment’s trend in Figure 11 shows the same trend we saw in Figure 9.
9 Conclusion
The importance of developing evaluation-efficient self-improving AI systems is substantial for its societal applications, yet this problem has remained unidentified and unaddressed. Motivated by the effectiveness of prompt optimization, which outperforms reinforcement fine-tuning methods, we formulate this problem as an iterative prompt optimization and address it.
We propose T-BoN BO, a simple, evaluation-efficient prompt optimization framework that leverages textual edits and Best-of-N search to guide prompt improvement. We prove that T-BoN BO BO statistically emulates UCB-BO, which is a provably evaluation-efficient algorithm. Experiments simulating ad alignment across eight creative-ad tasks demonstrate that T-BoN BO outperforms strong state-of-the-art baselines, including large Best-of-N and GEPA. We validate that these improvements are not attributed to the usage of specific persona dataset or trivial improvements from using poor initial prompts.
Our paper makes three key contributions to the computer science and marketing literature. First, it identifies evaluation efficiency as an essential objective for a societal self-improving system design. Second, it presents a simple yet provably efficient AI self-improving system design framework that can be applied to real-world societal systems, e.g., ad agencies, that incur high evaluation costs. Third, we theoretically prove that the presented framework is evaluation-efficient.
In sum, this work bridges classical Bayesian optimization and prompt-based self-improving AI systems, demonstrating that evaluation-efficient optimization in the prompt space is both theoretically tractable and practically impactful. As organizations increasingly rely on AI-generated content, methods like T-BoN BO can help managers achieve higher-quality outputs without proportionally increasing evaluation burdens.
Appendix A Prompts used for digital marketing
Best-of-N Gradient steps’ pairwise tournament prompt for ads. contents = [image1, image2, comparison_\_prompt]
comparison_\_prompt =
“You are evaluating two advertisement images for mobile Instagram ads. Which image would be more effective at engaging users and driving clicks?
RESPOND WITH ONLY THE NUMBER 1 OR 2. NOTHING ELSE.
1 = first image is better
2 = second image is better ”
Figure 12: $M_{\text{critic}}$’s pairwise tournament prompt for digital marketing example.
Meta reflection prompt for ads You are an expert at analyzing visual patterns in advertising performance. VISUAL ANALYSIS TASK:
I will show you images from the lowest-scoring and highest-scoring ad iterations. Your task is to identify specific visual patterns that distinguish effective from ineffective ads.
VISUAL EXAMPLES – BEST VS WORST PERFORMING:
RANK 1/10 WORST PERFORMING (Score: 2.14/5.0):
Prompt excerpt: <prompt>…
[Image]
RANK 2/10 LOWER HALF (Score: 2.52/5.0):
Prompt excerpt: <prompt>…
[Image]
$\ldots$ RANK 9/10 UPPER HALF (Score: 3.81/5.0):
Prompt excerpt: <prompt>…
[Image]
RANK 10/10 BEST PERFORMING (Score: 4.15/5.0):
Prompt excerpt: <prompt>…
[Image]
Based on your visual analysis, identify patterns that correlate with higher effectiveness scores: 1. Visual composition and framing differences 2. Lighting conditions and mood variations 3. Color palettes and visual tone patterns 4. Subject positioning and action effectiveness 5. Brand integration approaches 6. Environmental and atmospheric elements RESPONSE FORMAT:
Provide a structured analysis of visual patterns observed, focusing on what distinguishes high-performing from low-performing ads.
Figure 13: Meta reflection prompt for digital marketing example.
Prompt used for generating prompts for ads generation Generate a creative, structured and descriptive prompt for a generative AI model (e.g., Imagen) that will produce a brand-aligned, and scroll-stopping advertisement image suitable for an Instagram feed.
A successful prompt must be constructed using the following components and principles:
0. (Most important!) Prompt NEVER includes any description of kids. Kids are not allowed to generate in Imagen 4. Even if creative_brief includes description of "family" or "kids", never include kids in the prompt. Only create adults if humans are included.
1. Key Message (The "Why"): This is the foundational element that guides all other components. It defines the core idea or feeling the ad must communicate. Before writing the rest of the prompt, clearly articulate the message the ad will deliver. This message will act as the "North Star" for all subsequent creative choices. It may be desirable to put the message as the text overlay.
2. Core Components (The "What"):
- This should completely depend on the key message.
- Scene & Environment: Based on the key message, establish a clear, relatable setting that aligns with the brand’s lifestyle appeal.
- Action & Narrative: Based on the key message and the scene, describe a dynamic but clear action or interaction to create a sense of a captured moment and tell a micro-story.
- Composition & Framing: Specify the camera shot, angle, and framing (e.g., "low-angle shot," "dynamic medium shot," "close-up on the shoe").
3. Stylistic Qualities (The "How"):
- This should completely depend on the key message and the scene.
- Photography Style: Define the overall aesthetic (e.g., "photorealistic," "cinematic," "professional product photography," "lifestyle action shot").
- Lighting: Be specific about the lighting to set the mood (e.g., "warm golden hour light," "bright morning sun," "dramatic side-lighting").
- Color Palette & Tone: Guide the color scheme and emotional feel (e.g., "vibrant and energetic colors," "empowering and motivational tones," "clean and modern palette").
- Atmosphere & Feeling: Aim to evoke a specific feeling aligned with the brand (e.g., "a feeling of effortless performance," "an atmosphere of vibrant energy," "a sense of supreme comfort").
4. What to Avoid:
- Vagueness: Use specific, descriptive terms instead of "nice" or "good."
- Contradictory Elements: Ensure all elements work together harmoniously.
- Over-stuffing: Focus on a single, clear message without too many competing objects.
5. What to include:
- Logo: create a logo based on creative brief, and naturally place it.
- Text: If you are including a text message, prompt for "negative space for text overlay".
Figure 14: Prompt for generating the prompts for the initial 64 ads.
Prompt for generating the worst-of-64 in GreenBite (plant-based burger patty) ad scenario. **Key Message:** To visually communicate the core idea: "Experience the juicy, no-compromise satisfaction of a real burger, made better for you and the planet."
**Image Generation Prompt:** A dynamic, commercial lifestyle photograph for an Instagram feed ad.
**Core Components:**
* **Scene & Environment:** A stylish, modern outdoor patio during a late afternoon barbecue. The background is softly blurred, showing contemporary patio furniture, lush green potted plants, and the warm glow of string lights just beginning to turn on.
* **Subject & Action:** The focal point is a vibrant, stylish woman in her early thirties, captured in a candid moment of pure satisfaction. She is taking a bite of a huge, delicious-looking plant-based burger. Her eyes are closed in bliss, and a small, genuine smile plays on her lips, conveying an authentic "this is amazing" reaction. Juice from the burger is subtly visible on her hand to emphasize its succulence. She is dressed in a casual but chic linen shirt.
* **Product:** The GreenBite burger is the hero. It’s a thick, perfectly seared patty with grill marks, nestled in a glossy brioche bun. It’s layered with glistening melted cheddar cheese, crisp green lettuce, a thick slice of ripe red tomato, and caramelized onions spilling out slightly. The burger is partially wrapped in a simple, brown kraft paper sheet which has a small, clean logo on it: "GreenBite" in a modern sans-serif font next to a stylized green leaf.
* **Composition & Framing:** A medium close-up shot, angled slightly upwards to make the moment feel triumphant and empowering. The camera uses a shallow depth of field, keeping the woman’s expression and the burger in sharp focus while the background is beautifully out of focus (bokeh). The composition is balanced, with a clear area of soft background on the upper right side, creating negative space for a text overlay.
**Stylistic Qualities:**
* **Photography Style:** High-end, photorealistic commercial food photography with a warm, authentic lifestyle feel. The image should be incredibly sharp and detailed, especially the texture of the burger patty and the fresh ingredients.
* **Lighting:** Bathed in the warm, dreamy glow of the golden hour. The sunlight comes from the side, creating soft highlights and long, gentle shadows, emphasizing the textures of the food and giving her skin a radiant look.
* **Color Palette & Tone:** A rich and mouth-watering color palette. Saturated and vibrant, but natural. Warm tones of orange, gold, and red from the sun and the food are balanced by the deep greens of the background foliage and the crispness of the lettuce.
* **Atmosphere & Feeling:** The image radiates a feeling of guilt-free indulgence, vibrant well-being, and ultimate satisfaction. The atmosphere is confident, positive, and irresistibly appetizing, making the viewer crave that exact experience.
Prompt for generating the best-of-64 in GreenBite (plant-based burger patty) ad scenario. **1. Key Message (The "Why"):** To visually capture the peak moment of guilt-free indulgence, proving that a plant-based burger can deliver an unbelievably juicy, satisfying, and no-compromise flavor experience identical to a classic beef burger. The text overlay will read: "The No-Compromise Burger."
**2. Core Components (The "What"):**
* **Scene & Environment:** A vibrant, sun-drenched backyard barbecue scene, shot with a very shallow depth of field. The background is a beautiful, soft-focus bokeh of friends laughing (adults only, 25-40 age range), string lights, and lush greenery, suggesting a warm, sociable atmosphere without distracting from the main subject.
* **Action & Narrative:** A stylish woman in her early 30s, with a look of pure, blissful satisfaction, is taking the first, epic bite of a burger. Her eyes are closed in enjoyment. Her hands grip the burger firmly but gently. A single, perfect drop of sauce is starting to fall from the side of the burger, caught in mid-air, emphasizing its juiciness.
* **Composition & Framing:** A dramatic, macro close-up, food photography style. The shot is framed tightly on the lower half of the woman’s face and her hands holding the burger, making the burger the undeniable hero. The angle is slightly low to give the burger an epic, monumental feel. The right side of the frame has clean, out-of-focus background elements, creating intentional negative space for a text overlay.
**3. Stylistic Qualities (The "How"):**
* **Photography Style:** Hyper-realistic, professional commercial food photography. Every detail is rendered with extreme clarity: the char marks on the patty, the condensation on the lettuce, the glistening melt of the plant-based cheese, and the texture of the toasted brioche bun. Shot with a high-end DSLR camera and a macro lens.
* **Lighting:** Warm, radiant golden hour lighting. The sun acts as a dramatic backlight, creating a halo effect around the burger and the woman’s hands, and creating glistening specular highlights on the juicy patty and sauce. The lighting is bright, positive, and makes the food look incredibly appetizing.
* **Color Palette & Tone:** A mouth-watering and vibrant palette. Rich, savory browns of the seared "GreenBite" patty, deep reds of a ripe tomato slice, vibrant green of crisp lettuce, and a warm golden-yellow of the melted cheese and toasted bun. The overall tone is confident, delicious, and energetic.
* **Atmosphere & Feeling:** The atmosphere is one of ultimate satisfaction and pure, unadulterated foodie bliss. It feels warm, authentic, and completely satisfying, eliminating any doubt about the quality and taste of the plant-based option.
**4. Inclusions & Details:**
* **The Burger Details:** The GreenBite patty is thick, with a perfect, glistening sear and visible char marks, looking indistinguishable from high-quality ground beef. It’s topped with melting vegan cheddar cheese, a leaf of crisp butter lettuce, a thick slice of heirloom tomato, and a creamy aioli sauce.
* **Logo Placement:** The burger is partially wrapped at the base with a small square of branded, unbleached butcher paper. The "GreenBite" logo—a simple, modern green leaf icon next to the word "GreenBite" in a clean sans-serif font—is subtly visible on the paper.
* **Negative Space:** Ensure the composition leaves a clean, uncluttered area of soft-focus background on the right third of the image for text overlay.
Figure 15: Prompts for the best and worst ad among the initial 64 ads.
Prompt for generating the T-BoN $(T=10)$ ad for GreenBite (plant-based burger patty) "**Key Message:**
"The No-Compromise Burger"—to visually prove that a plant-based burger can be just as juicy, delicious, and satisfying as a traditional beef burger, making it the hero of the ultimate backyard BBQ.
*Core Components:**
* **Scene & Environment:** A vibrant, meticulously designed modern backyard patio, perfect for a relaxed, aspirational late-afternoon get-together. The background is a soft-focus composition of stylish outdoor furniture, lush green garden plants, and subtly visible, engaged friends enjoying themselves and partaking in the shared experience. Occasional hints of other complementary gourmet side dishes or refreshing beverages on the blurred patio table enhance the sense of a complete, desirable BBQ experience, creating a rich, warm, and inviting social atmosphere.
* **Action & Narrative:** A pair of adult hands are holding up a perfectly constructed, gourmet-style cheeseburger. The action captures either: (A) a moment of serene, authentic anticipation just before the first bite, with the subject’s face (in soft focus) conveying genuine contentment or subtle bliss, possibly with eyes gently closed, savoring the aroma; OR (B) the burger held out slightly towards the camera, inviting the viewer into the experience, with the subject’s soft-focus expression conveying a shared, authentic delight or playful ’aha!’ moment. The emphasis is on genuine human emotion and connection, avoiding any forced, exaggerated, or unnatural expressions (especially if a bite is implied). A single, glistening drop of rich, savory juice is visible, about to fall from the patty, powerfully emphasizing its irresistible succulence and indulgent quality without appearing messy. * **Central Subject:** The star of the image is the GreenBite burger. It’s a thick, plant-based patty with a deeply browned, flawlessly seared crust and visible grill marks, looking indistinguishable from premium ground beef. It’s topped with a layer of glistening, perfectly melted cheddar cheese, a crisp piece of bright green leaf lettuce, a thick, ruby-red slice of a heirloom tomato, and a few rings of sharp red onion, all between a fluffy, toasted brioche bun. * **Composition & Framing:** An enticing, slightly low-angle medium close-up shot where the GreenBite burger is the undeniable hero, prominently filling a significant portion of the frame and drawing the viewer’s eye directly to its appetizing qualities. The focus is tack-sharp on the GreenBite patty’s seared texture and the glistening juice. The hands holding the burger, and the subject’s face (if present), along with the backyard scene, are rendered in a beautiful, soft-focus bokeh, creating a strong emotional connection and aspirational appeal while ensuring the burger remains supremely appetizing. There is significant negative space with soft, blurred greenery in the upper portion of the frame for a text overlay.
**Stylistic Qualities:**
* **Photography Style:** High-end, photorealistic commercial food photography with a cinematic lifestyle feel. The image should be incredibly crisp, detailed, and textured. * **Lighting:** Bathed in the warm, dreamy glow of golden hour sunlight. The light source comes from the side, casting soft shadows that accentuate the burger’s shape and highlight the glistening textures of the melted cheese and juicy patty. A slight, elegant lens flare bleeds into the frame, adding to the warm, aspirational feel. * **Color Palette & Tone:** Rich, saturated, and mouth-watering. A color palette dominated by the warm browns of the seared patty, the vibrant orange of the cheese, the fresh greens of the lettuce, and the deep reds of the tomato, all set against the warm, earthy tones of the background.
* **Atmosphere & Feeling:** Evokes a powerful feeling of shared satisfaction, vibrant well-being, and guilt-free indulgence. The atmosphere is confidently positive, supremely delicious, and distinctly aspirational, conveying a desirable lifestyle of effortless enjoyment. **Brand Elements:** *
**Logo:** The simple, modern "GreenBite" logo (clean sans-serif font with a subtle leaf element) is integrated prominently yet naturally within the scene. It is elegantly printed on the branded wax paper partially wrapping the bottom of the burger, and also subtly visible on a branded napkin or a stylish sauce bottle on the blurred patio table in the background, reinforcing brand presence contextually within the aspirational setting. * In addition to the physical logo placements, a clean, prominent text overlay of the ’GreenBite’ brand name, and optionally a relevant tagline (e.g., ’Believe Your Bite’), should be integrated into the significant negative space in the upper portion of the frame. This overlay should be highly legible and reinforce clear brand recognition without detracting from the visual appeal."
Figure 16: Prompts for T-BoN ($T=10$)
Prompt for generating the GEPA ($T=10$) ad for GreenBite (plant-based burger patty) "**Key Message:**
Visually communicate the "no-compromise" promise of GreenBite by capturing the exact moment of pure, blissful satisfaction of a consumer realizing a plant-based burger can be just as juicy and delicious as its beef counterpart.
**Core Components:**
* **Scene & Environment:** A vibrant, sun-drenched modern backyard patio during a relaxed get-together of friends in their late 20s and early 30s. The setting is stylish and aspirational, featuring a natural wood deck, comfortable outdoor furniture, festoon string lights, and lush, green potted plants in the background.
**Action & Narrative:** The central focus is a woman with her eyes open, full of genuine, surprised satisfaction, capturing the peak moment of consumption. Her mouth is wide open, actively biting into the fully-loaded GreenBite burger. The patty is clearly visible within her mouth, emphasizing the powerful ’A-ha!’ moment of discovery and the sheer indulgence of the bite, directly communicating the ’no-compromise’ promise.
* **Composition & Framing:** Dynamic, slightly low-angle medium shot focusing on the woman and the burger. The burger is the hero; the depth of field is shallow, rendering the juicy, perfectly seared patty, melted plant-based cheddar, crisp lettuce, and glossy brioche bun in hyper-detailed, sharp focus. The background is beautifully out of focus with a pleasing bokeh effect. There is negative space in the upper left corner suitable for a text overlay.
**Stylistic Qualities:** * **Photography Style:** High-end, commercial lifestyle food photography. Polished, authentic, and incredibly photorealistic. Looks like an image from a top-tier food publication. * **Lighting:** Bathed in warm, cinematic golden hour sunlight. The light creates long, soft shadows and a beautiful, inviting glow, highlighting the textures of the food. Crucially, integrate a strong, warm sun flare or haze into the scene, creating a dramatic, aspirational glow that enhances the sense of satisfaction and makes the overall atmosphere feel more vibrant and special. * **Color Palette & Tone:** Rich, warm, and saturated colors. Deep browns of the seared patty, vibrant greens of the lettuce and surrounding plants, warm yellows and oranges from the cheese and sunlight. The overall tone is confident, energetic, and mouth-watering.
* **Atmosphere & Feeling:** A powerful feeling of guilt-free indulgence and vibrant satisfaction. The atmosphere is warm, happy, and effortlessly cool, capturing the joy of sharing great food with friends. **Inclusions:**
* **Logo:** Ensure the ’GreenBite’ logo is subtly yet clearly visible on the burger wrapper, or on a small, non-distracting element within the scene, to provide product context without pulling focus from the main action.
* **Text Overlay:** A prominent, clean white text overlay, featuring a strong, benefit-driven headline such as ’THE NO-COMPROMISE BURGER’ or ’BEYOND EXPECTATION.’, is strategically placed in the negative space (e.g., upper left corner). This headline should be clearly legible, visually impactful, and act as a primary visual element, prioritizing the core message and value proposition over just generic brand reinforcement."
Figure 17: Prompts for GEPA ($T=10$)
Appendix B Scenarios
Below are ten ad campaign scenarios with creative briefs, outlining the strategic and creative direction for each experimental scenario. The selected campaigns cover a diverse range of products across distinct categories. Each campaign addresses a unique consumer need—whether related to sustainability, luxury, technology, or lifestyle—ensuring comprehensive coverage of various market segments. From environmentally-conscious millennials to affluent, experience-driven travelers, the approach accounts for the full spectrum of consumer interests. By focusing on different consumer needs, the structure enables a holistic marketing strategy that captures both niche and broad demands, providing a well-rounded framework for the study.
B.1 Scenario 1: GreenBite Plant-Based Burger
  • Product: "GreenBite," a new plant-based burger patty.
  • Background: The market for plant-based food is growing, but many consumers remain skeptical about taste, believing they must sacrifice flavor for health.
  • The Challenge: Convince flexitarians that GreenBite offers the juicy, satisfying experience of a traditional beef burger with no compromise.
  • Core Insight: Consumers want to eat better for themselves and the planet, but they fear it means giving up the foods they love.
  • Single-Minded Proposition (SMP): The delicious, no-compromise burger experience that’s better for you and the planet.
  • Reasons to Believe (RTB): Made with savory pea protein; sears and tastes like real beef; free of soy and GMOs.
  • Desired Response: Think: "I can finally have a plant-based burger that tastes like the real thing.", Feel: Satisfied, vibrant, and guilt-free. Do: Purchase GreenBite for their next barbecue.
  • Brand Personality & Tone: Positive, confident, and mouth-watering.
B.2 Scenario 2: AuraSonics X1 Earbuds
  • Product: "AuraSonics X1," high-end, noise-canceling wireless earbuds.
  • Background: The rise of hybrid work and urban density has increased daily auditory overload, making focus a precious commodity.
  • The Challenge: Cut through the saturated audio market by positioning the X1 not as a gadget, but as an essential tool for mental clarity.
  • Core Insight: In a world of constant noise, true luxury is the ability to control your own soundscape.
  • Single-Minded Proposition (SMP): AuraSonics X1 creates a personal sanctuary for focus and immersion.
  • Reasons to Believe (RTB): Adaptive Active Noise Cancellation; studio-quality audio; all-day battery life; minimalist aluminum design.
  • Desired Response: Think: "This is the solution I need to escape the daily chaos." Feel: Calm, empowered, and sophisticated. Do: Visit the website to learn more.
  • Brand Personality & Tone: Minimalist, intelligent, and calm.
B.3 Scenario 3: Odyssey E-SUV
  • Product: The "Odyssey E-SUV," a new all-electric family SUV.
  • Background: Many families want to switch to electric vehicles but are concerned about sacrificing space, safety, and range for sustainability.
  • The Challenge: Position the Odyssey E-SUV as the first EV that meets all the practical and adventurous needs of a modern family, making the switch to electric feel like an upgrade, not a compromise.
  • Single-Minded Proposition (SMP): The future of family adventure is here: all-electric, zero compromise.
  • Reasons to Believe (RTB): 500km range; top-tier safety rating; spacious three-row seating; all-wheel drive capability.
  • Desired Response: Think: "An electric car can actually handle my family’s active lifestyle." Feel: Adventurous, optimistic, and secure. Do: Schedule a test drive.
  • Brand Personality & Tone: Inspiring, capable, and forward-thinking.
B.4 Scenario 4: Oasis Eco-Lodge
  • Product: "Oasis Eco-Lodge," a secluded, luxury resort operating in harmony with its natural surroundings.
  • Background: The luxury travel market often equates opulence with excess. A growing segment of affluent travelers seeks experiences that are both exclusive and responsible.
  • The Challenge: Redefine luxury as a seamless integration with nature, promising an experience that is restorative for both the guest and the environment.
  • Core Insight: For those who have everything, true luxury is not more things, but a deeper connection to something pure and serene.
  • Single-Minded Proposition (SMP): Rediscover tranquility in a luxury that respects nature.
  • Reasons to Believe (RTB): Secluded private bungalows; farm-to-table dining with locally sourced ingredients; carbon-neutral operations.
  • Desired Response: Think: "This is a truly special escape, not just another five-star hotel." Feel: Serene, exclusive, and rejuvenated. Do: Book a stay.
  • Brand Personality & Tone: Elegant, peaceful, and understated.
B.5 Scenario 5: Momentum Digital Banking
  • Product: "Momentum," a mobile-first banking app for freelancers and the gig economy.
  • Background: Traditional banks are not built for the fluctuating incomes and unique business needs of self-employed professionals.
  • The Challenge: Position Momentum as the essential financial co-pilot for the independent worker, simplifying complexity and providing stability.
  • Core Insight: Freelancers love their freedom but feel anxious about their financial instability. They need a tool that brings order to their financial chaos.
  • Single-Minded Proposition (SMP): Banking that is as flexible and entrepreneurial as you are.
  • Reasons to Believe (RTB): Automated tax-saving tools; integrated invoicing and payment tracking; instant business expense categorization.
  • Desired Response: Think: "Finally, a bank that gets the way I work." Feel: Empowered, organized, and financially confident. Do: Download the app and open an account.
  • Brand Personality & Tone: Modern, empowering, and simple.
B.6 Scenario 6: MindGarden Meditation App
  • Product: "MindGarden," a subscription-based meditation and mindfulness app.
  • Background: While many are interested in mindfulness for stress-relief, they are often intimidated by the practice, seeing it as difficult or time-consuming.
  • The Challenge: Make meditation feel accessible and achievable for absolute beginners, removing the barriers of time and perceived difficulty.
  • Core Insight: People want the benefits of mindfulness but are convinced they don’t have the time or ability to practice it correctly.
  • Single-Minded Proposition (SMP): Find your calm in just 5 minutes a day.
  • Reasons to Believe (RTB): Guided 5-minute sessions for specific needs (anxiety, focus); progress tracking to build a habit; simple, jargon-free instructions.
  • Desired Response: Think: "I can do this. 5 minutes is easy." Feel: Calm, supported, and hopeful. Do: Start a free trial.
  • Brand Personality & Tone: Gentle, approachable, and encouraging.
B.7 Scenario 7: Aeterno Watch
  • Product: The "Aeterno," a classic, automatic Swiss-made wristwatch with a heritage design.
  • Background: In an age of smartwatches that are obsolete in a few years, there is a renewed appreciation for objects with permanence and enduring value.
  • The Challenge: Reassert the relevance of the classic mechanical watch as a symbol of taste, craftsmanship, and timelessness.
  • Core Insight: In a disposable world, true status comes from owning something permanent that tells a story.
  • Single-Minded Proposition (SMP): A legacy on your wrist. Craftsmanship that transcends time.
  • Reasons to Believe (RTB): Swiss-made automatic movement; sapphire crystal glass; timeless design inspired by 1950s classics.
  • Desired Response: Think: "This is a beautiful object that I will own forever." Feel: Prestigious, sophisticated, and discerning. Do: Locate an authorized dealer.
  • Brand Personality & Tone: Elegant, timeless, and confident.
B.8 Scenario 8: SyncFlow B2B Software
  • Product: "SyncFlow," a project management and collaboration software platform for remote teams.
  • Background: Remote work has increased flexibility but also created challenges in team alignment, communication, and project visibility.
  • The Challenge: Cut through a crowded SaaS market by focusing on the outcome of "effortless collaboration" rather than just listing features.
  • Core Insight: Managers don’t want another tool to manage; they want the feeling of clarity and momentum that comes from a team working in perfect sync.
  • Single-Minded Proposition (SMP): Bring your remote team together for effortless collaboration and remarkable results.
  • Reasons to Believe (RTB): Centralized dashboards with real-time project status; integrated communication channels; automated workflows and reporting.
  • Desired Response: Think: "This could finally solve our remote work chaos and get everyone on the same page." Feel: Organized, in control, and successful. Do: Sign up for a team demo.
  • Brand Personality & Tone: Professional, efficient, and innovative.
Appendix C Extended experiment results
T-BoN BO
Aeterno Aurasonics Greenbite Mindgarden Momentum Oasis Odyssey Syncflow Average
Worst5-of-N 2.942 (0.025) 2.793 (0.023) 2.706 (0.024) 2.722 (0.024) 2.644 (0.024) 2.839 (0.025) 2.828 (0.025) 2.641 (0.023) 2.739
Steps 1 3.098 (0.028) 2.813 (0.022) 3.001 (0.027) 2.980 (0.027) 2.799 (0.027) 3.062 (0.028) 2.943 (0.027) 2.927 (0.023) 2.932
2 3.125 (0.026) 2.969 (0.027) 3.041 (0.026) 2.980 (0.027) 2.907 (0.025) 3.107 (0.026) 2.943 (0.027) 2.927 (0.023) 2.982
3 3.125 (0.026) 3.041 (0.025) 3.041 (0.026) 2.980 (0.027) 2.907 (0.025) 3.131 (0.019) 3.063 (0.019) 2.927 (0.023) 3.013
4 3.194 (0.029) 3.041 (0.025) 3.041 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.063 (0.019) 2.927 (0.023) 3.062
5 3.194 (0.029) 3.041 (0.025) 3.076 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.063 (0.019) 3.127 (0.022) 3.096
6 3.257 (0.023) 3.041 (0.025) 3.076 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.070 (0.022) 3.127 (0.022) 3.097
7 3.257 (0.023) 3.104 (0.028) 3.076 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.111
8 3.257 (0.023) 3.104 (0.028) 3.076 (0.026) 3.132 (0.022) 2.913 (0.026) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.112
9 3.257 (0.023) 3.104 (0.028) 3.076 (0.026) 3.132 (0.022) 2.926 (0.026) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.114
10 3.257 (0.023) 3.104 (0.028) 3.168 (0.025) 3.132 (0.022) 3.024 (0.025) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.155
Best-of-N 2.982 (0.025) 2.934 (0.027) 2.898 (0.025) 2.930 (0.028) 3.030 (0.027) 3.265 (0.022) 2.966 (0.027) 3.086 (0.023) 3.011
Table 2: Progress of T-BoN BO across 10 optimization steps from its starting point (Worst5-of-N) and comparison with the Best-of-N baseline. Each data point, in the mean (standard error) format, indicates the mean and standard error across 411 personas in the test persona set. The last column, labeled “Average”, reports the averages of the means across the eight scenarios.
GEPA
Aeterno Aurasonics Greenbite Mindgarden Momentum Oasis Odyssey Syncflow Average
Worst5-of-N 2.942 (0.025) 2.793 (0.023) 2.706 (0.024) 2.722 (0.024) 2.644 (0.024) 2.839 (0.025) 2.828 (0.025) 2.641 (0.023) 2.739
Steps 1 3.036 (0.025) 2.942 (0.025) 3.029 (0.025) 2.841 (0.024) 2.816 (0.024) 3.151 (0.024) 2.915 (0.025) 2.880 (0.025) 2.939
2 3.036 (0.025) 2.942 (0.025) 3.046 (0.024) 2.945 (0.023) 2.907 (0.023) 3.151 (0.024) 3.071 (0.028) 2.880 (0.025) 2.992
3 3.036 (0.025) 2.942 (0.025) 3.046 (0.024) 2.945 (0.023) 2.907 (0.023) 3.182 (0.025) 3.088 (0.026) 2.914 (0.018) 3.003
4 3.138 (0.025) 3.028 (0.022) 3.046 (0.024) 2.954 (0.024) 2.976 (0.026) 3.187 (0.023) 3.088 (0.026) 2.949 (0.028) 3.033
5 3.138 (0.025) 3.028 (0.022) 3.046 (0.024) 3.012 (0.022) 2.976 (0.026) 3.187 (0.023) 3.088 (0.026) 3.008 (0.026) 3.049
6 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.012 (0.022) 2.976 (0.026) 3.209 (0.020) 3.088 (0.026) 3.008 (0.026) 3.052
7 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.209 (0.020) 3.088 (0.026) 3.008 (0.026) 3.060
8 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.209 (0.020) 3.088 (0.026) 3.008 (0.026) 3.060
9 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.278 (0.023) 3.088 (0.026) 3.008 (0.026) 3.070
10 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.278 (0.023) 3.088 (0.026) 3.008 (0.026) 3.070
Best-of-N 2.982 (0.025) 2.934 (0.027) 2.898 (0.025) 2.930 (0.028) 3.030 (0.027) 3.265 (0.022) 2.966 (0.027) 3.086 (0.023) 3.011
Table 3: Progress of GEPA across 10 optimization steps from its starting point (Worst5-of-N) and comparison with the Best-of-N baseline. Each data point, in the mean (standard error) format, indicates the mean and standard error across 411 personas in the test persona set. The last column, labeled “Average”, reports the averages of the means across the eight scenarios.
Odyssey (Electric family SUV) Oasis (Secluded resort) Momentum (Mobile banking app) Aeterno (Luxury watch) Syncflow (Collaboration tool)
Worst5-of-64 Refer to caption Score: 2.828 Refer to caption Score: 2.839 Refer to caption Score: 2.644 Refer to caption Score: 2.942 Refer to caption Score: 2.641
Best-of-64 Refer to caption Score: 2.966 Refer to caption Score: 3.265 Refer to caption Score: 3.030 Refer to caption Score: 2.982 Refer to caption Score: 3.086
GEPA (T=10) Refer to caption Score: 3.088 Refer to caption Score: 3.278 Refer to caption Score: 2.976 Refer to caption Score: 3.157 Refer to caption Score: 3.008
T-BoN BO (T=10) Refer to caption Score: 3.107 Refer to caption Score: 3.324 Refer to caption Score: 3.024 Refer to caption Score: 3.257 Refer to caption Score: 3.127
Figure 18: T-BoN BO and baseline’s generated Ad images for five fictional brands: Odyssey electric family SUV (Scenario 3), Oasis eco-lodge (Scenario 4), Momentum Digital Banking (Scenario 5), Aeterno luxury watch (Scenario 7), and SyncFlow B2B software (Scenario 8)
Appendix D Proofs
D.1 Auxiliary lemmas
Lemma 3 (Uniform first-order expansion).
Let $g=\nabla\mu(x)$ and $h=\nabla\sigma(x)$. There exist $C>0$ and $\varepsilon_{0}>0$ such that for all $\varepsilon\in(0,\varepsilon_{0}]$ and all $u\in\mathbb{R}^{d}$,
$\mu(x+\varepsilon u)$ $=\mu(x)+\varepsilon\,g^{\top}u\;+\;r_{\mu}(\varepsilon,u),$ $|r_{\mu}(\varepsilon,u)|$ $\leq C\varepsilon^{2},$ (21)
$\sigma(x+\varepsilon u)$ $=\sigma(x)+\varepsilon\,h^{\top}u\;+\;r_{\sigma}(\varepsilon,u),$ $|r_{\sigma}(\varepsilon,u)|$ $\leq C\varepsilon^{2}.$ (22)
Proof.
By the mean-value theorem with Lipschitz gradients, for some $\xi$ between $x$ and $x+\varepsilon u$, $\mu(x+\varepsilon u)=\mu(x)+\nabla\mu(\xi)^{\top}(\varepsilon u).$ Hence $\mu(x+\varepsilon u)=\mu(x)+\varepsilon g^{\top}u+\varepsilon(\nabla\mu(\xi)-\nabla\mu(x))^{\top}u,$ and $|\nabla\mu(\xi)-\nabla\mu(x)|\leq L_{\mu}\|\xi-x\|\leq L_{\mu}\varepsilon$, giving $|r_{\mu}|\leq L_{\mu}\varepsilon^{2}$. The proof for $\sigma$ is identical with $L_{\sigma}$. Taking $C:=\max\{L_{\mu},L_{\sigma}\}$ yields the stated bounds with the same $C$. ∎
Lemma 4 (Max-stability under bounded perturbations).
Let $(f_{i})_{i\leq N}$ and $(g_{i})_{i\leq N}$ be real arrays. Let $i_{f}\in\arg\max_{i}f_{i}$ and $i_{g}\in\arg\max_{i}g_{i}$. If for some $\Delta\geq 0$,
$g_{i_{f}}\geq f_{i_{f}}-\Delta\quad\text{and}\quad f_{i_{g}}\geq g_{i_{g}}-\Delta,$ (23)
then $g_{i_{f}}\geq\max_{i}g_{i}-\Delta$ and $f_{i_{g}}\geq\max_{i}f_{i}-\Delta$. In particular, $g_{i_{f}}\geq\max_{i}g_{i}-\Delta$ implies that any maximizer of $f$ is a $\Delta$-approximate maximizer of $g$.
Proof.
Since $g_{i_{f}}\geq f_{i_{f}}-\Delta=\max_{i}f_{i}-\Delta\geq g_{i_{g}}-\Delta=\max_{i}g_{i}-\Delta$, the first claim holds. The second is symmetric. ∎
Lemma 5 (Spherical cap coverage).
For $v\in\mathbb{S}^{d-1}$ and $\eta\in(0,1)$, define $C(v,\eta):=\{u\in\mathbb{S}^{d-1}:v^{\top}u\geq 1-\eta\}$. Let $U_{1},\ldots,U_{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\rho$ as above. Then $\max_{i\leq N}v^{\top}U_{i}\to 1$ almost surely as $N\to\infty$. Moreover, for any deterministic sequence $\eta_{N}\downarrow 0$ with $N\cdot\rho\{u:v^{\top}u\geq 1-\eta_{N}\}\to\infty$, we have
$\mathbb{P}\!\left(\max_{i\leq N}v^{\top}U_{i}\geq 1-\eta_{N}\right)\to 1.$ (24)
Proof.
Let $C(\eta)=\{u\in\mathbb{S}^{d-1}:v^{\top}u\geq 1-\eta\}$ be a spherical cap. Since $\rho$ has a density bounded below, $\rho(C(\eta))\asymp\eta^{\frac{d-1}{2}}$ as $\eta\downarrow 0$. Then $\mathbb{P}(\max_{i}v^{\top}U_{i}<1-\eta)=(1-\rho(C(\eta)))^{N}\leq\exp(-N\rho(C(\eta))).$ For any fixed $\eta\in(0,1)$, $\sum_{N\geq 1}\mathbb{P}(\max_{i}v^{\top}U_{i}<1-\eta)\leq\sum_{N\geq 1}e^{-N\rho(C(\eta))}<\infty$, so by Borel–Cantelli, $\mathbb{P}(\max_{i}v^{\top}U_{i}<1-\eta\ \text{i.o.})=0$. Intersecting over a countable sequence $\eta_{m}\downarrow 0$ yields $\max_{i\leq N}v^{\top}U_{i}\to 1$ almost surely. Choosing $\eta_{N}\downarrow 0$ so that $N\rho(C(\eta_{N}))\to\infty$ yields the result, and taking $\eta=\eta_{N}\to 0$ proves the first statement. ∎
Lemma 6 (Near-maximum coupling in a thin band).
Let $M_{N}=\max_{i}\xi_{i}$ and define $v_{N}:=(g+M_{N}h)/\|g+M_{N}h\|$. Also, $S_{i}:=a_{i}+b_{i}M_{N}$, $Y_{i}^{(0)}:=a_{i}+b_{i}\xi_{i}$, $b_{i}:=\sigma(x)+\varepsilon h^{\top}U_{i}$. Fix any sequence $c_{N}\uparrow\infty$ with $c_{N}/q_{N}\to 0$ and set $\delta_{N}:=c_{N}/q_{N}$. Then with probability $\to 1$ there exists an index $j$ such that
$\xi_{j}\geq M_{N}-\delta_{N}\quad\text{and}\quad v_{N}^{\top}U_{j}\geq 1-\eta_{N},$ (25)
whenever $\eta_{N}\downarrow 0$ satisfies
$N\,\mathbb{P}(\xi\geq M_{N}-\delta_{N})\,\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}\to\infty.$ (26)
Moreover, for this $j$,
$S_{j}-Y_{j}^{(0)}=b_{j}(M_{N}-\xi_{j})\leq b_{\max}\,\delta_{N}.$ (27)
Proof.
Define $p_{\mathrm{cap},N}:=\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}$ and the event
$E_{N}:=\{M_{N}\geq q_{N}-\delta_{N}\}.$ (28)
By Lemma 1, $M_{N}-q_{N}\to 0$ in probability and $\delta_{N}\to 0$, hence $\mathbb{P}(E_{N})\to 1$.
Set the deterministic threshold $t_{N}:=q_{N}-2\delta_{N}$. On $E_{N}$ we have $M_{N}-\delta_{N}\geq q_{N}-2\delta_{N}=t_{N}$, hence by monotonicity of the tail,
$\mathbb{P}(\xi\geq t_{N})\;\geq\;\mathbb{P}(\xi\geq M_{N}-\delta_{N})\quad\text{on }E_{N}.$ (29)
Let
$K_{N}:=\sum_{i=1}^{N}\mathbf{1}\{\xi_{i}\geq t_{N}\}\sim\mathrm{Binomial}\!\big(N,\,p_{N}\big),\qquad p_{N}:=\mathbb{P}(\xi\geq t_{N}).$ (30)
A Chernoff bound gives
$\mathbb{P}\!\left(K_{N}<\tfrac{1}{2}Np_{N}\right)\;\leq\;\exp\!\left(-\tfrac{1}{8}Np_{N}\right).$ (31)
Consider
$L_{N}\;:=\;\sum_{i=1}^{N}\mathbf{1}\{\xi_{i}\geq t_{N}\}\,\mathbf{1}\{v_{N}^{\top}U_{i}\geq 1-\eta_{N}\}.$ (32)
[…]
$L_{N}\,\big|\,\{\xi_{i}\}\;\sim\;\mathrm{Binomial}\!\big(K_{N},\,p_{\mathrm{cap},N}\big),\qquad\mathbb{P}\big(L_{N}=0\,\big|\,\{\xi_{i}\}\big)\;=\;(1-p_{\mathrm{cap},N})^{K_{N}}\;\leq\;\exp\!\big(-K_{N}p_{\mathrm{cap},N}\big).$ (33)
We now bound $\mathbb{P}(L_{N}=0)$ by splitting on $\{K_{N}\geq\tfrac{1}{2}Np_{N}\}$ and $E_{N}$:
$\mathbb{P}(L_{N}=0)$ $\leq\mathbb{E}\!\left[\exp\!\big(-K_{N}p_{\mathrm{cap},N}\big)\,\mathbf{1}_{\{K_{N}\geq\frac{1}{2}Np_{N}\}\cap E_{N}}\right]\;+\;\mathbb{P}\!\left(K_{N}<\tfrac{1}{2}Np_{N}\right)\;+\;\mathbb{P}(E_{N}^{c})$
$\leq\mathbb{E}\!\left[\exp\!\big(-\tfrac{1}{2}Np_{N}\,p_{\mathrm{cap},N}\big)\,\mathbf{1}_{E_{N}}\right]\;+\;\exp\!\big(-\tfrac{1}{8}Np_{N}\big)\;+\;o(1)$
$\leq\mathbb{E}\!\left[\exp\!\big(-\tfrac{1}{2}Np_{N}\,p_{\mathrm{cap},N}\big)\right]\;+\;\exp\!\big(-\tfrac{1}{8}Np_{N}\big)\;+\;o(1).$ (34)
On $E_{N}$, (29) yields $p_{N}\geq\mathbb{P}(\xi\geq M_{N}-\delta_{N})$. By the lemma’s hypothesis,
$Z_{N}:=Np_{N}\,p_{\mathrm{cap},N}\;\geq\;N\,\mathbb{P}(\xi\geq M_{N}-\delta_{N})\,p_{\mathrm{cap},N}\;\xrightarrow{p}\;\infty.$
Since $0\leq e^{-\frac{1}{2}Z_{N}}\leq 1$, for any fixed $T>0$,
$\mathbb{E}\!\left[e^{-\frac{1}{2}Z_{N}}\right]\leq e^{-\frac{1}{2}T}+\mathbb{P}(Z_{N}\leq T)\;\longrightarrow\;e^{-\frac{1}{2}T}.$
Letting $T\to\infty$ gives $\mathbb{E}[e^{-\frac{1}{2}Z_{N}}]\to 0$. Moreover, $Z_{N}\leq Np_{N}$ since $p_{\mathrm{cap},N}\leq 1$. If $Np_{N}$ were bounded along a subsequence, then $Z_{N}$ would be bounded along that subsequence, contradicting $Z_{N}\xrightarrow{p}\infty$. Hence $Np_{N}\to\infty$, so $\exp(-\tfrac{1}{8}Np_{N})\to 0$. Therefore the second term in (34) vanishes, and $\mathbb{P}(L_{N}=0)\to 0$.
Therefore, with probability tending to one there exists $j$ such that $\xi_{j}\geq t_{N}$ and $v_{N}^{\top}U_{j}\geq 1-\eta_{N}$. Since $t_{N}\leq M_{N}-\delta_{N}$ on $E_{N}$, this $j$ also satisfies $\xi_{j}\geq M_{N}-\delta_{N}$. Finally, $S_{j}-Y_{j}^{(0)}=b_{j}(M_{N}-\xi_{j})\leq b_{\max}\delta_{N}$ holds deterministically by $M_{N}-\xi_{j}\leq\delta_{N}$ and the definition of $b_{\max}$. ∎
D.2 Proof of Theorem 2
Proof.
By Lemma 3,
$Y_{i}$ $=\mu(x)+\varepsilon g^{\top}u_{i}+r_{\mu}(\varepsilon,u_{i})+\big(\sigma(x)+\varepsilon h^{\top}u_{i}+r_{\sigma}(\varepsilon,u_{i})\big)\xi_{i}$
$=a_{i}+b_{i}\xi_{i}+R_{i},$ (35)
where $a_{i}:=\mu(x)+\varepsilon g^{\top}u_{i}$, $b_{i}:=\sigma(x)+\varepsilon h^{\top}u_{i}$, and $R_{i}:=r_{\mu}(\varepsilon,u_{i})+\xi_{i}r_{\sigma}(\varepsilon,u_{i}).$ Then $|R_{i}|\leq C\varepsilon^{2}(1+|\xi_{i}|)$ for all $i$. Define $S_{i}:=a_{i}+b_{i}M_{N}$, $b_{\max}:=\max_{1\leq i\leq N}|b_{i}|$, and $U_{i}:=u_{i}$.
Now let $Y_{i}^{(0)}:=a_{i}+b_{i}\xi_{i}$ and $i_{0}\in\arg\max_{i}Y_{i}^{(0)}$. From (35),
$Y_{i^{\star}}\geq Y_{i_{0}}\implies Y_{i^{\star}}^{(0)}\geq Y_{i_{0}}^{(0)}-|R_{i^{\star}}|-|R_{i_{0}}|.$
By the bound on $R_{i}$ and $|\xi_{i}|\leq M_{N}:=\max_{j}\xi_{j}$,
$Y_{i^{\star}}^{(0)}\geq\max_{i}Y_{i}^{(0)}-2C\varepsilon^{2}(1+M_{N}).$ (36)
Symmetrically, $\max_{i}Y_{i}^{(0)}\geq Y_{i^{\star}}^{(0)}\geq\max_{i}Y_{i}^{(0)}-2C\varepsilon^{2}(1+M_{N})$. By Lemma 4, $i^{\star}$ is a $2C\varepsilon^{2}(1+M_{N})$-approximate maximizer of $Y^{(0)}$.
Now let $j$ be the index from Lemma 6 so that $\xi_{j}\geq M_{N}-\delta_{N}$ and $v_{N}^{\top}U_{j}\geq 1-\eta_{N}$. By Lemma 3,
$\max_{i\leq N}S_{i}-S_{j}\;\leq\;\varepsilon\|g+M_{N}h\|\,(1-v_{N}^{\top}U_{j})\;+\;2C\varepsilon^{2}(1+M_{N})\;\leq\;\varepsilon\|g+M_{N}h\|\,\eta_{N}\;+\;2C\varepsilon^{2}(1+M_{N}).$
Also,
$S_{j}-S_{i^{\star}}=(S_{j}-Y_{j}^{(0)})+(Y_{j}^{(0)}-Y_{i^{\star}}^{(0)})+(Y_{i^{\star}}^{(0)}-S_{i^{\star}})\;\leq\;b_{\max}\delta_{N}\;+\;2C\varepsilon^{2}(1+M_{N})\;+\;\underbrace{(Y_{i^{\star}}^{(0)}-S_{i^{\star}})}_{\leq 0},$
where the last term is $\leq 0$ provided $b_{i}\geq 0$ for all $i$, which holds by Assumption A4 for $\varepsilon\leq\varepsilon_{0}:=\sigma(x)/(2\|h\|)$ when $\sigma(x)>0$ since $b_{i}=\sigma(x)+\varepsilon h^{\top}U_{i}\geq\sigma(x)-\varepsilon\|h\|\geq\tfrac{1}{2}\sigma(x)>0$. Therefore,
$\max_{i\leq N}S_{i}-S_{i^{\star}}\;\leq\;\varepsilon\|g+M_{N}h\|\,\eta_{N}\;+\;b_{\max}\delta_{N}\;+\;4C\varepsilon^{2}(1+M_{N}).$ (37)
Set:
$\Delta^{\prime}_{N}(\varepsilon)\;:=\;b_{\max}\,\frac{c_{N}}{q_{N}}\;+\;4C\varepsilon^{2}(1+M_{N}).$
By Lemma 3,
$S_{i}\;=\;\mu(x)+M_{N}\sigma(x)+\varepsilon(g+M_{N}h)^{\top}u_{i}+\tilde{r}_{i},\qquad|\tilde{r}_{i}|\leq C\varepsilon^{2}(1+M_{N}).$
Hence for any $i$,
$S_{i}-S_{i^{\star}}=\varepsilon\|g+M_{N}h\|\,(v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N})+(\tilde{r}_{i}-\tilde{r}_{i^{\star}}),$
so
$\big|\,(S_{i}-S_{i^{\star}})-\varepsilon\|g+M_{N}h\|\,(v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N})\,\big|\;\leq\;2C\varepsilon^{2}(1+M_{N}).$ (38)
Taking $i$ that attains $\max_{i}S_{i}$ and using (37) yields
$\max_{i\leq N}v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N}\;\leq\;\eta_{N}\;+\;\frac{\Delta^{\prime}_{N}(\varepsilon)}{\varepsilon\|g+M_{N}h\|}.$ (39)
Thus
[…]
Assume $\|h\|>0$ (the $\|h\|=0$ case is addressed in Remark 1). Using
$\|g+M_{N}h\|\;\geq\;M_{N}\|h\|-\|g\|$
and Lemma 1, $M_{N}\to\infty$ and $M_{N}\asymp q_{N}$. Also $b_{\max}\leq\sigma(x)+\varepsilon\|h\|+C\varepsilon^{2}$. Hence
$\frac{\Delta^{\prime}_{N}(\varepsilon)}{\varepsilon\|g+M_{N}h\|}=O_{p}\!\Big(\tfrac{c_{N}}{\varepsilon q_{N}^{2}\|h\|}\Big)+O\!\Big(\tfrac{\varepsilon}{\|h\|}\Big).$
Choosing, e.g., $c_{N}=\tfrac{1}{2}\log\log N$ and any $\eta_{N}\downarrow 0$ with
$N\,e^{c_{N}}\,\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}\to\infty$
makes the RHS of (40) equal to $o_{p}(1)+O(\varepsilon)$ for fixed $\varepsilon>0$; then let $\varepsilon\downarrow 0$.
With $N\to\infty$ first and fixed $\varepsilon>0$, (40) implies $1-v_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}0$. Then $\varepsilon\downarrow 0$ removes the $O(\varepsilon)$ residual, yielding $v_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}1$. Since $M_{N}-q_{N}\to 0$ in probability by Lemma 1, the unit vectors $v_{N}=(g+M_{N}h)/\|g+M_{N}h\|$ and $\tilde{v}_{N}:=(g+q_{N}h)/\|g+q_{N}h\|$ satisfy $\|v_{N}-\tilde{v}_{N}\|\to 0$ in probability. Hence $\widehat{u}_{N}\xrightarrow{p}\tilde{v}_{N}$, which is the claimed gradient direction of $A_{\beta_{N}}$ with $\beta_{N}=q_{N}$.
By Lemma 3,
$A_{\beta_{N}}(x+\Delta x^{(N)})=A_{\beta_{N}}(x)+\varepsilon\nabla A_{\beta_{N}}(x)^{\top}\widehat{u}_{N}+O(\varepsilon^{2}).$
Write $\nabla A_{\beta_{N}}(x)=g+q_{N}h$. Using $\tilde{v}_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}1$ gives $\nabla A_{\beta_{N}}(x)^{\top}\widehat{u}_{N}=\|\nabla A_{\beta_{N}}(x)\|\,\tilde{v}_{N}^{\top}\widehat{u}_{N}=\|\nabla A_{\beta_{N}}(x)\|(1-o_{p}(1)).$ Thus $A_{\beta_{N}}(x+\Delta x^{(N)})\geq A_{\beta_{N}}(x)+\varepsilon\|\nabla A_{\beta_{N}}(x)\|-o_{p}(\varepsilon),$ as claimed. ∎
Remark 1 (Edge cases).
(i) If $\|h\|=0$, then $\nabla A_{\beta_{N}}(x)=g$ and the direction reduces to $\nabla\mu(x)$. (ii) If $\sigma(x)=0$ but $\|h\|>0$, the result holds unchanged; only the linearization constants change. (iii) If $\sigma(x)=0$ and $\|h\|=0$, the model is locally deterministic and the problem is degenerate; selection aligns with $\nabla\mu(x)$ or is undefined if $g=0$.
Appendix E GP-UCB and its evaluation efficiency
Setup.
Let $f$ be a unknown, black-box function that maps a compact and convex feasible region $\mathcal{D}\subset\mathbb{R}^{d}$ to $\mathbb{R}$. At each time $t$, a sampling algorithm $\mathcal{A}$ selects a point $x_{t}\in\mathcal{D}$ to evaluate. This selection is adaptive, based on the history of previously chosen points and their observed outcomes. The evaluation at $x_{t}$ yields a noisy observation $y_{t}=f\left(x_{t}\right)+\epsilon_{t}$, where $\epsilon_{t}$ is a random noise term, where $\left\{\epsilon_{t}:t=1,2,\ldots\right\}$ are independent, sub-Gaussian random variables. The unknown function $f$ is assumed to be an element of a Reproducing Kernel Hilbert Space (RKHS), $\mathcal{N}_{\Psi}(\mathcal{D})$, induced by a stationary kernel $\Psi$. Functions within this RKHS possess a specific smoothness property, the nature of which is determined by the kernel used. The space $\mathcal{N}_{\Psi}(\mathcal{D})$ is endowed with an inner product and is formally defined as the set of functions
[…]
where the norm is given by: $\|g\|_{\mathcal{N}_{\Psi}(\mathcal{D})}:=\left(\sum_{j,l=1}^{\infty}c_{j}c_{l}\Psi\left(\mathbf{x}_{j}-\mathbf{x}_{l}\right)\right)^{\frac{1}{2}}$. For any two functions $g_{1}(\mathbf{x})=\sum_{j=1}^{\infty}a_{j}\Psi\left(\mathbf{x}-\mathbf{x}_{j}\right)$ and $g_{2}(\mathbf{x})=\sum_{l=1}^{\infty}b_{l}\Psi\left(\mathbf{x}-\mathbf{x}_{l}\right)$ in the space, their inner product is $\left\langle g_{1},g_{2}\right\rangle_{\mathcal{N}_{\Psi}(\mathcal{D})}:=\sum_{j,l=1}^{\infty}a_{j}b_{l}\Psi\left(\mathbf{x}_{j}-\mathbf{x}_{l}\right)$. A key feature of RKHS space is the reproducing property, $\langle g,\Psi(\mathbf{x}-\cdot)\rangle_{\mathcal{N}_{\Psi}(\mathcal{D})}=g(\mathbf{x})$, which holds for all $\mathbf{x}\in\mathcal{D}$. The norm of the function $f$ in this space is assumed to be bounded such that $\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B$ for some constant $B>0$.
The objectives.
The performance of $\mathcal{A}$ is often evaluated through two types of regret over a time horizon $T$. One is cumulative regret, defined as
$\mathcal{R}_{C}(T;f,\mathcal{A}):=\sum_{t=1}^{T}\left[\max_{\mathbf{x}\in\mathcal{D}}f(\mathbf{x})-f\left(\mathbf{x}_{t}\right)\right]$ (41)
The other is simple regret, defined as
$\mathcal{R}_{S}(T;f,\mathcal{A}):=\max_{\mathbf{x}\in\mathcal{D}}f(\mathbf{x})-f\left(\mathbf{x}^{(T)}\right)$ (42)
where $\mathbf{x}^{(T)}$ is the output of the algorithm $\mathcal{A}$ after taking $T$ function evaluations.
The lower bound.
scarlett2017lower showed that a lower bound exists on the regret of any algorithm for $f\in\mathcal{N}_{\Psi}(\mathcal{D})$. Specifically, for any constant $B>0$,
$\inf_{\mathcal{A}}\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}\mathbb{E}\left[\mathcal{R}_{C}(T;f,\mathcal{A})\right]=\begin{cases}\Omega(T^{\frac{\nu+d}{2\nu+d}}),&\text{ for Matérn kernels, }\\ \Omega(T^{\frac{1}{2}}\ln^{\frac{d}{2}}(T)),&\text{ for SE kernels. }\end{cases}$ (43)
$\inf_{\mathcal{A}}\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}T\left(\epsilon,f,\mathcal{A}\right)=\begin{cases}\Omega\left(\left(\frac{1}{\epsilon}\right)^{2+d/\nu}\right),&\text{ for Matérn kernels, }\\ \Omega\left(\frac{1}{\epsilon^{2}}\left(\log\frac{1}{\epsilon}\right)^{d/2}\right),&\text{ for SE kernels. }\end{cases}$ (44)
where Matérn kernels are defined as
$\Psi_{\mathrm{M}}\left(\mathbf{x}-\mathbf{x}^{\prime}\right):=\frac{1}{\Gamma(\nu)2^{\nu-1}}\left(\frac{2\sqrt{\nu}\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}}{\ell}\right)^{\nu}K_{\nu}\left(\frac{2\sqrt{\nu}\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}}{\ell}\right),\quad\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$ (45)
and SE kernels are defined as
$\Psi_{\mathrm{SE}}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)=\exp\left(-\frac{\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}^{2}}{2\ell^{2}}\right),\quad\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$ (46)
where $\nu>0$ is the smoothness parameter, $\ell>0$ is the length-scale parameter, $\Gamma(\cdot)$ is the gamma function, $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind of order $\nu$, and $\|\cdot\|_{2}$ denotes the Euclidean norm.
GP-UCB algorithm.
The Gaussian Process Upper Confidence Bound (GP-UCB) algorithm, introduced in the seminal work of srinivas2012information, operates from a Bayesian perspective by placing a Gaussian Process (GP) prior on the unknown objective function $f$. This GP is characterized by a zero-mean function and a symmetric, positive-definite covariance function, or kernel, $k:\mathcal{D}\times\mathcal{D}\mapsto\mathbb{R}$, where $k\left(\mathbf{x},\mathbf{x}^{\prime}\right)=\Psi\left(\mathbf{x}-\mathbf{x}^{\prime}\right)$ for all $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$. For any finite set of points $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{t}\}\subset\mathcal{D}$, the corresponding function values $(f(\mathbf{x}_{1}),\ldots,f(\mathbf{x}_{t}))^{\top}$ are assumed to follow a multivariate normal distribution with a zero mean vector and a covariance matrix with entries $\mathbf{K}_{t}]_{jl}=k(\mathbf{x}_{j},\mathbf{x}_{l})$. Under the assumption of independent, zero-mean Gaussian observation noise with variance $\sigma^{2}$, conditioning the GP prior on a set of observations $\mathbf{D}_{t}:=\{(\mathbf{x}_{j},y_{j}):j=1,\ldots,t\}$ yields a closed-form posterior distribution. The posterior mean and variance at any point $\mathbf{x}\in\mathcal{D}$ are given by:
$\mathbb{E}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]=\mathbf{k}_{t}^{\top}(\mathbf{x})\left(\mathbf{K}_{t}+\sigma^{2}\mathbf{I}_{t}\right)^{-1}\mathbf{y}_{t},\operatorname{Var}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]=k(\mathbf{x},\mathbf{x})-\mathbf{k}_{t}(\mathbf{x})^{\top}\left(\mathbf{K}_{t}+\sigma^{2}\mathbf{I}_{t}\right)^{-1}\mathbf{k}_{t}(\mathbf{x}),$ (47)
where $\mathbf{k}_{t}(\mathbf{x})$ is the vector of covariances between $\mathbf{x}$ and the observed points, $\mathbf{K}_{t}$ is the covariance matrix of the observed points, $\mathbf{y}_{t}$ is the vector of observed values, and $\mathbf{I}_{t}$ is the identity matrix. Inspired by upper confidence bound methods in the multi-armed bandit literature (Auer et al., 2002), GP-UCB employs an optimistic acquisition strategy to select subsequent evaluation points. At each step $t$, the next point $\mathbf{x}_{t+1}$ is chosen by maximizing an upper confidence bound on the function’s value $\mathbf{x}_{t+1}=\underset{\mathbf{x}\in\mathcal{D}}{\arg\max}\left(\mathbb{E}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]+\sqrt{\beta_{t}\operatorname{Var}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]}\right).$ This acquisition function naturally balances exploitation, driven by the posterior mean $\mathbb{E}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]$, with exploration, driven by the posterior standard deviation $\sqrt{\operatorname{Var}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]}$. The tunable parameter $\beta_{t}>0$ explicitly governs this trade-off, making its specification critical to the algorithm’s performance.
The evaluation efficiency of GP-UCB.
whitehouse2023sublinear, wang2023regret showed that GP-UCB srinivas2012information upper bound on the cumulative regret of any algorithm for $f\in\mathcal{N}_{\Psi}(\mathcal{D})$ that is almost nearly tight. Specifically, for any constant $B>0$, the GP-UCB algorithm $\mathcal{A}_{\text{GPUCB}}$ achieves
$\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}\mathbb{E}\left[\mathcal{R}_{C}(T;f,\mathcal{A}_{\text{GPUCB}})\right]=\begin{cases}O(T^{\frac{\nu+d}{2\nu+d}}),&\text{ for Matérn kernels, }\\ O(T^{\frac{1}{2}}\ln^{\frac{d}{2}}(T))&\text{ for SE kernels. }\end{cases}$ (48)
$\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}T(\epsilon,f,\mathcal{A}_{\mathrm{GPUCB}})=\begin{cases}O\left(\left(\frac{1}{\epsilon}\right)^{2+d/\nu}\right),&\text{for Matérn kernels,}\\ O\left(\frac{1}{\epsilon^{2}}\cdot\left(\log\frac{1}{\epsilon}\right)^{d+3}\right),&\text{for SE kernels.}\end{cases}$ (49)
wang2023regret’s approach to proving the regret optimality of GP-UCB follows the two-step framework. The first critical component is to construct a high-probability uniform error bound that quantifies the difference between the true objective function $f(x)$ and its posterior mean estimate $\mu_{t}(x)$ at any given step $t$. This bound takes the form $|f(x)-\mu_{t}(x)|\leq\sqrt{\beta_{t}}\sigma_{t}(x)$ for all $x\in\mathcal{D}$, where $\sigma_{t}(x)$ is the posterior standard deviation and $\beta_{t}$ is a carefully chosen exploration parameter. The second component involves bounding the cumulative sum of the instantaneous regrets, which, under the uniform error bound, can be related to the sum of the posterior variances $\sum_{t=1}^{T}\sigma_{t-1}^{2}(x_{t})$. This sum is, in turn, bounded by the maximal information gain, $\gamma_{T}$. The final regret bound is thus a function of $T$, $\beta_{T}$, and an upper bound on $\gamma_{T}$. wang2023regret employs tools from empirical process theory and decomposes the estimation error, $f(x)-\mu_{t}(x)$, into a bias term and a random error term. By leveraging the connection between Gaussian process regression and kernel ridge regression, along with properties of the Fourier transform for stationary kernels, they show that the bias term is bounded by $f||_{\mathcal{N}_{\Psi}(\mathcal{D})}\sigma_{t}(x)$. The more challenging random error term is handled by viewing it as an empirical process indexed by a class of functions and bounding its supremum. They bound the $\epsilon$-entropy of this function class, which allows for a high-probability bound on the random error. With the new uniform error bound, we can select a much smaller, dimension-independent exploration parameter $\beta_{t}$ that grows only logarithmically with $T$. Combining this sharper $\beta_{t}$ with the tightest known bounds on the maximal information gain $\gamma_{T}$ for Matérn and SE kernels, we can achieve the tight cumulative regret of GP-UCB.

### v2

@@
 Abstract
-Large Language Models (LLMs) have recently enabled self-improving AI, i.e., AI that iteratively generates, evaluates, and refine its own outcomes.
-Recent studies have shown that self-improving AI focusing on prompt optimization (i.e., context engineering) can outperform state-of-the-art reinforcement-learning fine-tuned LLMs.
+Large Language Models (LLMs) have recently enabled self-improving AI, i.e., AI that iteratively generates, evaluates, and refines its own outcomes.
+Recent studies have shown that self-improving AI focusing on prompt optimization (i.e., context engineering) can outperform state-of-the-art reinforcement learning fine-tuned LLMs.
 Here, their ‘performance’ is typically measured by query efficiency—the number of LLM-generated solution samples required to reach a specified performance threshold.
@@
 However, the difficulty of directly estimating suitable acquisition functions in LLMs’ minds makes this extension challenging.
-This paper overcomes this challenge by proving that the combination of the simple and widely used Best-of-N selection strategy and simple textual gradients (i.e., textual edits from a critic model) statistically emulates the behavior of the gradients on the canonical Upper Confidence Bound (UCB) acquisition function, which induces optimal exploration in terms of evaluation efficiency.
+This paper overcomes this challenge by proving that the combination of the simple and widely used Best-of-N selection strategy and simple textual gradients (i.e., textual edits from a critic model), statistically emulates the behavior of the gradients on the canonical Upper Confidence Bound (UCB) acquisition function, which induces optimal exploration in terms of evaluation efficiency.
 Based on this result, we propose TextGrad-Best-of-N Bayesian Optimization (T-BoN BO), a simple and eval-efficient language-space Bayesian optimization framework for AI self-improvement.
@@
 We briefly describe this paradigm using the example of ad-optimization; see the left panel of Figure 1.
-The process starts with a manager generating a set of ad creatives based on a creative brief.
+The process starts with generating a set of ad creatives based on a creative brief.
 This task is often done by ad agencies or marketers, who rely on their understanding of the product, theories of consumer behavior, market conditions, and feedback from prior ad campaigns.
@@
 Although effective in many settings, this human-driven workflow suffers from drawbacks.
-First, the speed with which firms can iterate over this process depends critically on how quickly humans can learn from the performance of prior solutions and propose new candidate solutions.
+First, the speed with which firms can iterate over this process depends critically on how quickly humans can generate new candidate solutions.
 Historically, generating new candidate solutions (or ads in the case of our example) has been time-consuming, and this step formed a key bottleneck in speeding up the process.
-Further, there exists no formal framework for how human agents can learn from the performance of prior solutions and leverage them to propose better candidates, and this step is often left to managerial intuition.
-Indeed, much of the research and managerial focus on data-driven business solutions has largely centered on the evaluation step, i.e., given a set of candidate solutions/options, how can we identify the most effective one – and solutions such as A/B tests and Multi-Armed Bandits (MAB) experiments have been proposed and tested (kohavi2020trustworthy; fiez_etal_2024); but how to leverage insights from prior rounds of evaluation systematically and improve the set of candidate solutions remains an unsolved problem.
-Ad generation (Step 1)(Weeks)Analysis (Step 3)(days)Ad evaluation (Step 2)(days)HumanBottleneckAd generation (Step 1)(seconds)Analysis (Step 3)(minutes–hours)Ad evaluation (Step 2)(days)AIBottleneck
+Further, there exists no formal framework for how human agents can analyze and learn from the performance of prior solutions and leverage them to propose better candidates, and this step is often left to managerial intuition.
+Indeed, much of the research and managerial focus on data-driven business solutions has largely centered on the evaluation step, i.e., given a set of candidate solutions/options, how can we identify the most effective one – and solutions such as A/B tests and Multi-Armed Bandits (MAB) experiments have been proposed and tested (kohavi2020trustworthy; fiez_etal_2024).
+Ad generation (Step 1)(Weeks)Analysis (Step 3)(Days)Ad evaluation (Step 2)(Days-weeks)HumanBottleneckAd generation (Step 1)(seconds)Analysis (Step 3)(minutes–hours)Ad evaluation (Step 2)(Days-weeks)AIBottleneck
 Figure 1: Comparison of human-driven and AI-driven ad self-improvement cycles in digital advertising.
@@
 First, candidate generation is fast and scalable with generative AI models.
-For example, in the case of digital ads, generative models can produce high-quality creative outputs, including ad copy, images, and complete campaign concepts that adhere to brand guidelines in a matter of minutes (jansen2024automated; hartmann2025power) (see right panel of Figure 1).
+For example, in the case of digital ads, generative models can produce high-quality creative outputs, including ad copy, images, and complete campaign concepts that adhere to brand guidelines in a matter of minutes (jansen2024automated; hartmann2025power).
 [^1] Indeed, a substantial share of creatives is now generated by AI, reflecting improvements in both quality and speed (coffee2025aiadbuying; Majic2025VibeMarketing).
 Second, Large language Models (LLMs) and related generative systems can now conduct sophisticated data analysis, summarizing performance patterns and identifying potential drivers of success (ferber2024context; guo2024ds; ghosh2024exploring; wiedemer2025videomodelszeroshotlearners), providing a pathway for analyzing historical data and improving candidate solutions.
-As such, with the new capabilities enabled by AI, the traditional bottlenecks of analysis and generation of candidate solutions are significantly reduced.
-Given these developments, there is immense interest in the development and deployment of self-improving AI (silver2025welcome), which automates the complete analysis–generation–evaluation loop at scale, allowing systems to update themselves continuously in response to emerging data (MantiaChatterjeeLee2025; chen2025xbench).
+As such, with the new capabilities enabled by AI, the time required for previous data analysis and generation of new candidate solutions are significantly reduced.
+Given these developments, there is immense interest in the development and deployment of AI-based automation of the complete generation–evaluation-analysis loop at scale, allowing systems to update themselves continuously in response to emerging data (See right panel of Figure 1).
+In this loop, generative models produce new solution candidates, data-science agents conduct evaluation of the solution candidates, and multi-modal reasoning LLM-based analytical agents interpret the evaluation results and synthesize insights to guide the next generation step.
+This type of automated cycle running system is often referred to as the self-improving AI or self-evolving AI (MantiaChatterjeeLee2025; chen2025xbench; silver2025welcome).
+This transformation has an important implication for practice: it dramatically lowers the marginal cost of iteration, making continuous experimentation feasible not only for large firms or organizations with dedicated analytics teams but also for small and medium-sized firms that historically lacked the creative and analytical capacity to run many cycles in parallel.
+When generation and analysis become cheap and largely automated, even resource-constrained organizations can maintain always-on optimization loops that adapt campaigns to evolving market conditions and consumer responses.
+As generation and analysis become increasingly automated and inexpensive in this self-improving loop, the bottleneck in data-centric decision making shifts from proposing and interpreting candidates to evaluating them in the real world.
+Even if analysis can be performed quickly and new candidate solutions can be generated almost instantaneously, each iteration still has to be validated by interacting with the real environment and collecting outcome data.
+The effective pace of the cycle is therefore governed by evaluation-related constraints such as available traffic, experimentation budgets, and the amount of data required to obtain statistically reliable comparisons among candidates, rather than by generation or analysis speed.
+For instance, in digital advertising, even when generating ad creatives via LLM queries is super fast, but evaluating an ad’s effectiveness still requires significant human feedback or large-scale experimentation (johnson2023inferno).
+A similar pattern arises in other scientific and engineering domains: protein structures or sequences can be proposed rapidly by models such as AlphaFold (jumper2021highly), but their functional properties must still be confirmed through slow and noisy wet-lab assays; and chip layouts can be generated in silico by systems such as AlphaChip (goldie2024chip), yet rigorous evaluation depends on physical fabrication and benchmarking.
+As generation and analysis become increasingly automated and inexpensive, the evaluation step thus emerges as the primary bottleneck that limits the overall cadence of self-improvement.
+As such, any societal and business AI system we design needs to be evaluation-efficient, i.e., reach a high-performing solution using as few costly evaluations as possible, rather than optimizing for a few inexpensive generations.
+1.2 Prompt optimization-based self-improving AI
+A natural question is, then, how the computer science literature has instantiated self-improving AI in concrete algorithmic frameworks.
+To date, a central line of success in self-improving AI has arisen in the domain of iterative prompt optimization, where AI agents iteratively refine their own prompts in response to performance feedback.
+agrawal2025gepa showed that iterative prompt optimization-based self-improving AI can be much more efficient than state-of-the-art model fine-tuning techniques such as GRPO (liu2024deepseek).
+Specifically, given any limited budget for the number of LLM generation query calls, their proposed iterative prompt optimization method, named GEPA (Genetic-Pareto), outperformed GRPO across their experiments.
+Those experiments included benchmarks such as complex questions across multiple Wikipedia articles, strict instruction-following under constraints, retrieval-augmented fact verification to support or refute claims, and maintaining answer quality while avoiding privacy leakage.
+In all those benchmarks, the primary efficiency metric was generation efficiency.
+i.e., the number of LLM query calls required to generate candidate answers to reach a target performance level.
+Likewise, prior literature on prompt-based self-improving AI (agrawal2025gepa; zhang2025agentic; wang2025maestro) has focused on measuring efficiency in terms of generation efficiency, not evaluation efficiency.
+Indeed, leading approaches are often specified as procedural heuristic search routines which finds near-optimal solution with small computations, e.g., population-based or random-search-style algorithms that treat prompts as individuals in an evolutionary process or as configurations in a heuristic pipeline.
+Specifically, they often maintain a pool or batch of prompts, repeatedly generate new variants via LLM-driven mutation, crossover, reflection, or trace-based resampling, score these candidates on held-out data or rollouts, retain the top-performing or Pareto-optimal prompts, and iterate this generate–score–select–mutate loop under a budget of LLM forward passes (10.5555/3692070.3692611; guo2024connecting; agrawal2025gepa; better_together).
+This mirrors the design philosophy of classical evolutionary and heuristic optimization, where the quality of the best solution is assessed as a function of the computation budget rather than the evaluation budget (nguyen2016understanding; ding2023quality; hu2024automated; liu2024evolution).
+Consequently, evaluation efficiency objective has been largely overlooked in the computer science literature.
+However, as we discussed earlier, in self-improving AI for real-world societal and business problems, the primary bottleneck is no longer generating new solutions efficiently (generation efficiency) but evaluating them efficiently (evaluation efficiency).
+This shift in the dominant constraint calls for rethinking prompt-optimization-based self-improving AI frameworks so that they explicitly treat evaluation budget as the scarce resource to be allocated and optimized over.
 1.3 Research agenda and Challenges
-In this paper, we study a prompt-optimization-based self-improving AI framework for societal and business problems, motivated by recent successes of prompt-optimization-based self-improving AI algorithms that outperform state-of-the-art reinforcement learning fine-tuning methods (agrawal2025gepa).
-While the use of modern generative AI systems has reduced the overhead associated with content generation and analysis (Steps 1 and 3), there are still three challenges that we need to overcome to design an efficient and effective self-improving AI that can automate decision-making in societal and business problems.
+Motivated by the effectiveness of iterative prompt-optimization-based self-improving AI and the overlooked importance of evaluation efficiency in existing frameworks, we now turn to our central question: how can we design a prompt-optimization-based self-improving AI framework that is evaluation-efficient?
+In addition to evaluation efficiency, we would like the framework to be easy to implement, theoretically grounded, and empirically perform well.
+However, realizing these desiderata in a unified framework is nontrivial.
+The key insight in this paper is that prompt-optimization-based self-improving AI follows a familiar pattern we saw in the optimization framework called Bayesian Optimization (BO) (frazier2018tutorial; srinivas2012information; papenmeier2025understanding).
+This insight is grounded in three observations: First, LLMs learn in-context from previous evaluations (brown2020language; coda2023meta; lu2024emergent; cahyawijaya2024llms), acting as Bayesian learners (xie2021explanation; wang2023large; wakayama2025context).
+Second, we iteratively choose a candidate prompt that generates a candidate solution, observe a noisy performance score associated with this solution, and then decide where to query next.
+Third, we have to be cognizant of three concerns: (1) we have no closed-form model linking prompts to scores, (2) we need to balance exploration and exploitation when choosing the next candidate solution, and (3) the goal is to optimize for the number of evaluations rather than candidate generation.
+BO is a principled framework for tackling such black-box optimization problems with such features.
+In BO, we build a function called the acquisition function and choose the point with the top of the acquisition function as the next point to evaluate.
+Among many variants of BO, Upper Confidence Bound (UCB)-BO (srinivas2012information), a version of BO with UCB acquisition function, is proven to be evaluation-efficient, in the sense that no algorithm can achieve better regret in terms of both simple regret and expected cumulative regret (whitehouse2023sublinear; wang2023regret).
+Figure 2 illustrates the procedure of UCB-BO.
+Figure 2: An example of a non-parallel UCB-BO procedure for four optimization steps.
+At each step, we choose the maximum (or multiple local maxima in the parallel UCB-BO) of the blue UCB function as the next point to evaluate.
+As you can see from Figure 2, the UCB acquisition function is often non-concave.
+To resolve this issue, in practice, we often use parallel gradient-based UCB-BO (wilson2018maximizing), where we 1) start from multiple initial points and take gradients to find a set of local optima, and 2) choose those local optima as the points to evaluate next.
+Parallel gradient-based BO has shown surprising empirical effectiveness for high-dimensional problems (papenmeier2025understanding).
+Building on the key insight, we propose to achieve evaluation efficiency by casting iterative prompt optimization-based self-improving AI as parallel gradient-based UCB Bayesian optimization in language space.
+At each iteration, we treat the prompts from the previous round as initialization points and take gradient steps to ascend the UCB acquisition function toward its local optima.
+We then select which prompts to evaluate from among these local optima.
+After each evaluation round, the LLM and its associated critic models implicitly update a surrogate belief over prompt quality via in-context learning and corresponding UCB, analogous to the posterior update in classical Bayesian optimization.
+However, there are two challenges we need to overcome.
   • $\bullet$
-    Challenge 1 – Evaluation-efficiency
-    First, in many real-world societal and business problems, the primary bottleneck is no longer generating new solutions efficiently (query efficiency) but evaluating them efficiently (evaluation efficiency).
-For instance, in the case of digital advertising, generating ad creatives using LLM queries is super fast, but evaluating an ad’s effectiveness requires significant human feedback or large-scale experimentation (johnson2023inferno).
-As such, any AI system that we design needs to be evaluation-efficient, i.e., reach a high-performing solution using as few costly evaluations as possible, rather than trying to optimize for a few inexpensive generations as possible.
-While there is a nascent but growing literature in computer science on self-improving AI systems, this literature mainly focuses on generation efficiency as the main bottleneck (agrawal2025gepa), and has largely overlooked evaluation efficiency.
+    Challenge 1– Lack of gradient functions
+    In prompt spaces, the notion of a gradient is ill-defined because they lack the inner-product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction).
+Thus, we first need to define a meaningful notion of a “gradient” in the domain of text.
+Although one might consider going around this challenge by embedding the prompt first, taking the gradient there, and reverting the prompt, this strategy runs into the well-known soft-to-discrete projection problem, where such reverted prompts break any clean correspondence between infinitesimal moves in embedding space and interpretable edits of the underlying natural-language prompt (cui2025automatic).
   • $\bullet$
-    Challenge 2 – Lack of gradient functions
-    Second, notice that we are essentially solving an optimization problem in language space, i.e., we want to take an input prompt, evaluate its performance, and improve it to reach a higher performance.
-[^2] Self-improving AI can thus be viewed as principled prompt optimization or context engineering since it involves improving textual inputs/prompts based on historic performance data (zhou2022large; khattab2023dspy; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro).
-In standard numerical settings, the natural way to solve such problems is to use gradient functions.
-However, in language spaces, the notion of a gradient is ill-defined because they lack the inner-product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction).
-Thus, we first need to define a meaningful notion of a “gradient” in the domain of text.
-  • $\bullet$
-    Challenge 3 – Balancing exploration and exploitation
-    Third, while it is tempting to move fully along the textual gradient that worked well previously, i.e., exploit the knowledge gained from prior prompts, this can be sub-optimal due to uncertainty about that knowledge.
-Thus, we also need to bake some exploration in the choice of gradient direction.
-Again, while there exist bandit-based approaches that balance exploration and exploitation in the case where we have a fixed candidate set and the exploration space is numerical (jamieson2014best; russo2016simple; russo2018tutorial; mason2020finding; simchi2023multi; fiez_etal_2024), these do not automatically translate to our setting since our goal is to explore in the language space and our exploration focuses on changing the set of candidate solutions rather than deciding how much to explore each option within a fixed candidate set.
-Thus, a key challenge is to account for uncertainty to guide the search over the language space.
-Canonical uncertainty measures, such as the Upper Confidence Bound (UCB) or the posterior distribution, are not readily available from off-the-shelf LLMs, making their direct application to language space non-trivial.
-Further, we note that self-consistency variance or embedding distances are poor proxies, and training a Gaussian Process over high-dimensional text embeddings is brittle, difficult to implement, expensive, and data-hungry.
-In this paper, our goal is to address the above challenges and build an evaluation-efficient self-improving AI that effectively balances exploration and exploitation in the language space.
-1.3 Our Framework and Algorithm
-We start with the observation that prompt optimization for self-improving AI follows a familiar pattern – we iteratively choose a candidate prompt that generates a candidate solution, observe a noisy performance score associated with this solution, and then decide where to query next while minimizing evaluations.
-Throughout, we have to be cognizant of three concerns: (1) we need to balance exploration and exploitation when choosing the next candidate solution, (2) we have no closed-form model linking prompts to scores, and (3) the main scarce resource is the number of evaluations rather than candidate generation.
-Our key insight is that these features of our problem are similar to the regime in which classic Bayesian Optimization (BO) operates successfully (frazier2018tutorial; srinivas2012information; papenmeier2025understanding).
-BO is a principled framework for tackling black-box optimization problems where evaluation efficiency is an important consideration.
-(See Web Appendix E for a detailed theoretical discussion of BO.) Thus, we seek to extend BO from numerical inputs to the space of natural language prompts to design an evaluation-efficient self-improving AI.
-At its core, BO assumes a surrogate model (e.g., Gaussian Process or Reproducing Hilbert Kernel Space) to approximate the objective function and provide uncertainty estimates of that approximation.
-From this surrogate, BO constructs a function often called the acquisition function and selects as the next candidate the point with the highest acquisition value.
-One of the most widely used BO methods is Bayesian Optimization with Upper Confidence Bound (UCB) acquisition function (srinivas2012information).
-In practice, since the global maximizer of the UCB acquisition function is often not available in a closed form, we typically use gradient-based BO with UCB acquisition function (wilson2018maximizing).
-Gradient-based BO methods have shown surprising effectiveness for high-dimensional problems when the gradient-based BO is adapted to avoid vanishing gradients, for example, through random multi-start (where each initialization yields a local maximum of UCB) and careful gradient sampling strategies (papenmeier2025understanding).
-However, the standard gradient-based UCB BO is not directly applicable to the language space, as the notion of a gradient is ill-defined because they lack the inner-product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction).
-Therefore, we first need to define a notion of gradients in the language space and devise an appropriate acquisition function in the text space.
-We now describe how we modify and adapt the BO framework to address the challenges discussed earlier.
-We propose a simple, evaluation-efficient iterative prompt-optimization framework that we call TextGrad-Best-of-N Bayesian Optimization (T-BoN BO) that addresses all three challenges discussed earlier.
-First, the BO framework naturally addresses the first challenge of the evaluation efficiency.
-It is known that BO with UCB acquisition function (srinivas2012information) is a theoretically evaluation-efficient algorithm, i.e., optimally balances exploration and exploitation, minimizing regret (whitehouse2023sublinear; wang2023regret).
-Specifically, for both expected cumulative regret and simple regret, the regret upper bound of BO with UCB acquisition function tightly matches the regret lower bound of scarlett2017lower.
-To solve the second challenge, T-BoN BO builds on the recently proposed TextGrad (textgrad) framework that adapt gradient descent to discrete text by asking an external auxiliary LLM critic to propose targeted edits that improve a candidate; applying these edits induces a local, directional update, i.e., a “textual gradient”, in language space.
+    Challenge 2 – Taking gradients for UCB
+    Second, even if we can define meaningful gradients in prompt space, we do not have direct access to the underlying UCB acquisition function whose gradient we would like to follow.
+The surrogate posterior over prompt quality that induces UCB lives implicitly in the LLM, rather than as an explicit Gaussian-process-style model that exposes closed-form posterior means, variances, and their derivatives.
+Consequently, standard gradient-based BO techniques that differentiate the UCB constructed from a known surrogate cannot be applied naively: we must design a mechanism that elicits UCB-like, optimism-in-the-face-of-uncertainty behavior using only black-box text interactions and critic feedback, without ever computing the UCB function or its gradient in closed form.
+In this paper, our goal is to address the above challenges and build an evaluation-efficient self-improving AI that optimally balances exploration and exploitation in the language space.
+1.4 Proposed Method
+Figure 3: Gradient steps of parallel gradient-based BO (Left image) and those of T-BoN BO (Right image). T-BoN BO extends parallel gradient-based BO to the language space.
+We propose a simple, evaluation-efficient iterative prompt-optimization framework, called TextGrad-Best-of-N Bayesian Optimization (T-BoN BO), that adapts parallel gradient-based UCB-BO for prompt optimization based self-improving AI (Figure 3), addressing both of the challenges discussed earlier.
+To solve the first challenge, T-BoN BO builds on the recently proposed TextGrad (textgrad) framework that adapts gradient descent to discrete text by asking an external auxiliary LLM critic to propose targeted edits that improve a candidate; applying these edits induces a local, directional update, i.e., a “textual gradient”, in language space.
 This makes gradient-style search feasible without continuous relaxations or mapping the LLM prompts to the embedding space.
+For instance, consider an ad creative generation prompt for an image model: “A simple photo of a plant-based burger on a white plate.” A critic LLM might suggest: “Make the burger appear delicious, and show it being enjoyed at a barbecue.” Incorporating this edit yields: “A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends.” This change is analogous to following a gradient direction—small, constructive, and aligned with improving performance according to the evaluation signal.
 textgrad show that textual gradients based on their approach enable consistent improvements across diverse scientific domains such as code-completion benchmarks, improving accuracy on graduate-level QA datasets, optimizing drug-like molecules, and refining radiotherapy planning for clinical objectives.
 These results establish textual gradients as an effective surrogate for numerical gradients when optimizing systems that operate in language space.
-To address the third challenge, T-BoN BO chooses the next prompt to evaluate via a series of textual gradients which are selected under the Best-of-N principle (snell2024scaling).
+To address the second challenge, T-BoN BO chooses the prompt to evaluate next via a series of textual gradients which are selected under the Best-of-N principle (snell2024scaling), resulting in the gradient direction we call Best-of-N gradient.
 Moving in the textual gradient direction suggested by the critic exploits the expected improvement implied by past results but ignores uncertainty in those expectations.
-T-BoN bakes in exploration into the choice of gradient direction by adopting an optimism-based rule in the spirit of the optimism-in-the-face-of-uncertainty principle: it samples multiple local textual edits and moves along the direction the LLM judges most promising, i.e., the Best-of-$N$ candidate.
-We show that Best-of-$N$ over locally sampled textual edits statistically emulates gradient ascent on a UCB acquisition function, so that Best-of-$N$ effectively acts as an implicit optimistic acquisition maximizer in language space.
+Therefore, it does not climb the UCB acquisition function.
+Best-of-N gradient bakes in exploration into the choice of gradient direction by adopting an optimism-based rule in the spirit of UCB’s optimism-in-the-face-of-uncertainty principle: it samples multiple local textual edits and moves along the direction the LLM judges most promising, i.e., the Best-of-$N$ candidate.
+That is, the Best-of-$N$ gradient is designed to simulate the implicit UCB acquisition gradient in language space.
 We now briefly describe our T-BoN BO algorithm.
 At each optimization iteration, in Step 1 (meta-reflection phase), we use the critic model to analyze the history of previously evaluated solutions and their scores, and update a meta-reflection that summarizes what has tended to work well or poorly.
-In Step 2 (Best-of-$N$ textual-gradient phase), we perform a sequence of textual gradient updates: for each gradient step, (i) we propose $N$ textual-gradient edits that transform the current prompt into $N$ candidate prompts, (ii) each candidate prompt is passed to a solution-generation model to produce a corresponding candidate solution, (iii) conditioned on the meta-reflections, the critic model predicts which candidate solution is most likely to perform best, and (iv) we set the prompt associated with this predicted-best solution as the new current prompt.
+In Step 2 (Best-of-$N$ textual-gradient phase), we perform a sequence of Best-of-N gradient updates from the current prompt.
+That is, for each gradient step, (i) we propose $N$ textual-gradient edits that transform the current prompt into $N$ candidate prompts, (ii) each candidate prompt is passed to a solution-generation model to produce a corresponding candidate solution, (iii) conditioned on the meta-reflections, the critic model predicts which candidate solution is most likely to perform best, and (iv) we set the prompt associated with this predicted-best solution as the new current prompt.
 After iterating Step 2, we obtain a refined prompt and its associated solution to evaluate.
-In Step 3 (evaluation phase), we evaluate this solution to obtain a scalar score, which is then used by T-BoN BO to update its history and meta-reflection for subsequent optimization steps.
+In Step 3 (evaluation phase), we evaluate this solution to get a scalar score, which T-BoN BO then uses to update its history and meta-reflection for subsequent optimization steps.
 1.5 Theoretical Guarantees and Empirical Evaluation
 We provide theoretical guarantees on the performance of our approach.
-We formally show that, under mild regularity conditions, the Best-of-$N$ selection over locally sampled textual edits induces, in probability, ascent directions of the UCB acquisition function with an exploration parameter $\beta_{N}=\Theta(\sqrt{\ln N})$ (Theorem 2).
-In other words, T-BoN BO emulates gradient-based UCB Bayesian Optimization in the implicit embedding space induced by the LLM, even though it operates purely in language space and does not construct an explicit surrogate or uncertainty model.
+We formally show that, under mild regularity conditions, the Best-of-$N$ gradient over locally sampled textual edits induces, in probability, ascent directions of the UCB acquisition function with an exploration parameter $\beta_{N}=\Theta(\sqrt{\ln N})$ (Theorem 2).
+In other words, T-BoN BO emulates parallel gradient-based UCB Bayesian Optimization in the implicit embedding space induced by the LLM, even though it operates purely in language space and does not construct an explicit surrogate or uncertainty model.
 Since UCB-BO is known to achieve sublinear regret and optimal evaluation-efficiency guarantees, T-BoN BO inherits these theoretical properties up to constant factors and approximation error stemming from textual gradients.
-In sum, we formally establish that T-BoN BO is theoretically equivalent to the gradient-based implementation of Bayesian Optimization with the Upper Confidence Bound (UCB) acquisition function (srinivas2012information; wilson2018maximizing), which is proven to be evaluation-efficient (whitehouse2023sublinear; wang2023regret).
-This equivalence establishes T-BoN BO as a valid, eval-efficient, gradient-based Bayesian optimization procedure in the language space.
-In addition to its theoretical efficiency, T-BoN BO is simple and easy to implement since does not require any explicit uncertainty modeling, embeddings, or Gaussian process surrogates.
+In sum, we formally establish that T-BoN BO is theoretically equivalent to the parallel gradient-based implementation of Bayesian Optimization with the Upper Confidence Bound (UCB) acquisition function (srinivas2012information; wilson2018maximizing), which is proven to be evaluation-efficient (whitehouse2023sublinear; wang2023regret).
+This equivalence establishes T-BoN BO as a valid, eval-efficient, parallel gradient-based Bayesian optimization procedure in the language space.
+In addition to its theoretical efficiency, T-BoN BO is simple and easy to implement since it does not require any explicit uncertainty modeling, embeddings, or Gaussian process surrogates.
 We now demonstrate the empirical performance of our approach using a series of numerical experiments.
@@
 Section 2 discusses the related literature.
-Section 3 describes the preliminaries related to the ideas discussed in this paper.
 Section 3 formalizes the AI system and its objective, along with self-improving algorithms.
-Section 5 introduces the main algorithm proposed in this paper, T-BoN BO, and its components.
-Section 6 proves that T-BoN BO theoretically emulates the gradient-based UCB Bayesian Optimization.
-Sections  7 and 8 empirically demonstrate T-BoN BO’s effectiveness on ad-alignment tasks versus strong baselines.
+Section  4 describes the concept of Best-of-N gradient and proves that it theoretically emulates the gradient of the UCB acquisition function.
+Section 5 introduces the main self-improving AI algorithm proposed in this paper, T-BoN BO.
+Sections 6 and 7 empirically demonstrate T-BoN BO’s effectiveness on ad-alignment tasks versus strong baselines.
 Finally, Section 8 concludes with a discussion of implications and limitations.
@@
 We now discuss each one below.
-The literature on self-improving AI systems has often been motivated by an observation: the progress in AI based only on human data is hitting limits, so AI systems must create and learn from their own data.
-In other words, powerful AI systems “should have their own stream of experience that progresses, like humans, over a long time-scale” (silver2025welcome).
-That is, we need to design the self-improving AI system that iteratively 1) generate data of outcomes, evaluate outcomes with grounded signals, and 2) self-tune the three knobs –models, tools, and prompts– in the system to achieve long-term objectives under real-world feedback (wang2025maestro).
-Here, tuning models is about altering language models’ input-output correspondences (huang-etal-2023-large; lee2023rlaif; acikgoz2025self); tuning tools is about specifying the tools the AI system has access to, including memory tools (hou2025model; ouyang2025reasoningbank; qiu2025alita); tuning prompts is about engineering prompts in the model’s context window, including system messages, task instructions, constraints, schemas, and few-shot exemplars (zhou2022large; khattab2023dspy; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro).
-In this paper, we study self-improving systems, focusing on tuning prompts, which have recently attracted attention after agrawal2025gepa showed that iterative prompt optimization can be more efficient than state-of-the-art model fine-tuning techniques such as GRPO (liu2024deepseek).
-Specifically, given any limited budget for the number of LLM generation query calls, their proposed iterative prompt optimization method, named GEPA (Genetic-Pareto), outperformed GRPO across their experiments.
-Those experiments included benchmarks such as complex questions across multiple Wikipedia articles, strict instruction-following under constraints, retrieval-augmented fact verification to support or refute claims, and maintaining answer quality while avoiding privacy leakage.
-In all those benchmarks, the number of LLM query calls for generating candidate answers was considered as the primary performance bottleneck.
-Likewise, prior literature on prompt-based self-improving AI (agrawal2025gepa; zhang2025agentic; wang2025maestro) has focused on measuring efficiency in terms of generation efficiency, i.e., minimizing the number of LLM-generated rollouts.
-On the other hand, our paper proposes and focuses on the evaluation efficiency objective.
-That is, instead of focusing on reducing the number of LLM outcome-generation rollouts, we aim to reduce the number of evaluations required to achieve an effective outcome.
+The literature on self-improving AI has often been motivated by an observation: the progress in AI based only on human data is hitting limits, so AI must create and learn from its own data.
+In other words, powerful AI “should have their own stream of experience that progresses, like humans, over a long time-scale” (silver2025welcome).
+That is, we need to design the self-improving AI that iteratively 1) generate data of outcomes, evaluate outcomes with grounded signals, and 2) self-tune the three knobs –models, tools, and prompts– in the system to achieve long-term objectives under real-world feedback (wang2025maestro).
+Here, tuning models is about altering language models’ input-output correspondences (huang-etal-2023-large; lee2023rlaif; acikgoz2025self); tuning tools is about specifying the tools the AI has access to, including memory tools (hou2025model; ouyang2025reasoningbank; qiu2025alita); tuning prompts is about engineering prompts in the model’s context window, including system messages, task instructions, constraints, schemas, and few-shot exemplars (zhou2022large; khattab2023dspy; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro).
+In this paper, we focus on studying self-improving systems based on tuning prompts.
 Our paper also relates to the literature on relating LLMs to Bayesian Optimization (BO).
 A series of papers focuses on applying LLM to enhance BO for numerical optimization problems.
-liularge focuses on leveraging LLMs’ contextual understanding and few-shot learning capabilities to improve BO’s efficiency, enabling better warm-starting and surrogate modeling; agliettifunbo uses an LLM to discover new acquisition functions for Bayesian optimization.
+liularge and cisse2025language focus on leveraging LLMs’ contextual understanding and few-shot learning capabilities to improve BO’s efficiency, enabling better warm-starting and surrogate modeling; agliettifunbo uses an LLM to discover new acquisition functions for Bayesian optimization.
 The LLM writes candidate acquisition function formulas as code, which are evaluated on various optimization problems.
 Conversely, singh2025mechanistic applies BO to enhance LLM’s internal behavior to boost the model’s zero-shot and few-shot performance.
-In contrast to these papers, our paper focuses on considering the prompt optimization problem that arises in LLM-based self-improving AI system design as a Bayesian Optimization problem.
+In contrast to these papers, our paper focuses on considering the prompt optimization problem that arises in formulating LLM-based self-improving AI system design as a Bayesian Optimization problem.
 In line with this focus, schneider2025hyperbandbased and kong2025metapromptoptimizationllmbasedsequential formulate prompt optimization among the fixed set of candidate prompts as Bayesian optimization.
 While schneider2025hyperbandbased utilizes ideas from adversarial bandits, kong2025metapromptoptimizationllmbasedsequential feeds each prompt text into an embedding model, and the GP surrogate operates on those fixed embedding features to guide selection.
-In contrast to these approaches, our work focuses on self-improving AI systems that actively generate new candidate prompts as the process proceeds, rather than pre-specifying a fixed set of candidate prompts.
+In contrast to these approaches, our work focuses on self-improving AI that actively generate new candidate prompts as the process proceeds, rather than pre-specifying a fixed set of candidate prompts.
 In terms of method evaluation benchmarks, our paper relates to the methodological literature on ad optimization in digital marketing.
@@
 The main difference between bandit-based papers and ours is the existence of a pre-defined set of candidate arms.
-Self-improving AI systems focus on creating new solutions by analyzing previous results.
+Self-improving AI focuses on creating new solutions by analyzing previous results.
 On the other hand, bandit methods aim to balance exploration of a predefined set of candidate ads optimally.
 More broadly, there is also a difference in problem scope.
-Self-improving AI systems iterate among three steps: generation, evaluation, and analysis.
+Self-improving AI iterates among three steps: generation, evaluation, and analysis.
 In contrast, bandit methods focus on efficiently executing the evaluation step.
 Indeed, an AI self-improvement framework’s evaluation step may employ a bandit method to improve efficiency further.
-3 Preliminaries
-3.1 AI system.
-An AI system $\Phi$ can be formally defined as a feed-forward modular workflow, which consists of one or more interconnected Large Language Model (LLM) components.
-In line with prior work (better_together; khattab2022demonstrate; miprov2; langprobe; agrawal2025gepa), we define an AI system $\Phi$ as a tuple $(M,C,\mathcal{X},\mathcal{Y})$.
-Here, $M=\langle M_{1},\ldots,M_{|M|}\rangle$ represents the set of LLM-powered components, $C$ is the control logic that orchestrates the workflow and sequences the tasks, and $\mathcal{X}$ and $\mathcal{Y}$ define the system input space and the system output space.
-Each component $M_{i}=(\pi_{i},\theta_{i},\mathcal{I}_{i},\mathcal{O}_{i})$ is an independent unit where $\pi_{i}$ is the prompt, $\theta_{i}$ represents the fixed weights of the underlying language model, and $\mathcal{I}_{i},\mathcal{O}_{i}$ specify the component’s input space and output space.
-We denote the prompts in the whole system $\Pi=\left\langle\pi_{1},\ldots,\pi_{|M|}\right\rangle$.
-As the prompts $\Pi$ are what we system designers can control for the system $\Phi$, we denote $\Phi$ as the functional $\Phi(\cdot)$ when we would like to emphasize system $\Phi$’s dependency on prompts $\Pi$.
-Accordingly, we call $\Phi(\Pi)$ the ‘system outcome’ corresponding to prompt $\Pi$.
-Note that $\Phi(\Pi)$ is itself a function that maps $\mathcal{X}$ to $\mathcal{Y}$, i.e., $\Phi(\Pi):\mathcal{X}\mapsto\mathcal{Y}$.
-Figure 2 illustrates an example AI system with five LLM-powered components.
-Here, its control logic $C=\{\mathcal{X}\rightarrow\left\{M_{1},M_{2}\right\},M_{1}\rightarrow M_{3},M_{2}\rightarrow M_{4},\left\{M_{3},M_{4}\right\}\rightarrow M_{5}\rightarrow\mathcal{Y}\}$ describes that the the system input $\mathcal{X}$ becomes the component input for $M_{1}$ and $M_{2}$, $M_{3}$’s input is $M_{1}$’s output, $M_{4}$’s input is $M_{2}$’s output, $M_{5}$’s input is $M_{3}$’s output and $M_{4}$’s output, and the final output $\mathcal{Y}$ is the output of $M_{5}$.
-The system $\Phi$’s input is the prompt $\Pi=\langle\pi_{1},\pi_{2},\pi_{3},\pi_{4},\pi_{5}\rangle$.
-[…]
-Figure 2: An example AI system with control logic $C=\{\mathcal{X}\rightarrow\left\{M_{1},M_{2}\right\},M_{1}\rightarrow M_{3},M_{2}\rightarrow M_{4},\left\{M_{3},M_{4}\right\}\rightarrow M_{5}\rightarrow\mathcal{Y}\}$ and prompt $\Pi=\langle\pi_{1},\pi_{2},\pi_{3},\pi_{4},\pi_{5}\rangle$.
-3.2 TextGrad: Textual gradients in language space.
-A central question for extending gradient-based Bayesian optimization to text is how to define a usable notion of “gradient” in language.
-A recent influential paper by (textgrad) addresses this problem by proposing a framework called the TextGrad; and it has has attracted broad interest across multiple research communities.
-Instead of computing numerical derivatives, TextGrad asks a critic LLM to propose targeted edits that would improve a candidate text.
-These edits act as local, directional signals, i.e., textual gradients, that guide search in a discrete language space.
-Importantly, this approach does not require embedding relaxations or external surrogate models, making it directly applicable to prompt optimization and other text-centric tasks.
-For instance, consider an ad creative generation prompt for an image model: “A simple photo of a plant-based burger on a white plate.” A critic LLM might suggest: “Make the burger appear delicious, and show it being enjoyed at a barbecue.” Incorporating this edit yields: “A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends.” This change is analogous to following a gradient direction—small, constructive, and aligned with improving performance according to the evaluation signal.
-Empirical studies provided in textgrad show that textual gradients enable consistent improvements across diverse scientific domains.
-In programming tasks, they raise success rates on challenging code-completion benchmarks.
-In knowledge-intensive reasoning, they improve accuracy on graduate-level QA datasets.
-In applied domains, they have been used to optimize drug-like molecules (via text encodings) and to refine radiotherapy planning prompts for clinical objectives.
-These results establish textual gradients as an effective surrogate for numerical gradients when optimizing systems that operate in and over language, justifying our use of gradient-style search in the discrete text space.
-3.3 Bayesian Optimization.
-Bayesian Optimization (BO) is a framework for maximizing an unknown and expensive-to-evaluate objective function under a limited evaluation budget.
-Formally, consider a black-box function $f:\mathcal{Z}\to\mathbb{R}$, where in our setting $f$ corresponds to the system-level objective $J(\Pi)$ over prompts.
-BO maintains a surrogate model, typically a Gaussian Process (GP) or another probabilistic regressor defined on $\mathcal{Z}$, that summarizes current beliefs about $f$.
-For any candidate $z\in\mathcal{Z}$, the surrogate yields a posterior mean $\mu(z)$ and an associated uncertainty measure $\sigma(z)$.
-The next evaluation point is selected by maximizing an acquisition function $A(z)$ constructed from $\mu(z)$ and $\sigma(z)$, rather than by directly maximizing $\mu(z)$.
-Acquisition functions encode an explicit trade-off between exploitation (preferring points with high predicted value) and exploration (preferring points with high uncertainty).
-A widely used choice is the Upper Confidence Bound (UCB) acquisition function is:
-$A_{\beta}(z)\;=\;\mu(z)\;+\;\beta\,\sigma(z),$ (1)
-where the exploration parameter $\beta>0$ controls the strength of the exploration term (srinivas2012information; wilson2018maximizing).
-In many applications, the maximizer of $A_{\beta}$ is not available in closed form, and one instead employs gradient-based BO, which performs local ascent on $A_{\beta}$ (often from multiple random initializations) to obtain approximate maximizers (wilson2018maximizing; papenmeier2025understanding).
-In gradient-based BO, each evaluation is followed by multiple gradient steps that identify local optima of the acquisition function.
-Gradient-based BO methods have shown surprising effectiveness for high-dimensional problems under random multi-starts and careful choice of gradients (papenmeier2025understanding).
-In our setting, we do not construct an explicit surrogate model; rather, we interpret $\mu(\cdot)$ and $\sigma(\cdot)$ as implicit beliefs realized within the LLM and design our method to emulate the behavior of gradient-based UCB in this implicit space.
-4 Overview of Self-Improving AI Algorithms
+3 Setup and Background
 3.1 The optimization objective.
-Consider a system designer’s decision making problem, where she chooses the prompt set $\Pi$ for the LLM-based AI system $\Phi$.
-[^4] We call $\Pi$ in a plural form ‘prompts’ or ‘prompt set’ instead of a singular form ‘prompt’, as there are often multiple prompt components in one system.
-See AI system paragraph in Section 3.1 for a detailed description of an AI system.
+Consider a system designer’s decision-making problem, where she chooses a prompt $\Pi$ for the LLM-based AI system $\Phi$.
 [^3] LLM-based AI systems can also be modified by tuning the LLM’s internal parameters using LLM fine-tuning methods such as GRPO (liu2024deepseek).
 In this paper, we focus on prompt optimization, which fixes those internal parameters throughout.
-We call $\Phi(\Pi)$ the ‘system outcome’ of $\Phi$ corresponding to a prompt set $\Pi$.
-In a digital marketing setting, $\Phi$ is an image generation AI, and $\Phi(\Pi)$ is the ad generated by the prompt set $\Pi$.
+We call $\Phi(\Pi)$ the ‘system outcome’ of $\Phi$ corresponding to a prompt $\Pi$.
+In a digital marketing setting, $\Phi$ is an image generation AI, and $\Phi(\Pi)$ is the ad generated by the prompt $\Pi$.
 As a system, $\Phi(\Pi)$ has inputs and corresponding outputs.
@@
 This motivates the use of an external LLM that can analyze prior data and propose directions for improvement without explicitly computing gradients, leading to self-improving AI algorithms we describe below that iteratively optimize for $J$.
-4.2 Critic LLM and Self-improving AI algorithm
-Given a scoring rule $r$, a self-improving AI algorithm $\mathcal{A}$ optimizes $J(\Pi)$ through iteratively analyzing the previous history, proposing new prompts, and observing their scores.
-This is often enabled by assuming an external critic LLM, denoted $M_{\text{critic}}$, that can analyze previous data and propose the next solution.
-Critic LLM.
-Specifically, a critic LLM is assumed to have the following capabilities:
-  • 1)
-    The critic LLM can look at $\Phi(\Pi_{1})$ and $\Phi(\Pi_{2})$ side by side and predict which one will be better.
-  • 2)
-    The critic LLM can read the history of past attempts and summarize the main lessons as a short list of rules or guidelines.
-  • 3)
-    The critic LLM can suggest a small set of prompt edits that are likely to improve the current prompt.
-In the digital-marketing example, these correspond to an LLM’s ability to (1) compare two ad creatives and predict which will perform better (calderon2025alternative), (2) examine past performance data and derive guidelines for effective ads in context (jiang2024many; ferber2024context), and (3) propose prompt improvements based on those guidelines (textgrad; li2025test).
-Self-improving AI algorithm
-Given an initial prompt defined as $\Pi_{0}$, a self-improving AI algorithm proceeds in discrete optimization iterations, indexed by $t=1,2,\ldots$.
-At the beginning of iteration $t$, the algorithm has access to the history of past prompts and their observed scores and then selects a new prompt set $\Pi_{t}$, deploys the corresponding system $\Phi(\Pi_{t})$, and receives feedback about its performance.
-At each algorithm iteration $t$, we define the score $s_{t}$ of prompts $\Pi_{t}$ as the empirical average reward at $t$, i.e.,
+3.2 Self-improving AI algorithm
+Given a scoring rule $r$, a self-improving AI algorithm $\mathcal{A}$ optimizes $J(\Pi):=\mathbb{E}[r(y(\Phi(\Pi),x))]$ through iteratively analyzing the previous history, proposing a new prompt, and observing its noisy score.
+Formally, starting from the initial prompt $\Pi_{0}$, a self-improving AI algorithm proceeds in discrete optimization iterations, indexed by $t=1,2,\ldots$; at the beginning of iteration $t$, the algorithm has access to the history of past prompts and their observed scores, then selects a new prompt $\Pi_{t}$, deploys the corresponding system outcome $\Phi(\Pi_{t})$, and receives a score feedback on its performance.
+We define the score $s_{t}$ of prompts $\Pi_{t}$ as the empirical average reward at $t$, i.e.,
 $s_{t}:=\frac{1}{L}\sum_{l=1}^{L}r\left(y\left(\Phi\left(\Pi_{t}\right),x_{t}^{(l)}\right)\right),\quad x_{t}^{(l)}\sim\mathcal{D}_{\mathcal{X}}\left(\Pi_{t}\right),$ (3)
-where $L$ denotes the number of input samples of $x_{t}^{(l)}$ we evaluate for each $t$. Note that we compare the performance of algorithms in terms of the progress of $s_{t}$ across iterations $t$.
-Given access to a history of past evaluations $H_{t-1}=\{(\Pi_{\tau},s_{\tau})\}_{i=\tau}^{t-1}$ where $s_{\tau}$ is the empirically evaluated score for the prompts $\Pi_{\tau}$, a self-improving AI algorithm is defined as an algorithm $\mathcal{A}$ iterates the following steps at each $t$ to optimize for $J(\Pi)$:
+where $L$ denotes the number of input samples $x_{t}^{(l)}$ we evaluate for each $t$. Note that we compare the performance of algorithms in terms of the progress of $s_{t}$ across iterations $t$.
+Definition 1 (Self-improving AI algorithm).
+Fix a system $\Phi$ and a scoring rule $r$. For each iteration $t\geq 1$, let the history of past evaluations be
+$H_{t-1}\;:=\;\{(\Pi_{\tau},c_{\tau},s_{\tau})\}_{\tau=0}^{t-1}$ (4)
+where $s_{\tau}$ is the empirically evaluated score of $\Pi_{\tau}$, defined as
+$s_{t}=\frac{1}{L}\sum_{l=1}^{L}r\bigl(y(\Phi(\Pi_{t}),x_{t}^{(l)})\bigr),\quad x_{t}^{(l)}\sim\mathcal{D}_{\mathcal{X}}(\Pi_{t}).$ (5)
+Then, a self-improving AI algorithm is defined as function $\mathcal{A}$ that, at each iteration $t$, maps the history $H_{t-1}$ to a new prompt $\Pi_{t}$, i.e.,
+$\Pi_{t}=\mathcal{A}_{t}(H_{t-1})$
+The goal of $\mathcal{A}$ is to choose the sequence $(\Pi_{t})_{t\geq 0}$ so as to optimize the objective $J(\Pi)$, equivalently to drive the sequence of empirical scores $(s_{t})_{t\geq 1}$ upward over iterations in expectation.
+Operationally, given access to the history $H_{t-1}$, a self-improving AI algorithm $\mathcal{A}$ goes through following three steps each iteration $t$:
   1. 1.
@@
   3. 3.
-    Evaluation: The system $\Phi(\Pi_{t})$ is deployed, and its performance is measured, yielding a score $s_{t}$. The new observation is added to the history: $H_{t}=H_{t-1}\cup\{(\Pi_{t},s_{t})\}$.
-[…]
-Figure 3: A self-improving AI algorithm implemented on the example AI system in Figure 2.
-Figure 3 illustrates how the example AI system described in Figure 2 is positioned in the self-improving AI.
-5 TextGrad-Best-of-N Bayesian Optimization (T-BoN BO)
-We now describe the self-improving AI framework proposed in this paper, TextGrad-Best-of-N Bayesian Optimization (T-BoN BO) (Algorithm 1).
-Along with the system to optimize $\Phi$ and the scoring rule $r$, it requires a critic model $M_{\text{critic}}$.
-Algorithm.
-T-BoN BO starts from an initial prompt $\Pi_{0}$.
-Initially at iteration $t=0$, it gets the initial system $c_{0}=\Phi(\Pi_{0})$ and evaluates it on $\mathcal{D}_{\mathcal{X}}(\Pi_{0})$ to obtain initial score $s_{0}=\mathrm{Eval}(c_{0},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$.
-The history and the reflection is initialized as $H_{0}=\{(\Pi_{0},c_{0},s_{0})\}$ and $R_{0}=\emptyset$.
-We first fix the hyperparameters: the number of candidates per gradient step ($N$), the number of gradient steps per iteration ($G$), and the number of iterations ($T$).
-Then at each iteration $t$, it first sets $\Pi_{t,0}$ as $\Pi_{t-1}$, the previous iteration’s prompt, and proceeds as follows:
-  • 1.
-    Best-of-N Textual Gradient steps. (Detailed in Section 5.1) Repeat steps 1) – 4) for $g=1,\ldots,G$:
+    Evaluation: The system outcome $\Phi(\Pi_{t})$ is deployed, and its performance is measured, yielding a score $s_{t}$.
+The new observation is added to the history: $H_{t}=H_{t-1}\cup\{(\Pi_{t},c_{t},s_{t})\}$.
+3.3 Bayesian Optimization (BO)
+As discussed in Section 1, we would like to formulate self-improving AI as parallel gradient-based UCB-BO.
+At its core, BO maximizes an unknown and expensive-to-evaluate objective function under a limited evaluation budget.
+Formally, consider a black-box function $f:\mathcal{Z}\to\mathbb{R}$, where in our setting $f$ corresponds to the system-level objective $J(\Pi)$ over prompts.
+BO maintains a surrogate model, typically a Gaussian Process (GP) or another probabilistic regressor defined on $\mathcal{Z}$, that summarizes current beliefs about $f$.
+For any candidate $z\in\mathcal{Z}$, the surrogate yields a posterior mean $\mu(z)$ and an associated uncertainty measure $\sigma(z)$.
+In BO, the following evaluation point is selected by choosing the point that maximizes an acquisition function $A(z)$ constructed from $\mu(z)$ and $\sigma(z)$, rather than by directly maximizing $\mu(z)$.
+Acquisition functions encode an explicit trade-off between exploitation (preferring points with high predicted value) and exploration (preferring points with high uncertainty).
+A widely used choice is the Upper Confidence Bound (UCB) acquisition function
+$A_{\beta}(z)\;=\;\mu(z)\;+\;\beta\,\sigma(z),$ (6)
+where the exploration parameter $\beta>0$ controls the strength of the exploration term (srinivas2012information).
+Intuitively, UCB acquisition function gives a positive bonus to unexplored areas in addition to the posterior mean.
+This allows us to follow the “optimism in the face of uncertainty” principle, which is known to trade off exploration and exploitation effectively.
+Indeed, UCB-BO is an optimal algorithm for achieving evaluation efficiency, in the sense that no algorithm can achieve better regret guarantees (wang2023regret; whitehouse2023sublinear) (See Web Appendix E for a detailed theoretical discussion of BO.)
+In practice, since the maxima of the UCB acquisition function are often not available in a closed form, we typically use parallel gradient-based BO with the UCB acquisition function (wilson2018maximizing).
+In parallel gradient-based BO, we start from multiple initial points and take gradients from them to climb up the UCB acquisition function to find its local maxima.
+We then choose the points to evaluate among the local maxima (Figure 3).
+In our setting, we do not construct an explicit $\mu(\cdot)$ and $\sigma(\cdot)$ (and their corresponding surrogate model and acquisition function) for the prompt space; instead, we interpret $\mu(\cdot)$ and $\sigma(\cdot)$ as implicit beliefs within the LLM and design our method to emulate parallel gradient-based UCB in this implicit space.
+Textual gradient (textgrad) is designed to elicit gradients for $\mu(\cdot)$ in the prompt space without explicitly modeling those beliefs within the LLM.
+However, what we want is not gradients for $\mu(\cdot)$; we want gradients for $A_{\beta}(z)$, the UCB acquisition function.
+4 Best-of-N gradient and Theory
+As discussed in Section 1, a core challenge in casting self-improving AI as parallel, gradient-based UCB Bayesian optimization in the space of prompts is the absence of an explicit acquisition function (e.g., UCB) and its gradient.
+As we discussed earlier, textual gradients (textgrad) can guide the prompt toward the highest expected improvement based on immediate feedback, as a textual gradient elicits the gradient of the posterior mean $\nabla\mu(x)$.
+However, following the posterior mean’s gradient rather than UCB’s is a purely exploitative strategy that often leads to suboptimal local maxima.
+In this section, we propose that a gradient direction we call Best-of-N gradient, which generates $N$ stochastic textual gradients and selecting the most promising candidate via a Best-of-$N$ mechanism, can elicit the gradient of the UCB acquisition function within LLM’s mind without explicitly constructing the LLM’s implicit posterior mean and the associated uncertainty.
+To operationalize the Best-of-N gradient, we should first formalize the requisite capabilities of the external Large Language Model, which we designate as the Critic LLM.
+4.1 Critic LLM.
+Let $\Pi$ denote a prompt and $c=\Phi(\Pi)$ denote the corresponding system outcome.
+We define the history of optimization up to iteration $t$ as $H_{t}=\{(\Pi_{\tau},c_{\tau},s_{\tau})\}_{\tau\leq t}$.
+Then we define the Critic LLM, $M_{\text{critic}}$, as an LLM that can conduct the operations of the following four distinct operators:
+  1. 1.
+    Pairwise Judge Operator: A function that compares two outcome $c_{1}$ and $c_{2}$, given a reflection $R$ and predicts which outcome will yield a higher true score:
+    $\textsc{PairwiseJudge}\big(c_{1},c_{2};R,M_{\text{critic}}\big)\in\{c_{1},c_{2}\}$ (7)
+  2. 2.
+    Reflection Operator: A function that compresses the entire history into a natural-language summary of rules or guidelines:
+    $R_{t}=\textsc{MetaReflect}(H_{t};M_{\text{critic}})$ (8)
+  3. 3.
+    Textual Gradient Operator: A stochastic function that generates a textual gradient $\delta$ (e.g., “Make the lighting warmer”) intended to improve the prompt:
+    $\delta\sim\nabla_{\text{text}}(\Pi;R)$ (9)
+  4. 4.
+    Rewrite Operator: A deterministic function that maps a prompt $\Pi$ and an edit $\delta$ to a new prompt $\Pi^{\prime}$:
+    $\Pi^{\prime}=\operatorname{Apply}(\Pi,\delta;M_{\text{critic}})$ (10)
+In the digital-marketing example, these correspond to an LLM’s ability to (1) compare two ad creatives and predict which will perform better (calderon2025alternative), (2) examine past performance data and derive guidelines for effective ads in context (jiang2024many; ferber2024context), (3) propose prompt improvements based on those guidelines (textgrad; li2025test), and (4) apply prompt improvements to get the new prompt.
+4.2 Best-of-N gradient
+The notion of a Best-of-$N$ gradient builds on the widely used Best-of-$N$ strategy for test-time alignment, where an LLM trades additional inference-time (test-time) compute for higher-quality outputs by sampling multiple candidate generations and then selecting the one that appears most aligned with a downstream objective (snell2024scaling; beirami2024theoretical).
+In the standard Best-of-$N$ strategy for prompting, an LLM generates $N$ candidate prompts, a generation model produces $N$ candidate outcomes, and a reward model selects the highest-scoring one.
+While effective for choosing a single output, our goal is to improve the prompt itself to ensure sustained performance improvement.
+To achieve this, we introduce the concept of the Best-of-N Gradient.
+Instead of sampling $N$ candidate prompts directly, we sample $N$ distinct textual gradients (directions of improvement).
+By applying these $N$ textual edits to the current prompt and selecting the LLM-as-a-judge-predicted most promising candidate solution (Best-of-N), we aim to identify the textual gradient direction for the implicit acquisition function.
+The procedure operates as follows:
   1. 1)
-      Textual gradient generation: From the current inner prompt $\Pi_{t,g-1}$, query the textual-gradient generation operator $\nabla_{\text{text}}(\Pi_{t,g-1};R_{t-1},M_{\text{critic}})$ independently $N$ times to obtain textual gradients $\{\delta^{(i)}_{t,g}\}_{i=1}^{N}$.
+    Textual gradient generation: Using $M_{\text{critic}}$, query the textual-gradient operator $N$ independent times to obtain a set of small edits:
+    $\delta^{(i)}\sim\nabla_{\text{text}}(\Pi;R,M_{\text{critic}}),\quad i=1,\dots,N.$ (11)
   2. 2)
-      Candidate prompt generation: Apply each textual gradient $\delta_{t,g}^{(i)}$ to $\Pi_{t,g-1}$ for all $i$ to get $\{\Pi_{t,g}^{(i)}\}_{i=1}^{N}$.
+    Candidate prompt generation: Apply each edit to the current prompt to generate $N$ candidate prompts:
+    $\Pi^{(i)}=\operatorname{Apply}(\Pi,\delta^{(i)};M_{\text{critic}}),\quad i=1,\dots,N.$ (12)
   3. 3)
-      Candidate system outcome generation: Produce the system outcomes $\{c_{t,g}^{(i)}:=\Phi(\Pi_{t,g}^{(i)})\}_{i=1}^{N}$.
+    Candidate system outcome generation: Generate the corresponding system outcomes for each candidate prompt:
+    $c^{(i)}=\Phi(\Pi^{(i)}),\quad i=1,\dots,N.$ (13)
   4. 4)
-      Best-of-N selection: Find the best of $\{(\Pi_{t,g}^{(i)},c_{t,g}^{(i)})\}_{i=1}^{N}$ and set the best one as $(\Pi_{t,g},c_{t,g})$.
-  • 2.
-    Evaluation.
-(Detailed in Section 5.2) Set $c_{t}=c_{t,G}$ and $\Pi_{t}=\Pi_{t,g}$.
-Evaluate $c_{t}$ to get the evaluation result $s_{t}$, i.e., $s_{t}=\mathrm{Eval}(c_{t},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$.
-Append $(\Pi_{t},c_{t},s_{t})$ to $H_{t-1}$ to obtain $H_{t}$.
-If $s_{t}>s_{t-1}$, keep $(\Pi_{t},c_{t})$.
-Otherwise, set $(\Pi_{t},c_{t})=(\Pi_{t-1},c_{t-1})$ and set $s_{t}=s_{t-1}$.
-[^6] This ‘gated’ acceptance/rollback was adopted from how textgrad implemented TextGrad.
-  • 3.
-    Meta-reflection. (Detailed in Section 5.3) Update the reflection $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$ for use at iteration t+1t{+}1.
-After $T$ optimization steps, $\{c_{t}\}_{t=1}^{T}$ and $\{s_{t}\}_{t=1}^{T}$ are reported as the algorithm’s optimization outcome.
-Note that T-BoN BO takes $N\times G$ generation of prompt/outcome candidates before one evaluation.
-This careful selection of a prompt/outcome candidate to evaluate is the key to evaluation efficiency in T-BoN BO: it deliberately trades inexpensive generation compute for fewer evaluations.
-[…]
-Figure 4: Summary of T-BoN BO: $T$ iterations, $G$ gradient steps, and Best-of-$N$ candidates at each gradient step.
-T-BoN BO deliberately trades inexpensive generation compute for fewer evaluations: each iteration generates $N\times G$ prompt/outcome candidates with no evaluation.
-Input: System $\Phi$; initial prompt $\Pi_{0}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$ ; Output: Optimized $\Pi_{T},c_{T},s_{T}$ 1 $c_{0}\leftarrow\Phi(\Pi_{0})$; 2 $s_{0}\leftarrow\text{Eval}(c_{0},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$; 3 $H_{0}\leftarrow\{(\Pi_{0},c_{0},s_{0})\}$, $R_{0}\leftarrow\emptyset$; 4 for $t=1$ to $T$ do 5 $\Pi_{t,0}\leftarrow\Pi_{t-1}$; 6 for $g=1$ to $G$ do 7    for $i=1$ to $N$ do 8       $\delta^{(i)}_{t,g}\leftarrow\nabla_{\text{text}}(\Pi_{t,g-1};R_{t-1},M_{\text{critic}})$ ; 9       $\Pi_{t,g}^{(i)}\leftarrow\operatorname{Apply}(\Pi_{t,g-1},\delta_{t,g}^{(i)})$ ; 10       $c_{t,g}^{(i)}\leftarrow\Phi(\Pi_{t,g}^{(i)})$ ; 11        12      end for 13    $(\Pi_{t,g},c_{t,g})\leftarrow\textsc{Best-of-N}(\{(\Pi_{t,g}^{(i)},c_{t,g}^{(i)})\},R_{t-1},M_{\text{critic}})$ 14   end for 15 $c_{t}\leftarrow c_{t,G}$, $\Pi_{t}\leftarrow\Pi_{t,G}$; 16 $s_{t}\leftarrow\text{Eval}(c_{t},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$; 17 $H_{t}\leftarrow H_{t-1}\cup\{(\Pi_{t},c_{t},s_{t})\}$; 18 if $s_{t}<s_{t-1}$ then 19    $\Pi_{t}\leftarrow\Pi_{t-1}$, $c_{t}\leftarrow c_{t-1}$, $s_{t}\leftarrow s_{t-1}$; 20     21   end if 22 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$; 23  24 end for return $\{c_{t}\}_{t=1}^{T}$ Algorithm 1 T-BoN BO
-Input: System $\Phi$; $\{\Pi_{0}^{j}\}_{j=1}^{J}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$; $J$ trajectories Output: Optimized triples $\{(\Pi^{j}_{T},c^{j}_{T},s^{j}_{T})\}_{j=1}^{J}$ 1 for $j=1$ to $J$ do 2 $c_{0}^{j}\leftarrow\Phi(\Pi_{0}^{j})$; 3 $s_{0}^{j}\leftarrow\text{Eval}(c_{0}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$; 4 $H_{0}\leftarrow\{(\Pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$, $R\leftarrow\emptyset$; 5  6 end for 7 8for $t=1$ to $T$ do 9 for $j=1$ to $J$ do 10    $\Pi^{j}_{t,0}\leftarrow\Pi^{j}_{t-1}$; 11    for $g=1$ to $G$ do 12       for $i=1$ to $N$ do 13          $\delta_{t,g}^{j,(i)}\leftarrow\nabla_{\text{text}}(\Pi_{t,g-1}^{j};R,M_{\text{critic}})$ ; 14          $\Pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}(\Pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)})$ ; 15          $c_{t,g}^{j,(i)}\leftarrow\Phi(\Pi_{t,g}^{j,(i)})$ ; 16           17         end for 18       $(\Pi_{t,g}^{j},c_{t,g}^{j})\leftarrow\textsc{Best-of-N}(\{(\Pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\},R_{t-1},M_{\text{critic}})$ 19      end for 20    $c_{t}^{j}\leftarrow c^{j}_{t,G}$, $\Pi_{t}^{j}\leftarrow\Pi_{t,G}^{j}$; 21    $s_{t}^{j}\leftarrow\text{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$; 22    $H_{t}\leftarrow H_{t-1}\cup\{(\Pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}$; 23    if $s_{t}^{j}<s_{t-1}^{j}$ then 24       $\Pi_{t}^{j}\leftarrow\Pi_{t-1}^{j}$, $c_{t}^{j}\leftarrow c_{t-1}^{j}$, $s_{t}^{j}\leftarrow s_{t-1}^{j}$; 25        26      end if 27     28   end for 29 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$; 30  31 end for return $\{c_{t}^{j_{t}^{*}}\}_{t=1}^{T}$, where $j_{t}^{*}:=\operatorname{argmax}_{j}s_{t}^{j}$ Algorithm 2 Parallel T-BoN BO
-Figure 5: T-BoN algorithm (Left) and Parallel T-BoN algorithm (Right).
-5.1 Best-of-N Textual Gradient steps.
-At each gradient step $g=1,\ldots,G$, the algorithm expands the current prompt $\Pi_{t,g-1}$ into a neighborhood of $N$ variants, then contracts back to a single survivor.
-The process begins with textual gradient generation: the critic model $M_{\text{critic}}$ produces $N$ improvement directions,
-$\delta^{(i)}_{t,g}\;\sim\;\nabla_{\text{text}}\!\big(\Pi_{t,g-1};\,R_{t-1},\,M_{\text{critic}}\big),\qquad i=1,\ldots,N,$ (5)
-where randomness is introduced by a nonzero sampling temperature and nucleus filtering (top-$p$ sampling) [^7] In the experiments, we used Gemini Flash 2.5 with temperature =0.8=0.8 and top $p=0.9$.
-, which ensures that the gradients are not identical but represent diverse, locally valid edits.
-Each textual gradient is then applied to form a candidate prompt
-$\Pi^{(i)}_{t,g}=\operatorname{Apply}\!\big(\Pi_{t,g-1},\,\delta^{(i)}_{t,g}\big)$, (6)
-with the rewrite operator configured for stability (low/zero temperature) so the edits remain coherent. The generator model then produces the system outcomes
-$c^{(i)}_{t,g}:=\Phi\!\big(\Pi^{(i)}_{t,g}\big)$. (7)
-The system outcome candidates are reduced to one winner by the Best-of-N choice operation.
-For the Best-of-N choice operation, following liu2025pairjudgermperformbestofn, we employ the pairwise tournament that uses critic model $M_{\text{critic}}$.
+    Best-of-N selection: Select the best candidate by iteratively applyng PairwiseJudge operation using $M_{\text{critic}}$ as LLM-as-a-judge.
+We define this selection as Best-of-N operation, which induces the index $i^{\star}$ as the best one among N:
+    $i^{\star}=\textsc{Best-of-N}\big(\{(\Pi^{(i)},c^{(i)})\}_{i=1}^{N};R,M_{\text{critic}}\big).$ (14)
+    The resulting prompt $\Pi^{(i^{\star})}$ represents the update step taken by the Best-of-N Gradient.
+For the Best-of-N selection, we employ the pairwise tournament (liu2025pairjudgermperformbestofn).
 First, the $N$ candidates are permuted to produce randomly chosen tournament matches.
-Each match compares two outcomes $(c^{(i)}_{t,g},c^{(j)}_{t,g})$ by querying $M_{\text{critic}}$ for a head-to-head judgment (For the prompt, please refer to Figure 12 in Appendix A).
+Each match compares two outcomes by querying $M_{\text{critic}}$ for a head-to-head judgment (For the prompt, please refer to Figure 11 in Appendix A).
 illustrates the prompt example for ad generation tasks.
-To mitigate noise, the comparison is repeated $K$ times (typically an even number to evenly swap orders to remove position bias (shi2025judgingjudgessystematicstudy), and break even by random; here we chose $K=4$), producing votes $\{V_{k}(i,j)\}_{k=1}^{K}$, where $V_{k}(i,j)=i$ if critic prefers $i$ over $j$, and vice versa.
-The winner of the pairwise match is then defined by the majority rule $\operatorname{argmax}_{u\in\{i,j\}}\sum_{k=1}^{K}\mathbf{1}_{\{V_{k}(i,j)=u\}}$.
-Winners advance round by round until only one candidate remains.
-The survivor $(\Pi_{t,g},c_{t,g})=(\Pi^{(i^{\star})}_{t,g},c^{(i^{\star})}_{t,g})$ is carried forward for the next gradient step $g+1$.
-5.2 Evaluation
-After completing $G$ gradient steps, we arrive at the $(\Pi_{t,G},c_{t,G})$, which we redefine as $(\Pi_{t},c_{t})$.
-We evaluate $(\Pi_{t},c_{t})$: recall from Section 4 that $(\Pi_{t},c_{t})$ is evaluated as
-$s_{t}\;:=\;\frac{1}{L}\sum_{l=1}^{L}r\!\left(y\left(c_{t},x_{t}^{(l)}\right)\right),\qquad x_{t}^{(l)}\sim\mathcal{D}_{\mathcal{X}}(\Pi_{t}),$ (8)
-where $r\!\left(y\left(c_{t},x_{t}^{(l)}\right)\right)$ is the evaluation of $c_{t}$ for the input $x_{t}^{(l)}$ and $L$ denotes the number of inputs evaluated.
-After evaluation, the triple $(\Pi_{t},c_{t},s_{t})$ is appended to the history $H_{t-1}$ to form $H_{t}$.
-Following textgrad, we adopt a simple acceptance rule that controls progress: if $s_{t}\geq s_{t-1}$, the new candidate is adopted; otherwise, the system rolls back to the previous iteration’s outcome, i.e., $(\Pi_{t-1},c_{t-1},s_{t-1})$.
-5.3 Meta-reflection.
-At the end of each iteration $t$, we compress the full history $H_{t}$ of prompts, generated system outcomes, and evaluation results into a natural language reflection $R_{t}$.
-$R_{t}$ consists of a small set of actionable rules.
-In the ad generation example, those rules may take the form “emphasize social context,” “avoid cluttered backgrounds,” or “prefer warm lighting.” This reflection serves as a global memory; in subsequent iterations, the critic LLM takes this reflection as a part of its context when generating textual gradients.
-(For the prompt used for meta-reflection, see Figure 13 in Appendix A).
-There are many ways to construct a reflection from past results.
-In our experiments, following li2025test, we instruct the critic model $M_{\text{critic}}$ to compare the five lowest- and five highest-performing outcomes from previous iterations (or, when fewer than ten are available, all outcomes) [^8] Limiting the number of outcomes in the critic model’s context window helps mitigate context rotting (hong2025context)., identify recurring visual and textual patterns, and distill them into a concise set of actionable rules.
-5.4 Parallel T-BoN BO.
-Although a local maximum found by gradient steps from a single starting point is often effective in practice, it may under-explore the ample prompt space.
-Following the practices (wilson2018maximizing; papenmeier2025understanding), we also introduce a parallel variant of T-BoN BO that runs $J$ optimization trajectories simultaneously (Algorithm 2).
-Here, each of $J$ trajectories starts from a different initial prompt (denoted as $\{\Pi_{0}^{j}\}_{j=1}^{J}$).
-After independent gradient steps and evaluations at each iteration, evaluations from all trajectories are merged into the shared history, and a single meta-reflection is produced.
-Different trajectories diversify the search, while the shared reflection ensures that discoveries made by one trajectory are immediately shared with the others.
-Combined with the careful choice of gradients in T-BoN BO, this parallel process ensures the broader coverage of the search space and faster convergence to high-quality solutions in high-dimensional BO (papenmeier2025understanding) with a slight loss of optimality in the worst case (wilson2018maximizing).
-At the end of $T$ optimization steps, the $J$-trajectory parallel T-BoN BO returns $\{(\Pi_{T}^{j},c_{T}^{j},s_{T}^{j})\}_{j=1}^{J}$.
-For the purpose of reporting, for each $t\leq T$, we select $j_{t}^{\ast}:=\operatorname{argmax}_{j}s_{t}^{j}$ and report its output $c_{t}^{j_{t}^{\ast}}$ as the step $t$’s outcome.
-That is, $\{c_{t}^{j_{t}^{\ast}}\}_{t=1}^{T}$ is reported as the final outcome.
-6 Theory: T-BoN BO’s Asymptotic Equivalence to UCB Bayesian Optimization
+To mitigate noise, the comparison is repeated $K$ times (typically an even number to evenly swap orders to remove positional bias (shi2024judging), and ties are broken by random.
+Winners advance round by round until only the final winner remains.
+Given the final winner’s index $i^{\star}$, we define $(\Pi^{j}_{t,g},c^{j}_{t,g}):=(\Pi^{j,(i^{\star})}_{t,g},c^{j,(i^{\star})}_{t,g})$ and carry it forward for the next gradient step $g+1$.
+4.3 Theory: Best-of-N gradient’s Asymptotic Equivalence to UCB gradient
 In this section, we prove that the Best-of-$N$ over textual gradients selects, in probability, the ascent direction of the UCB acquisition with exploration weight $\beta$ scaling as $\Theta(\sqrt{\ln N})$ (Theorem 2).
@@
 $A_{\beta_{N}}(x+\Delta x^{(N)})\;\geq\;A_{\beta_{N}}(x)\;+\;\varepsilon\,\|\nabla A_{\beta_{N}}(x)\|\;-\;o_{p}(\varepsilon).$ (21)
+5 TextGrad-Best-of-N Bayesian Optimization (T-BoN BO)
+We now describe the self-improving AI framework proposed in this paper: TextGrad-Best-of-N Bayesian Optimization (T-BoN BO).
+As discussed in Section 1, T-BoN BO is designed to implement parallel gradient-based UCB-BO in the discrete space of natural language (Figure 3).
+Specifically, T-BoN BO implements the parallel gradient-based UCB-BO that reuses the final gradient point as the initialization for the next iteration.
+Therefore, each initial point gives rise to a trajectory: a sequence of candidate points obtained by repeatedly taking gradient-ascent steps and evaluations starting from that initialization.
+[…]
+Figure 4: Visual summary of T-BoN BO (Algorithm 1) with $T$ iterations, $J$ trajectories, and $G$ Best-of-NN gradients per each iteration and trajectory.
+Algorithm.
+Throughout $T$ number of optimization iterations, T-BoN BO runs $J$ trajectories in parallel, starting from a set of initial prompts, denoted $\{\Pi_{0}^{j}\}_{j=1}^{J}$.
+At each iteration and a trajectory, we fix the hyperparameters: the number of candidates per gradient step ($N$) and the number of gradient steps per iteration ($G$).
+At iteration $t=0$, it starts from evaluating $\{\Pi_{0}^{j}\}_{j=1}^{J}$, i.e., the initial systems $\{c_{0}^{j}=\Phi(\Pi_{0}^{j})\}_{j=1}^{J}$, on $\{\mathcal{D}_{\mathcal{X}}(\Pi_{0}^{j})\}_{j=1}^{J}$ to obtain initial scores $\{s_{0}^{j}\}_{j=1}^{J}$.
+The history is then initialized as $H_{0}=\bigcup_{j=1}^{J}\{(\Pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$.
+The shared reflection is initialized as $R_{0}=\emptyset$.
+At each iteration $t$, for each trajectory $j\in\{1,\ldots,J\}$, we set $\Pi_{t,0}^{j}$ as $\Pi_{t-1}^{j}$ and proceed as follows:
+  • 1.
+    Best-of-N gradients. (Detailed in Section 4.2) For each trajectory $j$, repeat steps 1) – 4) for $g=1,\ldots,G$:
+    1. 1)
+      Textual gradient generation: From the current prompt $\Pi_{t,g-1}^{j}$, query the textual-gradient generation operator $\nabla_{\text{text}}(\Pi_{t,g-1}^{j};R_{t-1},M_{\text{critic}})$ independently $N$ times to obtain textual gradients $\{\delta^{j,(i)}_{t,g}\}_{i=1}^{N}$.
+    2. 2)
+      Candidate prompt generation: Apply each textual gradient $\delta_{t,g}^{j,(i)}$ to $\Pi_{t,g-1}^{j}$ for all $i$ to get $\{\Pi_{t,g}^{j,(i)}\}_{i=1}^{N}$.
+    3. 3)
+      Candidate system outcome generation: Produce the system outcomes $\{c_{t,g}^{j,(i)}:=\Phi(\Pi_{t,g}^{j,(i)})\}_{i=1}^{N}$.
+    4. 4)
+      Best-of-N selection: Find the best of $\{(\Pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\}_{i=1}^{N}$ using the critic and the shared reflection $R_{t-1}$, and set the winner as $(\Pi_{t,g}^{j},c_{t,g}^{j})$.
+  • 2.
+    Evaluation. For each trajectory $j$,
+    1. 1)
+      Set $c_{t}^{j}=c_{t,G}^{j}$ and $\Pi_{t}^{j}=\Pi_{t,G}^{j}$.
+    2. 2)
+      Evaluate $c_{t}^{j}$ to get the score $s_{t}^{j}=\mathrm{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}^{j}))$, where
+      $\mathrm{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}^{j}))\;:=\;\frac{1}{L}\sum_{l=1}^{L}r\!\left(y\left(c^{j}_{t},x_{t}^{j,(l)}\right)\right),\qquad x_{t}^{j,(l)}\sim\mathcal{D}_{\mathcal{X}}(\Pi^{j}_{t}),$ (22)
+    3. 3)
+      Given $\{(\Pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}_{j=1}^{J}$, append it to the shared history $H_{t-1}$ to form $H_{t}$.
+    4. 4)
+      If $s_{t}^{j}>s_{t-1}^{j}$, keep $(\Pi_{t}^{j},c_{t}^{j})$.
+Otherwise, revert trajectory $j$ to the previous state: $(\Pi_{t}^{j},c_{t}^{j})=(\Pi_{t-1}^{j},c_{t-1}^{j})$ and set $s_{t}^{j}=s_{t-1}^{j}$.
+    For 4), we follow textgrad and adopt a simple acceptance rule that controls progress: if $s_{t}\geq s_{t-1}$, the new candidate is adopted; otherwise, the system rolls back to the previous iteration’s outcome, i.e., $(\Pi_{t-1},c_{t-1},s_{t-1})$.
+After evaluation of all trajectories, $\{(\Pi^{j}_{t},c^{j}_{t},s^{j}_{t})\}_{j=1}^{J}$ are appended to the history $H_{t-1}$ to form $H_{t}$.
+  • 3.
+    Shared Meta-reflection.
+    At the end of each iteration $t$, we update the global reflection
+    $R_{t}\leftarrow\textsc{MetaReflect}(H_{t};M_{\text{critic}})$
+    using the aggregated history from all trajectories for use at iteration t+1t{+}1.
+Here, $\textsc{MetaReflect}(H_{t};M_{\text{critic}})$ compresses the full history $H_{t}$ of prompts, generated system outcomes, and evaluation results into a natural language reflection $R_{t}$.
+$R_{t}$ consists of a small set of actionable rules.
+In the ad generation example, those rules may take the form “emphasize social context,” “avoid cluttered backgrounds,” or “prefer warm lighting.” This reflection serves as a global memory; in subsequent iterations, the critic LLM takes this reflection as a part of its context when generating textual gradients.
+(For the prompt used for meta-reflection, see Figure 12 in Appendix A).
+    There are many ways to construct a reflection from past results.
+In our experiments, following li2025test, we instruct the critic model $M_{\text{critic}}$ to compare the five lowest- and five highest-performing outcomes from previous iterations (or, when fewer than ten are available, all outcomes) [^4] Limiting the number of outcomes in the critic model’s context window helps mitigate context rotting (hong2025context)., identify recurring visual and textual patterns, and distill them into a concise set of actionable rules.
+After $T$ optimization steps, the sequence of best-performing outcomes $\{c_{t}^{j_{t}^{\ast}}\}_{t=1}^{T}$, where $j_{t}^{\ast}:=\operatorname{argmax}_{j}s_{t}^{j}$, is reported as the algorithm’s result.
+Note that T-BoN BO takes $N\times G$ generation of prompt/outcome candidates before each evaluation.
+This careful selection of a prompt/outcome candidate to evaluate is the key to evaluation efficiency in T-BoN BO: it deliberately trades inexpensive generation compute for fewer evaluations.
+Input: System $\Phi$; $\{\Pi_{0}^{j}\}_{j=1}^{J}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$; $J$ trajectories Output: Optimized triples $\{(\Pi^{j}_{T},c^{j}_{T},s^{j}_{T})\}_{j=1}^{J}$ 1 for $j=1$ to $J$ do 2 $c_{0}^{j}\leftarrow\Phi(\Pi_{0}^{j})$; 3 $s_{0}^{j}\leftarrow\text{Eval}(c_{0}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$; 4 $H_{0}\leftarrow\{(\Pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$, $R\leftarrow\emptyset$; 5  6 end for 7 8for $t=1$ to $T$ do 9 for $j=1$ to $J$ do 10    $\Pi^{j}_{t,0}\leftarrow\Pi^{j}_{t-1}$; 11    for $g=1$ to $G$ do 12       for $i=1$ to $N$ do 13          $\delta_{t,g}^{j,(i)}\leftarrow\nabla_{\text{text}}(\Pi_{t,g-1}^{j};R,M_{\text{critic}})$ ; 14          $\Pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}(\Pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)})$ ; 15          $c_{t,g}^{j,(i)}\leftarrow\Phi(\Pi_{t,g}^{j,(i)})$ ; 16           17         end for 18       $(\Pi_{t,g}^{j},c_{t,g}^{j})\leftarrow\textsc{Best-of-N}(\{(\Pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\},R_{t-1},M_{\text{critic}})$; 19        20      end for 21    $c_{t}^{j}\leftarrow c^{j}_{t,G}$, $\Pi_{t}^{j}\leftarrow\Pi_{t,G}^{j}$; 22    $s_{t}^{j}\leftarrow\text{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$; 23    $H_{t}\leftarrow H_{t-1}\cup\{(\Pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}$; 24    if $s_{t}^{j}<s_{t-1}^{j}$ then 25       $\Pi_{t}^{j}\leftarrow\Pi_{t-1}^{j}$, $c_{t}^{j}\leftarrow c_{t-1}^{j}$, $s_{t}^{j}\leftarrow s_{t-1}^{j}$; 26        27      end if 28     29   end for 30 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$; 31  32 end for return $\{c_{t}^{j_{t}^{*}}\}_{t=1}^{T}$, where $j_{t}^{*}:=\operatorname{argmax}_{j}s_{t}^{j}$ Algorithm 1 T-BoN BO
 6 Case Study: Ad Optimization in Digital Marketing
@@
 6.1 Problem definition
-Formally, define the brand fine-tuned ad-generation AI system as $\Phi$, whose behavior is controlled by a prompt set $\Pi$ (Figure 6).
-For a given prompt set $\Pi$, we write $\Phi(\Pi)$ for the corresponding ad creative shown on the ad platform.
+Formally, define the brand fine-tuned ad-generation AI system as $\Phi$, whose behavior is controlled by a prompt $\Pi$ (Figure 5).
+For a given prompt $\Pi$, we write $\Phi(\Pi)$ for the corresponding ad creative shown on the ad platform.
 As a system, $\Phi(\Pi)$ maps the input space $\mathcal{X}$ (potential ad viewers on the platform) to a distribution over the output space $\mathcal{Y}$ (observable ad responses).
@@
 When a user $x\in\mathcal{X}$ is shown the ad creative $\Phi(\Pi)$, we observe an outcome $y\in\mathcal{Y}$, which is a realization of $y(\Phi(\Pi),x)$.
-As discussed earlier, the distribution of exposed users depends on the ad $\Phi(\Pi)$ and therefore on the prompt set $\Pi$.
+As discussed earlier, the distribution of exposed users depends on the ad $\Phi(\Pi)$ and therefore on the prompt $\Pi$.
 We denote by $\mathcal{D}_{\mathcal{X}}(\Pi)$ the unknown distribution of users $x$ who are shown the ad generated from $\Pi$ when the ad $\Phi(\Pi)$ is run on the platform.
@@
 which form the textual gradients. Each edit is then applied to the current prompt to obtain $N$ candidate prompts
-$\pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}\big(\pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)}\big),\qquad i=1,\dots,N.$ (19)
+$\pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}\big(\pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)};M_{\text{critic}}\big),\qquad i=1,\dots,N.$ (26)
 These textual gradients correspond to small, directed moves in language space that, according to the critic’s judgment, are likely to increase ad effectiveness.
@@
 Meta-reflection shared across trajectories.
-After all $J$ trajectories have been evaluated at iteration $t$, the algorithm updates the global meta-reflection $R_{t}=\textsc{MetaReflect}(H_{t})$, as in Section 5.3.
+After all $J$ trajectories have been evaluated at iteration $t$, the algorithm updates the global meta-reflection $R_{t}=\textsc{MetaReflect}(H_{t})$, as in Section 5.
 Concretely, $M_{\text{critic}}$ is prompted to review a subset of the best- and worst-performing ads and prompts in $H_{t}$ and to summarize the key patterns and rules that distinguish successful creatives from unsuccessful ones (see Figure 12 in Appendix A).
@@
 7 Experiments
-We validate the empirical effectiveness of the proposed T-BoN BO framework for a real-world digital marketing problem described in 7.
+We validate the empirical effectiveness of the proposed T-BoN BO framework for a real-world digital marketing problem described in 6.
 Specifically, we examine if and how much T-BoN BO improves outcomes under a fixed evaluation budget compared to state-of-the-art baselines.
@@
 to generate the initial 64 high-quality ads.
-(For prompts, see Figure 14 and 15 in Appendix A.) To maintain a rigorous baseline, we made the 64 initial prompts detailed enough to yield strong ads.
+(For prompts, see Figure 13 and 14 in Appendix A.) To maintain a rigorous baseline, we made the 64 initial prompts detailed enough to yield strong ads.
 This ensures that the algorithm’s subsequent improvements are not explained as trivial quality gains from having initial under-specified prompts.
@@
 7.5 Ablation study
-The main experiment in Section 8.1-8.4 evaluates the candidate ad at each iteration as the mean of scores across 411 personas (20% of the persona data in Twin-2k-500 (toubia2025twin)) that were injected into the LLM’s context.
+The main experiment in Section 7.1-7.4 evaluates the candidate ad at each iteration as the mean of scores across 411 personas (20% of the persona data in Twin-2k-500 (toubia2025twin)) that were injected into the LLM’s context.
 Each persona consists of a large amount of information, with 500 survey questions accounting for around 50,00050{,}000 tokens.

### v3

@@
-Bayesian Optimization in Language Space: An Eval-Efficient AI Self-Improvement Framework
+TextBO: Bayesian Optimization in Language Space for Eval-Efficient Self-Improving AI
 Abstract
-Large Language Models (LLMs) have recently enabled self-improving AI, i.e., AI that iteratively generates, evaluates, and refines its own outcomes.
-Recent studies have shown that self-improving AI focusing on prompt optimization (i.e., context engineering) can outperform state-of-the-art reinforcement learning fine-tuned LLMs.
-Here, their ‘performance’ is typically measured by query efficiency—the number of LLM-generated solution samples required to reach a specified performance threshold.
-However, in many societal applications, the primary limitation is not generating new solutions but evaluating them.
-For instance, evaluating an ad’s effectiveness requires significant human feedback, which is far more costly and time-consuming than generating a candidate ad.
-To optimize for the evaluation efficiency objective, a natural approach is to extend Bayesian Optimization (BO), a framework proven optimal for evaluation efficiency, to the language domain.
-However, the difficulty of directly estimating suitable acquisition functions in LLMs’ minds makes this extension challenging.
-This paper overcomes this challenge by proving that the combination of the simple and widely used Best-of-N selection strategy and simple textual gradients (i.e., textual edits from a critic model), statistically emulates the behavior of the gradients on the canonical Upper Confidence Bound (UCB) acquisition function, which induces optimal exploration in terms of evaluation efficiency.
-Based on this result, we propose TextGrad-Best-of-N Bayesian Optimization (T-BoN BO), a simple and eval-efficient language-space Bayesian optimization framework for AI self-improvement.
-We also empirically validate T-BoN BO by applying it to automated ad alignment tasks for persona distribution, demonstrating its superior performance compared to popular state-of-the-art baselines.
-In sum, T-BoN BO is a simple but principled framework for AI self-improving system design that bridges prompt optimization with classical Bayesian optimization.
-Keywords: Self-improving AI, Evaluation efficiency, Bayesian optimization, Context Engineering
+Large Language Models (LLMs) have enabled self-improving AI systems that iteratively generate, evaluate, and refine their outcomes.
+Recent studies show that prompt-optimization-based self-improvement can outperform state-of-the-art reinforcement-learning fine-tuning of LLMs, but performance is typically measured by generation efficiency.
+However, in many applications, the constraint is evaluation efficiency: obtaining reliable feedback is far more costly than generating candidates.
+To optimize for evaluation efficiency, we extend Upper Confidence Bound–Bayesian Optimization (UCB-BO), a framework known for optimal evaluation-efficiency guarantees, to the language domain.
+Doing so is challenging for two reasons: (i) gradients needed for UCB-BO are ill-defined in discrete prompt space; and (ii) UCB-style exploration relies on a surrogate model and acquisition function, which only live implicitly in the LLM.
+We overcome these challenges by proving that combining simple textual gradients (LLM-proposed local edits) with the Best-of-N selection strategy statistically emulates ascent along the gradient of the canonical UCB acquisition function.
+Based on this result, we propose TextBO, a simple, evaluation-efficient self-improving algorithm that operates purely in language space without explicit surrogates or calibrated uncertainty models.
+We empirically validate TextBO on automated ad-alignment tasks using a persona-induced preference distribution, demonstrating superior performance per evaluation compared to strong baselines such as Best-of-N and GEPA.
+We also evaluate TextBO’s Best‑of‑N multi‑step textual‑gradient mechanism on agentic AI benchmarks by augmenting GEPA with it and show that it performs better than standard GEPA.
+In sum, TextBO is a simple and principled framework for AI self-improving system design that bridges prompt optimization with classical Bayesian optimization.
+Keywords: Self-improving AI, Evaluation efficiency, Bayesian optimization, Context Engineering, Ad optimization, LLMs, Marketing.
 1 Introduction
-1.1 Self-improving AI
-Data-centric decision making has long relied on a human-driven cycle of historical data analysis, solution proposal, and experimentation.
-This human-driven iterative approach has supported a wide range of business decisions, ranging from website design (hauser2009website) to product recommendations (nandy2021b).
-We briefly describe this paradigm using the example of ad-optimization; see the left panel of Figure 1.
-The process starts with generating a set of ad creatives based on a creative brief.
-This task is often done by ad agencies or marketers, who rely on their understanding of the product, theories of consumer behavior, market conditions, and feedback from prior ad campaigns.
-This ad generation process typically takes weeks, and serves as the main bottleneck in the iterative ad-optimization process.
-Once we have a set of candidate ads, the manager moves to evaluation, i.e., feeds these ads into advertising platforms or consumer surveys to identify the best performing one, and this step usually takes a few days.
-Finally, to complete this feedback loop, the manager analyzes the performance of the ads from this campaign and uses the insights learned to generate a new slate of candidate ads; and the iterative process continues.
-Although effective in many settings, this human-driven workflow suffers from drawbacks.
-First, the speed with which firms can iterate over this process depends critically on how quickly humans can generate new candidate solutions.
-Historically, generating new candidate solutions (or ads in the case of our example) has been time-consuming, and this step formed a key bottleneck in speeding up the process.
-Further, there exists no formal framework for how human agents can analyze and learn from the performance of prior solutions and leverage them to propose better candidates, and this step is often left to managerial intuition.
-Indeed, much of the research and managerial focus on data-driven business solutions has largely centered on the evaluation step, i.e., given a set of candidate solutions/options, how can we identify the most effective one – and solutions such as A/B tests and Multi-Armed Bandits (MAB) experiments have been proposed and tested (kohavi2020trustworthy; fiez_etal_2024).
-Ad generation (Step 1)(Weeks)Analysis (Step 3)(Days)Ad evaluation (Step 2)(Days-weeks)HumanBottleneckAd generation (Step 1)(seconds)Analysis (Step 3)(minutes–hours)Ad evaluation (Step 2)(Days-weeks)AIBottleneck
+1.1 Self-improving AI for Data-driven Decision-making
+Data-centric decision making has long relied on an iterative, human-driven cycle of proposing solutions, testing them, and learning from the results.
+This paradigm underlies many business applications, ranging from website design (hauser2009website) to product recommendations (nandy2021b).
+We illustrate the cycle using ad optimization (left panel of Figure 1).
+An ad agency or marketer first generates a slate of creatives from a brief, drawing on product knowledge, theories of consumer behavior, market conditions, and feedback from prior campaigns.
+This generation step often takes weeks.
+The candidates are then evaluated (e.g., on advertising platforms or via consumer surveys) to identify the best performers, typically over days to weeks.
+Finally, the manager analyzes campaign outcomes and uses the insights to guide the next round of creative generation.
+This workflow is effective, but it is also constrained by two longstanding limitations.
+First, iteration speed depends heavily on how quickly humans can generate new candidate solutions, which has historically been the main bottleneck.
+Second, while evaluation has been studied extensively—with tools such as A/B testing and Multi-Armed Bandits (MAB) (kohavi2020trustworthy; fiez_etal_2024)—there is no comparable formal framework for how practitioners should systematically analyze results and translate them into improved candidates; in practice, this step is often driven by managerial intuition.
+Ad generation (Step 1)(Weeks)Analysis (Step 3)(Days)Ad evaluation (Step 2)(Days-weeks) Human-driven cycle BottleneckAd generation (Step 1)(seconds)Analysis (Step 3)(minutes–hours)Ad evaluation (Step 2)(Days-weeks) AI-driven cycle Bottleneck
 Figure 1: Comparison of human-driven and AI-driven ad self-improvement cycles in digital advertising.
-Recent advances in generative AI have begun to transform this human-driven paradigm by providing potential solutions to the issues discussed above.
-First, candidate generation is fast and scalable with generative AI models.
-For example, in the case of digital ads, generative models can produce high-quality creative outputs, including ad copy, images, and complete campaign concepts that adhere to brand guidelines in a matter of minutes (jansen2024automated; hartmann2025power).
-[^1] Indeed, a substantial share of creatives is now generated by AI, reflecting improvements in both quality and speed (coffee2025aiadbuying; Majic2025VibeMarketing).
-Second, Large language Models (LLMs) and related generative systems can now conduct sophisticated data analysis, summarizing performance patterns and identifying potential drivers of success (ferber2024context; guo2024ds; ghosh2024exploring; wiedemer2025videomodelszeroshotlearners), providing a pathway for analyzing historical data and improving candidate solutions.
-As such, with the new capabilities enabled by AI, the time required for previous data analysis and generation of new candidate solutions are significantly reduced.
-Given these developments, there is immense interest in the development and deployment of AI-based automation of the complete generation–evaluation-analysis loop at scale, allowing systems to update themselves continuously in response to emerging data (See right panel of Figure 1).
-In this loop, generative models produce new solution candidates, data-science agents conduct evaluation of the solution candidates, and multi-modal reasoning LLM-based analytical agents interpret the evaluation results and synthesize insights to guide the next generation step.
-This type of automated cycle running system is often referred to as the self-improving AI or self-evolving AI (MantiaChatterjeeLee2025; chen2025xbench; silver2025welcome).
-This transformation has an important implication for practice: it dramatically lowers the marginal cost of iteration, making continuous experimentation feasible not only for large firms or organizations with dedicated analytics teams but also for small and medium-sized firms that historically lacked the creative and analytical capacity to run many cycles in parallel.
-When generation and analysis become cheap and largely automated, even resource-constrained organizations can maintain always-on optimization loops that adapt campaigns to evolving market conditions and consumer responses.
-As generation and analysis become increasingly automated and inexpensive in this self-improving loop, the bottleneck in data-centric decision making shifts from proposing and interpreting candidates to evaluating them in the real world.
-Even if analysis can be performed quickly and new candidate solutions can be generated almost instantaneously, each iteration still has to be validated by interacting with the real environment and collecting outcome data.
-The effective pace of the cycle is therefore governed by evaluation-related constraints such as available traffic, experimentation budgets, and the amount of data required to obtain statistically reliable comparisons among candidates, rather than by generation or analysis speed.
-For instance, in digital advertising, even when generating ad creatives via LLM queries is super fast, but evaluating an ad’s effectiveness still requires significant human feedback or large-scale experimentation (johnson2023inferno).
-A similar pattern arises in other scientific and engineering domains: protein structures or sequences can be proposed rapidly by models such as AlphaFold (jumper2021highly), but their functional properties must still be confirmed through slow and noisy wet-lab assays; and chip layouts can be generated in silico by systems such as AlphaChip (goldie2024chip), yet rigorous evaluation depends on physical fabrication and benchmarking.
-As generation and analysis become increasingly automated and inexpensive, the evaluation step thus emerges as the primary bottleneck that limits the overall cadence of self-improvement.
-As such, any societal and business AI system we design needs to be evaluation-efficient, i.e., reach a high-performing solution using as few costly evaluations as possible, rather than optimizing for a few inexpensive generations.
-1.2 Prompt optimization-based self-improving AI
-A natural question is, then, how the computer science literature has instantiated self-improving AI in concrete algorithmic frameworks.
-To date, a central line of success in self-improving AI has arisen in the domain of iterative prompt optimization, where AI agents iteratively refine their own prompts in response to performance feedback.
-agrawal2025gepa showed that iterative prompt optimization-based self-improving AI can be much more efficient than state-of-the-art model fine-tuning techniques such as GRPO (liu2024deepseek).
-Specifically, given any limited budget for the number of LLM generation query calls, their proposed iterative prompt optimization method, named GEPA (Genetic-Pareto), outperformed GRPO across their experiments.
-Those experiments included benchmarks such as complex questions across multiple Wikipedia articles, strict instruction-following under constraints, retrieval-augmented fact verification to support or refute claims, and maintaining answer quality while avoiding privacy leakage.
-In all those benchmarks, the primary efficiency metric was generation efficiency.
-i.e., the number of LLM query calls required to generate candidate answers to reach a target performance level.
-Likewise, prior literature on prompt-based self-improving AI (agrawal2025gepa; zhang2025agentic; wang2025maestro) has focused on measuring efficiency in terms of generation efficiency, not evaluation efficiency.
-Indeed, leading approaches are often specified as procedural heuristic search routines which finds near-optimal solution with small computations, e.g., population-based or random-search-style algorithms that treat prompts as individuals in an evolutionary process or as configurations in a heuristic pipeline.
-Specifically, they often maintain a pool or batch of prompts, repeatedly generate new variants via LLM-driven mutation, crossover, reflection, or trace-based resampling, score these candidates on held-out data or rollouts, retain the top-performing or Pareto-optimal prompts, and iterate this generate–score–select–mutate loop under a budget of LLM forward passes (10.5555/3692070.3692611; guo2024connecting; agrawal2025gepa; better_together).
-This mirrors the design philosophy of classical evolutionary and heuristic optimization, where the quality of the best solution is assessed as a function of the computation budget rather than the evaluation budget (nguyen2016understanding; ding2023quality; hu2024automated; liu2024evolution).
-Consequently, evaluation efficiency objective has been largely overlooked in the computer science literature.
-However, as we discussed earlier, in self-improving AI for real-world societal and business problems, the primary bottleneck is no longer generating new solutions efficiently (generation efficiency) but evaluating them efficiently (evaluation efficiency).
-This shift in the dominant constraint calls for rethinking prompt-optimization-based self-improving AI frameworks so that they explicitly treat evaluation budget as the scarce resource to be allocated and optimized over.
+Recent advances in generative AI have begun to transform this human-driven paradigm by addressing the issues discussed above.
+Candidate generation is now fast and scalable: in digital advertising, AI models can produce high-quality ad copy, images, and campaign concepts that adhere to brand guidelines in minutes (jansen2024automated; hartmann2025power).
+[^1] A large share of ad creatives is now AI-generated, reflecting improvements in both quality and speed (coffee2025aiadbuying; Majic2025VibeMarketing).
+In parallel, LLM-based systems can autonomously conduct analyses by summarizing performance patterns and hypothesizing drivers of success (ferber2024context; guo2024ds; ghosh2024exploring; wiedemer2025videomodelszeroshotlearners).
+Together, these capabilities make it feasible to automate the full generation–evaluation–analysis loop at scale, allowing systems to update themselves continuously in response to new data (right panel of Figure 1).
+We refer to such automated, continuously updating systems as self-improving or self-evolving AI (MantiaChatterjeeLee2025; chen2025xbench; silver2025welcome).
+While there can be many types of self-improving AI systems, a central line of progress has been on iterative prompt optimization, where the AI agent repeatedly refines its own prompts based on performance feedback.
+Indeed, recent papers show that such approaches can often outperform reinforcement-learning based fine-tuning (agrawal2025gepa); see $\S$2 for a detailed discussion.
+The key performance objective that this stream of work focuses on is generation efficiency—the number of LLM calls (or candidate solution generations) needed to reach a target performance level.
+However, in most business settings, the key bottleneck is a limited evaluation budget rather than a generation budget.
+For instance, in digital advertising, generating ad creatives is now cheap and fast with modern LLMs/AI models, but ad evaluation, i.e., measuring whether the ad is effective or not, still requires significant human feedback in the form of field deployment (e.g., A/B tests or bandits in ad platforms) or consumer surveys.
+This is also true for other marketing decisions such as package design, product descriptions, and promotions – in all these cases, generating new candidate solutions is now easy, but evaluating the performance of each solution can take significant time, effort, and human feedback.
+Beyond business applications, this pattern is also common in many social and scientific domains: AI models like AlphaFold can propose protein structures quickly (jumper2021highly), but their functional properties must still be tested through slow and noisy wet-lab experiment; similarly, AlphaChip can generate chip layouts in silico (goldie2024chip), but performance ultimately depends on fabrication and benchmarking.
+In sum, the primary goal for self-improving AI systems in business and social settings should be evaluation efficiency.
+However, as discussed earlier, the existing work has largely focused on generation efficiency as the main objective.
 1.2 Research Agenda and Challenges
-Motivated by the effectiveness of iterative prompt-optimization-based self-improving AI and the overlooked importance of evaluation efficiency in existing frameworks, we now turn to our central question: how can we design a prompt-optimization-based self-improving AI framework that is evaluation-efficient?
-In addition to evaluation efficiency, we would like the framework to be easy to implement, theoretically grounded, and empirically perform well.
-However, realizing these desiderata in a unified framework is nontrivial.
-The key insight in this paper is that prompt-optimization-based self-improving AI follows a familiar pattern we saw in the optimization framework called Bayesian Optimization (BO) (frazier2018tutorial; srinivas2012information; papenmeier2025understanding).
-This insight is grounded in three observations: First, LLMs learn in-context from previous evaluations (brown2020language; coda2023meta; lu2024emergent; cahyawijaya2024llms), acting as Bayesian learners (xie2021explanation; wang2023large; wakayama2025context).
-Second, we iteratively choose a candidate prompt that generates a candidate solution, observe a noisy performance score associated with this solution, and then decide where to query next.
-Third, we have to be cognizant of three concerns: (1) we have no closed-form model linking prompts to scores, (2) we need to balance exploration and exploitation when choosing the next candidate solution, and (3) the goal is to optimize for the number of evaluations rather than candidate generation.
-BO is a principled framework for tackling such black-box optimization problems with such features.
-In BO, we build a function called the acquisition function and choose the point with the top of the acquisition function as the next point to evaluate.
-Among many variants of BO, Upper Confidence Bound (UCB)-BO (srinivas2012information), a version of BO with UCB acquisition function, is proven to be evaluation-efficient, in the sense that no algorithm can achieve better regret in terms of both simple regret and expected cumulative regret (whitehouse2023sublinear; wang2023regret).
-Figure 2 illustrates the procedure of UCB-BO.
-Figure 2: An example of a non-parallel UCB-BO procedure for four optimization steps.
-At each step, we choose the maximum (or multiple local maxima in the parallel UCB-BO) of the blue UCB function as the next point to evaluate.
-As you can see from Figure 2, the UCB acquisition function is often non-concave.
-To resolve this issue, in practice, we often use parallel gradient-based UCB-BO (wilson2018maximizing), where we 1) start from multiple initial points and take gradients to find a set of local optima, and 2) choose those local optima as the points to evaluate next.
-Parallel gradient-based BO has shown surprising empirical effectiveness for high-dimensional problems (papenmeier2025understanding).
-Building on the key insight, we propose to achieve evaluation efficiency by casting iterative prompt optimization-based self-improving AI as parallel gradient-based UCB Bayesian optimization in language space.
-At each iteration, we treat the prompts from the previous round as initialization points and take gradient steps to ascend the UCB acquisition function toward its local optima.
-We then select which prompts to evaluate from among these local optima.
-After each evaluation round, the LLM and its associated critic models implicitly update a surrogate belief over prompt quality via in-context learning and corresponding UCB, analogous to the posterior update in classical Bayesian optimization.
-However, there are two challenges we need to overcome.
-  • $\bullet$
-    Challenge 1– Lack of gradient functions
-    In prompt spaces, the notion of a gradient is ill-defined because they lack the inner-product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction).
-Thus, we first need to define a meaningful notion of a “gradient” in the domain of text.
-Although one might consider going around this challenge by embedding the prompt first, taking the gradient there, and reverting the prompt, this strategy runs into the well-known soft-to-discrete projection problem, where such reverted prompts break any clean correspondence between infinitesimal moves in embedding space and interpretable edits of the underlying natural-language prompt (cui2025automatic).
-  • $\bullet$
-    Challenge 2 – Taking gradients for UCB
-    Second, even if we can define meaningful gradients in prompt space, we do not have direct access to the underlying UCB acquisition function whose gradient we would like to follow.
-The surrogate posterior over prompt quality that induces UCB lives implicitly in the LLM, rather than as an explicit Gaussian-process-style model that exposes closed-form posterior means, variances, and their derivatives.
+In this paper, we seek to design an evaluation-efficient, theoretically grounded, and empirically effective framework for prompt-optimization-based self-improving AI.
+We define evaluation efficiency as: reaching high-performing solutions using as few costly evaluations as possible.
+Our key and novel insight is that iterative prompt-optimization-based self-improving AI can be cast as Bayesian Optimization (BO) in the language space.
+This recasting then allows us to inherit the key properties of classical BO, in particular, it’s evaluation efficiency.
+While intuitive and appealing, we nevertheless need to overcome technical challenges in order to translate BO to the language space.
+We now briefly summarize canonical BO and discuss the parallels between our problem and canonical BO and provide some intuition for why iterative prompt optimization can be viewed as BO in language space, and the describe the challenges that we need to address to fully leverage the BO framework for our problem.
+Bayesian Optimization (BO).
+Classical BO is a principled framework for tackling black-box optimization problems with two defining features: (1) There is no closed-form or analytical expression for the objective function or its derivative.
+In such cases, optimization has to rely on evaluating the function by sampling different points and getting a noisy response.
+(2) Function evaluation is costly.
+If evaluation is cheap, then this problem can be solved easily using methods such as grid search or numerical gradient estimation.
+However, if function evaluation is expensive, then it is important to minimize the number of samples drawn from the black box function.
+This is the regime in which BO is particularly effective (frazier2018tutorial).
+In BO, we maintain two models: (1) a surrogate model, that is used for approximating the objective function (given a history of prior evaluations).
+This model produces a posterior mean and an uncertainty estimate at each point in the space.
+(2) Acquisition model, that helps guide which point should be chosen for evaluation in the next iteration; this model trades-off predicted value and uncertainty (exploration-exploitation).
+Among many variants of BO, Upper Confidence Bound (UCB)-BO (srinivas2012information) is known to be evaluation-efficient, i.e., no algorithm can achieve better in terms of both simple regret and expected cumulative regret (whitehouse2023sublinear; wang2023regret).
+In practice, the UCB acquisition function is often non-concave, so practitioners usually use parallel gradient-based UCB-BO (wilson2018maximizing): start from multiple initial points, take gradient steps to reach several local optima, and evaluate those optima.
+Parallel gradient-based BO can be especially effective in high-dimensional settings (papenmeier2025understanding).
+Iterative prompt optimization as BO.
+The iterative-prompt optimization problem can be summarized as follows: in each iteration, the agent repeatedly selects a prompt, observes a noisy performance score, and decides what to try next.
+The overarching goal is to reach a high-performing prompt (or candidate solution) in as few evaluations as possible.
+We note that this problem shares some key features of the black-box optimization where BO typically performs well: the prompt-space is high-dimensional and there is no closed-form prompt-to-score model, exploration must be balanced with exploitation when we choose the next prompt to evaluate, and the goal is to maximize progress per costly evaluation.
+Building on this insight, we propose to cast iterative prompt optimization as parallel gradient-based UCB Bayesian optimization in language space: prompts from prior rounds serve as initializations, gradient-style updates move toward local optima of an implicit UCB acquisition, and in-context learning plays the role of posterior updating.
+Challenges.
+However, realizing this idea requires overcoming two key challenges.
+First, in prompt space, gradients are ill-defined because text lacks the inner product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction).
+Second, even if we can define a meaningful “textual gradient,” we do not have direct access to the UCB acquisition function (or its derivatives).
+The surrogate posterior over prompt quality that induces the UCB lives implicitly in the LLM, rather than as an explicit Gaussian-process-style model that exposes closed-form posterior means, variances, and their derivatives.
 Consequently, standard gradient-based BO techniques that differentiate the UCB constructed from a known surrogate cannot be applied naively: we must design a mechanism that elicits UCB-like, optimism-in-the-face-of-uncertainty behavior using only black-box text interactions and critic feedback, without ever computing the UCB function or its gradient in closed form.
-In this paper, our goal is to address the above challenges and build an evaluation-efficient self-improving AI that optimally balances exploration and exploitation in the language space.
-1.4 Proposed Method
-Figure 3: Gradient steps of parallel gradient-based BO (Left image) and those of T-BoN BO (Right image). T-BoN BO extends parallel gradient-based BO to the language space.
-We propose a simple, evaluation-efficient iterative prompt-optimization framework, called TextGrad-Best-of-N Bayesian Optimization (T-BoN BO), that adapts parallel gradient-based UCB-BO for prompt optimization based self-improving AI (Figure 3), addressing both of the challenges discussed earlier.
-To solve the first challenge, T-BoN BO builds on the recently proposed TextGrad (textgrad) framework that adapts gradient descent to discrete text by asking an external auxiliary LLM critic to propose targeted edits that improve a candidate; applying these edits induces a local, directional update, i.e., a “textual gradient”, in language space.
-This makes gradient-style search feasible without continuous relaxations or mapping the LLM prompts to the embedding space.
-For instance, consider an ad creative generation prompt for an image model: “A simple photo of a plant-based burger on a white plate.” A critic LLM might suggest: “Make the burger appear delicious, and show it being enjoyed at a barbecue.” Incorporating this edit yields: “A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends.” This change is analogous to following a gradient direction—small, constructive, and aligned with improving performance according to the evaluation signal.
-textgrad show that textual gradients based on their approach enable consistent improvements across diverse scientific domains such as code-completion benchmarks, improving accuracy on graduate-level QA datasets, optimizing drug-like molecules, and refining radiotherapy planning for clinical objectives.
-These results establish textual gradients as an effective surrogate for numerical gradients when optimizing systems that operate in language space.
-To address the second challenge, T-BoN BO chooses the prompt to evaluate next via a series of textual gradients which are selected under the Best-of-N principle (snell2024scaling), resulting in the gradient direction we call Best-of-N gradient.
-Moving in the textual gradient direction suggested by the critic exploits the expected improvement implied by past results but ignores uncertainty in those expectations.
-Therefore, it does not climb the UCB acquisition function.
-Best-of-N gradient bakes in exploration into the choice of gradient direction by adopting an optimism-based rule in the spirit of UCB’s optimism-in-the-face-of-uncertainty principle: it samples multiple local textual edits and moves along the direction the LLM judges most promising, i.e., the Best-of-$N$ candidate.
-That is, the Best-of-$N$ gradient is designed to simulate the implicit UCB acquisition gradient in language space.
-We now briefly describe our T-BoN BO algorithm.
-At each optimization iteration, in Step 1 (meta-reflection phase), we use the critic model to analyze the history of previously evaluated solutions and their scores, and update a meta-reflection that summarizes what has tended to work well or poorly.
+1.3 TextBO: Algorithm, Theory, and Numerical Experiments
+We propose TextBO, a simple evaluation-efficient prompt-optimization based self-improving AI framework that adapts parallel gradient-based UCB-BO to language space by addressing the above challenges.
+To address the first challenge, TextBO builds on TextGrad (textgrad), which adapts gradient descent to discrete text by using an auxiliary LLM critic to propose targeted edits; applying these edits induces a local, directional update (a “textual gradient”) in language space.
+For example, starting from “A simple photo of a plant-based burger on a white plate,” A critic LLM might suggest: “Make the burger appear delicious, and show it being enjoyed at a barbecue.” Incorporating this edit yields: “A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends.” Such edits serve as language-native analogues of gradient steps.
+To address the second challenge, TextBO selects gradient updates via the Best-of-N principle (snell2024scaling), producing what we call the Best-of-N gradient.
+A single textual gradient step exploits expected improvements implied by past results but does not account for uncertainty.
+Best-of-N gradient injects exploration into the acquisition by sampling multiple local edits and moving in the direction judged most promising by the critic—an optimism-based rule designed to emulate ascent on an implicit UCB acquisition gradient in language space.
+Algorithm.
+We now briefly describe our TextBO algorithm.
+Each iteration in the algorithm consists of three steps.
+In Step 1 (meta-reflection phase), we use the critic model to analyze the history of previously evaluated solutions and their scores, and update a meta-reflection that summarizes what has tended to work well or poorly.
 In Step 2 (Best-of-N textual-gradient phase), we perform a sequence of Best-of-N gradient updates from the current prompt.
 That is, for each gradient step, (i) we propose $N$ textual-gradient edits that transform the current prompt into $N$ candidate prompts, (ii) each candidate prompt is passed to a solution-generation model to produce a corresponding candidate solution, (iii) conditioned on the meta-reflections, the critic model predicts which candidate solution is most likely to perform best, and (iv) we set the prompt associated with this predicted-best solution as the new current prompt.
-After iterating Step 2, we obtain a refined prompt and its associated solution to evaluate.
-In Step 3 (evaluation phase), we evaluate this solution to get a scalar score, which T-BoN BO then uses to update its history and meta-reflection for subsequent optimization steps.
-1.5 Theoretical Guarantees and Empirical Evaluation
+After Step 2, we obtain a refined prompt and its associated solution to evaluate.
+In Step 3 (evaluation phase), we evaluate this solution to get a scalar score, which TextBO then uses to update its history and meta-reflection for subsequent optimization steps.
+Theory.
 We provide theoretical guarantees on the performance of our approach.
-We formally show that, under mild regularity conditions, the Best-of-$N$ gradient over locally sampled textual edits induces, in probability, ascent directions of the UCB acquisition function with an exploration parameter $\beta_{N}=\Theta(\sqrt{\ln N})$ (Theorem 2).
-In other words, T-BoN BO emulates parallel gradient-based UCB Bayesian Optimization in the implicit embedding space induced by the LLM, even though it operates purely in language space and does not construct an explicit surrogate or uncertainty model.
-Since UCB-BO is known to achieve sublinear regret and optimal evaluation-efficiency guarantees, T-BoN BO inherits these theoretical properties up to constant factors and approximation error stemming from textual gradients.
-In sum, we formally establish that T-BoN BO is theoretically equivalent to the parallel gradient-based implementation of Bayesian Optimization with the Upper Confidence Bound (UCB) acquisition function (srinivas2012information; wilson2018maximizing), which is proven to be evaluation-efficient (whitehouse2023sublinear; wang2023regret).
-This equivalence establishes T-BoN BO as a valid, eval-efficient, parallel gradient-based Bayesian optimization procedure in the language space.
-In addition to its theoretical efficiency, T-BoN BO is simple and easy to implement since it does not require any explicit uncertainty modeling, embeddings, or Gaussian process surrogates.
-We now demonstrate the empirical performance of our approach using a series of numerical experiments.
-Concretely, we instantiate T-BoN BO in a digital advertising setting, where the goal is to iteratively refine image-generation prompts for ad creatives so as to maximize ad responsiveness for a population of customers.
+Under mild regularity conditions, we prove that the Best-of-N gradient computed from locally sampled textual edits induces (in probability) an ascent direction of the UCB acquisition function with exploration parameter $\beta_{N}=\Theta(\sqrt{\ln N})$ (Theorem 1).
+Consequently, TextBO effectively implements parallel, gradient-based UCB Bayesian Optimization in the implicit embedding space induced by the LLM, despite operating purely in language space and without constructing an explicit surrogate or uncertainty model.
+It therefore inherits the sublinear-regret and evaluation-efficiency guarantees established for UCB-BO, up to constant factors and approximation error due to textual gradients.
+Numerical experiments.
+We consider two sets of numerical experiments to empirically demonstrate the performance of TextBO – (1) a advertising setting where the goal is to optimize ad creatives, and (2) a series of agentic AI benchmarks from the recent self-improving AI literature.
+In the ad optimization experiments, the goal is to iteratively refine image-generation prompts for ad creatives to maximize responsiveness for a population of customers.
 We test and measure how well and how fast a method optimizes for (i.e., aligns with) the preference distribution induced by LLM-based simulation of personas in the Digital Twins persona dataset called Twin-2k-500 (toubia2025twin).
-[^2] While LLM-based persona simulation can be biased and may not reflect real-world human behavior (li2025llm; peng2025mega), it still induces a valid preference distribution.
-As such, it can serve as a valid benchmark on which we can compare the performance of different AI-based optimization methods.
-We demonstrate that T-BoN BO significantly outperforms strong state-of-the-art baselines such as Best-of-N (snell2024scaling) and GEPA (agrawal2025gepa) in terms of evaluation efficiency, i.e., reaching fixed performance with a smaller number of steps.
-Specifically, we show that:
-  • $\bullet$
-    T-BoN BO, the proposed evaluation efficient framework, achieves nearly twice more improvements from Best-of-N baselines compared to GEPA (agrawal2025gepa), the state-of-the-art prompt optimization-based self-improving AI algorithm, which outperforms GRPO (liu2024deepseek), the state-of-the-art reinforcement learning fine-tuning algorithm.
-  • $\bullet$
-    T-BoN BO outperforms the state-of-the-art baseline even when we simulate preferences without any persona information, i.e., when the evaluator is a single LLM judge with no auxiliary context.
-Together, these results show that T-BoN BO can reach good performance both in settings where there is significant heterogeneity in the preferences of the target population as well as in settings with homogeneous preferences.
+We consider eight advertising scenarios and compare the performance of TextBO to two strong baselines – (1) Best-of-N (snell2024scaling), a popular and well-performing test-time alignment method, and (2) GEPA (agrawal2025gepa), a state-of-the-art prompt optimization based self-improving AI algorithm, which has been shown to outperform reinforcement learning fine-tuning algorithms such as GRPO (liu2024deepseek).
+We show that TextBO significantly outperforms both these baselines, across all eight scenarios.
+Further, in an ablation study, we show that these gains are not simply due to the persona data – even when we simulate preferences without any persona information, i.e., when the evaluator is a single LLM judge with no auxiliary context – TextBO outperforms baseline approaches.
+Our numerical experiments show that TextBO can reach good performance both in settings where there is significant heterogeneity in the preferences of the target population as well as in settings with homogeneous preferences.
+Next, we consider a series of agentic AI benchmark experiments in order to assess whether the same Best‑of‑N multi‑step update principle improves evaluation efficienc.
+Specifically, we consider the experiments used to evaluate GEPA in agrawal2025gepa – HotpotQA, HoVer, and Pupa.
+We then compare the performance of TextBO‑GEPA (GEPA augmented with Best‑of‑N gradient sampling and multiple gradient steps) to GEPA.
+Across all experiments, we show that TextBO-GEPA consistently outperforms GEPA.
+Together, these experiments establish TextBO as an effective and general purpose self-improving AI algorithm that can be used for data driven decision-making for both business decisions such as ad creatives, product descriptions, as well as agentic tasks.
+1.4 Contributions
 In sum, our paper makes four main contributions to the literature on AI, optimization, and data-driven decision making.
 First, from a conceptual perspective, we identify evaluation efficiency as the key objective to optimize in self-improving AI design in societal and business systems.
-Second, we propose T-BoN BO, a purely text-based self-improving AI framework that combines textual-gradient search with Best-of-$N$ selection to automate and optimize business and marketing decisions (e.g., ad creatives, product descriptions).
-Third, we provide theory showing that Best-of-$N$ over locally sampled textual edits selects, in probability, ascent directions of the UCB acquisition with exploration weight $\beta=\Theta(\sqrt{\ln N})$; thus, we prove that T-BoN BO emulates gradient-based UCB and inherits evaluation-efficiency guarantees.
-Finally, we demonstrate empirical gains on automated ad-alignment across eight scenarios: T-BoN BO improves faster and attains higher final scores compared to strong baselines, including Best-of-64 and GEPA, and remains effective both in settings with and without contextual information.
-The rest of the paper is organized as follows.
-Section 2 discusses the related literature.
-Section 3 formalizes the AI system and its objective, along with self-improving algorithms.
-Section  4 describes the concept of Best-of-N gradient and proves that it theoretically emulates the gradient of the UCB acquisition function.
-Section 5 introduces the main self-improving AI algorithm proposed in this paper, T-BoN BO.
-Sections 6 and 7 empirically demonstrate T-BoN BO’s effectiveness on ad-alignment tasks versus strong baselines.
-Finally, Section 8 concludes with a discussion of implications and limitations.
+Second, we propose TextBO, a purely text-based self-improving AI framework that combines textual-gradient search with Best-of-N selection to automate and optimize business and marketing decisions (e.g., ad creatives, product descriptions).
+TextBO is simple and easy to implement: it requires no explicit uncertainty modeling, embeddings, or Gaussian process surrogates.
+Third, we provide theory showing that Best-of-N over locally sampled textual edits selects, in probability, ascent directions of the UCB acquisition with exploration weight $\beta=\Theta(\sqrt{\ln N})$; thus, we prove that TextBO emulates gradient-based UCB and inherits evaluation-efficiency guarantees.
+Finally, we demonstrate empirical gains using a wide variety of numerical experiments, including automated ad-alignment across eight scenarios as well as the standard agentic AI benchmarks used in the self-improving AI literature – throughout, TextBO improves faster and attains higher final scores compared to strong baselines, and remains effective both in settings with and without contextual information.
 2 Related Literature
@@
 We now discuss each one below.
-The literature on self-improving AI has often been motivated by an observation: the progress in AI based only on human data is hitting limits, so AI must create and learn from its own data.
+The literature on self-improving AI has often been motivated by an important observation: the progress in AI based only on human data is hitting limits, so AI must create and learn from its own data.
 In other words, powerful AI “should have their own stream of experience that progresses, like humans, over a long time-scale” (silver2025welcome).
 That is, we need to design the self-improving AI that iteratively 1) generate data of outcomes, evaluate outcomes with grounded signals, and 2) self-tune the three knobs – models, tools, and prompts – in the system to achieve long-term objectives under real-world feedback (wang2025maestro).
-Here, tuning models is about altering language models’ input-output correspondences (huang-etal-2023-large; lee2023rlaif; acikgoz2025self); tuning tools is about specifying the tools the AI has access to, including memory tools (hou2025model; ouyang2025reasoningbank; qiu2025alita); tuning prompts is about engineering prompts in the model’s context window, including system messages, task instructions, constraints, schemas, and few-shot exemplars (zhou2022large; khattab2023dspy; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro).
-In this paper, we focus on studying self-improving systems based on tuning prompts.
+Here, tuning models is about altering language models’ input-output correspondences (huang-etal-2023-large; lee2023rlaif; acikgoz2025self); tuning tools is about specifying the tools the AI has access to, including memory tools (hou2025model; ouyang2025reasoningbank; qiu2025alita); tuning prompts is about engineering prompts in the model’s context window, including system messages, task instructions, constraints, schema, and few-shot exemplars (zhou2022large; khattab2023dspy; pryzant2023automatic; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro).
+In particular, iterative prompt-optimization based self-improving AI have become increasingly popular.
+In a recent paper, agrawal2025gepa demonstrate their method, GEPA (Genetic-Pareto) is able to outperform state-of-the-art model fine-tuning techniques such as GRPO (liu2024deepseek) under a fixed budget of LLM generation calls across a range of challenging tasks including multi-document QA, constrained instruction following, retrieval-augmented fact verification, and privacy-related objectives.
+At a high-level, both GEPA and other prompt-based self-improvement (agrawal2025gepa; zhang2025agentic; wang2025maestro) is specified as procedural heuristic search: maintain a pool of prompts, generate variants via LLM-driven mutation/crossover/reflection, score candidates on held-out data or rollouts, retain top performers, and iterate (10.5555/3692070.3692611; guo2024connecting; agrawal2025gepa; better_together).
+This mirrors classical heuristic and evolutionary optimization, where performance is assessed as a function of generation efficiency rather than evaluation efficiency (nguyen2016understanding; ding2023quality; hu2024automated; liu2024evolution).
+Yet, as discussed earlier, in many real-world deployments, the scarce resource is not candidate generation but evaluation.
+Thus, in this paper, we focus on developing an iterative prompt-optimization framework that treat evaluation calls as the budget to allocate and optimize.
 Our paper also relates to the literature on relating LLMs to Bayesian Optimization (BO).
-A series of papers focuses on applying LLM to enhance BO for numerical optimization problems.
+A series of papers focuses on applying LLMs to enhance BO for numerical optimization problems.
 liularge and cisse2025language focus on leveraging LLMs’ contextual understanding and few-shot learning capabilities to improve BO’s efficiency, enabling better warm-starting and surrogate modeling; agliettifunbo uses an LLM to discover new acquisition functions for Bayesian optimization.
 The LLM writes candidate acquisition function formulas as code, which are evaluated on various optimization problems.
-Conversely, singh2025mechanistic applies BO to enhance LLM’s internal behavior to boost the model’s zero-shot and few-shot performance.
-In contrast to these papers, our paper focuses on considering the prompt optimization problem that arises in formulating LLM-based self-improving AI system design as a Bayesian Optimization problem.
+singh2025mechanistic applies BO to enhance LLM’s internal behavior to boost the model’s zero-shot and few-shot performance.
+In contrast to these papers, we focus on the prompt optimization problem that arises when formulating LLM-based self-improving AI system design as a Bayesian Optimization problem.
 In line with this focus, schneider2025hyperbandbased and kong2025metapromptoptimizationllmbasedsequential formulate prompt optimization among the fixed set of candidate prompts as Bayesian optimization.
 While schneider2025hyperbandbased utilizes ideas from adversarial bandits, kong2025metapromptoptimizationllmbasedsequential feeds each prompt text into an embedding model, and the GP surrogate operates on those fixed embedding features to guide selection.
-In contrast to these approaches, our work focuses on self-improving AI that actively generate new candidate prompts as the process proceeds, rather than pre-specifying a fixed set of candidate prompts.
-In terms of method evaluation benchmarks, our paper relates to the methodological literature on ad optimization in digital marketing.
-A series of papers focused on the efficient execution of ad evaluation, i.e., iteratively improving ad choice efficiently given a fixed set of candidate ads.
-schwartz2017customer adopts Thompson Sampling with a hierarchical general linear model for a real display advertising campaign for a financial-services firm, which purchased display advertising across 59 websites.
+However, unlike these papers, we actively generate new candidate prompts as the process proceeds, rather than pre-specifying a fixed set of candidate prompts.
+Next, our paper relates to the literature on efficient ad evaluation in digital marketing using bandit methods.
+Here, the goal is to iteratively improving the ads served given a fixed set of candidate ads.
+schwartz2017customer adopts Thompson Sampling with a hierarchical general linear model for a real display advertising campaign for a financial-services firm.
 geng2020online partitions JD.com’s platform users into disjoint sub-populations and applies a contextual Thompson sampling algorithm to maximize the expected advertiser payoff.
@@
 ba2022advertising used parametric Thompson Sampling, where a parametric structure links features (media, audience attributes) to conversion likelihood to optimally allocate ad-media and target-audience combinations in high-dimensional settings with low success rates.
-The main difference between bandit-based papers and ours is the existence of a pre-defined set of candidate arms.
-Self-improving AI focuses on creating new solutions by analyzing previous results.
-On the other hand, bandit methods aim to balance exploration of a predefined set of candidate ads optimally.
-More broadly, there is also a difference in problem scope.
-Self-improving AI iterates among three steps: generation, evaluation, and analysis.
-In contrast, bandit methods focus on efficiently executing the evaluation step.
-Indeed, an AI self-improvement framework’s evaluation step may employ a bandit method to improve efficiency further.
-3 Setup and Background
-3.1 The optimization objective.
-Consider a system designer’s decision-making problem, where she chooses a prompt $\Pi$ for the LLM-based AI system $\Phi$.
+The main difference between these bandit-based papers and ours is the existence of a pre-defined set of candidate arms.
+Our paper focuses on creating new solutions by analyzing previous results.
+On the other hand, bandit methods aim to balance exploration of a pre-defined set of candidate ads optimally.
+The scope of our problem is also quite different.
+Our self-improving AI iterates over three steps–generation, evaluation, and analysis–whereas, bandits focus on efficiently executing the evaluation step.
+Indeed, one could employ a bandit method, such as best-arm-identification, in our evaluation step (instead of using A/B tests).
+Finally, our paper relates to the literature on ad content.
+Researchers have studied many different features of ad creatives and their effect on consumer response, e.g., creativity (rosengren2020meta), discount information (biswas_2025), emotionality of content (macinnis_etal_2002), and specific text phrases (rutz_etal_2017).
+However, a limitation of this literature is that each study is designed to test one specific ad feature in one specific advertising format/context (e.g., TV ads, search ads), and there is not much consensus on how various features and estimated effects should be combined to optimize the ad creative and whether these effects can be ported to other formats and audiences.
+In this paper, we move away from manually manipulating individual ad features and instead treat the ad-creative optimization problem as a high-dimensional optimization problem in the prompt-space; and our self-improving AI framework can be used to automate and optimize ad-creative generation at scale for any given combination product, ad-format, and audience.
+3 Problem Definition
+Let $\Phi$ denote a generative AI model (or more generally, an AI system) that, given a prompt $\pi\in\Pi$, produces a candidate solution $\Phi(\pi)$.
 [^2] LLM-based AI systems can also be modified by tuning the LLM’s internal parameters using LLM fine-tuning methods such as GRPO (liu2024deepseek).
 In this paper, we focus on prompt optimization, which fixes those internal parameters throughout.
-We call $\Phi(\Pi)$ the ‘system outcome’ of $\Phi$ corresponding to a prompt $\Pi$.
-In a digital marketing setting, $\Phi$ is an image generation AI, and $\Phi(\Pi)$ is the ad generated by the prompt $\Pi$.
-As a system, $\Phi(\Pi)$ has inputs and corresponding outputs.
-Specifically, $\Phi(\Pi)$ maps the input space $\mathcal{X}$ to a distribution over an output space $\mathcal{Y}$.
-For each $x\in\mathcal{X}$, we denote by $y(\Phi(\Pi),x)$ the random $\mathcal{Y}$-valued vector output of the system at input $x$.
-In a digital marketing setting, input $x\in\mathcal{X}$ is a platform user, and a realization of $y(\Phi(\Pi),x)$ represents that user’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\Pi)$.
-The distribution of inputs over $\mathcal{X}$ may depend on the prompts set $\Pi$.
-For example, in a digital marketing setting, a recommender system on the ad platform determines which users are shown the ad $\Phi(\Pi)$.
-To capture such dependences, we denote the input distribution over $\mathcal{X}$ by $\mathcal{D}_{\mathcal{X}}(\Pi)$.
+This candidate solution $\Phi(\pi)$, in conjunction with some environmental variables $x\in\mathcal{X}$ produces an output $y(\Phi(\pi),x)$, where $y$ is drawn from a distribution over an output space $\mathcal{Y}$.
+For example, in the ad creative optimization example from $\S$1, $\Phi$ denotes an image generation AI model (e.g., Gemini Flash), $\pi$ is an image generation prompt that reflects what the marketer wants the ad to depict, $\Phi(\pi)$ is the ad creative generated by prompt $\pi$, $x\in\mathcal{X}$ denotes a user on the ad platform and/orother contextual features that affect the ad’s effectiveness (e.g., seasonality or time of day), and a realization of $y(\Phi(\pi),x)$ represents that user/context $x$’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\pi)$.
+Throughout, we allow for the distribution over $\mathcal{X}$ to depend on the prompt $\pi$.
+This is important for practical applications where such dependencies are common; for example, in digital advertising platforms, recommendation systems often determine which users are shown the ad $\Phi(\pi)$ based on the ad content, which depends on $\pi$.
+To capture such dependencies, we denote the input distribution over $\mathcal{X}$ by $\mathcal{D}_{\mathcal{X}}(\pi)$.
+In a fully randomized test, the distribution over $\mathcal{X}$ is independent of $\pi$, and we can denote the distribution of $\mathcal{X}$ as $\mathcal{D}_{\mathcal{X}}$.
 To evaluate the output, we use a scoring rule: a function $r:\mathcal{Y}\to[0,1]$ that assigns a scalar effectiveness score to each $y\in\mathcal{Y}$.
 In the digital marketing example, $r$ is a marketer-defined function that maps observed user responses (clicks, conversions, spend) to a normalized scalar effectiveness score.
-Our optimization objective is to discover the prompt $\Pi^{*}$ that maximizes the objective function
+Then, our optimization objective is to discover the prompt $\pi^{*}$ that maximizes the objective function:
 $J(\pi):=\mathbb{E}[r(y(\Phi(\pi),x))].$ (1)
-Note that $J$ cannot be optimized by directly computing gradients, because the input distribution $\mathcal{D}_{\mathcal{X}}(\Pi)$ can shift with $\Pi$; even when $J(\Pi)$ is differentiable with respect to $\Pi$ and $y$’s randomness is not dependent on $\Pi$, we have:
+$J$ cannot be optimized by directly computing gradients, because the input distribution $\mathcal{D}_{\mathcal{X}}(\pi)$ can shift with $\pi$; even when $J(\pi)$ is differentiable with respect to $\pi$ and $y$’s randomness is not dependent on $\pi$, we have:
 $\nabla_{\pi}J(\pi)=\underbrace{\mathbb{E}\big[\nabla_{\pi}r(y(\Phi(\pi),x))\big]}_{\text{model term}}\;+\;\underbrace{\mathbb{E}\big[r(y(\Phi(\pi),x))\,\nabla_{\pi}\log d_{\pi}(x)\big]}_{\text{distribution term}},$ (2)
-where $d_{\Pi}$ is the density of $\mathcal{D}_{\mathcal{X}}(\Pi)$.
+where $d_{\pi}$ is the density of $\mathcal{D}_{\mathcal{X}}(\pi)$ (williams1992simple).
 Since $d_{\pi}$ and $\nabla_{\pi}\log d_{\pi}(x)$ are unobservable, we cannot compute gradients from observed rewards.
-This motivates the use of an external LLM that can analyze prior data and propose directions for improvement without explicitly computing gradients, leading to self-improving AI algorithms we describe below that iteratively optimize for $J$.
-3.2 Self-improving AI algorithm
-Given a scoring rule $r$, a self-improving AI algorithm $\mathcal{A}$ optimizes $J(\Pi):=\mathbb{E}[r(y(\Phi(\Pi),x))]$ through iteratively analyzing the previous history, proposing a new prompt, and observing its noisy score.
-Formally, starting from the initial prompt $\Pi_{0}$, a self-improving AI algorithm proceeds in discrete optimization iterations, indexed by $t=1,2,\ldots$; at the beginning of iteration $t$, the algorithm has access to the history of past prompts and their observed scores, then selects a new prompt $\Pi_{t}$, deploys the corresponding system outcome $\Phi(\Pi_{t})$, and receives a score feedback on its performance.
-We define the score $s_{t}$ of prompts $\Pi_{t}$ as the empirical average reward at $t$, i.e.,
+This motivates the use of an external LLM that can analyze prior data and propose directions for improvement without explicitly computing gradients, leading to self-improving AI algorithms we describe below that iteratively optimize for JJ.
+[^3] In fully randomized experiments, the distribution term drops out and only the model term remains.
+Self-improving AI algorithm:
+Given a scoring rule $r$, a self-improving AI algorithm $\mathcal{A}$ optimizes $J(\pi):=\mathbb{E}[r(y(\Phi(\pi),x))]$ by iteratively analyzing the previous history, proposing a new prompt, and observing its noisy score.
+Formally, starting from the initial prompt $\pi_{0}$, a self-improving AI algorithm proceeds in discrete optimization iterations, indexed by $t=1,2,\ldots$; at the beginning of iteration $t$, the algorithm has access to the history of past prompts and their observed scores, then selects a new prompt $\pi_{t}$, deploys the corresponding candidate solution $\Phi(\pi_{t})$, and receives a score feedback on its performance.
+We define the score $s_{t}$ of the prompt $\pi_{t}$ as the empirical average reward at $t$, i.e.,
 $s_{t}:=\frac{1}{L}\sum_{l=1}^{L}r\left(y\left(\Phi\left(\pi_{t}\right),x_{t}^{(l)}\right)\right),\quad x_{t}^{(l)}\sim\mathcal{D}_{\mathcal{X}}\left(\pi_{t}\right),$ (3)
-where $L$ denotes the number of input samples $x_{t}^{(l)}$ we evaluate for each $t$. Note that we compare the performance of algorithms in terms of the progress of $s_{t}$ across iterations $t$.
+where $L$ denotes the number of input samples $x_{t}^{(l)}$ we evaluate for each $t$.
+Note that we compare the performance of self-improving algorithms in terms of the progress of $s_{t}$ across iterations $t$.
 Definition 1 (Self-improving AI algorithm).
-Fix a system $\Phi$ and a scoring rule $r$. For each iteration $t\geq 1$, let the history of past evaluations be
-$H_{t-1}\;:=\;\{(\Pi_{\tau},c_{\tau},s_{\tau})\}_{\tau=0}^{t-1}$ (4)
-where $s_{\tau}$ is the empirically evaluated score of $\Pi_{\tau}$, defined as
+Fix a generative model $\Phi$ and a scoring rule $r$.
+For each iteration $t\geq 1$, let the history of past evaluations be $H_{t-1}\;:=\;\{(\pi_{\tau},\Phi(\pi_{\tau}),s_{\tau})\}_{\tau=0}^{t-1}$, where $s_{\tau}$ is the empirically evaluated score of $\pi_{\tau}$, defined as
 $s_{t}=\frac{1}{L}\sum_{l=1}^{L}r\bigl(y(\Phi(\pi_{t}),x_{t}^{(l)})\bigr),\quad x_{t}^{(l)}\sim\mathcal{D}_{\mathcal{X}}(\pi_{t}).$ (4)
 Then, a self-improving AI algorithm is defined as function $\mathcal{A}$ that, at each iteration $t$, maps the history $H_{t-1}$ to a new prompt $\pi_{t}$, i.e.,
-$\Pi_{t}=\mathcal{A}_{t}(H_{t-1})$
-The goal of $\mathcal{A}$ is to choose the sequence $(\Pi_{t})_{t\geq 0}$ so as to optimize the objective $J(\Pi)$, equivalently to drive the sequence of empirical scores $(s_{t})_{t\geq 1}$ upward over iterations in expectation.
+$\pi_{t}=\mathcal{A}_{t}(H_{t-1}).$
 Operationally, given access to the history $H_{t-1}$, a self-improving AI algorithm $\mathcal{A}$ goes through following three steps each iteration $t$:
   1. 1.
-    Analysis: The critic LLM analyzes the previous history $H_{t-1}$.
+    Analysis: Analyze the previous history $H_{t-1}$.
   2. 2.
-    Generation: The critic LLM proposes the next prompts $\Pi_{t}$ based on the analysis.
+    Generation: Propose the next prompts $\pi_{t}$ based on the analysis.
   3. 3.
     Evaluation: The system outcome $\Phi(\pi_{t})$ is deployed, and its performance is measured, yielding a score $s_{t}$.
-The new observation is added to the history: $H_{t}=H_{t-1}\cup\{(\Pi_{t},c_{t},s_{t})\}$.
-3.3 Bayesian Optimization (BO)
-As discussed in Section 1, we would like to formulate self-improving AI as parallel gradient-based UCB-BO.
-At its core, BO maximizes an unknown and expensive-to-evaluate objective function under a limited evaluation budget.
-Formally, consider a black-box function $f:\mathcal{Z}\to\mathbb{R}$, where in our setting $f$ corresponds to the system-level objective $J(\Pi)$ over prompts.
-BO maintains a surrogate model, typically a Gaussian Process (GP) or another probabilistic regressor defined on $\mathcal{Z}$, that summarizes current beliefs about $f$.
-For any candidate $z\in\mathcal{Z}$, the surrogate yields a posterior mean $\mu(z)$ and an associated uncertainty measure $\sigma(z)$.
-In BO, the following evaluation point is selected by choosing the point that maximizes an acquisition function $A(z)$ constructed from $\mu(z)$ and $\sigma(z)$, rather than by directly maximizing $\mu(z)$.
-Acquisition functions encode an explicit trade-off between exploitation (preferring points with high predicted value) and exploration (preferring points with high uncertainty).
-A widely used choice is the Upper Confidence Bound (UCB) acquisition function
-$A_{\beta}(z)\;=\;\mu(z)\;+\;\beta\,\sigma(z),$ (6)
-where the exploration parameter $\beta>0$ controls the strength of the exploration term (srinivas2012information).
-Intuitively, UCB acquisition function gives a positive bonus to unexplored areas in addition to the posterior mean.
-This allows us to follow the “optimism in the face of uncertainty” principle, which is known to trade off exploration and exploitation effectively.
-Indeed, UCB-BO is an optimal algorithm for achieving evaluation efficiency, in the sense that no algorithm can achieve better regret guarantees (wang2023regret; whitehouse2023sublinear) (See Web Appendix E for a detailed theoretical discussion of BO.)
-In practice, since the maxima of the UCB acquisition function are often not available in a closed form, we typically use parallel gradient-based BO with the UCB acquisition function (wilson2018maximizing).
-In parallel gradient-based BO, we start from multiple initial points and take gradients from them to climb up the UCB acquisition function to find its local maxima.
-We then choose the points to evaluate among the local maxima (Figure 3).
-In our setting, we do not construct an explicit $\mu(\cdot)$ and $\sigma(\cdot)$ (and their corresponding surrogate model and acquisition function) for the prompt space; instead, we interpret $\mu(\cdot)$ and $\sigma(\cdot)$ as implicit beliefs within the LLM and design our method to emulate parallel gradient-based UCB in this implicit space.
-Textual gradient (textgrad) is designed to elicit gradients for $\mu(\cdot)$ in the prompt space without explicitly modeling those beliefs within the LLM.
-However, what we want is not gradients for $\mu(\cdot)$; we want gradients for $A_{\beta}(z)$, the UCB acquisition function.
+The new observation is added to the history: $H_{t}=H_{t-1}\cup\{(\pi_{t},\Phi\left(\pi_{t}\right),s_{t})\}$.
+We will be more precise about how Steps 1 and 2 (analysis and generation) are implemented in our self-improving AI framework in $\S$5 and $\S$6.
+Optimization objective:
+In this paper, we formulate the goal of self-improving AI algorithms $\mathcal{A}$ as improving prompts $\left(\pi_{t}\right)_{t=1}^{T}$ over the evaluation budget $T$ to efficiently optimize the objective $J(\pi)$.
+Specifically, we discuss evaluation efficiency using two standard regret criteria.
+Let $\pi^{\star}\in\arg\max_{\pi\in\Pi}J(\pi)$.
+Then,
+  • $\bullet$
+    The (expected) cumulative regret is $\mathbb{E}\left[\sum_{t=1}^{T}\left(J\left(\pi^{\star}\right)-J\left(\pi_{t}\right)\right)\right]$.
+  • $\bullet$
+    The (expected) simple regret is $\mathbb{E}\left[J\left(\pi^{\star}\right)-\max_{1\leq t\leq T}J\left(\pi_{t}\right)\right]$.
+In the next section, we show how a self-improving AI algorithm can be viewed as UCB-based Bayesian Optimization, which is optimal under both regret criteria (whitehouse2023sublinear; wang2023regret).
+4 Our Approach: Recasting Iterative Prompt Optimization as Bayesian Optimization in Language Space
+As discussed in $\S$1, our key idea is to model prompt-based self-improving AI as a Bayesian Optimization (BO) problem in language space.
+At a high level, we want a self-improving AI algorithm that repeatedly chooses a prompt, observes a noisy scalar performance signal (e.g., ad effectiveness), and decides which prompt to try next under a limited evaluation budget.
+We note that our optimization problem has no closed-form expression mapping the prompt to a performance score and a key constraint is that evaluating the function is costly.
+This is precisely the regime where BO has been shown to be effective, i.e., black-box objective functions where function evaluation is costly (frazier2018tutorial; srinivas2012information; papenmeier2025understanding).
+Canonical UCB Bayesian Optimization.
+To fix ideas, we start with a quick overview of BO.
+Classical BO considers a black-box function $f:\mathcal{Z}\to\mathbb{R}$ defined over a continuous input space $\mathcal{Z}$.
+The algorithm maintains a surrogate model—typically a Gaussian Process or related probabilistic regressor—that, given a history of evaluations $H_{t-1}=\{(z_{\tau},y_{\tau})\}_{\tau=1}^{t-1}$, produces for each $z\in\mathcal{Z}$ a posterior mean $\mu_{t-1}(z)$ and an uncertainty estimate $\sigma_{t-1}(z)$.
+Rather than maximizing $\mu_{t-1}(z)$ directly, BO selects the next query by maximizing an acquisition function that trades off exploitation (high predicted value) and exploration (high uncertainty).
+A widely used acquisition is the Upper Confidence Bound (UCB)
+$A_{\beta_{t}}(z)\;=\;\mu_{t-1}(z)\;+\;\beta_{t}\,\sigma_{t-1}(z),$ (5)
+where $\beta_{t}>0$ controls the strength of exploration (srinivas2012information).
+Figure 2 illustrates how $A_{\beta_{t}}$ combines the surrogate mean and uncertainty.
+Intuitively, $A_{\beta_{t}}(z)$ implements an “optimism in the face of uncertainty” rule: points with large uncertainty receive a positive bonus even if their current mean is modest.
+Figure 2: Illustration of the upper confidence bound (UCB) acquisition function in Bayesian optimization.
+The unknown objective $f(z)$ (orange) is evaluated at a set of observed points ($\times$).
+A surrogate model yields the posterior mean $\mu(z)$ (green dashed) and $\sigma(z)$.
+The UCB acquisition $A_{\beta}(z)=\mu(z)+\beta\sigma(z)$ (blue) trades off exploitation and exploration by favoring locations with high predicted value and high uncertainty.
+Input: Evaluation budget $T$, exploration parameter schedule $\beta$, initial history $H_{0}=\emptyset$
+Output: Sequence of evaluation points $\{z_{t}\}_{t=1}^{T}$ and observations $\{y_{t}\}_{t=1}^{T}$
+1 for $t=1,\dots,T$ do
+2   Update surrogate model $p(f\mid H_{t-1})$ on past data
+3   Construct the UCB acquisition
+$A_{\beta}(z)\;=\;\mu_{t-1}(z)\;+\;\beta\,\sigma_{t-1}(z).$
+Select the next evaluation point
+$z_{t}\;\in\;\operatorname{argmax}_{z}A_{\beta}(z).$
+Evaluate $f$ at $z_{t}$ to obtain a noisy observation
+$y_{t}=f(z_{t})+\text{noise}.$
+Augment the data: $H_{t}\leftarrow H_{t-1}\cup\{(z_{t},y_{t})\}$
+4
+Algorithm 1 Canonical UCB Bayesian Optimization
+Algorithm 1 summarizes the resulting UCB-BO loop.
+At each iteration $t$, the surrogate is updated on $H_{t-1}$, the acquisition $A_{\beta_{t}}$ is constructed, the next point $z_{t}$ is chosen by maximizing $A_{\beta_{t}}$, the black-box function $f$ is evaluated at $z_{t}$ to obtain $y_{t}$, and the history is augmented.
+Under mild conditions, UCB-BO achieves sublinear regret and enjoys strong evaluation-efficiency guarantees: no algorithm can perform uniformly better in terms of regret scaling (srinivas2012information; whitehouse2023sublinear; wang2023regret).
+In many practical settings, the global maximizer of $A_{\beta_{t}}$ is not available in closed form, and the acquisition is highly non-convex.
+A standard remedy is parallel gradient-based BO (wilson2018maximizing; papenmeier2025understanding): one starts from multiple initial points, performs gradient ascent on $A_{\beta_{t}}$ from each initialization to find local maxima, and then evaluates $f$ at these local optima in parallel.
+This procedure has shown surprising empirical effectiveness in high-dimensional numerical optimization (papenmeier2025understanding).
+Mapping self-improving AI to UCB-BO.
+We now conceptually align our iterative prompt-optimization problem with the BO template above. The correspondence is:
+  • $\bullet$
+    The BO input space $\mathcal{Z}$ corresponds to the prompt space $\Pi$.
+  • $\bullet$
+    The black-box function $f(z)$ corresponds to the system-level objective $f(\pi)\equiv J(\pi):=\mathbb{E}[r(y(\Phi(\pi),x))]$, where $x\sim\mathcal{D}_{\mathcal{X}}(\pi)$ and $r$ is the scoring rule defined in § 3.
+  • $\bullet$
+    The evaluation of $f(\pi)$ is the costly step: it requires deploying $\Phi(\pi)$, interacting with the environment (e.g., ad platform users), and computing $s_{t}$ as an empirical estimate of $J(\pi)$.
+  • $\bullet$
+    The cheap computation is everything the LLM-based system can do without new evaluations: analyzing past history, proposing textual edits, or internally comparing candidate ads.
+If we had an explicit surrogate model over prompts, providing $\mu_{t}(\pi)$ and $\sigma_{t}(\pi)$ for each $\pi\in\Pi$, we could in principle construct a UCB acquisition
+$A_{\beta_{t}}(\pi)\;=\;\mu_{t}(\pi)\;+\;\beta_{t}\,\sigma_{t}(\pi),$ (6)
+and then run parallel gradient-based BO over $\Pi$: start from multiple initial prompts, ascend the gradient $\nabla_{\pi}A_{\beta_{t}}(\pi)$ to obtain local optima, and evaluate $J(\pi)$ only at those local optima.
+This would give us a self-improving AI algorithm that is, by construction, evaluation-efficient.
+However, directly instantiating this plan in language space runs into two central challenges (originally discussed in $\S$1).
+Challenge 1: No native gradients in prompt space.
+Prompts are discrete sequences of tokens, not elements of a Euclidean vector space.
+As such, prompt space lacks the inner-product structure of a Hilbert space that underlies standard notions of directions, norms, and gradients (young1988introduction).
+Expressions such as $\nabla_{\pi}A_{\beta_{t}}(\pi)$ are therefore not well-defined in the usual calculus sense.
+A seemingly natural workaround is to embed prompts into a continuous space, take gradients there, and then “project back” to text:
+$\pi\;\xrightarrow{\text{embed}}\;e=\mathsf{E}(\pi)\in\mathbb{R}^{d}\;\xrightarrow{\text{gradient step}}\;e^{\prime}\;\xrightarrow{\text{decode}}\;\pi^{\prime}=\mathsf{E}^{-1}(e^{\prime}).$ (7)
+where $\mathsf{E}$ is the prompt embedding.
+Such a strategy is indeed theoretically plausible thanks to the invertibility of prompt embedding (nikolaou2025language).
+However, in practice, it encounters the well-known soft-to-discrete projection problem: learning the inverse prompt embedding ($\mathsf{E}^{-1}$), i.e., learning to decode an embedding into a prompt, is highly challenging (cui2025automatic).
+As a result, we cannot reliably implement gradient ascent in prompt space by differentiating through embeddings.
+Challenge 2: No explicit UCB or uncertainty model.
+Even if we had a notion of gradient in language space, we would still face a second obstacle: the UCB acquisition function itself is not explicitly available.
+In standard BO, the surrogate model exposes both a posterior mean $\mu_{t}(\cdot)$ and an uncertainty estimate $\sigma_{t}(\cdot)$, from which we build $A_{\beta_{t}}(\cdot)$ and compute its gradient.
+In our setting, by contrast, there is no explicit Gaussian process or parametric surrogate over prompts.
+The relevant beliefs about prompt quality and uncertainty live implicitly inside the LLM and its critic: they are encoded in how the model generalizes from past evaluations, not in an externally accessible $\mu_{t}$ or $\sigma_{t}$.
+Consequently, we cannot simply write down $A_{\beta_{t}}(\pi)$, evaluate it at arbitrary prompts, or compute $\nabla_{\pi}A_{\beta_{t}}(\pi)$ in closed form.
+Naive proxies for uncertainty (e.g., self-consistency variance, embedding distances) tend to be poorly calibrated or brittle, and training a separate Gaussian Process over high-dimensional text embeddings is often expensive, data-hungry, and empirically unstable.
+Textual gradients as a partial solution.
+Recent work on textual gradients offers a way to tackle the first challenge.
+The TextGrad framework (textgrad) shows that one can approximate gradient ascent on an implicit objective by asking an auxiliary LLM critic to propose local natural-language edits that are predicted to improve a scalar score.
+Applying these edits to the current prompt yields a sequence of locally improving prompts, which empirically and theoretically behaves as gradient ascent on the posterior mean $\mu(\pi)$ in an implicit embedding space thanks to bi-Lipschitzness of LLM’s prompt embedding (nikolaou2025language; tang2025understanding).
+However, textual gradients as in textgrad are inherently exploitative: they aim to climb the posterior mean $\mu(\pi)$, ignoring epistemic uncertainty $\sigma(\pi)$.
+They, therefore, correspond to gradient ascent on $\mu(\pi)$, not on the UCB acquisition $A_{\beta_{t}}(\pi)=\mu(\pi)+\beta_{t}\sigma(\pi)$.
+To recover the evaluation-efficiency guarantees of UCB-BO, we need an update rule whose effective search direction behaves like $\nabla_{\pi}A_{\beta_{t}}(\pi)$, even though we never observe $\sigma(\pi)$ or $\nabla_{\pi}\sigma(\pi)$ explicitly.
+Our approach.
+The central technical idea of this paper is to address Challenge 1 and Challenge 2 jointly by combining textual gradients with a Best-of-N selection mechanism.
+Rather than differentiating an explicit acquisition function, we (i) use a critic LLM to generate multiple stochastic textual edits around the current prompt, (ii) apply these edits to obtain nearby candidate prompts and system outcomes, and (iii) ask the critic, conditioned on past history, to pick the most promising candidate in a Best-of-N fashion.
+In $\S$5, we formalize this procedure as a Best-of-N gradient and show that, under mild regularity conditions, the direction selected by this mechanism converges (in probability) to the ascent direction of a UCB-style acquisition function in the LLM’s implicit embedding space, with exploration parameter $\beta_{N}=\Theta(\sqrt{\ln N})$ (Theorem 1).
+This theoretical result gives us a principled way to emulate parallel gradient-based UCB-BO entirely in language space using only black-box text interactions.
+We then leverage this equivalence to design a concrete self-improving AI algorithm in $\S$6, TextBO, that is evaluation-efficient, simple to implement, and readily deployable in real-world applications such as digital advertising.
 5 Best-of-N gradient and Theory
-As discussed in Section 1, a core challenge in casting self-improving AI as parallel, gradient-based UCB Bayesian optimization in the space of prompts is the absence of an explicit acquisition function (e.g., UCB) and its gradient.
-As we discussed earlier, textual gradients (textgrad) can guide the prompt toward the highest expected improvement based on immediate feedback, as a textual gradient elicits the gradient of the posterior mean $\nabla\mu(x)$.
-However, following the posterior mean’s gradient rather than UCB’s is a purely exploitative strategy that often leads to suboptimal local maxima.
-In this section, we propose that a gradient direction we call Best-of-N gradient, which generates $N$ stochastic textual gradients and selecting the most promising candidate via a Best-of-$N$ mechanism, can elicit the gradient of the UCB acquisition function within LLM’s mind without explicitly constructing the LLM’s implicit posterior mean and the associated uncertainty.
-To operationalize the Best-of-N gradient, we should first formalize the requisite capabilities of the external Large Language Model, which we designate as the Critic LLM.
-4.1 Critic LLM.
-Let $\Pi$ denote a prompt and $c=\Phi(\Pi)$ denote the corresponding system outcome.
-We define the history of optimization up to iteration $t$ as $H_{t}=\{(\Pi_{\tau},c_{\tau},s_{\tau})\}_{\tau\leq t}$.
-Then we define the Critic LLM, $M_{\text{critic}}$, as an LLM that can conduct the operations of the following four distinct operators:
+As discussed in $\S$4, our aim is to implement UCB-style Bayesian Optimization in prompt space using only black-box access to an LLM critic.
+The theoretical challenge is to show that the combination of (i) textual gradients and (ii) Best-of-N selection actually induces gradient ascent on a UCB acquisition function in the LLM’s implicit embedding space.
+This section formalizes that idea and proves our main result.
+We proceed in three steps.
+In $\S$5.1, we formalize the role of the critic LLM, which provides the primitive operations—pairwise judgments, textual gradients, and meta-reflections—that enable search in language space.
+$\S$5.2 then introduces the Best-of-N gradient, our key mechanism for turning these critic capabilities into a stochastic, locally improving update rule.
+Finally, in $\S$5.3, we prove that Best-of-N gradients asymptotically emulate gradient ascent on a UCB acquisition function in the LLM’s implicit embedding space.
+This establishes the theoretical justification for the TextBO framework developed in the next section.
+5.1 Critic LLM
+The use of a separate LLM as a critic is now a standard pattern in the prompt-optimization and self-improving AI literature: critic models are routinely used to provide structured feedback, propose improvements, or act as reward models and judges that guide search over prompts and outputs (zhou2022large; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro; shi2024judging).
+We therefore treat the critic LLM as a natural and practically realistic component that supplies the basic operations needed to implement Bayesian-optimization-style updates in language space.
+Recall that we denote a prompt as $\pi$ and the corresponding candidate solution (e.g., an ad creative) as $\Phi(\pi)$, the optimization history up to iteration $t$ as $H_{t}=\{(\pi_{\tau},\Phi(\pi_{\tau}),s_{\tau})\}_{\tau\leq t}$, where $s_{\tau}$ is the evaluated score of $\pi_{\tau}$.
+Then, let $M_{\text{critic}}$ denote the critic LLM, which can perform the following four operations:
   1. 1.
-    Pairwise Judge Operator: A function that compares two outcome $c_{1}$ and $c_{2}$, given a reflection $R$ and predicts which outcome will yield a higher true score:
-    $\textsc{PairwiseJudge}\big(c_{1},c_{2};R,M_{\text{critic}}\big)\in\{c_{1},c_{2}\}$ (7)
+    Pairwise Judge Operator. Given two prompts $\pi^{(1)}$ and $\pi^{(2)}$, and their corresponding outcomes $\Phi(\pi^{(1)})$ and $\Phi(\pi^{(2)})$, predict which is better given reflection $R$:
+    $\textsc{PairwiseJudge}(\Phi(\pi^{(1)}),\Phi(\pi^{(2)});R,M_{\text{critic}})\in\{\Phi(\pi^{(1)}),\Phi(\pi^{(2)})\}.$ (8)
   2. 2.
-    Reflection Operator: A function that compresses the entire history into a natural-language summary of rules or guidelines:
-    $R_{t}=\textsc{MetaReflect}(H_{t};M_{\text{critic}})$ (8)
+    Meta-reflection Operator. Compress the entire history into a small set of natural-language rules or guidelines:
+    $R_{t}=\textsc{MetaReflect}(H_{t};M_{\text{critic}}).$ (9)
   3. 3.
-    Textual Gradient Operator: A stochastic function that generates a textual gradient $\delta$ (e.g., “Make the lighting warmer”) intended to improve the prompt:
-    $\delta\sim\nabla_{\text{text}}(\Pi;R)$ (9)
+    Textual Gradient Operator. Stochastically generate a local edit $\delta$ intended to improve $\pi$ under reflection $R$:
+    $\delta\sim\nabla_{\text{text}}(\pi;R,M_{\text{critic}}).$ (10)
   4. 4.
-    Rewrite Operator: A deterministic function that maps a prompt $\Pi$ and an edit $\delta$ to a new prompt $\Pi^{\prime}$:
-    $\Pi^{\prime}=\operatorname{Apply}(\Pi,\delta;M_{\text{critic}})$ (10)
-In the digital-marketing example, these correspond to an LLM’s ability to (1) compare two ad creatives and predict which will perform better (calderon2025alternative), (2) examine past performance data and derive guidelines for effective ads in context (jiang2024many; ferber2024context), (3) propose prompt improvements based on those guidelines (textgrad; li2025test), and (4) apply prompt improvements to get the new prompt.
+    Apply Operator. Apply an edit $\delta$ to obtain a new prompt:
+    $\pi^{\prime}=\operatorname{Apply}(\pi,\delta;M_{\text{critic}}).$ (11)
+These four operators map directly onto well-documented LLM capabilities in the prompt-optimization and evaluative-feedback literature.
+The PairwiseJudge operator aligns with the growing use of LLMs as reliable comparison judges for creative and reasoning tasks (shi2024judging; calderon2025alternative).
+The MetaReflect operator captures reflection-based prompt-improvement methods that distill past outcomes into actionable guidelines (e.g., in the ad optimization case, it could suggest that “social settings outperform isolated product shots") (jiang2024many; ferber2024context; wang2025maestro).
+The TextualGradient operator corresponds to LLM-driven proposals for targeted, locally improving edits (e.g., “make the lighting warmer and emphasize people enjoying the product”), as formalized in textgrad and applied across domains such as scientific reasoning and code refinement (li2025test).
+Finally, the Apply operator reflects standard LLM-based prompt editing widely used in prompt engineering (zhou2022large; khattab2023dspy).
+Together, these capabilities make an LLM critic a natural, practically grounded component for enabling BO-style search over prompt space.
 5.2 Best-of-N gradient
-The notion of a Best-of-$N$ gradient builds on the widely used Best-of-$N$ strategy for test-time alignment, where an LLM trades additional inference-time (test-time) compute for higher-quality outputs by sampling multiple candidate generations and then selecting the one that appears most aligned with a downstream objective (snell2024scaling; beirami2024theoretical).
-In the standard Best-of-$N$ strategy for prompting, an LLM generates $N$ candidate prompts, a generation model produces $N$ candidate outcomes, and a reward model selects the highest-scoring one.
-While effective for choosing a single output, our goal is to improve the prompt itself to ensure sustained performance improvement.
-To achieve this, we introduce the concept of the Best-of-N Gradient.
-Instead of sampling $N$ candidate prompts directly, we sample $N$ distinct textual gradients (directions of improvement).
-By applying these $N$ textual edits to the current prompt and selecting the LLM-as-a-judge-predicted most promising candidate solution (Best-of-N), we aim to identify the textual gradient direction for the implicit acquisition function.
-The procedure operates as follows:
+The notion of a Best-of-N gradient builds on the widely used Best-of-N strategy for test-time alignment, where an LLM trades additional inference-time compute for higher-quality outputs by sampling multiple candidates and then selecting the one that appears most aligned with a downstream objective (snell2024scaling; beirami2024theoretical).
+In standard Best-of-N prompting, an LLM generates $N$ candidate prompts, a generation model produces $N$ candidate outcomes, and a reward model selects the highest-scoring one.
+While effective for choosing a single output, our goal here is to improve the prompt itself.
+To accomplish this, we introduce the concept of Best-of-N gradient.
+Rather than sampling $N$ full prompts, we sample $N$ distinct textual gradients, apply them to the current prompt, and then select the candidate whose predicted outcome is judged most promising.
+This selection defines a localized “direction of improvement” in prompt space, and can be viewed as an acquisition function.
+Formally, given a current prompt $\pi$ and reflection $R$, a single Best-of-N gradient step proceeds as follows:
   1. 1)
-    Textual gradient generation: Using $M_{\text{critic}}$, query the textual-gradient operator $N$ independent times to obtain a set of small edits:
-    $\delta^{(i)}\sim\nabla_{\text{text}}(\Pi;R,M_{\text{critic}}),\quad i=1,\dots,N.$ (11)
+    Textual gradient generation. Use $M_{\text{critic}}$ to generate $N$ independent textual edits:
+    $\delta^{(i)}\sim\nabla_{\text{text}}(\pi;R,M_{\text{critic}}),\qquad i=1,\dots,N.$ (12)
   2. 2)
-    Candidate prompt generation: Apply each edit to the current prompt to generate $N$ candidate prompts:
-    $\Pi^{(i)}=\operatorname{Apply}(\Pi,\delta^{(i)};M_{\text{critic}}),\quad i=1,\dots,N.$ (12)
+    Candidate prompt generation. Apply each edit to produce $N$ candidate prompts:
+    $\pi^{(i)}=\operatorname{Apply}(\pi,\delta^{(i)};M_{\text{critic}}).$ (13)
   3. 3)
-    Candidate system outcome generation: Generate the corresponding system outcomes for each candidate prompt:
-    $c^{(i)}=\Phi(\Pi^{(i)}),\quad i=1,\dots,N.$ (13)
+    Candidate system outcome generation. Produce the associated system outcomes:
+    $\Phi(\pi^{(i)}),\qquad i=1,\dots,N.$ (14)
   4. 4)
-    Best-of-N selection: Select the best candidate by iteratively applyng PairwiseJudge operation using $M_{\text{critic}}$ as LLM-as-a-judge.
-We define this selection as Best-of-N operation, which induces the index $i^{\star}$ as the best one among N:
-    $i^{\star}=\textsc{Best-of-N}\big(\{(\Pi^{(i)},c^{(i)})\}_{i=1}^{N};R,M_{\text{critic}}\big).$ (14)
-    The resulting prompt $\Pi^{(i^{\star})}$ represents the update step taken by the Best-of-N Gradient.
-For the Best-of-N selection, we employ the pairwise tournament (liu2025pairjudgermperformbestofn).
-First, the $N$ candidates are permuted to produce randomly chosen tournament matches.
-Each match compares two outcomes by querying $M_{\text{critic}}$ for a head-to-head judgment (For the prompt, please refer to Figure 11 in Appendix A).
-illustrates the prompt example for ad generation tasks.
-To mitigate noise, the comparison is repeated $K$ times (typically an even number to evenly swap orders to remove positional bias (shi2024judging), and ties are broken by random.
-Winners advance round by round until only the final winner remains.
-Given the final winner’s index $i^{\star}$, we define $(\Pi^{j}_{t,g},c^{j}_{t,g}):=(\Pi^{j,(i^{\star})}_{t,g},c^{j,(i^{\star})}_{t,g})$ and carry it forward for the next gradient step $g+1$.
+    Best-of-N selection. Identify the most promising candidate using the critic:
+    $i^{\star}=\textsc{Best-of-N}\big(\{(\pi^{(i)},\Phi(\pi^{(i)})\}_{i=1}^{N};R,M_{\text{critic}}\big).$ (15)
+    For the Best-of-N selection, we employ a pairwise tournament (liu2025pairjudgermperformbestofn).
+The $N$ candidates are first randomly permuted into tournament buckets.
+Each match compares two outcomes using the PairwiseJudge operator, with the comparison repeated $K$ times and the order of presentation swapped to mitigate positional bias (shi2024judging) (see the prompt template in Figure A3 in Web Appendix C.3 for the ad-optimization case).
+Ties are broken at random.
+Winners advance round by round until a single final winner remains; its associated prompt and outcome are then carried forward to the next gradient step.
+The resulting prompt $\pi^{(i^{\star})}$ defines the update made by the Best-of-N gradient for a given step.
+Intuitively, each textual edit represents a noisy proposal for a local improvement direction; Best-of-N filters these proposals through an optimism-based rule, selecting the direction most likely to improve performance under the critic LLM’s current understanding.
 5.3 Theory: Best-of-N gradient’s Asymptotic Equivalence to UCB gradient
-In this section, we prove that the Best-of-$N$ over textual gradients selects, in probability, the ascent direction of the UCB acquisition with exploration weight $\beta$ scaling as $\Theta(\sqrt{\ln N})$ (Theorem 2).
+We now prove that the Best-of-N over textual gradients selects, in probability, the ascent direction of the UCB acquisition with exploration weight $\beta$ scaling as $\Theta(\sqrt{\ln N})$ (Theorem 1).
 What allows us to do rigorous theoretical analysis is that LLM prompt embeddings are invertible and bi-Lipschitz: LLM’s prompt embeddings are known to be injective, and therefore invertible (nikolaou2025language); LLM embeddings are Lipschitz (tang2025understanding), implying they are bi-Lipschitz.
 Hence, small textual edits can be represented as small changes in the embedding space and vice versa.
-This enables us to map changes in prompts to changes in embeddings in Euclidean space, and later bring the theoretical guarantees back in the prompt space.
-Intuitive explanation.
-Denote $x$ to be an embedding of a prompt. Pick a small radius $\varepsilon>0$ around $x$ and sample $N$ nearby edits $x_{i}=x+\varepsilon u_{i}$ with $u_{i}$ on the unit sphere. Score each point
-$Y_{i}=\mu(x_{i})+\xi_{i}\,\sigma(x_{i}),\quad\xi_{i}\text{ i.i.d. sub-Gaussian and independent of }\{u_{i}\}.$ (15)
-Let $M_{N}=\max_{i}\xi_{i}$ and $q_{N}=F^{-1}(1-1/N)$. Since $M_{N}\simeq q_{N}$ and $q_{N}=\Theta(\sqrt{\ln N})$, Best-of-$N$ acts like picking the sampled $x_{i}$ with the largest
-$\mu(x_{i})+q_{N}\,\sigma(x_{i}).$ (16)
-On the small sphere around $x$ the function $A_{\beta}(x^{\prime})=\mu(x^{\prime})+\beta\,\sigma(x^{\prime})$ has a single best direction.
-Call it $v_{\beta}$.
+Consequently, small textual edits correspond to small perturbations in the embedding space, allowing us to analyze Best-of-N gradient steps in $\mathbb{R}^{d}$ and then translate the theoretical guarantees back to the prompt space.
+5.3.1 Intuitive Explanation
+Denote $e$ to be an embedding of a prompt.
+Pick a small radius $\varepsilon>0$ around $e$ and sample $N$ nearby edits $e_{i}=e+\varepsilon u_{i}$ with $u_{i}$ on the unit sphere.
+Next, denote the critic LLM’s ($M_{critic}$ from $\S$5.1) internal scoring of each edit, $Y_{i}$, as follows:
+$Y_{i}=\mu(e_{i})+\xi_{i}\,\sigma(e_{i}),$ (16)
+where $\mu$ is the (implicit) posterior mean, $\sigma$ is the (implicit) epistemic uncertainty, and $\xi_{i}$ is an i.i.d.
+sub-Gaussian noise variable independent of $u_{i}$.
+We note that $Y_{i}$ is in line with the BO scoring function from $\S$4.
+Let $M_{N}=\max_{i}\xi_{i}$ and let $q_{N}=F^{-1}(1-1/N)$ be the (11/N)(1-1/N)-quantile of the noise distribution.
+For Gaussian noise, classical extreme-value theory implies that $M_{N}$ concentrates around $q_{N}$ and that $q_{N}=\Theta(\sqrt{\ln N})$.
+In other words, as $N$ grows large, the largest noise term among $\{\xi_{i}\}_{i=1}^{N}$ behaves like $q_{N}$ up to lower-order fluctuations.
+Consequently, the index $i^{\star}$ selected by Best-of-N approximately maximizes
+$\mu(e_{i})+q_{N}\,\sigma(e_{i}).$ (17)
+Consequently, for large $N$, selecting the Best-of-N index $i^{\star}$ is asymptotically equivalent to selecting the maximizer of UCB-style scores with exploration weight $q_{N}$.
+Now consider a small sphere around $e$ in embedding space and the UCB acquisition $A_{\beta}(e^{\prime})=\mu(e^{\prime})+\beta\,\sigma(e^{\prime})$.
+On this sphere, there is a single dominant ascent direction for $A_{\beta}$, which we denote by $v_{\beta}$ (the direction in which $A_{\beta}$ increases the fastest to first order).
 Draw a narrow cone around $v_{\beta}$.
-Then, with a large enough $N$, at least one $u_{i}$ lands in that cone with high probability.
-The winner comes from that cone.
-Set $\beta=q_{N}$.
-Then Best-of-$N$ behaves like UCB with weight $\beta$: it picks the sampled point most similar to the $A_{\beta}$ best direction.
-Rigorous derivation.
-Let $\mathcal{M}$ be the set of syntactically valid prompts.
-Suppose that there exists a text embedding $\mathsf{E}:\mathcal{M}\to\mathbb{R}^{d}$, where $x\;:=\;\mathsf{E}(\Pi)\in\mathbb{R}^{d}$ is the embedding of a prompt $\Pi\in\mathcal{M}$.
+For sufficiently large $N$, at least one of the sampled directions $u_{i}$ falls inside this cone with high probability, and the Best-of-N winner $e_{i^{\star}}$ comes from this cone.
+Thus, among the locally sampled perturbations $e_{i}=e+\varepsilon u_{i}$, the Best-of-N rule selects a candidate whose direction is (with high probability, for large enough $N$) very close to the ascent direction of the UCB acquisition function with $\beta_{N}:=q_{N}$.
+5.3.2 Rigorous Derivation
+Let $\Pi$ be the set of syntactically valid prompts.
+Suppose that there exists a text embedding $\mathsf{E}:\Pi\to\mathbb{R}^{d}$, where $e\;:=\;\mathsf{E}(\pi)\in\mathbb{R}^{d}$ is the embedding of a prompt $\pi\in\Pi$.
 We first state the assumptions motivated by how language model embeddings behave in practice, and then state the main theorem, Theorem 1.
-At each gradient step of T-BoN BO, we sample edits that correspond to $u_{1},\ldots,u_{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\rho$ on the embedding space $\subset\mathbb{R}^{d}$, where $\rho$ is assumed to have a density bounded below by a positive constant with respect to surface measure.
+At each gradient step of TextBO, we sample edits that correspond to $u_{1},\ldots,u_{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\rho$ on the embedding space $\subset\mathbb{R}^{d}$, where $\rho$ is assumed to have a density bounded below by a positive constant with respect to surface measure.
 Assumption 1 implies that small textual edits induce near‑linear moves in $\mathbb{R}^{d}$, consistent with the recent findings that prompt embeddings preserve phrase similarity under small changes (reimers2019sentence; he2025position; tang2025understanding; nikolaou2025language).
 Assumption 1 (Embedding map and edit realization).
-For the current prompt $\Pi$ with $x=\mathsf{E}(\Pi)$, constants $C<\infty$ and $\varepsilon_{0}>0$ such that for every unit vector $u\in\mathbb{R}^{d}$ and every $\varepsilon\in(0,\varepsilon_{0}]$ the edit operator $\mathrm{Apply}$ realizes a first–order move:
-$\mathsf{E}\!\big(\mathrm{Apply}(\Pi,\varepsilon u)\big)=x+\varepsilon u+r(\varepsilon,u),\qquad\sup_{\|u\|=1}\|r(\varepsilon,u)\|\leq C\,\varepsilon^{2}.$ (17)
-Next, the assumption 2 is a standard regularity/technical condition: let’s call a function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $C^{1,1}$ if it is continuously differentiable and its gradient is Lipschitz: $\exists L<\infty$ such that $\|\nabla f(y)-\nabla f(z)\|\leq L\|y-z\|$ for all $y,z$.
+For the current prompt $\pi$ with $e=\mathsf{E}(\pi)$, constants $C<\infty$ and $\varepsilon_{0}>0$ such that for every unit vector $u\in\mathbb{R}^{d}$ and every $\varepsilon\in(0,\varepsilon_{0}]$ the edit operator $\mathrm{Apply}$ realizes a first–order move:
+$\mathsf{E}\!\big(\mathrm{Apply}(\pi,\varepsilon u)\big)=e+\varepsilon u+r(\varepsilon,u),\qquad\sup_{\|u\|=1}\|r(\varepsilon,u)\|\leq C\,\varepsilon^{2}.$ (18)
+Next, Assumption 2 is a standard regularity condition: function $f:\mathbb{R}^{d}\to\mathbb{R}$ is of $C^{1,1}$ if it is continuously differentiable and its gradient is Lipschitz: $\exists L<\infty$ such that $\|\nabla f(y)-\nabla f(z)\|\leq L\|y-z\|$ for all $y,z$.
 Assumption 2 ($\mu,\sigma$ are of $C^{1,1}$).
-Let $\mu(x)$ and $\sigma(x)$ denote the mean and standard deviation of the evaluation variable $Y$ at embedding $x$.
+Let $\mu(e)$ and $\sigma(e)$ denote the mean and standard deviation of the critic LLM’s internal score of embedding $e$.
 We assume $\mu,\sigma$ are $C^{1,1}$ locally.
-We use $g=\nabla\mu(x)$ and $h=\nabla\sigma(x)$.
-Assumption 3 specifies a heteroscedastic sub-Gaussian model for candidate scores (with i.i.d.
+We use $g=\nabla\mu(e)$ and $h=\nabla\sigma(e)$.
+Assumption 3 specifies a heteroskedastic sub-Gaussian model for candidate scores (with i.i.d.
 mean-zero, unit-variance, unbounded-support noise independent of the sampled directions) and an ideal Best-of-N choice oracle.
 Assumption 3 (Noise model and choice oracle).
-Let $\{\xi_{i}\}_{i=1}^{N}$ be i.i.d. mean-zero, unit-variance, Gaussian random variables with unbounded support, independent of $\{u_{i}\}$. For candidates $x_{i}:=x+\varepsilon u_{i}$ we observe
-$Y_{i}\;=\;\mu(x_{i})\;+\;\sigma(x_{i})\,\xi_{i},$ (18)
-and the Best-of-$N$ oracle returns $i^{\star}\in\arg\max_{i\in[N]}Y_{i}$.
+Let $\{\xi_{i}\}_{i=1}^{N}$ be i.i.d.
+mean-zero, unit-variance, Gaussian random variables with unbounded support, independent of $\{u_{i}\}$.
+For candidates $e_{i}:=x+\varepsilon u_{i}$, the critic LLM observes its internal scoring
+$Y_{i}\;=\;\mu(e_{i})\;+\;\sigma(e_{i})\,\xi_{i},$ (19)
+and the Best-of-N selection oracle (often built using pairwise tournaments in practice; see $\S$5.2) returns corresponding $i^{\star}\in\arg\max_{i\in[N]}Y_{i}$.
 Assumption 4 (Small-step regime and limit order).
@@
 Under Assumptions 1-4, define $\beta_{N}:=q_{N}$ from Lemma 1. Let $i^{\star}\in\arg\max_{i}Y_{i}$ and set $\widehat{u}_{N}:=u_{i^{\star}}$. Define the step
-$\Delta x^{(N)}\;:=\;\mathsf{E}\!\big(\mathrm{Apply}(\Pi,\varepsilon\widehat{u}_{N})\big)-x\;=\;\varepsilon\widehat{u}_{N}+r(\varepsilon,\widehat{u}_{N}).$ (19)
+$\Delta e^{(N)}\;:=\;\mathsf{E}\!\big(\mathrm{Apply}(\pi,\varepsilon\widehat{u}_{N})\big)-e\;=\;\varepsilon\widehat{u}_{N}+r(\varepsilon,\widehat{u}_{N}).$ (20)
 Then, with the limit order $N\to\infty$ first and $\varepsilon\downarrow 0$ second,
-$\frac{\Delta x^{(N)}}{\|\Delta x^{(N)}\|}\;=\;\widehat{u}_{N}\;\xrightarrow{p}\;\frac{\nabla A_{\beta_{N}}(x)}{\|\nabla A_{\beta_{N}}(x)\|}\;=\;\frac{g+\beta_{N}h}{\|g+\beta_{N}h\|}.$ (20)
+$\frac{\Delta e^{(N)}}{\|\Delta e^{(N)}\|}\;=\;\widehat{u}_{N}\;\xrightarrow{p}\;\frac{\nabla A_{\beta_{N}}(e)}{\|\nabla A_{\beta_{N}}(e)\|}\;=\;\frac{g+\beta_{N}h}{\|g+\beta_{N}h\|}.$ (21)
 Moreover,
-$A_{\beta_{N}}(x+\Delta x^{(N)})\;\geq\;A_{\beta_{N}}(x)\;+\;\varepsilon\,\|\nabla A_{\beta_{N}}(x)\|\;-\;o_{p}(\varepsilon).$ (21)
-5 TextGrad-Best-of-N Bayesian Optimization (T-BoN BO)
-We now describe the self-improving AI framework proposed in this paper: TextGrad-Best-of-N Bayesian Optimization (T-BoN BO).
-As discussed in Section 1, T-BoN BO is designed to implement parallel gradient-based UCB-BO in the discrete space of natural language (Figure 3).
-Specifically, T-BoN BO implements the parallel gradient-based UCB-BO that reuses the final gradient point as the initialization for the next iteration.
-Therefore, each initial point gives rise to a trajectory: a sequence of candidate points obtained by repeatedly taking gradient-ascent steps and evaluations starting from that initialization.
+$A_{\beta_{N}}(e+\Delta e^{(N)})\;\geq\;A_{\beta_{N}}(e)\;+\;\varepsilon\,\|\nabla A_{\beta_{N}}(e)\|\;-\;o_{p}(\varepsilon).$ (22)
+Thus, each Best-of-N gradient step asymptotically performs a first-order improvement on an implicit UCB acquisition function, with exploration strength $\beta_{N}$ controlled entirely by $N$.
+The full proof is given in Web Appendix B.
+Theorem 1 provides the central conceptual and technical contribution of this paper: it shows that a procedure that looks like a purely heuristic rule in language space implicitly implements the same optimism-driven search direction as UCB BO in the model’s embedding space.
+This is novel and important because it removes the main obstacle to bringing UCB BO’s evaluation-efficiency guarantees into prompt optimization: we never construct an explicit surrogate model, never compute $\sigma(\cdot)$, and never differentiate a closed-form acquisition function, yet the effective update direction converges to the UCB gradient direction.
+Moreover, the result establishes a new use case for the ubiquitous Best-of-N principle as an interpretable exploration mechanism, where $N$ becomes a principled “exploration knob” for the exploration weight $\beta_{N}=\Theta(\sqrt{\ln N})$.
+In short, Theorem 1 is what turns Best-of-N gradient from an empirical recipe into a principled BO-style optimizer: it justifies viewing our language-space self-improvement loop as parallel gradient-based UCB-BO and is the reason TextBO can inherit BO’s evaluation-efficiency benefits.
+6 TextBO Algorithm
+We now describe the self-improving AI algorithm proposed in this paper: TextBO.
+TextBO is designed to implement parallel gradient-based UCB-BO in the space of natural language prompts.
+Concretely, TextBO adopts the parallel gradient-based UCB-BO scheme that reuses the final gradient point from one iteration as the initialization for the next.
+Thus, each initial prompt gives rise to a trajectory: a sequence of candidate prompts and outcomes generated by repeatedly taking gradient-ascent steps and then evaluating the resulting system outcome.
+The overview of the TextBO algorithm is shown in Figure 3.
+For each iteration $t\in[1,T]$ and each trajectory $j\in[1,J]$, we evaluate one prompt; to determine the one prompt to evaluate, we take $G$ steps of Best-of-N gradients beforehand, without any evaluation.
+We present the pseudocode of the TextBO algorithm (Algorithm 2) and provide the line-by-line description of the details of the procedure below.
 […]
-Figure 4: Visual summary of T-BoN BO (Algorithm 1) with $T$ iterations, $J$ trajectories, and $G$ Best-of-NN gradients per each iteration and trajectory.
-Algorithm.
-Throughout $T$ number of optimization iterations, T-BoN BO runs $J$ trajectories in parallel, starting from a set of initial prompts, denoted $\{\Pi_{0}^{j}\}_{j=1}^{J}$.
-At each iteration and a trajectory, we fix the hyperparameters: the number of candidates per gradient step ($N$) and the number of gradient steps per iteration ($G$).
-At iteration $t=0$, it starts from evaluating $\{\Pi_{0}^{j}\}_{j=1}^{J}$, i.e., the initial systems $\{c_{0}^{j}=\Phi(\Pi_{0}^{j})\}_{j=1}^{J}$, on $\{\mathcal{D}_{\mathcal{X}}(\Pi_{0}^{j})\}_{j=1}^{J}$ to obtain initial scores $\{s_{0}^{j}\}_{j=1}^{J}$.
-The history is then initialized as $H_{0}=\bigcup_{j=1}^{J}\{(\Pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$.
-The shared reflection is initialized as $R_{0}=\emptyset$.
-At each iteration $t$, for each trajectory $j\in\{1,\ldots,J\}$, we set $\Pi_{t,0}^{j}$ as $\Pi_{t-1}^{j}$ and proceed as follows:
+Figure 3: A pictorial overview of TextBO (Algorithm 2) with $T$ iterations, $J$ trajectories, and $G$ Best-of-N gradient steps per iteration and trajectory.
+Input: System $\Phi$; $\{\pi_{0}^{j}\}_{j=1}^{J}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$; $J$ trajectories Output: Optimized triples $\{(\pi^{j}_{T},c^{j}_{T},s^{j}_{T})\}_{j=1}^{J}$ 1 for $j=1$ to $J$ do 2 $c_{0}^{j}\leftarrow\Phi(\pi_{0}^{j})$ 3 $s_{0}^{j}\leftarrow\text{Eval}(c_{0}^{j},\mathcal{D}_{\mathcal{X}}(\pi_{0}))$ 4 $H_{0}\leftarrow\{(\pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$, $R\leftarrow\emptyset$ 5  6for $t=1$ to $T$ do 7 for $j=1$ to $J$ do 8    $\pi^{j}_{t,0}\leftarrow\pi^{j}_{t-1}$ 9    for $g=1$ to $G$ do 10       for $i=1$ to $N$ do 11          $\delta_{t,g}^{j,(i)}\leftarrow\nabla_{\text{text}}(\pi_{t,g-1}^{j};R,M_{\text{critic}})$ 12          $\pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}(\pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)})$ 13          $c_{t,g}^{j,(i)}\leftarrow\Phi(\pi_{t,g}^{j,(i)})$ 14           15       $(\pi_{t,g}^{j},c_{t,g}^{j})\leftarrow\textsc{Best-of-N}(\{(\pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\},R_{t-1},M_{\text{critic}})$ 16        17    $c_{t}^{j}\leftarrow c^{j}_{t,G}$, $\pi_{t}^{j}\leftarrow\pi_{t,G}^{j}$ 18    $s_{t}^{j}\leftarrow\text{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\pi_{t}))$ 19    $H_{t}\leftarrow H_{t-1}\cup\{(\pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}$ 20    if $s_{t}^{j}<s_{t-1}^{j}$ then 21       $\pi_{t}^{j}\leftarrow\pi_{t-1}^{j}$, $c_{t}^{j}\leftarrow c_{t-1}^{j}$, $s_{t}^{j}\leftarrow s_{t-1}^{j}$ 22        23 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$ 24  return $\{c_{t}^{j_{t}^{*}}\}_{t=1}^{T}$, where $j_{t}^{*}:=\operatorname{argmax}_{j}s_{t}^{j}$ Algorithm 2 TextBO
+Initialization ($t=0$).
+At iteration $t=0$, we first evaluate all initial prompts.
+For each trajectory $j=1,\dots,J$, we first generate the initial system outcome $\Phi(\pi_{0}^{j})$.
+Next, we evaluate it on the corresponding input distribution $\mathcal{D}_{\mathcal{X}}(\pi_{0}^{j})$ to obtain a scalar score $s_{0}^{j}$.
+The shared history is then initialized as $H_{0}\;=\;\bigcup_{j=1}^{J}\{(\pi_{0}^{j},\Phi(\pi_{0}^{j}),s_{0}^{j})\}$, and and the shared reflection is initialized as $R_{0}=\emptyset$.
+Per-iteration update ($t\geq 1$).
+At each optimization iteration $t\in\{1,\dots,T\}$, the algorithm updates all trajectories in parallel.
+For each trajectory $j\in\{1,\dots,J\}$, we set the starting prompt for this iteration to be the final prompt from the previous iteration, i.e., $\pi_{t,0}^{j}\leftarrow\pi_{t-1}^{j}$.
+We then perform the following three steps.
   • 1.
-    Best-of-N gradients. (Detailed in Section 4.2) For each trajectory $j$, repeat steps 1) – 4) for $g=1,\ldots,G$:
+    Best-of-N gradients phase.
+(Detailed in $\S$5.2.) For each trajectory $j$, we take $G$ successive Best-of-N textual-gradient steps, which do not involve any evaluation.
+That is, for $j=1,\ldots J$, for $g=1,\dots,G$:
     1. 1)
-      Textual gradient generation: From the current prompt $\Pi_{t,g-1}^{j}$, query the textual-gradient generation operator $\nabla_{\text{text}}(\Pi_{t,g-1}^{j};R_{t-1},M_{\text{critic}})$ independently $N$ times to obtain textual gradients $\{\delta^{j,(i)}_{t,g}\}_{i=1}^{N}$.
+      Textual gradient generation. From the current prompt $\pi_{t,g-1}^{j}$, query the textual-gradient operator
+      $\delta_{t,g}^{j,(i)}\sim\nabla_{\text{text}}(\pi_{t,g-1}^{j};R_{t-1},M_{\text{critic}})$
+      independently $N$ times to obtain textual gradients $\{\delta_{t,g}^{j,(i)}\}_{i=1}^{N}$.
     2. 2)
-      Candidate prompt generation: Apply each textual gradient $\delta_{t,g}^{j,(i)}$ to $\Pi_{t,g-1}^{j}$ for all $i$ to get $\{\Pi_{t,g}^{j,(i)}\}_{i=1}^{N}$.
+      Candidate prompt generation. Apply each textual gradient to the current prompt $\pi_{t,g-1}^{j}$ to obtain candidate prompts $\{\pi_{t,g}^{j,(i)}\}_{i=1}^{N}$:
+      $\pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}(\pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)}),\qquad i=1,\dots,N.$
     3. 3)
-      Candidate system outcome generation: Produce the system outcomes $\{c_{t,g}^{j,(i)}:=\Phi(\Pi_{t,g}^{j,(i)})\}_{i=1}^{N}$.
+      Candidate system outcome generation. For each candidate prompt $\pi_{t,g}^{j,(i)}$, generate the corresponding system outcome:
+      $\Phi(\pi_{t,g}^{j,(i)}),\qquad i=1,\dots,N.$
     4. 4)
-      Best-of-N selection: Find the best of $\{(\Pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\}_{i=1}^{N}$ using the critic and the shared reflection $R_{t-1}$, and set the winner as $(\Pi_{t,g}^{j},c_{t,g}^{j})$.
+      Best-of-N selection.
+Using the critic model and the shared reflection $R_{t-1}$, select the candidate that is predicted to be the most promising by the critic model among $\{(\pi_{t,g}^{j,(i)},\Phi(\pi_{t,g}^{j,(i)})\}_{i=1}^{N}$, and set the winner as $(\pi_{t,g}^{j},\Phi(\pi_{t,g}^{j}))$.
+    After $G$ gradient steps, trajectory $j$ has an updated prompt-outcome pair $(\pi_{t,G}^{j},\Phi(\pi_{t,G}^{j}))$ for this iteration.
+    Geometrically, as illustrated in Figure 4, at iteration $t$ the prompts $\{\pi_{t-1}^{j}\}_{j=1}^{J}$ (equivalently, the initial points $\{\pi_{t,0}^{j}\}_{j=1}^{J}$) act as starting locations in the implicit UCB landscape over prompt space.
+For each trajectory $j$, the $G$ Best-of-N gradient steps climb this landscape, producing the updated prompts $\{\pi_{t,G}^{j}\}_{j=1}^{J}$ that approximate local maxima of the UCB acquisition function.
+[^4] For $\{\pi_{t,G}^{j}\}_{j=1}^{J}$ to be considered local optima in the usual gradient-based BO sense, $G$ should be large enough.
+These local optima are then passed to the evaluation phase (Phase 2).
+    Figure 4: Illustration of how TextBO (Algorithm 2) implements Best-of-N gradients phase at iteration $t$.
+The prompts identified as local optima in this phase are evaluated in the later evaluation phase.
   • 2.
-    Evaluation. For each trajectory $j$,
+    Evaluation phase. For each trajectory $j$:
     1. 1)
-      Set $c_{t}^{j}=c_{t,G}^{j}$ and $\Pi_{t}^{j}=\Pi_{t,G}^{j}$.
+      Set the trajectory’s candidate for evaluation to the final output of the Best-of-N gradient phase:
+      $\pi_{t}^{j}:=\pi_{t,G}^{j}$
     2. 2)
-      Evaluate $c_{t}^{j}$ to get the score $s_{t}^{j}=\mathrm{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}^{j}))$, where
-      $\mathrm{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}^{j}))\;:=\;\frac{1}{L}\sum_{l=1}^{L}r\!\left(y\left(c^{j}_{t},x_{t}^{j,(l)}\right)\right),\qquad x_{t}^{j,(l)}\sim\mathcal{D}_{\mathcal{X}}(\Pi^{j}_{t}),$ (22)
+      Evaluate $c_{t}^{j}$ to obtain the scalar score
+      $s_{t}^{j}$ $:=\mathrm{Eval}(\Phi(\pi_{t}^{j}),\mathcal{D}_{\mathcal{X}}(\pi_{t}^{j}))=\frac{1}{L}\sum_{l=1}^{L}r\!\left(y\!\left(\Phi(\pi_{t}^{j}),x_{t}^{j,(l)}\right)\right),\;\;\textrm{where}\;\;x_{t}^{j,(l)}\sim\mathcal{D}_{\mathcal{X}}(\pi_{t}^{j}).$ (23)
+      Note that this is the step where the evaluation happens; and in each iteration, we conduct $J$ evaluations – one for each trajectory.
     3. 3)
-      Given $\{(\Pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}_{j=1}^{J}$, append it to the shared history $H_{t-1}$ to form $H_{t}$.
+      Append the new triple $(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})$ to the shared history:
+      $H_{t}\leftarrow H_{t-1}\cup\{(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})\}.$
     4. 4)
-      If $s_{t}^{j}>s_{t-1}^{j}$, keep $(\Pi_{t}^{j},c_{t}^{j})$.
-Otherwise, revert trajectory $j$ to the previous state: $(\Pi_{t}^{j},c_{t}^{j})=(\Pi_{t-1}^{j},c_{t-1}^{j})$ and set $s_{t}^{j}=s_{t-1}^{j}$.
-    For 4), we follow textgrad and adopt a simple acceptance rule that controls progress: if $s_{t}\geq s_{t-1}$, the new candidate is adopted; otherwise, the system rolls back to the previous iteration’s outcome, i.e., $(\Pi_{t-1},c_{t-1},s_{t-1})$.
-After evaluation of all trajectories, $\{(\Pi^{j}_{t},c^{j}_{t},s^{j}_{t})\}_{j=1}^{J}$ are appended to the history $H_{t-1}$ to form $H_{t}$.
+      Acceptance rule. If the new score does not improve, revert this trajectory to its previous state:
+      $\text{if }s_{t}^{j}\leq s_{t-1}^{j}\text{ then }(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})\leftarrow(\pi_{t-1}^{j},\Phi(\pi_{t-1}^{j}),s_{t-1}^{j}).$
+    This acceptance rule follows textgrad: it enforces monotone non-decreasing scores along each trajectory by rolling back any non-improving update.
   • 3.
-    Shared Meta-reflection.
-    At the end of each iteration $t$, we update the global reflection
-    $R_{t}\leftarrow\textsc{MetaReflect}(H_{t};M_{\text{critic}})$
-    using the aggregated history from all trajectories for use at iteration t+1t{+}1.
-Here, $\textsc{MetaReflect}(H_{t};M_{\text{critic}})$ compresses the full history $H_{t}$ of prompts, generated system outcomes, and evaluation results into a natural language reflection $R_{t}$.
-$R_{t}$ consists of a small set of actionable rules.
-In the ad generation example, those rules may take the form “emphasize social context,” “avoid cluttered backgrounds,” or “prefer warm lighting.” This reflection serves as a global memory; in subsequent iterations, the critic LLM takes this reflection as a part of its context when generating textual gradients.
-(For the prompt used for meta-reflection, see Figure 12 in Appendix A).
-    There are many ways to construct a reflection from past results.
-In our experiments, following li2025test, we instruct the critic model $M_{\text{critic}}$ to compare the five lowest- and five highest-performing outcomes from previous iterations (or, when fewer than ten are available, all outcomes) [^4] Limiting the number of outcomes in the critic model’s context window helps mitigate context rotting (hong2025context)., identify recurring visual and textual patterns, and distill them into a concise set of actionable rules.
-After $T$ optimization steps, the sequence of best-performing outcomes $\{c_{t}^{j_{t}^{\ast}}\}_{t=1}^{T}$, where $j_{t}^{\ast}:=\operatorname{argmax}_{j}s_{t}^{j}$, is reported as the algorithm’s result.
-Note that T-BoN BO takes $N\times G$ generation of prompt/outcome candidates before each evaluation.
-This careful selection of a prompt/outcome candidate to evaluate is the key to evaluation efficiency in T-BoN BO: it deliberately trades inexpensive generation compute for fewer evaluations.
-Input: System $\Phi$; $\{\Pi_{0}^{j}\}_{j=1}^{J}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$; $J$ trajectories Output: Optimized triples $\{(\Pi^{j}_{T},c^{j}_{T},s^{j}_{T})\}_{j=1}^{J}$ 1 for $j=1$ to $J$ do 2 $c_{0}^{j}\leftarrow\Phi(\Pi_{0}^{j})$; 3 $s_{0}^{j}\leftarrow\text{Eval}(c_{0}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$; 4 $H_{0}\leftarrow\{(\Pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$, $R\leftarrow\emptyset$; 5  6 end for 7 8for $t=1$ to $T$ do 9 for $j=1$ to $J$ do 10    $\Pi^{j}_{t,0}\leftarrow\Pi^{j}_{t-1}$; 11    for $g=1$ to $G$ do 12       for $i=1$ to $N$ do 13          $\delta_{t,g}^{j,(i)}\leftarrow\nabla_{\text{text}}(\Pi_{t,g-1}^{j};R,M_{\text{critic}})$ ; 14          $\Pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}(\Pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)})$ ; 15          $c_{t,g}^{j,(i)}\leftarrow\Phi(\Pi_{t,g}^{j,(i)})$ ; 16           17         end for 18       $(\Pi_{t,g}^{j},c_{t,g}^{j})\leftarrow\textsc{Best-of-N}(\{(\Pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\},R_{t-1},M_{\text{critic}})$; 19        20      end for 21    $c_{t}^{j}\leftarrow c^{j}_{t,G}$, $\Pi_{t}^{j}\leftarrow\Pi_{t,G}^{j}$; 22    $s_{t}^{j}\leftarrow\text{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$; 23    $H_{t}\leftarrow H_{t-1}\cup\{(\Pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}$; 24    if $s_{t}^{j}<s_{t-1}^{j}$ then 25       $\Pi_{t}^{j}\leftarrow\Pi_{t-1}^{j}$, $c_{t}^{j}\leftarrow c_{t-1}^{j}$, $s_{t}^{j}\leftarrow s_{t-1}^{j}$; 26        27      end if 28     29   end for 30 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$; 31  32 end for return $\{c_{t}^{j_{t}^{*}}\}_{t=1}^{T}$, where $j_{t}^{*}:=\operatorname{argmax}_{j}s_{t}^{j}$ Algorithm 1 T-BoN BO
-6 Case Study: Ad Optimization in Digital Marketing
-[…]
-Figure 5: An ad generation AI system $\Phi$ and prompt $\Pi=\langle\pi\rangle$.
-[…]
-Figure 6: A self-improving AI implemented on the ad generation AI system in Figure 5.
-In this section, we present a case study on how ad platforms can use T-BoN BO for ad optimization and automate the ad generation, analysis, and evaluation cycle.
-Given an a priori fine-tuned image generation AI model that is aligned with its client’s brand, the objective is to efficiently refine the prompts for ad generation to better align with the users’ preference distribution.
-6.1 Problem definition
-Formally, define the brand fine-tuned ad-generation AI system as $\Phi$, whose behavior is controlled by a prompt $\Pi$ (Figure 5).
-For a given prompt $\Pi$, we write $\Phi(\Pi)$ for the corresponding ad creative shown on the ad platform.
-As a system, $\Phi(\Pi)$ maps the input space $\mathcal{X}$ (potential ad viewers on the platform) to a distribution over the output space $\mathcal{Y}$ (observable ad responses).
-For each $x\in\mathcal{X}$, we denote by $y(\Phi(\Pi),x)$ the random $\mathcal{Y}$-valued output of the system at input $x$.
-In a digital marketing setting, $x$ is a platform user, and a realization of $y(\Phi(\Pi),x)$ represents that user’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\Pi)$.
-When a user $x\in\mathcal{X}$ is shown the ad creative $\Phi(\Pi)$, we observe an outcome $y\in\mathcal{Y}$, which is a realization of $y(\Phi(\Pi),x)$.
-As discussed earlier, the distribution of exposed users depends on the ad $\Phi(\Pi)$ and therefore on the prompt $\Pi$.
-We denote by $\mathcal{D}_{\mathcal{X}}(\Pi)$ the unknown distribution of users $x$ who are shown the ad generated from $\Pi$ when the ad $\Phi(\Pi)$ is run on the platform.
-To evaluate system performance, a marketer pre-defines a scoring rule $r:\mathcal{Y}\to[0,1]$ that maps observed user responses (clicks, conversions, spend) to a normalized effectiveness score.
-The ad optimization objective of the agency for a given client and campaign can then be written as
-$J(\Pi)\;:=\;\mathbb{E}\big[\,r\big(y(\Phi(\Pi),x)\big)\big],$ (23)
-where the expectation is taken over $x\sim\mathcal{D}_{\mathcal{X}}(\Pi)$ and any additional randomness in $y(\Phi(\Pi),x)$.
-The self–improving ad system built by the agency can therefore be viewed as an iterative algorithm that searches over prompts $\pi$ in order to maximize $J(\pi)$ while minimizing the number of costly ad evaluations on the platform (Figure 6).
-At iteration $t$, a candidate prompt $\pi_{t}$ is selected, the corresponding ad creative $\Phi(\pi_{t})$ is generated, and the campaign is (conceptually) exposed to a batch of users $\{x_{t}^{(l)}\}_{l=1}^{L}\sim\mathcal{D}_{\mathcal{X}}(\pi_{t})$.
-The empirical evaluation
-$s_{t}\;:=\;\frac{1}{L}\sum_{l=1}^{L}r\big(y(\Phi(\pi_{t}),x_{t}^{(l)})\big)$ (24)
-serves as a noisy estimate of $J(\pi_{t})$ and is fed back into the optimization algorithm (T-BoN BO) to update the history $H_{t}$ and the meta–reflection $R_{t}$.
-6.2 Application of T-BoN BO
-We now describe how the general T-BoN BO framework in Section 5 is instantiated for ad agency’s ad optimization problem.
-Specifically, we focus on the parallel version of T-BoN BO (Algorithm 1), where $J$ optimization trajectories start from different initial prompts and run simultaneously.
-Initialization.
-The decision-maker/platform starts from initial prompts $\{\pi_{0}^{j}\}_{j=1}^{J}$, i.e., prompts already in use or high-quality baselines provided by humans.
-(e.g., “A simple, photorealistic image of a plant-based burger patty.
-[^5] Note that in the experiments in Section 7, we use Worst5-of-N prompts as initial starting points, which are higher-quality, much more detailed initial prompts; See Appendix A.) For each initial prompt $\pi_{0}^{j}$, the system generates an ad $\Phi(\pi_{0}^{j})$ and obtains an initial score $s_{0}^{j}$ by deploying the ad and observing outcomes.
-These tuples $\{(\pi_{0}^{j},\Phi(\pi_{0}^{j}),s_{0}^{j})\}_{j=1}^{J}$ form the initial history $H_{0}$ for the parallel T-BoN BO procedure.
-Best-of-NN textual-gradient steps.
-Within each optimization iteration $t$, T-BoN BO refines each trajectory’s prompt $\pi_{t-1}^{j}$ through $G$ Best-of-N textual-gradient steps (Figure 4).
-For a given trajectory $j$ and inner step $g$, the critic model $M_{\text{critic}}$ (a large language model such as Gemini 2.5 Flash) receives as context: (i) the current prompt $\pi_{t,g-1}^{j}$, (ii) the current meta-reflection $R_{t-1}$, and (iii) a short description of the brand and campaign scenario.
-Conditioned on this context, it proposes $N$ small natural-language edits
-$\delta_{t,g}^{j,(1)},\ldots,\delta_{t,g}^{j,(N)}\;\sim\;\nabla_{\text{text}}(\pi_{t,g-1}^{j};R_{t-1},M_{\text{critic}}),$ (25)
-which form the textual gradients. Each edit is then applied to the current prompt to obtain $N$ candidate prompts
-$\pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}\big(\pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)};M_{\text{critic}}\big),\qquad i=1,\dots,N.$ (26)
-These textual gradients correspond to small, directed moves in language space that, according to the critic’s judgment, are likely to increase ad effectiveness.
-For example, conditioned on the reflection $R_{t-1}$ (e.g., “social occasions and visible grill marks tend to perform better than plain pack shots”), the critic $M_{\text{critic}}$ may propose the following textual-gradient candidates:
-  1. 1.
-    Emphasize social context and enjoyment: “Show the burger being enjoyed at an outdoor barbecue with friends, not alone on a plate.”
-  2. 2.
-    Highlight sensory appeal and indulgence: “Make the patty look extra juicy with clear grill marks, a toasted bun, and melty toppings.”
-  3. 3.
-    Make sustainability salient but secondary: “Subtly include eco-friendly cues (like a small ‘plant-based’ tag or greenery in the background) without overpowering the food.”
-Each of these textual gradients $\delta^{(i)}$ is then applied to the original prompt to produce candidate prompts. For example, applying the first gradient yields the prompt:
-‘‘A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends, everyone smiling and reaching for food.’’
-Applying the second gradient yields the prompt:
-‘‘A close-up, photorealistic shot of a plant-based burger with deep grill marks, a toasted brioche bun, melty vegan cheese, fresh toppings, and steam rising to suggest warmth and juiciness.’’
-To select the next prompt that will go forward to the next gradient step among $N$ candidate prompts $\{\pi_{t,g}^{j,(i)}\}_{i=1}^{N}$, T-BoN BO applies the Best-of-$N$ choice rule over the $N$ corresponding candidate ads $\{c_{t,g}^{j,(i)}\}_{i=1}^{N}$.
-Following liu2025pairjudgermperformbestofn, we implement Best-of-$N$ via a pairwise tournament: in each match, the critic model $M_{\text{critic}}$ compares two creatives side by side and predicts which will perform better, given the campaign goals and the current meta-reflection $R_{t-1}$ (see Figure 11 in Appendix A).
-After multiple randomized pairwise comparisons, the winner $c_{t,g}^{j}$ and its corresponding prompt $\pi_{t,g}^{j}$ are carried forward to the next gradient step.
-Repeating this expansion-and-selection process for $g=1,\dots,G$ yields the refined prompt $\pi_{t}^{j}$ and ad $\Phi(\pi_{t}^{j})$ for trajectory $j$ at iteration $t$.
-Evaluation as the costly step.
-Once all $G$ gradient steps are completed, T-BoN BO performs a single expensive evaluation for each trajectory by deploying the ad $\Phi(\pi_{t}^{j})$ and computing the score $s_{t}^{j}$.
-In the real ad-agency use case, this evaluation step would correspond to running the creative in an A/B test, multi-armed bandit experiment, or survey-based study, and mapping the resulting performance metrics to a scalar score via the scoring rule $r$.
-The new triple $(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})$ is appended to the history $H_{t}$, and a simple acceptance rule is applied: if $s_{t}^{j}$ fails to improve upon $s_{t-1}^{j}$, the trajectory rolls back to $(\pi_{t-1}^{j},\Phi(\pi_{t-1}^{j}),s_{t-1}^{j})$.
-Meta-reflection shared across trajectories.
-After all $J$ trajectories have been evaluated at iteration $t$, the algorithm updates the global meta-reflection $R_{t}=\textsc{MetaReflect}(H_{t})$, as in Section 5.
-Concretely, $M_{\text{critic}}$ is prompted to review a subset of the best- and worst-performing ads and prompts in $H_{t}$ and to summarize the key patterns and rules that distinguish successful creatives from unsuccessful ones (see Figure 12 in Appendix A).
-For example, the reflection may emphasize that scenes with social eating and visible grill marks perform better than isolated product shots, or that emphasizing “satisfying” and “juicy” yields higher effectiveness scores.
-This reflection is then injected into the context of subsequent textual-gradient generation and Best-of-$N$ comparisons, allowing all trajectories to share accumulated knowledge even though they explore different regions of the prompt space.
-Reported outcome.
-After $T$ iterations, the system returns the sequence of best-performing ads $\{c_{t}^{j_{t}^{\ast}}\}_{t=1}^{T}$, where $j_{t}^{\ast}=\arg\max_{j}s_{t}^{j}$ is the trajectory that achieves the highest score at iteration $t$.
-7 Experiments
-We validate the empirical effectiveness of the proposed T-BoN BO framework for a real-world digital marketing problem described in 6.
-Specifically, we examine if and how much T-BoN BO improves outcomes under a fixed evaluation budget compared to state-of-the-art baselines.
-7.1 Evaluation Module using Persona data and LLM-based Simulation
-To evaluate how well an ad performs on the target population and improve prompts, we need a target population and a way to measure the preferences of that population for a given ad.
-There are many ways to do this, such as field experiments, consumer surveys, or simulated responses from a pre-defined target population.
-In our experiments, we use LLM-generated preferences over the Twin-2k-500 persona distribution provided by toubia2025twin.
-An important advantage of using this persona distribution is that it allows us to test the performance of different algorithms fast without the need for large-scale field experiments or expensive surveys.
-[^6] Note that naive field deployment cannot be used to test the relative performance of algorithms since an algorithm’s measured advantage would be a mix of its own effect and distributional effects induced by the ad platform’s ad targeting systems.
-This dataset includes the data of 2,058 U.S.
-adults answering 500 survey questions spanning demographics, personality, cognitive ability, economic preferences, classic heuristics-and-biases tasks, and a pricing study.
-These survey questions of each persona are passed to a multi-modal LLM (Gemini 2.5 Flash) as context so that the LLM can simulate the ad effectiveness for the persona.
-[^7] While LLM-based persona simulation is often biased and does not reflect real-world human behavior (li2025llm; peng2025mega), it still induces a valid preference distribution for which, as a benchmark, we would like to see how well a method can optimize.
-We divided the Twin-2k-500 personas into 80% (1647 personas) for the training set and 20% (411 personas) for the test set.
-Every time we evaluate an ad during the algorithm (i.e., training stage), we randomly sample 200 personas from the training set and simulate the ad’s effectiveness for them.
-Every time we want to test an ad, such as a final ad output from the algorithm, we simulate its effectiveness for all the test set personas.
-7.2 Experiment Setup
-We consider eight synthetic ad campaign scenarios that cover a diverse range of products across distinct categories, defined by the creative brief that outlines the strategic and creative direction for each scenario.
-(For the creative brief for each scenario, see Appendix B.)
+    Shared meta-reflection update.
+    At the end of iteration $t$, after all $J$ trajectories have been evaluated and the history updated, we recompute the global reflection:
+    $R_{t}\leftarrow\textsc{MetaReflect}(H_{t};M_{\text{critic}}).$
+    The operator $\textsc{MetaReflect}(H_{t};M_{\text{critic}})$ compresses the full history $H_{t}$—prompts, system outcomes, and scores—into a concise natural-language reflection $R_{t}$.
+The reflection consists of a small set of actionable rules summarizing cross-trajectory patterns.
+In the ad-generation setting, such rules might take the form “emphasize social context,” “avoid cluttered backgrounds,” or “prefer warm lighting.” This reflection serves as a global memory: in subsequent iterations, the critic LLM conditions on $R_{t}$ when producing textual gradients and performing Best-of-N judgments.
+    There are many ways to construct such reflections.
+In our experiments, following li2025test, we instruct $M_{\text{critic}}$ to compare the five lowest- and five highest-performing outcomes from previous iterations (or all available outcomes if fewer than ten exist) [^5] Limiting the number of outcomes in the critic model’s context window helps mitigate context rotting (hong2025context)., identify recurring visual and textual patterns, and distill these patterns into a small number of generalizable rules.
+After $T$ optimization iterations, we report as the algorithm’s output the sequence of best-performing outcomes
+$\{\Phi(\pi_{t}^{j_{t}^{*}})\}_{t=1}^{T},\qquad j_{t}^{\ast}:=\operatorname{argmax}_{j}s_{t}^{j},$
+i.e., the best trajectory at each iteration, according to the evaluation scores.
+Note that, in each iteration, TextBO generates $N\times G$ prompt/outcome candidates for each trajectory before conducting one evaluation.
+This deliberate trade-off is the key to its evaluation efficiency: the algorithm uses inexpensive generation and critic-based comparisons to identify promising directions, so that each actual evaluation is performed on a small set of carefully selected, high-potential candidates.
+7 Case Study: Digital Ad Optimization
+We now instantiate our framework in a digital advertising setting, showing how an advertiser (or ad platform acting on behalf of an advertiser) can use TextBO to automate the ad generation–analysis–evaluation loop.
+Here, the goal is to refine the ad-generation prompts so that the resulting creatives are better aligned with a target population’s preference distribution, under a limited budget of costly ad evaluations.
+We specialize the general setup from $\S$3 to ad optimization as follows:
+  • $\bullet$
+    $\Phi$ is an ad-generation AI system that can generate ads given a prompt $\pi$.
+[^6] In practice, this image-generation model can be fine-tuned for a particular brand and/or customized to the platform.
+Our approach takes any pre-training as given.
+  • $\bullet$
+    For a prompt $\pi$, $\Phi(\pi)$ is the resulting ad creative deployed on the platform.
+  • $\bullet$
+    $\mathcal{X}$ is the population of platform users, and $\mathcal{D}_{\mathcal{X}}(\pi)$ is the (unknown) distribution of users who are exposed to $\Phi(\pi)$ when this ad is run.
+  • $\bullet$
+    $\mathcal{Y}$ is the space of observed ad responses (clicks, conversions, spend). For each user $x\in\mathcal{X}$, $y(\Phi(\pi),x)$ is the random ad response when $x$ is shown ad $\Phi(\pi)$.
+  • $\bullet$
+    $r:\mathcal{Y}\to[0,1]$ is a scoring rule that maps responses to a normalized effectiveness score (e.g., a weighted function of click, conversion, and spend).
+The campaign objective is the same as the general objective introduced in $\S$3: $J(\pi)=\mathbb{E}\big[\,r\big(y(\Phi(\pi),x)\big)\big]$, where the expectation taken over $x\sim\mathcal{D}_{\mathcal{X}}(\pi)$ and the stochasticity in $y(\Phi(\pi),x)$.
+7.1 Ad Campaign Scenarios
+We consider eight synthetic ad campaign scenarios that cover a diverse range of products across distinct categories, each defined by a creative brief that outlines the strategic and creative direction (see Web Appendix C.1 for the full creative briefs):
   • Scenario 1: “GreenBite,” a new plant-based burger patty.
@@
   • Scenario 5: “Momentum,” a mobile-first banking app for freelancers and the gig economy.
-  • Scenario 6: “MindGarden," a subscription-based meditation and mindfulness app.
+  • Scenario 6: “MindGarden,” a subscription-based meditation and mindfulness app.
   • Scenario 7: “Aeterno,” a classic, automatic Swiss-made wristwatch with a heritage design.
   • Scenario 8: “SyncFlow,” a project management and collaboration software platform for remote teams.
-For each scenario, from its creative brief, we generated 64 high-quality initial prompts that were used as input to the state-of-the-art image generation AI model named Imagen 4  [^8] See Vertex AI’s description (https://console.cloud.google.com/vertex-ai/publishers/google/model-garden/imagen-4.0-generate-preview-06-06) for details.
-to generate the initial 64 high-quality ads.
-(For prompts, see Figure 13 and 14 in Appendix A.) To maintain a rigorous baseline, we made the 64 initial prompts detailed enough to yield strong ads.
-This ensures that the algorithm’s subsequent improvements are not explained as trivial quality gains from having initial under-specified prompts.
-That is, the measured improvements are attributed to the algorithm’s ability to align ads with the underlying persona distribution better.
-When simulating an ad’s effectiveness for a persona, we ask the LLM (with the prompt in Figure 7) what the ad’s “effectiveness” [^9] The vagueness of the term “effectiveness” constitutes an evaluation distribution induced by LLM-persona based on what LLM ‘thinks’ about the term, which we target to align the algorithms with.
-That is, as long as we consistently query the same LLM to predict “effectiveness” using the same prompt, such vagueness is justified.
-for the persona will be, with the score of [1, 2, 3, 4, 5].
-Here, the score of one is the least effective, and the score of five is the most effective.
-As the LLM’s internal assignment of probability for the scores is available as ‘logprobs’ [^10] See Vertex AI’s logprobs description (https://cloud.google.com/vertex-ai/generative-ai/docs/model-reference/inference#logprobs) for details., we use them to calculate the mean score, and consider it as the evaluation score of the ad for the persona (which is a fraction number like 3.41), instead of considering the actual response score from the LLM (which is an integer number among 1-5).
-We then take the average of the evaluation scores for the personas to get the final evaluation score of the ad.
-Persona prompt for simulating ad effectiveness. PERSONA DATA: $\{$
-Which part of the United States do you currently live in? [Answer]
-What is the highest level of schooling or degree that you have completed? [Answer]
-$\ldots$ (Many other survey questions and answers) $\ldots$
+7.2 Implementation Details
+There are five main components of the algorithm: (1) initial prompt ($\pi_{0}$), and (2) Ad-generation module ($\Phi$), (3) Critic LLM and its four operators defined in $\S$5.1, (4) the population distribution $\mathcal{D}_{\mathcal{X}}(\pi)$ and the evaluation operator $\mathrm{Eval}()$, as defined in Equation (23).
+We now describe how each of these components is implemented in our experiments.
+Initial prompts ($\pi_{0}^{j}$).
+For each scenario, we first created 64 high-quality prompts from its creative brief using a meta-prompt.
+In Web Appendix C.2, see Figure A1 for the details of the meta-prompt and Figure A2 for an example of two prompts in scenario 1.
+Among those 64 ads, we chose five as initial prompts; we will describe the details of the choice rule in $\S$7.3.
+Ad-generation module ($\Phi$).
+We utilize a state-of-the-art image generation model, Google Imagen 4, [^7] See Vertex AI’s description (https://console.cloud.google.com/vertex-ai/publishers/google/model-garden/imagen-4.0-generate-preview-06-06) for details.
+Imagen 4 delivers high-fidelity, photorealistic text-to-image generation at up to 2K resolution, and excels in rendering intricate details such as realistic textures, lighting, and camera effects.
+Its features include typography accuracy for legible in-image text and precise adherence to complex prompts.
+Critic LLM and its Operators.
+Unless otherwise noted, we use Gemini 2.5 Flash  [^8] See Vertex AI’s description (https://docs.cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-5-flash) for details.
+as the multimodal critic model $M_{\text{critic}}$ and implement the four operators discussed in $\S$5.1 using the meta-prompts described in Web Appendix C.3: (1) PairwiseJudge operation; see meta-prompt in Figure A3, (2) MetaReflect operation; see meta-prompt in Figure A4, (3) Textual gradient operation $\nabla_{\text{text}}$; see meta-prompt in Figure A5, and (4) Rewrite operation Apply; see meta-prompt in Figure A6.
+Evaluation module using persona data and LLM-based simulation.
+We now describe how we implement the evaluation operator $\mathrm{Eval}$ in this case study.
+To evaluate how well an ad performs on the target population and to update prompts, we need (i) a target population ($\mathcal{D}_{\mathcal{X}}(\pi)$.) and (ii) a way to measure that population’s preferences for a given ad ($\mathrm{Eval}(\cdot)$).
+In principle, there are three ways define the evaluation operator – (1) field experiments/deployment in a real ad platform, (2) consumer surveys, and (3) simulated responses from a pre-defined target population.
+The former two options (field deployment and consumer surveys) are costly in time and resources, and can require large samples to identify effective differences across ads given low CTRs in digital ad settings; see (lewis2015unfavorable; johnson2023inferno).
+Thus, using these approaches for algorithm evaluation is not ideal.
 […]
-$\ldots$ (Many other survey questions and answers) $\ldots$
-AD IMAGE: [image] TASK: Return only one item from ["1","2","3","4","5"] for ad effectiveness.
-Effective Score Scale Definition:
-1: Extremely Unlikely. The persona would actively ignore or be annoyed by this ad.
-2: Unlikely. The persona would likely scroll past without a second thought.
-3: Mediocre. It is hard to decide whether the personal would click or don’t click.
-4: Likely. The persona is intrigued and has a good chance of clicking to learn more.
-5: Extremely Likely. The persona is the ideal target; a click is almost certain.
-No explanation. Just the score.
-Figure 7: Persona simulation prompt for the digital marketing problem.
-Among the initial ads, we identified the best initial ad, which we call Best-of-64, and the worst five initial ads, which we call Worst5-of-64, by running a best-arm identification bandit algorithm (russo2016simple) by alternating the best-arm identification objective and worst 5-arm identification objective for evaluations of 5,000 sequential random samples from the training persona set.
+We split the Twin-2k-500 personas into 80% (1,647 personas) for training and 20% (411 personas) for testing.
+During algorithm training, each evaluation of an ad $\Phi(\pi)$ proceeds as follows: we randomly sample 200 personas from the training set and simulate ad effectiveness for each persona via the LLM.
+When we test a final ad (e.g., after running TextBO), we simulate its effectiveness for all personas in the test set.
+We note that important advantage of our approach: using this persona distribution with LLM-generated preferences enables fast and controlled comparisons of different algorithms without large-scale field experiments or expensive surveys.
+[^10] Naive field deployment cannot be used to test the relative performance of algorithms, since an algorithm’s measured advantage would then mix its own effect with distributional effects induced by the ad platform’s targeting systems.
+Indeed, our approach can serve as a template for algorithm comparisons in many other settings where researchers require evaluations that reflect the variance/heterogeneity of human preferences.
+Further, we note that while LLM-based preference simulation may not reflect true human preferences (li2025llm; peng2025mega), it still induces a well-defined preference distribution on which we can benchmark optimization methods and algorithms.
+Specifically, kang2025llmpersonassubstitutefield proves that swapping humans for personas is equivalent to changing the evaluation population (e.g., from the New York population to the Jakarta population), under two benchmark hygiene conditions: (i) methods observe only the aggregate outcome (called aggregate-only observation condition) and (ii) evaluation depends only on the submitted artifact and not on the algorithm’s identity or provenance (called algorithm-blind evaluation condition).
+Furthermore, they show that simply increasing the size of the persona dataset is sufficient to guarantee that persona simulation becomes as useful a benchmark as field experimentation.
+Nevertheless, we also provide evaluations of our approach that do not leverage persona datasets; see $\S$8.4 for details.
+7.3 Application of TextBO
+We now describe how the TextBO algorithm is instantiated for the ad-optimization problem.
+Here, we choose the hyperparameters as: iterations $T=10$; trajectories $J=5$; gradient steps $G=5$; and candidates per gradient step $N=5$.
+Initialization.
+As described in $\S$7.2, for each scenario, we first generate 64 high-quality prompts from its creative brief using a meta-prompt.
+Among the 64 ads for each scenario, we identify the best initial ad, Best-of-64 (which will be used as one of the baselines but is not directly used in our algorithm; see $\S$8.1 for details), and the worst five, Worst5-of-64, by running a best-arm identification bandit algorithm (russo2016simple).
+Specifically, we alternate between a best-arm identification objective and a worst 5-arm identification objective over 5,000 sequential random samples from the training persona set.
+These Worst5-of-64 ads serve as starting points $\{\pi_{0}^{j}\}_{j=1}^{5}$ for TextBO procedure.
+This starting point allows us to attribute observed gains to the algorithms rather than to favorable starting prompts.
+[^11] To maintain a rigorous baseline, we make these 64 prompts detailed enough to yield strong creatives.
+This ensures that subsequent improvements are not merely due to trivial gains from starting with an under-specified prompts; instead, improvements reflect better alignment with the underlying target audience distribution.
+Let $\{s_{0}^{j}\}_{j=1}^{J}$ denote the evaluation scores for these five ads.
+Then, the tuples $\{(\pi_{0}^{j},\Phi(\pi_{0}^{j}),s_{0}^{j})\}_{j=1}^{J}$ form the initial history $H_{0}$ for the algorithm.
+Best-of-N textual-gradient steps.
+Within each optimization iteration $t$, TextBO refines each trajectory’s prompt $\pi_{t-1}^{j}$ through five steps of Best-of-N textual-gradients before evaluation (see Figure 3).
+We follow the generic procedure from $\S$6, but here specialize the critic’s context and edits to the ad domain.
+[…]
+  • $\bullet$
+    Emphasize social context and enjoyment: “Show the burger being enjoyed at an outdoor barbecue with friends, not alone on a plate.”
+  • $\bullet$
+    Highlight sensory appeal and indulgence: “Make the patty look extra juicy with clear grill marks, a toasted bun, and melty toppings.”
+  • $\bullet$
+    Make sustainability salient but secondary: “Subtly include eco-friendly cues (like a small ‘plant-based’ tag or greenery in the background) without overpowering the food.”
+We then query these edits for the operation Apply operation to an initial prompt (see Figure A6 in Web Appendix C.3 for the meta-prompt) may yield candidates such as:
+‘‘A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends, everyone smiling and reaching for food.’’
+‘‘A close-up, photorealistic shot of a plant-based burger with deep grill marks, a toasted brioche bun, melty vegan cheese, fresh toppings, and steam rising to suggest warmth and juiciness.’’
+‘‘A photorealistic image of a sizzling plant-based burger with rich grill marks, served on a rustic wooden table with subtle greenery in the background and a small plant-based tag, conveying indulgence without sacrificing sustainability.’’
+To select which candidate to keep for the next gradient step, TextBO applies the Best-of-N rule from $\S$5.2 to the ads $\{\Phi(\pi_{t,g}^{j,(i)})\}_{i=1}^{N}$.
+Specifically, following liu2025pairjudgermperformbestofn, we implement Best-of-N via a pairwise tournament: in each match, the critic compares two creatives side by side and predicts which will perform better using operator PairwiseJudge, given the campaign goals and the current reflection $R_{t-1}$ (see Figure A3 in Web Appendix C.3 for the meta-prompt).
+The tournament winner prompt and its corresponding ad, $\left(\pi_{t,g}^{j},\Phi\left(\pi_{t,g}^{j}\right)\right)$, are carried forward to the next gradient step.
+Repeating this for $g=1,\dots,G$ yields the final refined prompt $\pi_{t}^{j}$ and ad $\Phi(\pi_{t}^{j})$ for trajectory $j$ at iteration $t$.
+Evaluation as the costly step.
+After the $G$ gradient steps, TextBO performs a single expensive evaluation per trajectory (each of which consists of assessing the ad for 200 randomly sampled personas from the training persona set) by deploying $\Phi(\pi_{t}^{j})$ to get the evaluation result $s_{t}^{j}$ using the evaluation operator $\mathrm{Eval}$ from Equation (23) (instantiated using the evaluation module discussed in $\S$7.2).
+The new triple $(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})$ is appended to $H_{t}$, and the usual acceptance rule is applied: if $s_{t}^{j}\leq s_{t-1}^{j}$, the trajectory reverts to $(\pi_{t-1}^{j},\Phi(\pi_{t-1}^{j}),s_{t-1}^{j})$, ensuring non-decreasing scores along each trajectory.
+Meta-reflection shared across trajectories.
+After all $J$ trajectories have been evaluated at iteration $t$, the algorithm updates the global meta-reflection $R_{t}=\textsc{MetaReflect}(H_{t};M_{\text{critic}})$.
+Concretely, $M_{\text{critic}}$ is prompted to review a subset of the best- and worst-performing ads and prompts in $H_{t}$ and to summarize the key patterns that distinguish successful creatives from unsuccessful ones (see Figure A4 in Web Appendix C.3 for the meta-prompt).
+For example, it may infer that scenes with social eating and visible grill marks outperform isolated pack shots, or that emphasizing “satisfying” and “juicy” tends to improve effectiveness.
+This reflection is then injected into the context of subsequent textual-gradient generation and Best-of-N comparisons, allowing trajectories to share accumulated knowledge even as they explore different regions of the prompt space.
+Reported outcome.
+After $T=10$ iterations, the system reports the sequence of best-performing ads $\{\Phi\left(\pi_{t}^{j_{t}^{*}}\right)\}_{t=1}^{T}$, where $j_{t}^{\ast}=\arg\max_{j}s_{t}^{j}$ is the trajectory that achieved the highest score at iteration $t$.
+8 Experiments for the Digital Ad optimization
+In this section, we discuss the numerical performance of TextBO for the ad-optimization case (discussed in §7).
+In $\S$8.1, we first discuss the baseline algorithms we compare against, and then in $\S$8.2, we describe the experiment design.
+We present the main results in $\S$8.3 and then finally end with an ablation study in $\S$8.4.
 8.1 Baselines
-As baselines, we consider the performance of two popular state-of-the-art algorithms.
-One is the Best-of-64 (snell2024scaling) baseline, which is one of the most popular and well-performing test-time alignment methods that has been reported to outperform reinforcement learning fine-tuning methods under large enough $N$ (gao2023scaling; mudgal2023controlled; eisenstein2023helping; beirami2024theoretical).
-Specifically, the test persona set-based evaluation result for the Best-of-64 ad (identified above in Section 7.2 using training persona set) was used as the Best-of-64 baseline.
-We also evaluated the ads in Worst5-of-64 in the same way and took the mean of their scores to obtain the Worst5-of-64 baseline.
-Another baseline is GEPA (Genetic-Pareto) (agrawal2025gepa).
-As discussed earlier, this reflective and evolutionary prompt evolution algorithm has recently demonstrated superior performance compared to GRPO (liu2024deepseek), a state-of-the-art reinforcement fine-tuning technique.
-GEPA keeps a set of candidate prompts, and samples a prompt to mutate based on its evaluation result for a problem (in our case, a persona) set they call the Pareto set.
-A Pareto set stores ‘important’ problems.
-Since we randomly sample personas from the training persona set, we cannot store specific personas and retrieve them later.
-Therefore, we instead use the version of GEPA that 1) randomly samples a persona and 2) mutates all candidate prompts equally and throws away the old prompts to avoid unbounded growth of the candidate set.
-For the experiment, we run five-trajectory ($J=5$) parallel T-BoN BO and GEPA starting from Worst5-of-64, i.e., the worst five initial prompts we identified in Section 7.2.
-This starting point allows us to attribute observed gains to the algorithms rather than to favorable starting prompts.
-We test if T-BoN BO outperforms the Best-of-N baseline and GEPA baseline throughout the optimization process.
-For both T-BoN BO and GEPA, we ran 10 optimization steps, evaluating 200 randomly sampled personas per step (2,000 total persona evaluations).
-We used a multi-modal LLM (Gemini-2.5-Flash) to serve as the critic model for both T-BoN BO and GEPA.
+We consider two state-of-the-art algorithms in the AI literature as baselines – (1) Best-of-N (snell2024scaling) and (2) GEPA (agrawal2025gepa).
+Best-of-N is one of the most popular and well-performing test-time alignment methods that has been reported to outperform reinforcement learning fine-tuning methods under large enough $N$ (gao2023scaling; mudgal2023controlled; eisenstein2023helping; beirami2024theoretical).
+For any given scenario, this algorithm samples $N$ candidate outputs from a fixed generator and returns the single candidate with the highest predicted score under a judge/reward model, trading extra inference-time compute for higher expected quality (snell2024scaling; beirami2024theoretical).
+In our setting, recall that we generate a set of 64 initial ads (see $\S$7.2 on the details of the initial ad set generation).
+Then, we use the training dataset of personas and the Eval() operator in conjunction witha best-arm identification algorithm to identify the Best-of-64 ad (as described in $\S$7.3).
+Finally, we evaluate Best-of-64’s performance on the testing persona dataset; this score serves as our Best-of-64 baseline for that scenario.
+GEPA is a reflective and evolutionary prompt optimization algorithm that has demonstrated superior performance compared to state-of-the-art reinforcement fine-tuning techniques such as GRPO (liu2024deepseek).
+In our implementation of GEPA, we use the same set of initial prompts, critic models, ad generation modules, and evaluation operators we use for TextBO (see $\S$7.2 for details).
+Further, for each of the eight scenarios, we start GEPA with the same Worst5-of-64 prompts as TextBO, i.e., the worst five initial prompts we identified in $\S$7.2.
+We also run GEPA for the same number of iterations ($T=10$).
+Here, keeping the components of the algorithms and the starting points the same across both algorithms ensures a fair comparison.
+Notably, when we adapt GEPA to the digital ad optimization problem setup (where we cannot store and retrieve problem instances (e.g., personas) freely anytime we want as in agentic AI benchmarks of agrawal2025gepa), it effectively reduces to TextBO with $G=1$ and N=1N=1121212In $\S$9, we adapt the principles of Best-of-N gradient multiple gradient steps of TextB for GEPA and test it for agentic AI benchmark experiments (agrawal2025gepa) for which GEPA was tested.; see Web Appendix D.1 for the implementation details of the adapted GEPA.
+8.2 Experiment Design
+Our experiment is designed to test whether we can start from ex-post least-aligned ads (Worst5-of-64) and apply TextBO to outperform the ex-post best-aligned ad (Best-of-64).
+Here, “alignment” is measured by the evaluation score induced by our persona-based evaluation environment (see $\S$7.2).
+This design choice matters because the Best-of-64 baseline is already a strong, ex-post selection benchmark: it is the best-performing ad within a reasonably rich, human-like initial pool of 64 creatives generated from the same creative brief (snell2024scaling; beirami2024theoretical).
+By initializing optimization from the opposite end of the same pool (Worst5-of-64), we attribute improvements to the algorithm’s performance rather than to a favorable starting prompt.
+Relatedly, we are also interested in comparing the performance of TextBO with GEPA.
+To do so, we also need to ensure that the comparison between GEPA and TextBO is fair and both use similar starting points and algorithm components.
+Therefore, for each scenario, we proceed as follows:
+  • $\bullet$
+    Construct a common initial candidate pool.
+Using the scenario’s creative brief, we generate 64 detailed prompts and their corresponding ad images.
+All algorithms start with this initial pool of prompts and ads.
+  • $\bullet$
+    Define ex-post baselines and initializations within the pool.
+Using the training persona set and the fixed evaluation operator $\mathrm{Eval}(\cdot)$, we identify (i) the single best ad in the pool (Best-of-64) and (ii) the five worst ads (Worst5-of-64) via best-arm identification (see $\S$7.3).
+We use Worst5-of-64 as the starting set $\{\pi_{0}^{j}\}_{j=1}^{5}$ for both TextBO and GEPA, while Best-of-64 serves as the ex-post “best-aligned” benchmark.
+  • $\bullet$
+    Hold the algorithm components and evaluation environment fixed across methods.
+All methods use the same (i) scenarios and initial pool, (ii) ad generator $\Phi$, (iii) critic model $M_{\text{critic}}$, and (iv) persona-based evaluation operator.
+This ensures that differences in outcomes are attributable to the optimization procedure, rather than the evaluation environment.
+  • $\bullet$
+    Evaluate under a common budget and report out-of-sample performance.
+During optimization steps in GEPA and TextBO, each costly evaluation uses the same protocol (randomly sampled 200 personas from the training set; see $\S$7.2).
+For reporting, we evaluate the resulting ads on the held-out 411 test persona set and compare progress trajectories and final performance across methods.
 8.3 Results and Analysis
-Trend analysis
-Figure 8: Progress comparison of T-BoN BO and GEPA with Best-of-64 baseline.
-T-BoN BO implements parallel T-BoN BO with $J=5$.
+We now present the main results from the experiments.
+Note that while the algorithm is designed to minimize regret, we cannot directly plot regret here since we don’t have true $\pi^{\star}$ (or the the prompt that maximizes the score), a priori, for any of the scenarios.
+Therefore, throughout, we report experiment results in terms of empirical performance scores.
+Figure 5: Progress comparison of TextBO and GEPA with Best-of-64 baseline.
+TextBO implements parallel TextBO with $J=5$.
 The performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the testing persona set.
-Since it is the mean of means, we don’t plot the standard errors.
-Figure 8 plots how much T-BoN BO, GEPA, and Best-of-N improved their performance from the initial Worst5-of-N score across eight scenarios.
-(For the detailed data used to generate this plot, refer to Table 2 and 3 in Appendix C.) At each step (the $x$-axis), the performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the testing persona set.
-For example, the step-2 score is the average of the eight step-2 mean evaluation scores.
-As the figure illustrates, both T-BoN BO and GEPA improve consistently across steps, with T-BoN BO demonstrating faster early gains and achieving a significantly better final performance than GEPA.
-Notably, by the third step on average, which is equivalent to 600 persona evaluations, T-BoN BO and GEPA both outperform the Best-of-64 baseline.
-Statistical analysis
-Model
-$\text{score}=\alpha+\beta_{1}\mathbf{1}\{t=10\}+\beta_{2}\mathbf{1}\{\mathrm{TBoN}\}\,\mathbf{1}\{t=10\}+\text{persona FE}+\text{scenario FE}+\varepsilon$
+Main results and trend analysis:
+Figure 5 shows the results of our numerical experiments over 10 iterations: it plots how much TextBO and GEPA improved their performance from the initial Worst5-of-N score across eight scenarios.
+(For the detailed data used to generate this plot, see Tables A1 and A2 in Web Appendix D.2.) At each step (the $x$-axis), the performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the test persona dataset.
+[^13] Best-of-N remains constant because, as discussed earlier, it is simply the best among the initial pool of ads and this baseline does not involve generation of new ads.
+For example, the step 2 score is the average of the eight step 2 mean evaluation scores.
+Both TextBO and GEPA improve consistently across steps, with TextBO demonstrating faster early gains and achieving a significantly better final performance than GEPA.
+Notably, by the third step on average, which is equivalent to 600 persona evaluations, both TextBO and GEPA outperform the Best-of-64 baseline.
+Further, by step 8, both algorithms have reached stable performance.
 Effect Estimate Std. err T-stat
 $\beta_{1}$: GEPA gain $(t=10)$ from Worst5-of-64 0.3168*** 0.0325 9.752
-$\beta_{2}$: T-BoN BO gain - GEPA gain $(t=10)$ from Worst5-of-64 0.0743*** 0.0248 2.991
-$\beta_{1}+\beta_{2}$: T-BoN BO gain $(t=10)$ from Worst5-of-64 0.3911*** 0.0355 11.011
+$\beta_{2}$: TextBO gain - GEPA gain $(t=10)$ from Worst5-of-64 0.0743*** 0.0248 2.991
+$\beta_{1}+\beta_{2}$: TextBO gain $(t=10)$ from Worst5-of-64 0.3911*** 0.0355 11.011
 Summary statistics
@@
   • * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
-Table 1: Two-way fixed effect (TWFE) analysis result with persona and scenario fixed effects with two-way clustered SEs (cameron2011robust).
-Standard error of $\beta_{1}+\beta_{2}$ was computed using Delta method.
-In addition to the aggregate trend analysis provided in Figure 8, we now provide statistical evidence that T-BoN BO indeed outperforms the GEPA.
-For this, we conducted a Two-way Fixed Effect (TWFE) analysis using T-BoN BO and GEPA’s final ($t=10$) scores to compare the effectiveness of each method rigorously.
-Specifically, we consider the following TWFE model:
-$\text{score}=\alpha+\beta_{1}\mathbf{1}\{t=10\}+\beta_{2}\mathbf{1}\{\mathrm{TBoN}\}\,\mathbf{1}\{t=10\}+\text{persona FE}+\text{scenario FE}+\varepsilon.$ (27)
-where $\mathbf{1}\{t=10\}$ is the indicator variable of whether it is $t=0$ (GEPA and T-BoN BO’s initial starting point, i.e., Worst5-of-64 score) or $t=10$ (GEPA and T-BoN BO’s final score), $\beta_{1}$ is the GEPA’s gain from Worst5-of-64 after $t=10$, $\beta_{2}$ is the T-BoN BO’s advantage over GEPA after $t=10$, and $\beta_{1}+\beta_{2}$ is the T-BoN BO’s gain from Worst5-of-64 after $t=10$.
-The unobserved residual $\varepsilon$ is assumed to follow the normal distribution.
-To remove potential correlation among unobserved residuals within the same persona or scenario, we conducted a cluster-robust Two-way Fixed Effect (TWFE) analysis with two-way clustering (cameron2011robust; cameron2015practitioner).
-The result of TWFE analysis is shown in Table 1.
-As we can see, both of T-BoN BO’s gain and GEPA ’s gain from the starting point (Worst5-of-64) after $T=10$ are statistically significant.
-In addition, we conclude that T-BoN BO has a significant 23.5% more gain in score compared to GEPA’s gain.
-Qualitative analysis
-The significant gains of GEPA and T-BoN BO can be interpreted as demonstrating their ability to align with the persona distribution.
-However, such significance can also be easily gained by starting from a poor starting point, i.e., a poor Worst5-of-N prompt and corresponding ad image.
-One way to check whether this happens is to compare the quality of the prompts; we provide them in Appendix A to demonstrate that this is not the case.
-Another qualitative approach is to compare the generated ad images and verify whether the differences among ads are in quality or in the message and representation.
-Figure 9 compares the ad images of the baseline methods.
-As we can see, there is no clear visual quality difference among those ad images; the differences are mainly in the ad’s message and how it is represented.
-This supports the claim that T-BoN BO and GEPA improvements from Worst5-of-N demonstrated in Figure 8 are attributed to alignment, not trivial quality gains from correcting poor or under-specified prompts.
+Table 1: Two-way fixed effect (TWFE) analysis result with persona and scenario fixed effects with two-way clustered SEs. Standard error of $\beta_{1}+\beta_{2}$ is computed using the Delta method.
+Statistical analysis:
+Figure 5 shows the mean of means (i.e., scores are averaged over personas in test dataset and then over the eight scenarios).
+Since both the set of personas and scenarios are common across all algorithms, we cannot directly plot standard errors.
+Therefore, we now present a statistical analysis using the raw scores at the persona-scenario level and examine whether the performance of TextBO is statistically significant different from that of GEPA.
+We consider the Two-way Fixed Effect (TWFE) model:
+$\text{score(persona, scenario, t)}=\alpha+\beta_{1}\mathbf{1}\{t=10\}+\beta_{2}\mathbf{1}\{\mathrm{TextBO}\}\,\mathbf{1}\{t=10\}+\text{persona FE}+\text{scenario FE}+\varepsilon,$
+where $\mathbf{1}\{t=10\}$ is the indicator variable of whether it is $t=0$ (GEPA and TextBO’s initial starting point, i.e., Worst5-of-64 score) or $t=10$ (GEPA and TextBO’s final score), $\beta_{1}$ is the GEPA’s gain from Worst5-of-64 after $t=10$, $\beta_{2}$ is the TextBO’s advantage over GEPA after $t=10$, and $\beta_{1}+\beta_{2}$ is the TextBO’s gain from Worst5-of-64 after $t=10$.
+Further, we control for persona and scenario fixed effects to account for persistent differences in personas and scenarios (across both algorithms).
+The standard errors are two-way clustered (cameron2011robust; cameron2015practitioner).
+The results from this TWFE model are shown in Table 1.
+We see that both TextBO’s and GEPA show statistically significant gains from the starting point (Worst5-of-64) after ten iterations.
+In addition, at $T=10$, TextBO shows a 23.5% more gain in score compared to GEPA.
+Together, these findings provide two key takeaways: (1) both GEPA and TextBO are able to self-improve and generate ads better aligned with the target population, and (2) the performance of TextBO is significantly better than that of GEPA for the same number of evaluations, and this difference is both statistically significant and meaningful in magnitude.
+Qualitative analysis:
+So far, we have shown that TextBO is able to self-improve and generate ads that align well with the persona distribution within a few iterations.
+However, one could argue that this can be simply due to a weak set of initial prompts, i.e., a poor set of Worst5-of-N prompts and corresponding ad images.
+We examine this alternative explanation by examining the prompts and ad creatives/images.
+First, one natural way to rule out this explanation is to compare the quality of the prompts at the starting point and end point of the algorithm, and show that the prompts at the starting point are grammatically correct, meaningful, and sensible and that the differences in the prompts at $t=0$ and $t=10$ stem from differences in the message and focus of the prompt rather than grammar/structure.
+We confirm that this in indeed the case by comparing the prompts at the starting and end points of the algorithm for all the scenarios and trajectories.
+For example, see the prompt for the Worst-of-64 prompt for Scenario 1 in Figure A2 in Web Appendix C.2 and the prompt for TextBO at the last step in Figure A9 in Web Appendix D.3.
+A second way to rule this out this alternative explanation is to compare the generated ad images and verify that the differences among ads across algorithms and (over the iterations, within a algorithm) stem from the message and representation, and not from the image quality.
+To that end, in Figure 6, we show the ad images at the start and end points for both GEPA and TextBO, as well the Best-of-N baseline (for three scenarios).
+There is no clear visual quality differences among these ad images; the differences are mainly in the ads’ message and how it is represented.
+This further suggests that the improvements from TextBO (and GEPA) over the Worst5-of-N starting point can be attributed to alignment, and are not due to trivial quality gains from correcting poor or under-specified initial prompts.
 GreenBite (Plant-based patty) MindGarden (Meditation app) AuraSonics X1 (Noise-canceling earbuds)
@@
 GEPA (after T=10) Refer to caption Score: 3.046 Refer to caption Score: 3.068 Refer to caption Score: 3.028
-T-BoN BO (after T=10) Refer to caption Score: 3.168 Refer to caption Score: 3.132 Refer to caption Score: 3.104
-Figure 9: T-BoN BO and baseline’s generated Ad images for three fictional brands: GreenBite burger patty (Scenario 1), MindGarden meditation app (Scenario 6), and AuraSonics noise-canceling earbuds (Scenario 2).
-For the Ad images of the other five brand scenarios, see Figure 17 in Appendix C.
-Figure 10: An ablation study experimenting T-BoN BO with $J=5$ and Best-of-64 for Gemini 2.5 Flash instead of persona distribution with temperature 0.
-(GEPA was excluded, as there is no problem-dependent heterogeneity.) The performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the testing persona set.
-Again, since it is the mean of means, we don’t plot the standard deviations.
+TextBO (after T=10) Refer to caption Score: 3.168 Refer to caption Score: 3.132 Refer to caption Score: 3.104
+Figure 6: TextBO and baselines’ generated ad images for three fictional brands: GreenBite burger patty (Scenario 1), MindGarden meditation app (Scenario 6), and AuraSonics noise-canceling earbuds (Scenario 2).
+For the ad images of the other five scenarios, see Figure A8 in Web Appendix D.2.
+Figure 7: Ablation study with TextBO with $J=5$ and Best-of-64 for Gemini 2.5 Flash as the evaluator, with temperature 0.
+The performance score is the average of eight scenarios’ mean evaluation score at that step.
 8.4 Ablation Study
-The main experiment in Section 7.1-7.4 evaluates the candidate ad at each iteration as the mean of scores across 411 personas (20% of the persona data in Twin-2k-500 (toubia2025twin)) that were injected into the LLM’s context.
-Each persona consists of a large amount of information, with 500 survey questions accounting for around 50,00050{,}000 tokens.
-Would T-BoN BO still outperform strong state-of-the-art baselines when the LLM receives less persona context information?
-To test this, we ran a no-persona ablation experiment in which the evaluator is a single LLM (Gemini 2.5 Flash) with no auxiliary persona context.
-Concretely, at each iteration, we present only the ad image candidate to LLM judge (Gemini 2.5 Flash) together with the same 1–5 effectiveness rubric, set temperature to 0, and compute the same scalar score as the log-probability–weighted expectation over $\{1,\dots,5\}$ as before.
+So far, in all our experiments, we used the persona dataset for evaluation and the goal was to optimize the ads serve to the target population captured by the persona dataset (toubia2025twin).
+However, a natural question here is whether TextBO can still perform well if we simply use a different target population/evaluator without access to extensive persona information during training.
+(Recall that each persona consists of a large amount of information, with 500 survey questions accounting for around 50,00050{,}000 tokens.) Therefore, we run a no-persona ablation experiment in which the evaluator is a single LLM (Gemini 2.5 Flash) with no auxiliary persona context.
+Concretely, at each iteration, we present only the ad image candidate to a LLM judge (Gemini 2.5 Flash) together with the same 1–5 effectiveness rubric, set the temperature to 0, and compute the same scalar score as the log-probability–weighted expectation over $\{1,\dots,5\}$ as before.
 [^14] With temperature =0=0, the judge’s output is deterministic for a fixed image and rubric; the logprob vector is still available, so the expected score is stable across repeated calls.
-This collapses the evaluation to a deterministic mapping and removes heterogeneity from personas and sampling, allowing us to test whether T-BoN BO’s performance depends on specific persona information-related setup.
-To ensure there is no leakage of preference, we used GPT-5 instead of Gemini 2.5 Flash for generating meta-reflection, generating textual gradients, and making Best-of-N choices.
-We again compare T-BoN BO (J=5J{=}5 trajectories) against the Best-of-64 initial baseline; we exclude GEPA, which requires multiple personas (multiple problems, in GEPA’s language.)
-Figure 10 plots T-BoN BO across 10 optimization steps starting from the Worst5-of-64 prompts under the ablation setup.
-As in Figure 8, the main experiment’s results, the value shown in Figure 10 is the mean of that step’s scores over eight scenarios.
-As can see in Figure 10, T-BoN BO again outperforms Best-of-64 baseline from optimization step 3.
-Except for the fact that in general the evaluation scores without persona information are more optimistic than scores with persona information for both T-BoN BO and Best-of-64, the ablation experiment’s trend in Figure 10 shows the same trend we saw in Figure 8.
+This collapses the evaluation to a deterministic mapping between ads and scores; it removes heterogeneity from personas and stochasticity from sampling, allowing us to test whether TextBO’s performance depends on specific persona information-related setup.
+To ensure that there is no leakage of preferences, we used GPT-5 instead of Gemini 2.5 Flash for critic model $M_{critic}$ (i.e., the operators used for generating meta-reflection, generating textual gradients, and making the Best-of-N choices.).
+We again compare TextBO (J=5J{=}5 trajectories) against the Best-of-64 initial baseline.
+The results from this exercise are shown in Figure 7 plots TextBO across 10 optimization steps starting from the Worst5-of-64 prompts under the ablation setup.
+[^15] Here, Worst5-of-64 and Best-of-64 are different from those of the main experiments, as the preference distribution is different.
+As before, we see that TextBO outperforms Best-of-64 baseline from optimization step 3 onwards, and the overall trend in Figure 7 is similar to that in Figure 5.
+The main difference is that the evaluation scores without persona information are more optimistic than scores with persona information for both TextBO and Best-of-64.
+This is understandable, since peng2025mega find that persona information does not significantly improve LLMs’ overall ability to predict preferences and mainly adds variance to the predictions.
+Since variance makes the learning problem harder, it is natural that the main experiment, which uses personas, shows slower improvement than the pure-LLM case here.
+Overall, these findings suggest that when learning the target population’s preferences is difficult, the method may require more iterations and larger training samples to optimize effectively.
+9 Experiments for Agentic AI Benchmarks
+In $\S$8, we saw that TextBO significantly outperforms GEPA for ad optimization experiments.
+Recall that the adapted version of GEPA in the ads settings effectively reduced to TextBO with $N=1$ and $G=1$.
+As such, one of the main takeaways from those experiments is that Best-of-N gradient sampling and multiple gradient steps can effectively improve the performance of self-improving AI under limited evaluation budget.
+[…]
+9.1 Experiment Design
+To rigorously evaluate GEPA and TextBO-GEPA, we closely replicate three agentic AI benchmarks and their corresponding experimental setups from agrawal2025gepa.161616agrawal2025gepa also included IFBench ifbench_bench, for which they report that none of the algorithm show improvement.
+Therefore we exclude it in our analysis.
+In all the experiments, we used Qwen3-8B (qwen3_technical_report), the same model and setup used in agrawal2025gepa (decoding temperature of 0.6, top-p of 0.95, and top-k of 20 for training as well as inference).
+As reported in agrawal2025gepa, GEPA saturates after $7,500$ evaluations (rollouts), so we conduct experiments up to $7,500$ evaluations.
+We now briefly describe the three benchmark tasks:
+HotpotQAhotpotqa_bench provides HotpotQA, a set of diverse, explainable multi-hop question answering tasks comprising approximately 113k Wikipedia-based examples.
+This task requires reasoning over multiple supporting documents and providing sentence-level supporting facts to enable intense supervision.
+We follow the implementation details described in agrawal2025gepa, except for the retriever module, which was not specified in the paper.
+So we conservatively set it to BM25 (robertson2009probabilistic), one of the most widely used retrievers.
+We use 150 examples for training, 300 for validation, and 300 for testing.
+HoVerhover_bench provides HoVer, a set of many-hop fact-extraction and claim-verification tasks, constructed from the same 113k Wikipedia-based examples used in HotpotQA.
+It tests complex reasoning that requires aggregating evidence from multiple sentences across disparate documents, often involving diverse reasoning graphs and long-range dependencies.
+Again, we follow the implementation details described in agrawal2025gepa, except for the retriever module, which was again not specified in the paper, and was set to BM25 (robertson2009probabilistic).
+Again, we use 150 examples for training, 300 for validation, and 300 for testing.
+PUPApapillon_bench provides PUPA, a set of privacy-conscious delegation tasks which address real-world user queries by orchestrating an ensemble of trusted and untrusted models.
+The primary objective is to maintain a high response quality, comparable to that of proprietary models, while minimizing exposure of personally identifiable information (PII) to the untrusted backend.
+Again, we follow the implementation details described in agrawal2025gepa, except that we use case-insensitive substring matching instead of LLM-as-a-judge to check whether each PII exists in the output.
+We use 111 training examples, 111 for validation, and 221 for testing.
+9.2 Results and Analysis
+(a) HotpotQA, Qwen3 8B
+(b) HoVer, Qwen3 8B
+(c) PUPA, Qwen3 8B
+Figure 8: Experiment results for three agentic AI benchmarks considered in agrawal2025gepa.
+As more evaluations are conducted, TextBO-GEPA (GEPA with Best-of-N gradient sampling and multiple gradient steps) consistently outperform vanilla GEPA.
+Figure 8 and Table 2 present the results of agentic AI benchmark experiments.
+We plot the improvement from their initial scores for both TextBO-GPEA and GEPA.
+Relative to the baseline, TextBO-GEPA delivers larger gains than GEPA across all datasets.
+On HotpotQA, the improvement rises from +1.00 to +3.00 points, which is a $200\%$ increase in the baseline-relative gain.
+On Hover, the gain increases from +5.00 to +7.67 points $(+53.4\%)$.
+On PUPA, it grows from +10.28 to +15.95 points $(+55.2\%)$.
+Overall, the aggregate improvement increases from +5.43 to +8.87 points, corresponding to a $63.4\%$ increase in baseline-relative improvement.
+Considering that language model evaluation benchmarks often have standard deviations well below 1 point, and 95% CI sizes often around 1–2 points (madaan2024quantifying),these results illustrate that TextBO-GEPA consistently significantly outperforms GEPA across the agentic AI benchmarks.
+Model HotpotQA Hover PUPA Aggregate Improvement
+Qwen3-8B
+Baseline 56.67 41.33 66.58 54.86
+GEPA 57.67 46.33 76.86 60.29 +5.43
+TextBO-GEPA 59.67 49.00 82.53 63.73 +8.87
+Table 2: Results on HotpotQA, Hover, and PUPA for Qwen3-8B after  7,500 evaluations. Aggregate is the mean across the three datasets; Improvement is vs. Baseline aggregate.
 10 Conclusion
-The importance of developing evaluation-efficient self-improving AI systems is substantial for its societal applications, yet this problem has remained unidentified and unaddressed.
-Motivated by the effectiveness of prompt optimization, which outperforms reinforcement fine-tuning methods, we formulate this problem as an iterative prompt optimization and address it.
-We propose T-BoN BO, a simple, evaluation-efficient prompt optimization framework that leverages textual edits and Best-of-N search to guide prompt improvement.
-We prove that T-BoN BO BO statistically emulates UCB-BO, which is a provably evaluation-efficient algorithm.
-Experiments simulating ad alignment across eight creative-ad tasks demonstrate that T-BoN BO outperforms strong state-of-the-art baselines, including large Best-of-N and GEPA.
-We validate that these improvements are not attributed to the usage of specific persona dataset or trivial improvements from using poor initial prompts.
-Our paper makes three key contributions to the computer science and marketing literature.
-First, it identifies evaluation efficiency as an essential objective for a societal self-improving system design.
-Second, it presents a simple yet provably efficient AI self-improving system design framework that can be applied to real-world societal systems, e.g., ad agencies, that incur high evaluation costs.
-Third, we theoretically prove that the presented framework is evaluation-efficient.
-In sum, this work bridges classical Bayesian optimization and prompt-based self-improving AI systems, demonstrating that evaluation-efficient optimization in the prompt space is both theoretically tractable and practically impactful.
-As organizations increasingly rely on AI-generated content, methods like T-BoN BO can help managers achieve higher-quality outputs without proportionally increasing evaluation burdens.
-Appendix A Prompts used for digital marketing
-Best-of-N Gradient steps’ pairwise tournament prompt for ads. contents = [image1, image2, comparison_\_prompt]
-comparison_\_prompt =
-“You are evaluating two advertisement images for mobile Instagram ads. Which image would be more effective at engaging users and driving clicks?
-RESPOND WITH ONLY THE NUMBER 1 OR 2. NOTHING ELSE.
-1 = first image is better
-2 = second image is better ”
-Figure 11: $M_{\text{critic}}$’s pairwise tournament prompt for digital marketing example.
-Meta reflection prompt for ads You are an expert at analyzing visual patterns in advertising performance. VISUAL ANALYSIS TASK:
-I will show you images from the lowest-scoring and highest-scoring ad iterations. Your task is to identify specific visual patterns that distinguish effective from ineffective ads.
-VISUAL EXAMPLES – BEST VS WORST PERFORMING:
-RANK 1/10 WORST PERFORMING (Score: 2.14/5.0):
-Prompt excerpt: <prompt>…
-[Image]
-RANK 2/10 LOWER HALF (Score: 2.52/5.0):
-Prompt excerpt: <prompt>…
-[Image]
-$\ldots$ RANK 9/10 UPPER HALF (Score: 3.81/5.0):
-Prompt excerpt: <prompt>…
-[Image]
-RANK 10/10 BEST PERFORMING (Score: 4.15/5.0):
-Prompt excerpt: <prompt>…
-[Image]
-Based on your visual analysis, identify patterns that correlate with higher effectiveness scores: 1.
-Visual composition and framing differences 2.
-Lighting conditions and mood variations 3.
-Color palettes and visual tone patterns 4.
-Subject positioning and action effectiveness 5.
-Brand integration approaches 6.
-Environmental and atmospheric elements RESPONSE FORMAT:
-Provide a structured analysis of visual patterns observed, focusing on what distinguishes high-performing from low-performing ads.
-Figure 12: Meta reflection prompt for digital marketing example.
-Prompt used for generating prompts for ads generation Generate a creative, structured and descriptive prompt for a generative AI model (e.g., Imagen) that will produce a brand-aligned, and scroll-stopping advertisement image suitable for an Instagram feed.
+Evaluation-efficient self-improving AI is increasingly important in societal and business settings where real-world feedback (e.g., human responses, field experiments) is slow and costly, even as candidate generation and analysis have become cheap.
+In this paper, we address this gap by proposing TextBO, a simple evaluation-efficient prompt-optimization framework that combines TextGrad-style local textual updates with Best-of-N gradient selection.
+Our main theoretical result shows that Best-of-N over locally sampled textual edits induces, in probability, ascent directions of a UCB-style acquisition function.
+This provides a principled bridge between language-space self-improvement and classical UCB Bayesian optimization: TextBO emulates parallel gradient-based UCB-BO in the implicit embedding space induced by the LLM, while operating purely in language space and without constructing an explicit surrogate or calibrated uncertainty model.
+Empirically, we find that TextBO improves evaluation efficiency in ad-creative prompt optimization, and our ablations indicate that the gains are not driven by a particular persona dataset or by starting from unusually weak prompts.
+We additionally find that the same Best‑of‑N multi‑step mechanism strengthens GEPA in agentic benchmark settings: TextBO‑GEPA exhibits larger improvements than GEPA across evaluation steps on agentic AI benchmarks.
+Overall, this work makes four contributions to literature on AI, optimization, and marketing.
+First, we elevate evaluation efficiency as a first-order design objective for self-improving AI deployed in real-world decisions.
+Second, we introduce TextBO, a simple and practical language-space optimization procedure that combines textual gradients with Best-of-N selection.
+Third, we provide theory showing that Best-of-N textual-gradient steps implement optimism-driven UCB-style exploration in an implicit embedding space, yielding evaluation-efficiency guarantees.
+Fourth, we demonstrate empirical gains relative to strong baselines in both the digital advertising case study and the agentic AI benchmarks.
+From a managerial perspective, evaluation-efficient self-improving AI can sharply reduce the marginal cost of iteration in data-driven decision making and enable even small businesses and resource constrained organizations to adopt AI tools (in ad optimization, promotion design, and other marketing activities).
+As generation and analysis become automated, our framework offers firms the tools and ability to sustain continuous optimization loops that adapt campaigns to evolving market conditions and consumer responses.
+Appendix A GP-UCB and Its Evaluation Efficiency
+Setup.
+Let $f$ be a unknown, black-box function that maps a compact and convex feasible region $\mathcal{D}\subset\mathbb{R}^{d}$ to $\mathbb{R}$.
+At each time $t$, a sampling algorithm $\mathcal{A}$ selects a point $x_{t}\in\mathcal{D}$ to evaluate.
+This selection is adaptive, based on the history of previously chosen points and their observed outcomes.
+The evaluation at $x_{t}$ yields a noisy observation $y_{t}=f\left(x_{t}\right)+\epsilon_{t}$, where $\epsilon_{t}$ is a random noise term, where $\left\{\epsilon_{t}:t=1,2,\ldots\right\}$ are independent, sub-Gaussian random variables.
+The unknown function $f$ is assumed to be an element of a Reproducing Kernel Hilbert Space (RKHS), $\mathcal{N}_{\Psi}(\mathcal{D})$, induced by a stationary kernel $\Psi$.
+Functions within this RKHS possess a specific smoothness property, the nature of which is determined by the kernel used.
+The space $\mathcal{N}_{\Psi}(\mathcal{D})$ is endowed with an inner product and is formally defined as the set of functions
+[…]
+where the norm is given by: $\|g\|_{\mathcal{N}_{\Psi}(\mathcal{D})}:=\left(\sum_{j,l=1}^{\infty}c_{j}c_{l}\Psi\left(\mathbf{x}_{j}-\mathbf{x}_{l}\right)\right)^{\frac{1}{2}}$.
+For any two functions $g_{1}(\mathbf{x})=\sum_{j=1}^{\infty}a_{j}\Psi\left(\mathbf{x}-\mathbf{x}_{j}\right)$ and $g_{2}(\mathbf{x})=\sum_{l=1}^{\infty}b_{l}\Psi\left(\mathbf{x}-\mathbf{x}_{l}\right)$ in the space, their inner product is $\left\langle g_{1},g_{2}\right\rangle_{\mathcal{N}_{\Psi}(\mathcal{D})}:=\sum_{j,l=1}^{\infty}a_{j}b_{l}\Psi\left(\mathbf{x}_{j}-\mathbf{x}_{l}\right)$.
+A key feature of RKHS space is the reproducing property, $\langle g,\Psi(\mathbf{x}-\cdot)\rangle_{\mathcal{N}_{\Psi}(\mathcal{D})}=g(\mathbf{x})$, which holds for all $\mathbf{x}\in\mathcal{D}$.
+The norm of the function $f$ in this space is assumed to be bounded such that $\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B$ for some constant $B>0$.
+The objectives.
+The performance of $\mathcal{A}$ is often evaluated through two types of regret over a time horizon $T$. One is cumulative regret, defined as
+$\mathcal{R}_{C}(T;f,\mathcal{A}):=\sum_{t=1}^{T}\left[\max_{\mathbf{x}\in\mathcal{D}}f(\mathbf{x})-f\left(\mathbf{x}_{t}\right)\right]$ (A1)
+The other is simple regret, defined as
+$\mathcal{R}_{S}(T;f,\mathcal{A}):=\max_{\mathbf{x}\in\mathcal{D}}f(\mathbf{x})-f\left(\mathbf{x}^{(T)}\right)$ (A2)
+where $\mathbf{x}^{(T)}$ is the output of the algorithm $\mathcal{A}$ after taking $T$ function evaluations.
+The lower bound.
+scarlett2017lower showed that a lower bound exists on the regret of any algorithm for $f\in\mathcal{N}_{\Psi}(\mathcal{D})$. Specifically, for any constant $B>0$,
+$\inf_{\mathcal{A}}\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}\mathbb{E}\left[\mathcal{R}_{C}(T;f,\mathcal{A})\right]=\begin{cases}\Omega(T^{\frac{\nu+d}{2\nu+d}}),&\text{ for Matérn kernels, }\\ \Omega(T^{\frac{1}{2}}\ln^{\frac{d}{2}}(T)),&\text{ for SE kernels.
+}\end{cases}$ (A3)
+$\inf_{\mathcal{A}}\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}T\left(\epsilon,f,\mathcal{A}\right)=\begin{cases}\Omega\left(\left(\frac{1}{\epsilon}\right)^{2+d/\nu}\right),&\text{ for Matérn kernels, }\\ \Omega\left(\frac{1}{\epsilon^{2}}\left(\log\frac{1}{\epsilon}\right)^{d/2}\right),&\text{ for SE kernels.
+}\end{cases}$ (A4)
+where Matérn kernels are defined as
+$\Psi_{\mathrm{M}}\left(\mathbf{x}-\mathbf{x}^{\prime}\right):=\frac{1}{\Gamma(\nu)2^{\nu-1}}\left(\frac{2\sqrt{\nu}\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}}{\ell}\right)^{\nu}K_{\nu}\left(\frac{2\sqrt{\nu}\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}}{\ell}\right),\quad\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$ (A5)
+and SE kernels are defined as
+$\Psi_{\mathrm{SE}}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)=\exp\left(-\frac{\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}^{2}}{2\ell^{2}}\right),\quad\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$ (A6)
+where $\nu>0$ is the smoothness parameter, $\ell>0$ is the length-scale parameter, $\Gamma(\cdot)$ is the gamma function, $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind of order $\nu$, and $\|\cdot\|_{2}$ denotes the Euclidean norm.
+GP-UCB algorithm.
+The Gaussian Process Upper Confidence Bound (GP-UCB) algorithm, introduced in the seminal work of srinivas2012information, operates from a Bayesian perspective by placing a Gaussian Process (GP) prior on the unknown objective function $f$.
+This GP is characterized by a zero-mean function and a symmetric, positive-definite covariance function, or kernel, $k:\mathcal{D}\times\mathcal{D}\mapsto\mathbb{R}$, where $k\left(\mathbf{x},\mathbf{x}^{\prime}\right)=\Psi\left(\mathbf{x}-\mathbf{x}^{\prime}\right)$ for all $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$.
+For any finite set of points $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{t}\}\subset\mathcal{D}$, the corresponding function values $(f(\mathbf{x}_{1}),\ldots,f(\mathbf{x}_{t}))^{\top}$ are assumed to follow a multivariate normal distribution with a zero mean vector and a covariance matrix with entries $\mathbf{K}_{t}]_{jl}=k(\mathbf{x}_{j},\mathbf{x}_{l})$.
+Under the assumption of independent, zero-mean Gaussian observation noise with variance $\sigma^{2}$, conditioning the GP prior on a set of observations $\mathbf{D}_{t}:=\{(\mathbf{x}_{j},y_{j}):j=1,\ldots,t\}$ yields a closed-form posterior distribution.
+The posterior mean and variance at any point $\mathbf{x}\in\mathcal{D}$ are given by:
+$\mathbb{E}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]=\mathbf{k}_{t}^{\top}(\mathbf{x})\left(\mathbf{K}_{t}+\sigma^{2}\mathbf{I}_{t}\right)^{-1}\mathbf{y}_{t},\operatorname{Var}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]=k(\mathbf{x},\mathbf{x})-\mathbf{k}_{t}(\mathbf{x})^{\top}\left(\mathbf{K}_{t}+\sigma^{2}\mathbf{I}_{t}\right)^{-1}\mathbf{k}_{t}(\mathbf{x}),$ (A7)
+where $\mathbf{k}_{t}(\mathbf{x})$ is the vector of covariances between $\mathbf{x}$ and the observed points, $\mathbf{K}_{t}$ is the covariance matrix of the observed points, $\mathbf{y}_{t}$ is the vector of observed values, and $\mathbf{I}_{t}$ is the identity matrix.
+Inspired by upper confidence bound methods in the multi-armed bandit literature (Auer et al., 2002), GP-UCB employs an optimistic acquisition strategy to select subsequent evaluation points.
+At each step $t$, the next point $\mathbf{x}_{t+1}$ is chosen by maximizing an upper confidence bound on the function’s value $\mathbf{x}_{t+1}=\underset{\mathbf{x}\in\mathcal{D}}{\arg\max}\left(\mathbb{E}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]+\sqrt{\beta_{t}\operatorname{Var}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]}\right).$ This acquisition function naturally balances exploitation, driven by the posterior mean $\mathbb{E}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]$, with exploration, driven by the posterior standard deviation $\sqrt{\operatorname{Var}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]}$.
+The tunable parameter $\beta_{t}>0$ explicitly governs this trade-off, making its specification critical to the algorithm’s performance.
+The evaluation efficiency of GP-UCB.
+whitehouse2023sublinear, wang2023regret showed that GP-UCB srinivas2012information regret upper bound for both expected cumulative regret and simple regret almost tightly matches the lower bound provided by scarlett2017lower.
+Specifically, for any constant $B>0$, the GP-UCB algorithm $\mathcal{A}_{\text{GPUCB}}$ achieves
+$\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}\mathbb{E}\left[\mathcal{R}_{C}(T;f,\mathcal{A}_{\text{GPUCB}})\right]=\begin{cases}O(T^{\frac{\nu+d}{2\nu+d}}),&\text{ for Matérn kernels, }\\ O(T^{\frac{1}{2}}\ln^{\frac{d}{2}}(T))&\text{ for SE kernels.
+}\end{cases}$ (A8)
+$\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}T(\epsilon,f,\mathcal{A}_{\mathrm{GPUCB}})=\begin{cases}O\left(\left(\frac{1}{\epsilon}\right)^{2+d/\nu}\right),&\text{for Matérn kernels,}\\ O\left(\frac{1}{\epsilon^{2}}\cdot\left(\log\frac{1}{\epsilon}\right)^{d+3}\right),&\text{for SE kernels.}\end{cases}$ (A9)
+wang2023regret’s approach to proving the regret optimality of GP-UCB follows the two-step framework.
+The first critical component is to construct a high-probability uniform error bound that quantifies the difference between the true objective function $f(x)$ and its posterior mean estimate $\mu_{t}(x)$ at any given step $t$.
+This bound takes the form $|f(x)-\mu_{t}(x)|\leq\sqrt{\beta_{t}}\sigma_{t}(x)$ for all $x\in\mathcal{D}$, where $\sigma_{t}(x)$ is the posterior standard deviation and $\beta_{t}$ is a carefully chosen exploration parameter.
+The second component involves bounding the cumulative sum of the instantaneous regrets, which, under the uniform error bound, can be related to the sum of the posterior variances $\sum_{t=1}^{T}\sigma_{t-1}^{2}(x_{t})$.
+This sum is, in turn, bounded by the maximal information gain, $\gamma_{T}$.
+The final regret bound is thus a function of $T$, $\beta_{T}$, and an upper bound on $\gamma_{T}$.
+wang2023regret employs tools from empirical process theory and decomposes the estimation error, $f(x)-\mu_{t}(x)$, into a bias term and a random error term.
+By leveraging the connection between Gaussian process regression and kernel ridge regression, along with properties of the Fourier transform for stationary kernels, they show that the bias term is bounded by $f||_{\mathcal{N}_{\Psi}(\mathcal{D})}\sigma_{t}(x)$.
+The more challenging random error term is handled by viewing it as an empirical process indexed by a class of functions and bounding its supremum.
+They bound the $\epsilon$-entropy of this function class, which allows for a high-probability bound on the random error.
+With the new uniform error bound, we can select a much smaller, dimension-independent exploration parameter $\beta_{t}$ that grows only logarithmically with $T$.
+Combining this sharper $\beta_{t}$ with the tightest known bounds on the maximal information gain $\gamma_{T}$ for Matérn and SE kernels, we can achieve the tight cumulative regret of GP-UCB.
+Appendix B Proofs
+B.1 Auxiliary lemmas
+Lemma 2 (Uniform first-order expansion).
+Let $g=\nabla\mu(e)$ and $h=\nabla\sigma(e)$. There exist $C>0$ and $\varepsilon_{0}>0$ such that for all $\varepsilon\in(0,\varepsilon_{0}]$ and all $u\in\mathbb{R}^{d}$,
+$\mu(e+\varepsilon u)$ $=\mu(e)+\varepsilon\,g^{\top}u\;+\;r_{\mu}(\varepsilon,u),$ $|r_{\mu}(\varepsilon,u)|$ $\leq C\varepsilon^{2},$ (A10)
+$\sigma(e+\varepsilon u)$ $=\sigma(e)+\varepsilon\,h^{\top}u\;+\;r_{\sigma}(\varepsilon,u),$ $|r_{\sigma}(\varepsilon,u)|$ $\leq C\varepsilon^{2}.$ (A11)
+Proof.
+By the mean-value theorem with Lipschitz gradients, for some $\xi$ between $e$ and $e+\varepsilon u$, $\mu(e+\varepsilon u)=\mu(e)+\nabla\mu(\xi)^{\top}(\varepsilon u).$ Hence $\mu(e+\varepsilon u)=\mu(e)+\varepsilon g^{\top}u+\varepsilon(\nabla\mu(\xi)-\nabla\mu(e))^{\top}u,$ and $|\nabla\mu(\xi)-\nabla\mu(e)|\leq L_{\mu}\|\xi-e\|\leq L_{\mu}\varepsilon$, giving $|r_{\mu}|\leq L_{\mu}\varepsilon^{2}$.
+The proof for $\sigma$ is identical with $L_{\sigma}$.
+Taking $C:=\max\{L_{\mu},L_{\sigma}\}$ yields the stated bounds with the same $C$.
+∎
+Lemma 3 (Max-stability under bounded perturbations).
+Let $(f_{i})_{i\leq N}$ and $(g_{i})_{i\leq N}$ be real arrays. Let $i_{f}\in\arg\max_{i}f_{i}$ and $i_{g}\in\arg\max_{i}g_{i}$. If for some $\Delta\geq 0$,
+$g_{i_{f}}\geq f_{i_{f}}-\Delta\quad\text{and}\quad f_{i_{g}}\geq g_{i_{g}}-\Delta,$ (A12)
+then $g_{i_{f}}\geq\max_{i}g_{i}-\Delta$ and $f_{i_{g}}\geq\max_{i}f_{i}-\Delta$.
+In particular, $g_{i_{f}}\geq\max_{i}g_{i}-\Delta$ implies that any maximizer of $f$ is a $\Delta$-approximate maximizer of $g$.
+Proof.
+Since $g_{i_{f}}\geq f_{i_{f}}-\Delta=\max_{i}f_{i}-\Delta\geq g_{i_{g}}-\Delta=\max_{i}g_{i}-\Delta$, the first claim holds. The second is symmetric. ∎
+Lemma 4 (Spherical cap coverage).
+For $v\in\mathbb{S}^{d-1}$ and $\eta\in(0,1)$, define $C(v,\eta):=\{u\in\mathbb{S}^{d-1}:v^{\top}u\geq 1-\eta\}$.
+Let $U_{1},\ldots,U_{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\rho$ as above.
+Then $\max_{i\leq N}v^{\top}U_{i}\to 1$ almost surely as $N\to\infty$.
+Moreover, for any deterministic sequence $\eta_{N}\downarrow 0$ with $N\cdot\rho\{u:v^{\top}u\geq 1-\eta_{N}\}\to\infty$, we have
+$\mathbb{P}\!\left(\max_{i\leq N}v^{\top}U_{i}\geq 1-\eta_{N}\right)\to 1.$ (A13)
+Proof.
+Let $C(\eta)=\{u\in\mathbb{S}^{d-1}:v^{\top}u\geq 1-\eta\}$ be a spherical cap.
+Since $\rho$ has a density bounded below, $\rho(C(\eta))\asymp\eta^{\frac{d-1}{2}}$ as $\eta\downarrow 0$.
+Then $\mathbb{P}(\max_{i}v^{\top}U_{i}<1-\eta)=(1-\rho(C(\eta)))^{N}\leq\exp(-N\rho(C(\eta))).$ For any fixed $\eta\in(0,1)$, $\sum_{N\geq 1}\mathbb{P}(\max_{i}v^{\top}U_{i}<1-\eta)\leq\sum_{N\geq 1}e^{-N\rho(C(\eta))}<\infty$, so by Borel–Cantelli, $\mathbb{P}(\max_{i}v^{\top}U_{i}<1-\eta\ \text{i.o.})=0$.
+Intersecting over a countable sequence $\eta_{m}\downarrow 0$ yields $\max_{i\leq N}v^{\top}U_{i}\to 1$ almost surely.
+Choosing $\eta_{N}\downarrow 0$ so that $N\rho(C(\eta_{N}))\to\infty$ yields the result, and taking $\eta=\eta_{N}\to 0$ proves the first statement.
+∎
+Lemma 5 (Near-maximum coupling in a thin band).
+Let $M_{N}=\max_{i}\xi_{i}$ and define $v_{N}:=(g+M_{N}h)/\|g+M_{N}h\|$.
+Also, $S_{i}:=a_{i}+b_{i}M_{N}$, $Y_{i}^{(0)}:=a_{i}+b_{i}\xi_{i}$, $b_{i}:=\sigma(e)+\varepsilon h^{\top}U_{i}$.
+Fix any sequence $c_{N}\uparrow\infty$ with $c_{N}/q_{N}\to 0$ and set $\delta_{N}:=c_{N}/q_{N}$.
+Then with probability $\to 1$ there exists an index $j$ such that
+$\xi_{j}\geq M_{N}-\delta_{N}\quad\text{and}\quad v_{N}^{\top}U_{j}\geq 1-\eta_{N},$ (A14)
+whenever $\eta_{N}\downarrow 0$ satisfies
+$N\,\mathbb{P}(\xi\geq M_{N}-\delta_{N})\,\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}\to\infty.$ (A15)
+Moreover, for this $j$,
+$S_{j}-Y_{j}^{(0)}=b_{j}(M_{N}-\xi_{j})\leq b_{\max}\,\delta_{N}.$ (A16)
+Proof.
+Define $p_{\mathrm{cap},N}:=\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}$ and the event
+$E_{N}:=\{M_{N}\geq q_{N}-\delta_{N}\}.$ (A17)
+By Lemma 1, $M_{N}-q_{N}\to 0$ in probability and $\delta_{N}\to 0$, hence $\mathbb{P}(E_{N})\to 1$.
+Set the deterministic threshold $t_{N}:=q_{N}-2\delta_{N}$. On $E_{N}$ we have $M_{N}-\delta_{N}\geq q_{N}-2\delta_{N}=t_{N}$, hence by monotonicity of the tail,
+$\mathbb{P}(\xi\geq t_{N})\;\geq\;\mathbb{P}(\xi\geq M_{N}-\delta_{N})\quad\text{on }E_{N}.$ (A18)
+Let
+$K_{N}:=\sum_{i=1}^{N}\mathbf{1}\{\xi_{i}\geq t_{N}\}\sim\mathrm{Binomial}\!\big(N,\,p_{N}\big),\qquad p_{N}:=\mathbb{P}(\xi\geq t_{N}).$ (A19)
+A Chernoff bound gives
+$\mathbb{P}\!\left(K_{N}<\tfrac{1}{2}Np_{N}\right)\;\leq\;\exp\!\left(-\tfrac{1}{8}Np_{N}\right).$ (A20)
+Consider
+$L_{N}\;:=\;\sum_{i=1}^{N}\mathbf{1}\{\xi_{i}\geq t_{N}\}\,\mathbf{1}\{v_{N}^{\top}U_{i}\geq 1-\eta_{N}\}.$ (A21)
+[…]
+$L_{N}\,\big|\,\{\xi_{i}\}\;\sim\;\mathrm{Binomial}\!\big(K_{N},\,p_{\mathrm{cap},N}\big),\qquad\mathbb{P}\big(L_{N}=0\,\big|\,\{\xi_{i}\}\big)\;=\;(1-p_{\mathrm{cap},N})^{K_{N}}\;\leq\;\exp\!\big(-K_{N}p_{\mathrm{cap},N}\big).$ (A22)
+We now bound $\mathbb{P}(L_{N}=0)$ by splitting on $\{K_{N}\geq\tfrac{1}{2}Np_{N}\}$ and $E_{N}$:
+$\mathbb{P}(L_{N}=0)$ $\leq\mathbb{E}\!\left[\exp\!\big(-K_{N}p_{\mathrm{cap},N}\big)\,\mathbf{1}_{\{K_{N}\geq\frac{1}{2}Np_{N}\}\cap E_{N}}\right]\;+\;\mathbb{P}\!\left(K_{N}<\tfrac{1}{2}Np_{N}\right)\;+\;\mathbb{P}(E_{N}^{c})$
+$\leq\mathbb{E}\!\left[\exp\!\big(-\tfrac{1}{2}Np_{N}\,p_{\mathrm{cap},N}\big)\,\mathbf{1}_{E_{N}}\right]\;+\;\exp\!\big(-\tfrac{1}{8}Np_{N}\big)\;+\;o(1)$
+$\leq\mathbb{E}\!\left[\exp\!\big(-\tfrac{1}{2}Np_{N}\,p_{\mathrm{cap},N}\big)\right]\;+\;\exp\!\big(-\tfrac{1}{8}Np_{N}\big)\;+\;o(1).$ (A23)
+On $E_{N}$, (A18) yields $p_{N}\geq\mathbb{P}(\xi\geq M_{N}-\delta_{N})$. By the lemma’s hypothesis,
+$Z_{N}:=Np_{N}\,p_{\mathrm{cap},N}\;\geq\;N\,\mathbb{P}(\xi\geq M_{N}-\delta_{N})\,p_{\mathrm{cap},N}\;\xrightarrow{p}\;\infty.$
+Since $0\leq e^{-\frac{1}{2}Z_{N}}\leq 1$, for any fixed $T>0$,
+$\mathbb{E}\!\left[e^{-\frac{1}{2}Z_{N}}\right]\leq e^{-\frac{1}{2}T}+\mathbb{P}(Z_{N}\leq T)\;\longrightarrow\;e^{-\frac{1}{2}T}.$
+Letting $T\to\infty$ gives $\mathbb{E}[e^{-\frac{1}{2}Z_{N}}]\to 0$.
+Moreover, $Z_{N}\leq Np_{N}$ since $p_{\mathrm{cap},N}\leq 1$.
+If $Np_{N}$ were bounded along a subsequence, then $Z_{N}$ would be bounded along that subsequence, contradicting $Z_{N}\xrightarrow{p}\infty$.
+Hence $Np_{N}\to\infty$, so $\exp(-\tfrac{1}{8}Np_{N})\to 0$.
+Therefore the second term in (A23) vanishes, and $\mathbb{P}(L_{N}=0)\to 0$.
+Therefore, with probability tending to one there exists $j$ such that $\xi_{j}\geq t_{N}$ and $v_{N}^{\top}U_{j}\geq 1-\eta_{N}$.
+Since $t_{N}\leq M_{N}-\delta_{N}$ on $E_{N}$, this $j$ also satisfies $\xi_{j}\geq M_{N}-\delta_{N}$.
+Finally, $S_{j}-Y_{j}^{(0)}=b_{j}(M_{N}-\xi_{j})\leq b_{\max}\delta_{N}$ holds deterministically by $M_{N}-\xi_{j}\leq\delta_{N}$ and the definition of $b_{\max}$.
+∎
+B.2 Proof of Theorem 1
+Proof.
+By Lemma 2,
+$Y_{i}$ $=\mu(e)+\varepsilon g^{\top}u_{i}+r_{\mu}(\varepsilon,u_{i})+\big(\sigma(e)+\varepsilon h^{\top}u_{i}+r_{\sigma}(\varepsilon,u_{i})\big)\xi_{i}$
+$=a_{i}+b_{i}\xi_{i}+R_{i},$ (A24)
+where $a_{i}:=\mu(e)+\varepsilon g^{\top}u_{i}$, $b_{i}:=\sigma(e)+\varepsilon h^{\top}u_{i}$, and $R_{i}:=r_{\mu}(\varepsilon,u_{i})+\xi_{i}r_{\sigma}(\varepsilon,u_{i}).$ Then $|R_{i}|\leq C\varepsilon^{2}(1+|\xi_{i}|)$ for all $i$.
+Define $S_{i}:=a_{i}+b_{i}M_{N}$, $b_{\max}:=\max_{1\leq i\leq N}|b_{i}|$, and $U_{i}:=u_{i}$.
+Now let $Y_{i}^{(0)}:=a_{i}+b_{i}\xi_{i}$ and $i_{0}\in\arg\max_{i}Y_{i}^{(0)}$. From (A24),
+$Y_{i^{\star}}\geq Y_{i_{0}}\implies Y_{i^{\star}}^{(0)}\geq Y_{i_{0}}^{(0)}-|R_{i^{\star}}|-|R_{i_{0}}|.$
+By the bound on $R_{i}$ and $|\xi_{i}|\leq M_{N}:=\max_{j}\xi_{j}$,
+$Y_{i^{\star}}^{(0)}\geq\max_{i}Y_{i}^{(0)}-2C\varepsilon^{2}(1+M_{N}).$ (A25)
+Symmetrically, $\max_{i}Y_{i}^{(0)}\geq Y_{i^{\star}}^{(0)}\geq\max_{i}Y_{i}^{(0)}-2C\varepsilon^{2}(1+M_{N})$.
+By Lemma 3, $i^{\star}$ is a $2C\varepsilon^{2}(1+M_{N})$-approximate maximizer of $Y^{(0)}$.
+Now let $j$ be the index from Lemma 5 so that $\xi_{j}\geq M_{N}-\delta_{N}$ and $v_{N}^{\top}U_{j}\geq 1-\eta_{N}$. By Lemma 2,
+$\max_{i\leq N}S_{i}-S_{j}\;\leq\;\varepsilon\|g+M_{N}h\|\,(1-v_{N}^{\top}U_{j})\;+\;2C\varepsilon^{2}(1+M_{N})\;\leq\;\varepsilon\|g+M_{N}h\|\,\eta_{N}\;+\;2C\varepsilon^{2}(1+M_{N}).$
+Also,
+$S_{j}-S_{i^{\star}}=(S_{j}-Y_{j}^{(0)})+(Y_{j}^{(0)}-Y_{i^{\star}}^{(0)})+(Y_{i^{\star}}^{(0)}-S_{i^{\star}})\;\leq\;b_{\max}\delta_{N}\;+\;2C\varepsilon^{2}(1+M_{N})\;+\;\underbrace{(Y_{i^{\star}}^{(0)}-S_{i^{\star}})}_{\leq 0},$
+where the last term is $\leq 0$ provided $b_{i}\geq 0$ for all $i$, which holds by Assumption A4 for $\varepsilon\leq\varepsilon_{0}:=\sigma(e)/(2\|h\|)$ when $\sigma(e)>0$ since $b_{i}=\sigma(e)+\varepsilon h^{\top}U_{i}\geq\sigma(e)-\varepsilon\|h\|\geq\tfrac{1}{2}\sigma(e)>0$.
+Therefore,
+$\max_{i\leq N}S_{i}-S_{i^{\star}}\;\leq\;\varepsilon\|g+M_{N}h\|\,\eta_{N}\;+\;b_{\max}\delta_{N}\;+\;4C\varepsilon^{2}(1+M_{N}).$ (A26)
+Set:
+$\Delta^{\prime}_{N}(\varepsilon)\;:=\;b_{\max}\,\frac{c_{N}}{q_{N}}\;+\;4C\varepsilon^{2}(1+M_{N}).$
+By Lemma 2,
+$S_{i}\;=\;\mu(e)+M_{N}\sigma(e)+\varepsilon(g+M_{N}h)^{\top}u_{i}+\tilde{r}_{i},\qquad|\tilde{r}_{i}|\leq C\varepsilon^{2}(1+M_{N}).$
+Hence for any $i$,
+$S_{i}-S_{i^{\star}}=\varepsilon\|g+M_{N}h\|\,(v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N})+(\tilde{r}_{i}-\tilde{r}_{i^{\star}}),$
+so
+$\big|\,(S_{i}-S_{i^{\star}})-\varepsilon\|g+M_{N}h\|\,(v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N})\,\big|\;\leq\;2C\varepsilon^{2}(1+M_{N}).$ (A27)
+Taking $i$ that attains $\max_{i}S_{i}$ and using (A26) yields
+$\max_{i\leq N}v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N}\;\leq\;\eta_{N}\;+\;\frac{\Delta^{\prime}_{N}(\varepsilon)}{\varepsilon\|g+M_{N}h\|}.$ (A28)
+Thus
+[…]
+Assume $\|h\|>0$ (the $\|h\|=0$ case is addressed in Remark 1). Using
+$\|g+M_{N}h\|\;\geq\;M_{N}\|h\|-\|g\|$
+and Lemma 1, $M_{N}\to\infty$ and $M_{N}\asymp q_{N}$. Also $b_{\max}\leq\sigma(e)+\varepsilon\|h\|+C\varepsilon^{2}$. Hence
+$\frac{\Delta^{\prime}_{N}(\varepsilon)}{\varepsilon\|g+M_{N}h\|}=O_{p}\!\Big(\tfrac{c_{N}}{\varepsilon q_{N}^{2}\|h\|}\Big)+O\!\Big(\tfrac{\varepsilon}{\|h\|}\Big).$
+Choosing, e.g., $c_{N}=\tfrac{1}{2}\log\log N$ and any $\eta_{N}\downarrow 0$ with
+$N\,e^{c_{N}}\,\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}\to\infty$
+makes the RHS of (A29) equal to $o_{p}(1)+O(\varepsilon)$ for fixed $\varepsilon>0$; then let $\varepsilon\downarrow 0$.
+With $N\to\infty$ first and fixed $\varepsilon>0$, (A29) implies $1-v_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}0$.
+Then $\varepsilon\downarrow 0$ removes the $O(\varepsilon)$ residual, yielding $v_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}1$.
+Since $M_{N}-q_{N}\to 0$ in probability by Lemma 1, the unit vectors $v_{N}=(g+M_{N}h)/\|g+M_{N}h\|$ and $\tilde{v}_{N}:=(g+q_{N}h)/\|g+q_{N}h\|$ satisfy $\|v_{N}-\tilde{v}_{N}\|\to 0$ in probability.
+Hence $\widehat{u}_{N}\xrightarrow{p}\tilde{v}_{N}$, which is the claimed gradient direction of $A_{\beta_{N}}$ with $\beta_{N}=q_{N}$.
+By Lemma 2,
+$A_{\beta_{N}}(e+\Delta e^{(N)})=A_{\beta_{N}}(e)+\varepsilon\nabla A_{\beta_{N}}(e)^{\top}\widehat{u}_{N}+O(\varepsilon^{2}).$
+Write $\nabla A_{\beta_{N}}(e)=g+q_{N}h$.
+Using $\tilde{v}_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}1$ gives $\nabla A_{\beta_{N}}(e)^{\top}\widehat{u}_{N}=\|\nabla A_{\beta_{N}}(e)\|\,\tilde{v}_{N}^{\top}\widehat{u}_{N}=\|\nabla A_{\beta_{N}}(e)\|(1-o_{p}(1)).$ Thus $A_{\beta_{N}}(e+\Delta e^{(N)})\geq A_{\beta_{N}}(e)+\varepsilon\|\nabla A_{\beta_{N}}(e)\|-o_{p}(\varepsilon),$ as claimed.
+∎
+Remark 1 (Edge cases).
+If $\|h\|=0$, then $\nabla A_{\beta_{N}}(e)=g$ and the direction reduces to $\nabla\mu(e)$. The Bayesian Optimization setup does not allow $\sigma(e)=0$.
+Appendix C Appendix for Section 7
+C.1 Creative Briefs for the Scenarios
+Below are the creative briefs for the eight ad campaigns.
+Each creative brief outlines the strategic and creative direction for each scenario.
+The selected campaigns cover a diverse range of products across distinct categories.
+Each campaign addresses a unique consumer need—whether related to sustainability, luxury, technology, or lifestyle—ensuring comprehensive coverage of various market segments.
+From environmentally-conscious millennials to affluent, experience-driven travelers, the approach accounts for the full spectrum of consumer interests.
+By focusing on different consumer needs, the structure enables a holistic marketing strategy that captures both niche and broad demands, providing a well-rounded framework for the study.
+C.1.1 Scenario 1: GreenBite Plant-Based Burger
+  • $\bullet$
+    Product: "GreenBite," a new plant-based burger patty.
+  • $\bullet$
+    Background: The market for plant-based food is growing, but many consumers remain skeptical about taste, believing they must sacrifice flavor for health.
+  • $\bullet$
+    The Challenge: Convince flexitarians that GreenBite offers the juicy, satisfying experience of a traditional beef burger with no compromise.
+  • $\bullet$
+    Core Insight: Consumers want to eat better for themselves and the planet, but they fear it means giving up the foods they love.
+  • $\bullet$
+    Single-Minded Proposition (SMP): The delicious, no-compromise burger experience that’s better for you and the planet.
+  • $\bullet$
+    Reasons to Believe (RTB): Made with savory pea protein; sears and tastes like real beef; free of soy and GMOs.
+  • $\bullet$
+    Desired Response: Think: "I can finally have a plant-based burger that tastes like the real thing.", Feel: Satisfied, vibrant, and guilt-free. Do: Purchase GreenBite for their next barbecue.
+  • $\bullet$
+    Brand Personality & Tone: Positive, confident, and mouth-watering.
+C.1.2 Scenario 2: AuraSonics X1 Earbuds
+  • $\bullet$
+    Product: "AuraSonics X1," high-end, noise-canceling wireless earbuds.
+  • $\bullet$
+    Background: The rise of hybrid work and urban density has increased daily auditory overload, making focus a precious commodity.
+  • $\bullet$
+    The Challenge: Cut through the saturated audio market by positioning the X1 not as a gadget, but as an essential tool for mental clarity.
+  • $\bullet$
+    Core Insight: In a world of constant noise, true luxury is the ability to control your own soundscape.
+  • $\bullet$
+    Single-Minded Proposition (SMP): AuraSonics X1 creates a personal sanctuary for focus and immersion.
+  • $\bullet$
+    Reasons to Believe (RTB): Adaptive Active Noise Cancellation; studio-quality audio; all-day battery life; minimalist aluminum design.
+  • $\bullet$
+    Desired Response: Think: "This is the solution I need to escape the daily chaos." Feel: Calm, empowered, and sophisticated. Do: Visit the website to learn more.
+  • $\bullet$
+    Brand Personality & Tone: Minimalist, intelligent, and calm.
+C.1.3 Scenario 3: Odyssey E-SUV
+  • $\bullet$
+    Product: The "Odyssey E-SUV," a new all-electric family SUV.
+  • $\bullet$
+    Background: Many families want to switch to electric vehicles but are concerned about sacrificing space, safety, and range for sustainability.
+  • $\bullet$
+    The Challenge: Position the Odyssey E-SUV as the first EV that meets all the practical and adventurous needs of a modern family, making the switch to electric feel like an upgrade, not a compromise.
+  • $\bullet$
+    Single-Minded Proposition (SMP): The future of family adventure is here: all-electric, zero compromise.
+  • $\bullet$
+    Reasons to Believe (RTB): 500km range; top-tier safety rating; spacious three-row seating; all-wheel drive capability.
+  • $\bullet$
+    Desired Response: Think: "An electric car can actually handle my family’s active lifestyle." Feel: Adventurous, optimistic, and secure. Do: Schedule a test drive.
+  • $\bullet$
+    Brand Personality & Tone: Inspiring, capable, and forward-thinking.
+C.1.4 Scenario 4: Oasis Eco-Lodge
+  • $\bullet$
+    Product: "Oasis Eco-Lodge," a secluded, luxury resort operating in harmony with its natural surroundings.
+  • $\bullet$
+    Background: The luxury travel market often equates opulence with excess. A growing segment of affluent travelers seeks experiences that are both exclusive and responsible.
+  • $\bullet$
+    The Challenge: Redefine luxury as a seamless integration with nature, promising an experience that is restorative for both the guest and the environment.
+  • $\bullet$
+    Core Insight: For those who have everything, true luxury is not more things, but a deeper connection to something pure and serene.
+  • $\bullet$
+    Single-Minded Proposition (SMP): Rediscover tranquility in a luxury that respects nature.
+  • $\bullet$
+    Reasons to Believe (RTB): Secluded private bungalows; farm-to-table dining with locally sourced ingredients; carbon-neutral operations.
+  • $\bullet$
+    Desired Response: Think: "This is a truly special escape, not just another five-star hotel." Feel: Serene, exclusive, and rejuvenated. Do: Book a stay.
+  • $\bullet$
+    Brand Personality & Tone: Elegant, peaceful, and understated.
+C.1.5 Scenario 5: Momentum Digital Banking
+  • $\bullet$
+    Product: "Momentum," a mobile-first banking app for freelancers and the gig economy.
+  • $\bullet$
+    Background: Traditional banks are not built for the fluctuating incomes and unique business needs of self-employed professionals.
+  • $\bullet$
+    The Challenge: Position Momentum as the essential financial co-pilot for the independent worker, simplifying complexity and providing stability.
+  • $\bullet$
+    Core Insight: Freelancers love their freedom but feel anxious about their financial instability. They need a tool that brings order to their financial chaos.
+  • $\bullet$
+    Single-Minded Proposition (SMP): Banking that is as flexible and entrepreneurial as you are.
+  • $\bullet$
+    Reasons to Believe (RTB): Automated tax-saving tools; integrated invoicing and payment tracking; instant business expense categorization.
+  • $\bullet$
+    Desired Response: Think: "Finally, a bank that gets the way I work." Feel: Empowered, organized, and financially confident. Do: Download the app and open an account.
+  • $\bullet$
+    Brand Personality & Tone: Modern, empowering, and simple.
+C.1.6 Scenario 6: MindGarden Meditation App
+  • $\bullet$
+    Product: "MindGarden," a subscription-based meditation and mindfulness app.
+  • $\bullet$
+    Background: While many are interested in mindfulness for stress-relief, they are often intimidated by the practice, seeing it as difficult or time-consuming.
+  • $\bullet$
+    The Challenge: Make meditation feel accessible and achievable for absolute beginners, removing the barriers of time and perceived difficulty.
+  • $\bullet$
+    Core Insight: People want the benefits of mindfulness but are convinced they don’t have the time or ability to practice it correctly.
+  • $\bullet$
+    Single-Minded Proposition (SMP): Find your calm in just 5 minutes a day.
+  • $\bullet$
+    Reasons to Believe (RTB): Guided 5-minute sessions for specific needs (anxiety, focus); progress tracking to build a habit; simple, jargon-free instructions.
+  • $\bullet$
+    Desired Response: Think: "I can do this. 5 minutes is easy." Feel: Calm, supported, and hopeful. Do: Start a free trial.
+  • $\bullet$
+    Brand Personality & Tone: Gentle, approachable, and encouraging.
+C.1.7 Scenario 7: Aeterno Watch
+  • $\bullet$
+    Product: The "Aeterno," a classic, automatic Swiss-made wristwatch with a heritage design.
+  • $\bullet$
+    Background: In an age of smartwatches that are obsolete in a few years, there is a renewed appreciation for objects with permanence and enduring value.
+  • $\bullet$
+    The Challenge: Reassert the relevance of the classic mechanical watch as a symbol of taste, craftsmanship, and timelessness.
+  • $\bullet$
+    Core Insight: In a disposable world, true status comes from owning something permanent that tells a story.
+  • $\bullet$
+    Single-Minded Proposition (SMP): A legacy on your wrist. Craftsmanship that transcends time.
+  • $\bullet$
+    Reasons to Believe (RTB): Swiss-made automatic movement; sapphire crystal glass; timeless design inspired by 1950s classics.
+  • $\bullet$
+    Desired Response: Think: "This is a beautiful object that I will own forever." Feel: Prestigious, sophisticated, and discerning. Do: Locate an authorized dealer.
+  • $\bullet$
+    Brand Personality & Tone: Elegant, timeless, and confident.
+C.1.8 Scenario 8: SyncFlow B2B Software
+  • $\bullet$
+    Product: "SyncFlow," a project management and collaboration software platform for remote teams.
+  • $\bullet$
+    Background: Remote work has increased flexibility but also created challenges in team alignment, communication, and project visibility.
+  • $\bullet$
+    The Challenge: Cut through a crowded SaaS market by focusing on the outcome of "effortless collaboration" rather than just listing features.
+  • $\bullet$
+    Core Insight: Managers don’t want another tool to manage; they want the feeling of clarity and momentum that comes from a team working in perfect sync.
+  • $\bullet$
+    Single-Minded Proposition (SMP): Bring your remote team together for effortless collaboration and remarkable results.
+  • $\bullet$
+    Reasons to Believe (RTB): Centralized dashboards with real-time project status; integrated communication channels; automated workflows and reporting.
+  • $\bullet$
+    Desired Response: Think: "This could finally solve our remote work chaos and get everyone on the same page." Feel: Organized, in control, and successful. Do: Sign up for a team demo.
+  • $\bullet$
+    Brand Personality & Tone: Professional, efficient, and innovative.
+C.2 Prompts for Initial Ad Generation
+Meta-prompt used for generating prompts for initial 64 ads generation CREATIVE BRIEF:
+{creative_\_brief}
+Based on this creative brief, generate a creative, structured, and descriptive prompt for a generative AI model (e.g., Imagen) that will produce a brand-aligned and scroll-stopping advertisement image suitable for an Instagram feed.
 A successful prompt must be constructed using the following components and principles:
@@
 - Text: If you are including a text message, prompt for "negative space for text overlay".
-Figure 13: Prompt for generating the prompts for the initial 64 ads.
-Prompt for generating the worst-of-64 in GreenBite (plant-based burger patty) ad scenario.
+Figure A1: Meta-prompt for generating the prompts for the initial 64 ads.
+Worst-of-64 prompt for GreenBite (plant-based burger patty) ad scenario.
 **Key Message:** To visually communicate the core idea: "Experience the juicy, no-compromise satisfaction of a real burger, made better for you and the planet."
@@
 The atmosphere is confident, positive, and irresistibly appetizing, making the viewer crave that exact experience.
-Prompt for generating the best-of-64 in GreenBite (plant-based burger patty) ad scenario.
+Best-of-64 prompt for GreenBite (plant-based burger patty) ad scenario.
 **1.
@@
 * **Negative Space:** Ensure the composition leaves a clean, uncluttered area of soft-focus background on the right third of the image for text overlay.
-Figure 14: Prompts for the best and worst ad among the initial 64 ads.
-Prompt for generating the T-BoN $(T=10)$ ad for GreenBite (plant-based burger patty) "**Key Message:**
+Figure A2: Prompts for the best and worst ad among the initial 64 ads generated using the meta-prompt in Figure A1.
+C.3 Prompts used for Critic LLM Operators
+Best-of-N Gradient steps’ pairwise tournament prompt for ads. contents = [image1, image2, comparison_\_prompt]
+comparison_\_prompt =
+“You are evaluating two advertisement images for mobile Instagram ads. Which image would be more effective at engaging users and driving clicks?
+RESPOND WITH ONLY THE NUMBER 1 OR 2. NOTHING ELSE.
+1 = first image is better
+2 = second image is better ”
+Figure A3: $M_{\text{critic}}$’s pairwise tournament prompt for digital marketing example.
+Meta reflection prompt for ads You are an expert at analyzing visual patterns in advertising performance. VISUAL ANALYSIS TASK:
+I will show you images from the lowest-scoring and highest-scoring ad iterations. Your task is to identify specific visual patterns that distinguish effective from ineffective ads.
+VISUAL EXAMPLES – BEST VS WORST PERFORMING:
+RANK 1/10 WORST PERFORMING (Score: 2.14/5.0):
+Prompt excerpt: <prompt>…
+[Image]
+RANK 2/10 LOWER HALF (Score: 2.52/5.0):
+Prompt excerpt: <prompt>…
+[Image]
+$\ldots$ RANK 9/10 UPPER HALF (Score: 3.81/5.0):
+Prompt excerpt: <prompt>…
+[Image]
+RANK 10/10 BEST PERFORMING (Score: 4.15/5.0):
+Prompt excerpt: <prompt>…
+[Image]
+Based on your visual analysis, identify patterns that correlate with higher effectiveness scores: 1.
+Visual composition and framing differences 2.
+Lighting conditions and mood variations 3.
+Color palettes and visual tone patterns 4.
+Subject positioning and action effectiveness 5.
+Brand integration approaches 6.
+Environmental and atmospheric elements RESPONSE FORMAT:
+Provide a structured analysis of visual patterns observed, focusing on what distinguishes high-performing from low-performing ads.
+Figure A4: Meta reflection prompt for digital marketing example.
+Textual gradient generating meta-prompt for ads (at $t>0$) You are an expert at optimizing image generation prompts for advertising effectiveness. CURRENT PROMPT TO IMPROVE:
+{current_prompt}
+PERFORMANCE ANALYSIS FROM PREVIOUS ITERATIONS:
+{performance_analysis}
+TASK: Based on the performance analysis above, generate specific, actionable improvements to make the current prompt more effective.
+Focus on implementing the successful visual patterns identified in the analysis while avoiding the ineffective elements.
+Provide 3–5 specific, implementable suggestions for improvement.
+(It can be addition of a new prompt part, deletion of existing prompt part, or rewriting a prompt part).
+Each suggestion should reference insights from the performance analysis.
+RESPONSE FORMAT: 1.
+[Specific improvement suggestion based on performance analysis] 2.
+[Specific improvement suggestion based on performance analysis] 3.
+$\ldots$ Be concrete and actionable.
+Focus on changes that will meaningfully improve ad effectiveness based on the analysis.
+Textual gradient generating meta-prompt for ads (at $t=0$) You are an expert at optimizing image generation prompts for advertising effectiveness. CURRENT PROMPT TO IMPROVE:
+{current_prompt}
+TASK: Generate specific, actionable improvements to make this ad prompt more effective.
+Focus on elements that will increase engagement and appeal to the target audience.
+Consider improvements in: 1.
+Visual composition and framing 2.
+Emotional appeal and messaging 3.
+Color palette and lighting 4.
+Subject positioning and action 5.
+Brand integration and logo placement 6.
+Text overlay space and readability Provide 3–5 specific, implementable suggestions for improvement.
+Each suggestion should be concrete and actionable.
+RESPONSE FORMAT: 1.
+[Specific improvement suggestion based on performance analysis] 2.
+[Specific improvement suggestion based on performance analysis] 3.
+$\ldots$ Be concise but specific.
+Focus on changes that will meaningfully improve ad effectiveness.
+Figure A5: Meta prompts for improving ad image-generation prompts (with and without performance analysis).
+Meta-prompt for applying textual gradients You are an expert at revising image generation prompts based on improvement suggestions. ORIGINAL PROMPT:
+{current_prompt}
+IMPROVEMENT SUGGESTIONS:
+{gradient}
+TASK: Rewrite the prompt incorporating the improvement suggestions while maintaining the core message and structure.
+REQUIREMENTS: Keep the prompt structure and format similar to the original Integrate the improvement suggestions naturally Maintain coherence and readability Ensure the prompt is optimized for image generation Keep the prompt length reasonable (not too long) OUTPUT CONSTRAINT:
+Return ONLY the revised prompt, no explanations or additional text.
+Figure A6: Meta-prompt for applying improvement suggestions to revise an image-generation ad prompt.
+C.4 Prompt for Evaluation Based on Persona
+Persona prompt for simulating ad effectiveness. SYSTEM: $\{$
+You are an AI assistant.
+Your task is to answer the TASK as if you are the individual described in the ‘Persona Profile’ (which contains their past survey responses).
+Remain consistent with the persona’s past survey responses and stated characteristics.
+Carefully follow any instructions provided for the new question, including formatting requirements.
+PERSONA DATA: $\{$
+Which part of the United States do you currently live in?
+Question Type: Single Choice
+Options:
+1 - Northeast (PA, NY, NJ, RI, CT, MA, VT, NH, ME)
+2 - Midwest (ND, SD, NE, KS, MN, IA, MO, WI, IL, MI, IN, OH)
+3 - South (TX, OK, AR, LA, KY, TN, MS, AL, WV, DC, MD, DE, VA, NC, SC, GA, FL)
+4 - West (WA, OR, ID, MT, WY, CA, NV, UT, CO, AZ, NM)
+5 - Pacific (HI, AK)
+Answer: 2 - Midwest (ND, SD, NE, KS, MN, IA, MO, WI, IL, MI, IN, OH)
+What is the highest level of schooling or degree that you have completed?
+Question Type: Single Choice
+Options:
+1 - Less than high school
+2 - High school graduate
+3 - Some college, no degree
+4 - Associate’s degree
+5 - College graduate/some postgrad
+6 - Postgraduate
+Answer: 3 - Some college, no degree
+$\ldots$ (Many other survey questions and answers) $\ldots$
+Suppose you were given \$5 and had to offer to another (anonymous) person a way to split the money.
+The other person can either accept or reject your offer.
+If the other person accepts your offer, you would each receive the amount you proposed.
+If the other person rejects your offer, you would both receive \$0.
+How much would you offer to the other person?
+Question Type: Single Choice
+Options:
+1 - \$0
+2 - \$1
+3 - \$2
+4 - \$3
+5 - \$4
+6 - \$5
+Answer: 3 - \$2
+$\ldots$ (Many other survey questions and answers) $\ldots$
+AD IMAGE: [image] TASK: Return only one item from ["1","2","3","4","5"] for ad effectiveness.
+Effective Score Scale Definition:
+1: Extremely Unlikely. The persona would actively ignore or be annoyed by this ad.
+2: Unlikely. The persona would likely scroll past without a second thought.
+3: Mediocre. It is hard to decide whether the personal would click or don’t click.
+4: Likely. The persona is intrigued and has a good chance of clicking to learn more.
+5: Extremely Likely. The persona is the ideal target; a click is almost certain.
+No explanation. Just the score.
+Figure A7: Prompt for simulating the effectiveness of a given ad-persona combination.
+Appendix D Appendix for Section 8
+D.1 Implementation Details of GEPA for digital ad optimization
+Implementation overview
+GEPA [agrawal2025gepa] is an iterative, prompt-optimization-based, self-improving AI method.
+The key idea of GEPA is to “combine reflection with evolution”.
+Here, an evolution is the same concept as taking a textual gradient in textgrad, and reflection is the same reflection TextBO (\$6) uses.
+GEPA iterates over three steps in each optimization iteration: Step 1) choosing a prompt and evaluating it; Step 2) updating the meta-reflection; and Step 3) improving the prompt.
+GEPA optimizes the prompt for a set of problem instances.
+In standard agentic benchmarks considered in agrawal2025gepa, a problem instance is a single task in a benchmark (together with whatever information is needed to evaluate it) that can be stored and revisited.
+In our digital ad setting, a problem instance corresponds to an individual ad viewer (a persona): each evaluation asks how effective a given ad is for a specific ad viewer (i.e., an ad–ad viewer combination), and the overall score for an ad is computed by averaging effectiveness ratings over a random sample of ad viewers (personas).
+Unlike the agentic benchmarks, in ad optimization, ad platforms do not have the ability to store and retrieve a particular ad viewer at will.
+Instead, ad platforms (or managers) test an ad to a population of ad viewers and receive an aggregated evaluation signal (e.g., an average effectiveness rating across a sampled set of personas), which prevents us from maintaining instance-level records, such as which prompts win on which ad viewer (in our case, a persona).
+Therefore, to apply GEPA for digital ad optimization, we need to adapt it.
+For Step 1), in agentic AI experiments considered in §9 and agrawal2025gepa), where we can store, retrieve, or evaluate individual problem instances), it uses instance-level information to (i) maintain a Pareto archive of which prompts in the candidate prompt set currently win on which problem instances, (ii) stochastically sample a non-dominated candidate prompt (favoring candidates that win on more instances), and (iii) evaluate that sampled candidate by rolling it out on a minibatch of problem instances.
+On the other hand, in the digital ad optimization experiment (§7), where algorithms have access only to aggregated evaluation results, this step’s procedure simplifies to that of TextBO: evaluate the prompt across the entire test persona set.
+In Step 2), we employ the same meta-reflection procedure as used for TextBO.
+For Step 3), we take one step of evolution (or equivalently, one step of textual gradient) of the prompt we evaluated in Step 1) and accept it only if it has a better evaluation than the original prompt.
+For agentic AI experiments considered in $\S$9 and agrawal2025gepa) where we have access to individual problem instances, when the offspring prompt is accepted, we append it to the candidate prompt set and don’t throw away the original prompt from the candidate set.
+On the other hand, in digital ad optimization in $\S$7, where we cannot maintain the instance-level Pareto archive and thus cannot construct the frequency-weighted, non-dominated sampling distribution, we discard the original prompt from the candidate set.
+This is because, if we were to keep appending accepted prompts, the candidate prompt set would grow without bound, and the evaluation budget per prompt would be uniformly diluted.
+Therefore, when the offspring prompt is accepted, we discard the original prompt, keeping the candidate prompt set size constant.
+Notably, after adapting GEPA to the digital ad optimization problem setup, where we cannot store and retrieve personas freely, it effectively reduces to TextBO with $G=1$ and $N=1$; we describe the pseudocode of the adapted version of GEPA in Algorithm 3.
+Shared components with TextBO.
+To ensure a fair comparison, we keep the non-algorithmic components identical to those used for TextBO (§ 7.2): (i) the same scenario-specific initial pool of 64 prompts generated from the creative brief, (ii) the same ad-generation module $\Phi$ (Imagen 4), (iii) the same critic model $M_{\text{critic}}$ (Gemini 2.5 Flash) and the same meta-prompts for reflection and prompt rewriting, and (iv) the same persona-based evaluation operator $\mathrm{Eval}(\cdot)$ induced by Twin-2k-500 personas and LLM-scored effectiveness ratings.
+Candidate prompt pool (“population”).
+In the ad setting, each prompt $\pi$ produces a single ad image $\Phi(\pi)$ and a single scalar score $s=\mathrm{Eval}(\Phi(\pi),\mathcal{D}_{\mathcal{X}})$.
+As we described earlier, we implement GEPA with a fixed-size candidate pool (population) of size $J=5$.
+We initialize this pool using the same Worst5-of-64 prompts described in $\S$7.3.
+For each candidate $j\in\{1,\dots,J\}$ we maintain a triple $(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})$ where $s_{t}^{j}$ is the empirical evaluation score.
+Evaluation
+Each time we evaluate an ad image $\Phi(\pi)$ during training, we follow the same protocol as in § 7.2: we randomly sample 200 personas from the training persona set and compute the mean effectiveness score across those personas (using the same 1–5 rubric and the same logprob-to-expectation conversion).
+This yields the scalar feedback used by GEPA for selection.
+For reporting final performance (plots/tables), we evaluate selected ads on the held-out test personas.
+Meta-reflection update.
+After collecting evaluations, GEPA constructs a global meta-reflection $R_{t}$ using the same MetaReflect operator as TextBO (§ 6 and § 7.2).
+Concretely, the critic is prompted to compare a subset of the best- and worst-performing creatives observed so far (subject to context-length limits) and to distill recurring patterns into a small set of actionable rules (e.g., “prefer warm lighting,” “show social context,” “avoid clutter”).
+This reflection is then used to guide subsequent prompt evolution steps.
+Evolution (mutation) step and acceptance rule.
+Given a current candidate prompt $\pi_{t}^{j}$ and reflection $R_{t}$, we generate a single offspring prompt by (i) sampling a targeted textual improvement using the critic (a TextGrad-style local edit [textgrad]) and (ii) rewriting the prompt accordingly:
+$\delta_{t}^{j}\sim\nabla_{\text{text}}(\pi_{t}^{j};R_{t},M_{\text{critic}}),\qquad\tilde{\pi}_{t}^{j}=\operatorname{Apply}(\pi_{t}^{j},\delta_{t}^{j}).$
+We then generate the corresponding offspring creative $\Phi(\tilde{\pi}_{t}^{j})$ and evaluate it to obtain $\tilde{s}_{t}^{j}=\mathrm{Eval}(\Phi(\tilde{\pi}_{t}^{j}),\mathcal{D}_{\mathcal{X}})$.
+We use an elitist acceptance rule: if $\tilde{s}_{t}^{j}>s_{t}^{j}$, the offspring replaces the parent in the pool; otherwise the parent is kept.
+This keeps the pool size constant across iterations (preventing unbounded growth of candidates under a fixed evaluation budget) while ensuring that each candidate lineage is non-decreasing in its accepted scores.
+Full procedure and hyperparameters.
+As in TextBO, we run GEPA for $T=10$ iterations with pool size $J=5$ per scenario.
+At each iteration, we evolve each of the $J$ candidates once (one offspring per candidate), evaluate the offspring using 200 sampled training personas, apply the acceptance rule above, and then update the shared meta-reflection.
+For comparison plots and tables, we track the best-performing candidate in the pool at each iteration, $j_{t}^{\star}:=\arg\max_{j}s_{t}^{j}$, and report the corresponding ad $\Phi(\pi_{t}^{j_{t}^{\star}})$.
+Input: System $\Phi$; initial prompts $\{\pi_{0}^{j}\}_{j=1}^{J}$; critic $M_{\text{critic}}$; iterations $T$; pool size $J$
+Output: Best ads $\{\Phi(\pi_{t}^{j_{t}^{\star}})\}_{t=0}^{T}$, where $j_{t}^{\star}=\arg\max_{j}s_{t}^{j}$
+1 for $j=1$ to $J$ do
+2 $c_{0}^{j}\leftarrow\Phi(\pi_{0}^{j})$
+3 $s_{0}^{j}\leftarrow\text{Eval}(c_{0}^{j},\mathcal{D}_{\mathcal{X}})$
+[…]
+5
+6for $t=0$ to T1T-1 do
+7 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t};M_{\text{critic}})$
+8 for $j=1$ to $J$ do
+9    $\delta_{t}^{j}\leftarrow\nabla_{\text{text}}(\pi_{t}^{j};R_{t},M_{\text{critic}})$
+10    $\tilde{\pi}_{t+1}^{j}\leftarrow\operatorname{Apply}(\pi_{t}^{j},\delta_{t}^{j})$
+11    $\tilde{c}_{t+1}^{j}\leftarrow\Phi(\tilde{\pi}_{t+1}^{j})$
+12    $\tilde{s}_{t+1}^{j}\leftarrow\text{Eval}(\tilde{c}_{t+1}^{j},\mathcal{D}_{\mathcal{X}})$
+13    if $\tilde{s}_{t+1}^{j}>s_{t}^{j}$ then
+14       $(\pi_{t+1}^{j},c_{t+1}^{j},s_{t+1}^{j})\leftarrow(\tilde{\pi}_{t+1}^{j},\tilde{c}_{t+1}^{j},\tilde{s}_{t+1}^{j})$
+15
+16    else
+17       $(\pi_{t+1}^{j},c_{t+1}^{j},s_{t+1}^{j})\leftarrow(\pi_{t}^{j},c_{t}^{j},s_{t}^{j})$
+18
+19 $H_{t+1}\leftarrow H_{t}\cup\{(\pi_{t+1}^{j},c_{t+1}^{j},s_{t+1}^{j})\}_{j=1}^{J}$
+20
+return $\{\Phi(\pi_{t}^{j_{t}^{\star}})\}_{t=0}^{T}$
+Algorithm 3 GEPA (adapted for ad optimization)
+D.2 Detailed Experiment Results
+TextBO
+Aeterno Aurasonics Greenbite Mindgarden Momentum Oasis Odyssey Syncflow Average
+Worst5-of-N 2.942 (0.025) 2.793 (0.023) 2.706 (0.024) 2.722 (0.024) 2.644 (0.024) 2.839 (0.025) 2.828 (0.025) 2.641 (0.023) 2.739
+Steps 1 3.098 (0.028) 2.813 (0.022) 3.001 (0.027) 2.980 (0.027) 2.799 (0.027) 3.062 (0.028) 2.943 (0.027) 2.927 (0.023) 2.932
+2 3.125 (0.026) 2.969 (0.027) 3.041 (0.026) 2.980 (0.027) 2.907 (0.025) 3.107 (0.026) 2.943 (0.027) 2.927 (0.023) 2.982
+3 3.125 (0.026) 3.041 (0.025) 3.041 (0.026) 2.980 (0.027) 2.907 (0.025) 3.131 (0.019) 3.063 (0.019) 2.927 (0.023) 3.013
+4 3.194 (0.029) 3.041 (0.025) 3.041 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.063 (0.019) 2.927 (0.023) 3.062
+5 3.194 (0.029) 3.041 (0.025) 3.076 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.063 (0.019) 3.127 (0.022) 3.096
+6 3.257 (0.023) 3.041 (0.025) 3.076 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.070 (0.022) 3.127 (0.022) 3.097
+7 3.257 (0.023) 3.104 (0.028) 3.076 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.111
+8 3.257 (0.023) 3.104 (0.028) 3.076 (0.026) 3.132 (0.022) 2.913 (0.026) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.112
+9 3.257 (0.023) 3.104 (0.028) 3.076 (0.026) 3.132 (0.022) 2.926 (0.026) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.114
+10 3.257 (0.023) 3.104 (0.028) 3.168 (0.025) 3.132 (0.022) 3.024 (0.025) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.155
+Best-of-N 2.982 (0.025) 2.934 (0.027) 2.898 (0.025) 2.930 (0.028) 3.030 (0.027) 3.265 (0.022) 2.966 (0.027) 3.086 (0.023) 3.011
+Table A1: Progress of TextBO across 10 optimization steps from its starting point (Worst5-of-N) and comparison with the Best-of-N baseline.
+Each data point, in the mean (standard error) format, indicates the mean and standard error across 411 personas in the test persona set.
+The last column, labeled “Average”, reports the averages of the means across the eight scenarios.
+GEPA
+Aeterno Aurasonics Greenbite Mindgarden Momentum Oasis Odyssey Syncflow Average
+Worst5-of-N 2.942 (0.025) 2.793 (0.023) 2.706 (0.024) 2.722 (0.024) 2.644 (0.024) 2.839 (0.025) 2.828 (0.025) 2.641 (0.023) 2.739
+Steps 1 3.036 (0.025) 2.942 (0.025) 3.029 (0.025) 2.841 (0.024) 2.816 (0.024) 3.151 (0.024) 2.915 (0.025) 2.880 (0.025) 2.939
+2 3.036 (0.025) 2.942 (0.025) 3.046 (0.024) 2.945 (0.023) 2.907 (0.023) 3.151 (0.024) 3.071 (0.028) 2.880 (0.025) 2.992
+3 3.036 (0.025) 2.942 (0.025) 3.046 (0.024) 2.945 (0.023) 2.907 (0.023) 3.182 (0.025) 3.088 (0.026) 2.914 (0.018) 3.003
+4 3.138 (0.025) 3.028 (0.022) 3.046 (0.024) 2.954 (0.024) 2.976 (0.026) 3.187 (0.023) 3.088 (0.026) 2.949 (0.028) 3.033
+5 3.138 (0.025) 3.028 (0.022) 3.046 (0.024) 3.012 (0.022) 2.976 (0.026) 3.187 (0.023) 3.088 (0.026) 3.008 (0.026) 3.049
+6 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.012 (0.022) 2.976 (0.026) 3.209 (0.020) 3.088 (0.026) 3.008 (0.026) 3.052
+7 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.209 (0.020) 3.088 (0.026) 3.008 (0.026) 3.060
+8 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.209 (0.020) 3.088 (0.026) 3.008 (0.026) 3.060
+9 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.278 (0.023) 3.088 (0.026) 3.008 (0.026) 3.070
+10 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.278 (0.023) 3.088 (0.026) 3.008 (0.026) 3.070
+Best-of-N 2.982 (0.025) 2.934 (0.027) 2.898 (0.025) 2.930 (0.028) 3.030 (0.027) 3.265 (0.022) 2.966 (0.027) 3.086 (0.023) 3.011
+Table A2: Progress of GEPA across 10 optimization steps from its starting point (Worst5-of-N) and comparison with the Best-of-N baseline.
+Each data point, in the mean (standard error) format, indicates the mean and standard error across 411 personas in the test persona set.
+The last column, labeled “Average”, reports the averages of the means across the eight scenarios.
+Odyssey (Electric family SUV) Oasis (Secluded resort) Momentum (Mobile banking app) Aeterno (Luxury watch) Syncflow (Collaboration tool)
+Worst5-of-64 Refer to caption Score: 2.828 Refer to caption Score: 2.839 Refer to caption Score: 2.644 Refer to caption Score: 2.942 Refer to caption Score: 2.641
+Best-of-64 Refer to caption Score: 2.966 Refer to caption Score: 3.265 Refer to caption Score: 3.030 Refer to caption Score: 2.982 Refer to caption Score: 3.086
+GEPA (T=10) Refer to caption Score: 3.088 Refer to caption Score: 3.278 Refer to caption Score: 2.976 Refer to caption Score: 3.157 Refer to caption Score: 3.008
+TextBO (T=10) Refer to caption Score: 3.107 Refer to caption Score: 3.324 Refer to caption Score: 3.024 Refer to caption Score: 3.257 Refer to caption Score: 3.127
+Figure A8: TextBO and baseline’s generated Ad images for five fictional brands: Odyssey electric family SUV (Scenario 3), Oasis eco-lodge (Scenario 4), Momentum Digital Banking (Scenario 5), Aeterno luxury watch (Scenario 7), and SyncFlow B2B software (Scenario 8)
+D.3 Final Prompts in TextBO and GEPA
+Prompt for generating the TextBO $(T=10)$ ad for GreenBite (plant-based burger patty) "**Key Message:**
 "The No-Compromise Burger"—to visually prove that a plant-based burger can be just as juicy, delicious, and satisfying as a traditional beef burger, making it the hero of the ultimate backyard BBQ.
@@
 This overlay should be highly legible and reinforce clear brand recognition without detracting from the visual appeal."
-Figure 15: Prompts for T-BoN ($T=10$)
+Figure A9: Prompts for TextBO ($T=10$)
 Prompt for generating the GEPA ($T=10$) ad for GreenBite (plant-based burger patty) "**Key Message:**
@@
 This headline should be clearly legible, visually impactful, and act as a primary visual element, prioritizing the core message and value proposition over just generic brand reinforcement."
-Figure 16: Prompts for GEPA ($T=10$)
-Appendix B Scenarios
-Below are ten ad campaign scenarios with creative briefs, outlining the strategic and creative direction for each experimental scenario.
-The selected campaigns cover a diverse range of products across distinct categories.
-Each campaign addresses a unique consumer need—whether related to sustainability, luxury, technology, or lifestyle—ensuring comprehensive coverage of various market segments.
-From environmentally-conscious millennials to affluent, experience-driven travelers, the approach accounts for the full spectrum of consumer interests.
-By focusing on different consumer needs, the structure enables a holistic marketing strategy that captures both niche and broad demands, providing a well-rounded framework for the study.
-B.1 Scenario 1: GreenBite Plant-Based Burger
-  • Product: "GreenBite," a new plant-based burger patty.
-  • Background: The market for plant-based food is growing, but many consumers remain skeptical about taste, believing they must sacrifice flavor for health.
-  • The Challenge: Convince flexitarians that GreenBite offers the juicy, satisfying experience of a traditional beef burger with no compromise.
-  • Core Insight: Consumers want to eat better for themselves and the planet, but they fear it means giving up the foods they love.
-  • Single-Minded Proposition (SMP): The delicious, no-compromise burger experience that’s better for you and the planet.
-  • Reasons to Believe (RTB): Made with savory pea protein; sears and tastes like real beef; free of soy and GMOs.
-  • Desired Response: Think: "I can finally have a plant-based burger that tastes like the real thing.", Feel: Satisfied, vibrant, and guilt-free. Do: Purchase GreenBite for their next barbecue.
-  • Brand Personality & Tone: Positive, confident, and mouth-watering.
-B.2 Scenario 2: AuraSonics X1 Earbuds
-  • Product: "AuraSonics X1," high-end, noise-canceling wireless earbuds.
-  • Background: The rise of hybrid work and urban density has increased daily auditory overload, making focus a precious commodity.
-  • The Challenge: Cut through the saturated audio market by positioning the X1 not as a gadget, but as an essential tool for mental clarity.
-  • Core Insight: In a world of constant noise, true luxury is the ability to control your own soundscape.
-  • Single-Minded Proposition (SMP): AuraSonics X1 creates a personal sanctuary for focus and immersion.
-  • Reasons to Believe (RTB): Adaptive Active Noise Cancellation; studio-quality audio; all-day battery life; minimalist aluminum design.
-  • Desired Response: Think: "This is the solution I need to escape the daily chaos." Feel: Calm, empowered, and sophisticated. Do: Visit the website to learn more.
-  • Brand Personality & Tone: Minimalist, intelligent, and calm.
-B.3 Scenario 3: Odyssey E-SUV
-  • Product: The "Odyssey E-SUV," a new all-electric family SUV.
-  • Background: Many families want to switch to electric vehicles but are concerned about sacrificing space, safety, and range for sustainability.
-  • The Challenge: Position the Odyssey E-SUV as the first EV that meets all the practical and adventurous needs of a modern family, making the switch to electric feel like an upgrade, not a compromise.
-  • Single-Minded Proposition (SMP): The future of family adventure is here: all-electric, zero compromise.
-  • Reasons to Believe (RTB): 500km range; top-tier safety rating; spacious three-row seating; all-wheel drive capability.
-  • Desired Response: Think: "An electric car can actually handle my family’s active lifestyle." Feel: Adventurous, optimistic, and secure. Do: Schedule a test drive.
-  • Brand Personality & Tone: Inspiring, capable, and forward-thinking.
-B.4 Scenario 4: Oasis Eco-Lodge
-  • Product: "Oasis Eco-Lodge," a secluded, luxury resort operating in harmony with its natural surroundings.
-  • Background: The luxury travel market often equates opulence with excess. A growing segment of affluent travelers seeks experiences that are both exclusive and responsible.
-  • The Challenge: Redefine luxury as a seamless integration with nature, promising an experience that is restorative for both the guest and the environment.
-  • Core Insight: For those who have everything, true luxury is not more things, but a deeper connection to something pure and serene.
-  • Single-Minded Proposition (SMP): Rediscover tranquility in a luxury that respects nature.
-  • Reasons to Believe (RTB): Secluded private bungalows; farm-to-table dining with locally sourced ingredients; carbon-neutral operations.
-  • Desired Response: Think: "This is a truly special escape, not just another five-star hotel." Feel: Serene, exclusive, and rejuvenated. Do: Book a stay.
-  • Brand Personality & Tone: Elegant, peaceful, and understated.
-B.5 Scenario 5: Momentum Digital Banking
-  • Product: "Momentum," a mobile-first banking app for freelancers and the gig economy.
-  • Background: Traditional banks are not built for the fluctuating incomes and unique business needs of self-employed professionals.
-  • The Challenge: Position Momentum as the essential financial co-pilot for the independent worker, simplifying complexity and providing stability.
-  • Core Insight: Freelancers love their freedom but feel anxious about their financial instability. They need a tool that brings order to their financial chaos.
-  • Single-Minded Proposition (SMP): Banking that is as flexible and entrepreneurial as you are.
-  • Reasons to Believe (RTB): Automated tax-saving tools; integrated invoicing and payment tracking; instant business expense categorization.
-  • Desired Response: Think: "Finally, a bank that gets the way I work." Feel: Empowered, organized, and financially confident. Do: Download the app and open an account.
-  • Brand Personality & Tone: Modern, empowering, and simple.
-B.6 Scenario 6: MindGarden Meditation App
-  • Product: "MindGarden," a subscription-based meditation and mindfulness app.
-  • Background: While many are interested in mindfulness for stress-relief, they are often intimidated by the practice, seeing it as difficult or time-consuming.
-  • The Challenge: Make meditation feel accessible and achievable for absolute beginners, removing the barriers of time and perceived difficulty.
-  • Core Insight: People want the benefits of mindfulness but are convinced they don’t have the time or ability to practice it correctly.
-  • Single-Minded Proposition (SMP): Find your calm in just 5 minutes a day.
-  • Reasons to Believe (RTB): Guided 5-minute sessions for specific needs (anxiety, focus); progress tracking to build a habit; simple, jargon-free instructions.
-  • Desired Response: Think: "I can do this. 5 minutes is easy." Feel: Calm, supported, and hopeful. Do: Start a free trial.
-  • Brand Personality & Tone: Gentle, approachable, and encouraging.
-B.7 Scenario 7: Aeterno Watch
-  • Product: The "Aeterno," a classic, automatic Swiss-made wristwatch with a heritage design.
-  • Background: In an age of smartwatches that are obsolete in a few years, there is a renewed appreciation for objects with permanence and enduring value.
-  • The Challenge: Reassert the relevance of the classic mechanical watch as a symbol of taste, craftsmanship, and timelessness.
-  • Core Insight: In a disposable world, true status comes from owning something permanent that tells a story.
-  • Single-Minded Proposition (SMP): A legacy on your wrist. Craftsmanship that transcends time.
-  • Reasons to Believe (RTB): Swiss-made automatic movement; sapphire crystal glass; timeless design inspired by 1950s classics.
-  • Desired Response: Think: "This is a beautiful object that I will own forever." Feel: Prestigious, sophisticated, and discerning. Do: Locate an authorized dealer.
-  • Brand Personality & Tone: Elegant, timeless, and confident.
-B.8 Scenario 8: SyncFlow B2B Software
-  • Product: "SyncFlow," a project management and collaboration software platform for remote teams.
-  • Background: Remote work has increased flexibility but also created challenges in team alignment, communication, and project visibility.
-  • The Challenge: Cut through a crowded SaaS market by focusing on the outcome of "effortless collaboration" rather than just listing features.
-  • Core Insight: Managers don’t want another tool to manage; they want the feeling of clarity and momentum that comes from a team working in perfect sync.
-  • Single-Minded Proposition (SMP): Bring your remote team together for effortless collaboration and remarkable results.
-  • Reasons to Believe (RTB): Centralized dashboards with real-time project status; integrated communication channels; automated workflows and reporting.
-  • Desired Response: Think: "This could finally solve our remote work chaos and get everyone on the same page." Feel: Organized, in control, and successful. Do: Sign up for a team demo.
-  • Brand Personality & Tone: Professional, efficient, and innovative.
-Appendix C Extended experiment results
-T-BoN BO
-Aeterno Aurasonics Greenbite Mindgarden Momentum Oasis Odyssey Syncflow Average
-Worst5-of-N 2.942 (0.025) 2.793 (0.023) 2.706 (0.024) 2.722 (0.024) 2.644 (0.024) 2.839 (0.025) 2.828 (0.025) 2.641 (0.023) 2.739
-Steps 1 3.098 (0.028) 2.813 (0.022) 3.001 (0.027) 2.980 (0.027) 2.799 (0.027) 3.062 (0.028) 2.943 (0.027) 2.927 (0.023) 2.932
-2 3.125 (0.026) 2.969 (0.027) 3.041 (0.026) 2.980 (0.027) 2.907 (0.025) 3.107 (0.026) 2.943 (0.027) 2.927 (0.023) 2.982
-3 3.125 (0.026) 3.041 (0.025) 3.041 (0.026) 2.980 (0.027) 2.907 (0.025) 3.131 (0.019) 3.063 (0.019) 2.927 (0.023) 3.013
-4 3.194 (0.029) 3.041 (0.025) 3.041 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.063 (0.019) 2.927 (0.023) 3.062
-5 3.194 (0.029) 3.041 (0.025) 3.076 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.063 (0.019) 3.127 (0.022) 3.096
-6 3.257 (0.023) 3.041 (0.025) 3.076 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.070 (0.022) 3.127 (0.022) 3.097
-7 3.257 (0.023) 3.104 (0.028) 3.076 (0.026) 3.132 (0.022) 2.907 (0.025) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.111
-8 3.257 (0.023) 3.104 (0.028) 3.076 (0.026) 3.132 (0.022) 2.913 (0.026) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.112
-9 3.257 (0.023) 3.104 (0.028) 3.076 (0.026) 3.132 (0.022) 2.926 (0.026) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.114
-10 3.257 (0.023) 3.104 (0.028) 3.168 (0.025) 3.132 (0.022) 3.024 (0.025) 3.324 (0.028) 3.107 (0.027) 3.127 (0.022) 3.155
-Best-of-N 2.982 (0.025) 2.934 (0.027) 2.898 (0.025) 2.930 (0.028) 3.030 (0.027) 3.265 (0.022) 2.966 (0.027) 3.086 (0.023) 3.011
-Table 2: Progress of T-BoN BO across 10 optimization steps from its starting point (Worst5-of-N) and comparison with the Best-of-N baseline.
-Each data point, in the mean (standard error) format, indicates the mean and standard error across 411 personas in the test persona set.
-The last column, labeled “Average”, reports the averages of the means across the eight scenarios.
+Figure A10: Prompts for GEPA ($T=10$)
+Appendix E Appendix for Section 9
+E.1 Implementation Details of GEPA and TextBO-GEPA for agentic AI experiments
 GEPA
-Aeterno Aurasonics Greenbite Mindgarden Momentum Oasis Odyssey Syncflow Average
-Worst5-of-N 2.942 (0.025) 2.793 (0.023) 2.706 (0.024) 2.722 (0.024) 2.644 (0.024) 2.839 (0.025) 2.828 (0.025) 2.641 (0.023) 2.739
-Steps 1 3.036 (0.025) 2.942 (0.025) 3.029 (0.025) 2.841 (0.024) 2.816 (0.024) 3.151 (0.024) 2.915 (0.025) 2.880 (0.025) 2.939
-2 3.036 (0.025) 2.942 (0.025) 3.046 (0.024) 2.945 (0.023) 2.907 (0.023) 3.151 (0.024) 3.071 (0.028) 2.880 (0.025) 2.992
-3 3.036 (0.025) 2.942 (0.025) 3.046 (0.024) 2.945 (0.023) 2.907 (0.023) 3.182 (0.025) 3.088 (0.026) 2.914 (0.018) 3.003
-4 3.138 (0.025) 3.028 (0.022) 3.046 (0.024) 2.954 (0.024) 2.976 (0.026) 3.187 (0.023) 3.088 (0.026) 2.949 (0.028) 3.033
-5 3.138 (0.025) 3.028 (0.022) 3.046 (0.024) 3.012 (0.022) 2.976 (0.026) 3.187 (0.023) 3.088 (0.026) 3.008 (0.026) 3.049
-6 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.012 (0.022) 2.976 (0.026) 3.209 (0.020) 3.088 (0.026) 3.008 (0.026) 3.052
-7 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.209 (0.020) 3.088 (0.026) 3.008 (0.026) 3.060
-8 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.209 (0.020) 3.088 (0.026) 3.008 (0.026) 3.060
-9 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.278 (0.023) 3.088 (0.026) 3.008 (0.026) 3.070
-10 3.157 (0.023) 3.028 (0.022) 3.046 (0.024) 3.068 (0.027) 2.976 (0.026) 3.278 (0.023) 3.088 (0.026) 3.008 (0.026) 3.070
-Best-of-N 2.982 (0.025) 2.934 (0.027) 2.898 (0.025) 2.930 (0.028) 3.030 (0.027) 3.265 (0.022) 2.966 (0.027) 3.086 (0.023) 3.011
-Table 3: Progress of GEPA across 10 optimization steps from its starting point (Worst5-of-N) and comparison with the Best-of-N baseline.
-Each data point, in the mean (standard error) format, indicates the mean and standard error across 411 personas in the test persona set.
-The last column, labeled “Average”, reports the averages of the means across the eight scenarios.
-Odyssey (Electric family SUV) Oasis (Secluded resort) Momentum (Mobile banking app) Aeterno (Luxury watch) Syncflow (Collaboration tool)
-Worst5-of-64 Refer to caption Score: 2.828 Refer to caption Score: 2.839 Refer to caption Score: 2.644 Refer to caption Score: 2.942 Refer to caption Score: 2.641
-Best-of-64 Refer to caption Score: 2.966 Refer to caption Score: 3.265 Refer to caption Score: 3.030 Refer to caption Score: 2.982 Refer to caption Score: 3.086
-GEPA (T=10) Refer to caption Score: 3.088 Refer to caption Score: 3.278 Refer to caption Score: 2.976 Refer to caption Score: 3.157 Refer to caption Score: 3.008
-T-BoN BO (T=10) Refer to caption Score: 3.107 Refer to caption Score: 3.324 Refer to caption Score: 3.024 Refer to caption Score: 3.257 Refer to caption Score: 3.127
-Figure 17: T-BoN BO and baseline’s generated Ad images for five fictional brands: Odyssey electric family SUV (Scenario 3), Oasis eco-lodge (Scenario 4), Momentum Digital Banking (Scenario 5), Aeterno luxury watch (Scenario 7), and SyncFlow B2B software (Scenario 8)
-Appendix D Proofs
-D.1 Auxiliary lemmas
-Lemma 3 (Uniform first-order expansion).
-Let $g=\nabla\mu(x)$ and $h=\nabla\sigma(x)$. There exist $C>0$ and $\varepsilon_{0}>0$ such that for all $\varepsilon\in(0,\varepsilon_{0}]$ and all $u\in\mathbb{R}^{d}$,
-$\mu(x+\varepsilon u)$ $=\mu(x)+\varepsilon\,g^{\top}u\;+\;r_{\mu}(\varepsilon,u),$ $|r_{\mu}(\varepsilon,u)|$ $\leq C\varepsilon^{2},$ (28)
-$\sigma(x+\varepsilon u)$ $=\sigma(x)+\varepsilon\,h^{\top}u\;+\;r_{\sigma}(\varepsilon,u),$ $|r_{\sigma}(\varepsilon,u)|$ $\leq C\varepsilon^{2}.$ (29)
-Proof.
-By the mean-value theorem with Lipschitz gradients, for some $\xi$ between $x$ and $x+\varepsilon u$, $\mu(x+\varepsilon u)=\mu(x)+\nabla\mu(\xi)^{\top}(\varepsilon u).$ Hence $\mu(x+\varepsilon u)=\mu(x)+\varepsilon g^{\top}u+\varepsilon(\nabla\mu(\xi)-\nabla\mu(x))^{\top}u,$ and $|\nabla\mu(\xi)-\nabla\mu(x)|\leq L_{\mu}\|\xi-x\|\leq L_{\mu}\varepsilon$, giving $|r_{\mu}|\leq L_{\mu}\varepsilon^{2}$.
-The proof for $\sigma$ is identical with $L_{\sigma}$.
-Taking $C:=\max\{L_{\mu},L_{\sigma}\}$ yields the stated bounds with the same $C$.
-∎
-Lemma 4 (Max-stability under bounded perturbations).
-Let $(f_{i})_{i\leq N}$ and $(g_{i})_{i\leq N}$ be real arrays. Let $i_{f}\in\arg\max_{i}f_{i}$ and $i_{g}\in\arg\max_{i}g_{i}$. If for some $\Delta\geq 0$,
-$g_{i_{f}}\geq f_{i_{f}}-\Delta\quad\text{and}\quad f_{i_{g}}\geq g_{i_{g}}-\Delta,$ (30)
-then $g_{i_{f}}\geq\max_{i}g_{i}-\Delta$ and $f_{i_{g}}\geq\max_{i}f_{i}-\Delta$.
-In particular, $g_{i_{f}}\geq\max_{i}g_{i}-\Delta$ implies that any maximizer of $f$ is a $\Delta$-approximate maximizer of $g$.
-Proof.
-Since $g_{i_{f}}\geq f_{i_{f}}-\Delta=\max_{i}f_{i}-\Delta\geq g_{i_{g}}-\Delta=\max_{i}g_{i}-\Delta$, the first claim holds. The second is symmetric. ∎
-Lemma 5 (Spherical cap coverage).
-For $v\in\mathbb{S}^{d-1}$ and $\eta\in(0,1)$, define $C(v,\eta):=\{u\in\mathbb{S}^{d-1}:v^{\top}u\geq 1-\eta\}$.
-Let $U_{1},\ldots,U_{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\rho$ as above.
-Then $\max_{i\leq N}v^{\top}U_{i}\to 1$ almost surely as $N\to\infty$.
-Moreover, for any deterministic sequence $\eta_{N}\downarrow 0$ with $N\cdot\rho\{u:v^{\top}u\geq 1-\eta_{N}\}\to\infty$, we have
-$\mathbb{P}\!\left(\max_{i\leq N}v^{\top}U_{i}\geq 1-\eta_{N}\right)\to 1.$ (31)
-Proof.
-Let $C(\eta)=\{u\in\mathbb{S}^{d-1}:v^{\top}u\geq 1-\eta\}$ be a spherical cap.
-Since $\rho$ has a density bounded below, $\rho(C(\eta))\asymp\eta^{\frac{d-1}{2}}$ as $\eta\downarrow 0$.
-Then $\mathbb{P}(\max_{i}v^{\top}U_{i}<1-\eta)=(1-\rho(C(\eta)))^{N}\leq\exp(-N\rho(C(\eta))).$ For any fixed $\eta\in(0,1)$, $\sum_{N\geq 1}\mathbb{P}(\max_{i}v^{\top}U_{i}<1-\eta)\leq\sum_{N\geq 1}e^{-N\rho(C(\eta))}<\infty$, so by Borel–Cantelli, $\mathbb{P}(\max_{i}v^{\top}U_{i}<1-\eta\ \text{i.o.})=0$.
-Intersecting over a countable sequence $\eta_{m}\downarrow 0$ yields $\max_{i\leq N}v^{\top}U_{i}\to 1$ almost surely.
-Choosing $\eta_{N}\downarrow 0$ so that $N\rho(C(\eta_{N}))\to\infty$ yields the result, and taking $\eta=\eta_{N}\to 0$ proves the first statement.
-∎
-Lemma 6 (Near-maximum coupling in a thin band).
-Let $M_{N}=\max_{i}\xi_{i}$ and define $v_{N}:=(g+M_{N}h)/\|g+M_{N}h\|$.
-Also, $S_{i}:=a_{i}+b_{i}M_{N}$, $Y_{i}^{(0)}:=a_{i}+b_{i}\xi_{i}$, $b_{i}:=\sigma(x)+\varepsilon h^{\top}U_{i}$.
-Fix any sequence $c_{N}\uparrow\infty$ with $c_{N}/q_{N}\to 0$ and set $\delta_{N}:=c_{N}/q_{N}$.
-Then with probability $\to 1$ there exists an index $j$ such that
-$\xi_{j}\geq M_{N}-\delta_{N}\quad\text{and}\quad v_{N}^{\top}U_{j}\geq 1-\eta_{N},$ (32)
-whenever $\eta_{N}\downarrow 0$ satisfies
-$N\,\mathbb{P}(\xi\geq M_{N}-\delta_{N})\,\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}\to\infty.$ (33)
-Moreover, for this $j$,
-$S_{j}-Y_{j}^{(0)}=b_{j}(M_{N}-\xi_{j})\leq b_{\max}\,\delta_{N}.$ (34)
-Proof.
-Define $p_{\mathrm{cap},N}:=\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}$ and the event
-$E_{N}:=\{M_{N}\geq q_{N}-\delta_{N}\}.$ (35)
-By Lemma 1, $M_{N}-q_{N}\to 0$ in probability and $\delta_{N}\to 0$, hence $\mathbb{P}(E_{N})\to 1$.
-Set the deterministic threshold $t_{N}:=q_{N}-2\delta_{N}$. On $E_{N}$ we have $M_{N}-\delta_{N}\geq q_{N}-2\delta_{N}=t_{N}$, hence by monotonicity of the tail,
-$\mathbb{P}(\xi\geq t_{N})\;\geq\;\mathbb{P}(\xi\geq M_{N}-\delta_{N})\quad\text{on }E_{N}.$ (36)
-Let
-$K_{N}:=\sum_{i=1}^{N}\mathbf{1}\{\xi_{i}\geq t_{N}\}\sim\mathrm{Binomial}\!\big(N,\,p_{N}\big),\qquad p_{N}:=\mathbb{P}(\xi\geq t_{N}).$ (37)
-A Chernoff bound gives
-$\mathbb{P}\!\left(K_{N}<\tfrac{1}{2}Np_{N}\right)\;\leq\;\exp\!\left(-\tfrac{1}{8}Np_{N}\right).$ (38)
-Consider
-$L_{N}\;:=\;\sum_{i=1}^{N}\mathbf{1}\{\xi_{i}\geq t_{N}\}\,\mathbf{1}\{v_{N}^{\top}U_{i}\geq 1-\eta_{N}\}.$ (39)
-[…]
-$L_{N}\,\big|\,\{\xi_{i}\}\;\sim\;\mathrm{Binomial}\!\big(K_{N},\,p_{\mathrm{cap},N}\big),\qquad\mathbb{P}\big(L_{N}=0\,\big|\,\{\xi_{i}\}\big)\;=\;(1-p_{\mathrm{cap},N})^{K_{N}}\;\leq\;\exp\!\big(-K_{N}p_{\mathrm{cap},N}\big).$ (40)
-We now bound $\mathbb{P}(L_{N}=0)$ by splitting on $\{K_{N}\geq\tfrac{1}{2}Np_{N}\}$ and $E_{N}$:
-$\mathbb{P}(L_{N}=0)$ $\leq\mathbb{E}\!\left[\exp\!\big(-K_{N}p_{\mathrm{cap},N}\big)\,\mathbf{1}_{\{K_{N}\geq\frac{1}{2}Np_{N}\}\cap E_{N}}\right]\;+\;\mathbb{P}\!\left(K_{N}<\tfrac{1}{2}Np_{N}\right)\;+\;\mathbb{P}(E_{N}^{c})$
-$\leq\mathbb{E}\!\left[\exp\!\big(-\tfrac{1}{2}Np_{N}\,p_{\mathrm{cap},N}\big)\,\mathbf{1}_{E_{N}}\right]\;+\;\exp\!\big(-\tfrac{1}{8}Np_{N}\big)\;+\;o(1)$
-$\leq\mathbb{E}\!\left[\exp\!\big(-\tfrac{1}{2}Np_{N}\,p_{\mathrm{cap},N}\big)\right]\;+\;\exp\!\big(-\tfrac{1}{8}Np_{N}\big)\;+\;o(1).$ (41)
-On $E_{N}$, (36) yields $p_{N}\geq\mathbb{P}(\xi\geq M_{N}-\delta_{N})$. By the lemma’s hypothesis,
-$Z_{N}:=Np_{N}\,p_{\mathrm{cap},N}\;\geq\;N\,\mathbb{P}(\xi\geq M_{N}-\delta_{N})\,p_{\mathrm{cap},N}\;\xrightarrow{p}\;\infty.$
-Since $0\leq e^{-\frac{1}{2}Z_{N}}\leq 1$, for any fixed $T>0$,
-$\mathbb{E}\!\left[e^{-\frac{1}{2}Z_{N}}\right]\leq e^{-\frac{1}{2}T}+\mathbb{P}(Z_{N}\leq T)\;\longrightarrow\;e^{-\frac{1}{2}T}.$
-Letting $T\to\infty$ gives $\mathbb{E}[e^{-\frac{1}{2}Z_{N}}]\to 0$.
-Moreover, $Z_{N}\leq Np_{N}$ since $p_{\mathrm{cap},N}\leq 1$.
-If $Np_{N}$ were bounded along a subsequence, then $Z_{N}$ would be bounded along that subsequence, contradicting $Z_{N}\xrightarrow{p}\infty$.
-Hence $Np_{N}\to\infty$, so $\exp(-\tfrac{1}{8}Np_{N})\to 0$.
-Therefore the second term in (41) vanishes, and $\mathbb{P}(L_{N}=0)\to 0$.
-Therefore, with probability tending to one there exists $j$ such that $\xi_{j}\geq t_{N}$ and $v_{N}^{\top}U_{j}\geq 1-\eta_{N}$.
-Since $t_{N}\leq M_{N}-\delta_{N}$ on $E_{N}$, this $j$ also satisfies $\xi_{j}\geq M_{N}-\delta_{N}$.
-Finally, $S_{j}-Y_{j}^{(0)}=b_{j}(M_{N}-\xi_{j})\leq b_{\max}\delta_{N}$ holds deterministically by $M_{N}-\xi_{j}\leq\delta_{N}$ and the definition of $b_{\max}$.
-∎
-D.2 Proof of Theorem 2
-Proof.
-By Lemma 3,
-$Y_{i}$ $=\mu(x)+\varepsilon g^{\top}u_{i}+r_{\mu}(\varepsilon,u_{i})+\big(\sigma(x)+\varepsilon h^{\top}u_{i}+r_{\sigma}(\varepsilon,u_{i})\big)\xi_{i}$
-$=a_{i}+b_{i}\xi_{i}+R_{i},$ (42)
-where $a_{i}:=\mu(x)+\varepsilon g^{\top}u_{i}$, $b_{i}:=\sigma(x)+\varepsilon h^{\top}u_{i}$, and $R_{i}:=r_{\mu}(\varepsilon,u_{i})+\xi_{i}r_{\sigma}(\varepsilon,u_{i}).$ Then $|R_{i}|\leq C\varepsilon^{2}(1+|\xi_{i}|)$ for all $i$.
-Define $S_{i}:=a_{i}+b_{i}M_{N}$, $b_{\max}:=\max_{1\leq i\leq N}|b_{i}|$, and $U_{i}:=u_{i}$.
-Now let $Y_{i}^{(0)}:=a_{i}+b_{i}\xi_{i}$ and $i_{0}\in\arg\max_{i}Y_{i}^{(0)}$. From (42),
-$Y_{i^{\star}}\geq Y_{i_{0}}\implies Y_{i^{\star}}^{(0)}\geq Y_{i_{0}}^{(0)}-|R_{i^{\star}}|-|R_{i_{0}}|.$
-By the bound on $R_{i}$ and $|\xi_{i}|\leq M_{N}:=\max_{j}\xi_{j}$,
-$Y_{i^{\star}}^{(0)}\geq\max_{i}Y_{i}^{(0)}-2C\varepsilon^{2}(1+M_{N}).$ (43)
-Symmetrically, $\max_{i}Y_{i}^{(0)}\geq Y_{i^{\star}}^{(0)}\geq\max_{i}Y_{i}^{(0)}-2C\varepsilon^{2}(1+M_{N})$.
-By Lemma 4, $i^{\star}$ is a $2C\varepsilon^{2}(1+M_{N})$-approximate maximizer of $Y^{(0)}$.
-Now let $j$ be the index from Lemma 6 so that $\xi_{j}\geq M_{N}-\delta_{N}$ and $v_{N}^{\top}U_{j}\geq 1-\eta_{N}$. By Lemma 3,
-$\max_{i\leq N}S_{i}-S_{j}\;\leq\;\varepsilon\|g+M_{N}h\|\,(1-v_{N}^{\top}U_{j})\;+\;2C\varepsilon^{2}(1+M_{N})\;\leq\;\varepsilon\|g+M_{N}h\|\,\eta_{N}\;+\;2C\varepsilon^{2}(1+M_{N}).$
-Also,
-$S_{j}-S_{i^{\star}}=(S_{j}-Y_{j}^{(0)})+(Y_{j}^{(0)}-Y_{i^{\star}}^{(0)})+(Y_{i^{\star}}^{(0)}-S_{i^{\star}})\;\leq\;b_{\max}\delta_{N}\;+\;2C\varepsilon^{2}(1+M_{N})\;+\;\underbrace{(Y_{i^{\star}}^{(0)}-S_{i^{\star}})}_{\leq 0},$
-where the last term is $\leq 0$ provided $b_{i}\geq 0$ for all $i$, which holds by Assumption A4 for $\varepsilon\leq\varepsilon_{0}:=\sigma(x)/(2\|h\|)$ when $\sigma(x)>0$ since $b_{i}=\sigma(x)+\varepsilon h^{\top}U_{i}\geq\sigma(x)-\varepsilon\|h\|\geq\tfrac{1}{2}\sigma(x)>0$.
-Therefore,
-$\max_{i\leq N}S_{i}-S_{i^{\star}}\;\leq\;\varepsilon\|g+M_{N}h\|\,\eta_{N}\;+\;b_{\max}\delta_{N}\;+\;4C\varepsilon^{2}(1+M_{N}).$ (44)
-Set:
-$\Delta^{\prime}_{N}(\varepsilon)\;:=\;b_{\max}\,\frac{c_{N}}{q_{N}}\;+\;4C\varepsilon^{2}(1+M_{N}).$
-By Lemma 3,
-$S_{i}\;=\;\mu(x)+M_{N}\sigma(x)+\varepsilon(g+M_{N}h)^{\top}u_{i}+\tilde{r}_{i},\qquad|\tilde{r}_{i}|\leq C\varepsilon^{2}(1+M_{N}).$
-Hence for any $i$,
-$S_{i}-S_{i^{\star}}=\varepsilon\|g+M_{N}h\|\,(v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N})+(\tilde{r}_{i}-\tilde{r}_{i^{\star}}),$
-so
-$\big|\,(S_{i}-S_{i^{\star}})-\varepsilon\|g+M_{N}h\|\,(v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N})\,\big|\;\leq\;2C\varepsilon^{2}(1+M_{N}).$ (45)
-Taking $i$ that attains $\max_{i}S_{i}$ and using (44) yields
-$\max_{i\leq N}v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N}\;\leq\;\eta_{N}\;+\;\frac{\Delta^{\prime}_{N}(\varepsilon)}{\varepsilon\|g+M_{N}h\|}.$ (46)
-Thus
-[…]
-Assume $\|h\|>0$ (the $\|h\|=0$ case is addressed in Remark 1). Using
-$\|g+M_{N}h\|\;\geq\;M_{N}\|h\|-\|g\|$
-and Lemma 1, $M_{N}\to\infty$ and $M_{N}\asymp q_{N}$. Also $b_{\max}\leq\sigma(x)+\varepsilon\|h\|+C\varepsilon^{2}$. Hence
-$\frac{\Delta^{\prime}_{N}(\varepsilon)}{\varepsilon\|g+M_{N}h\|}=O_{p}\!\Big(\tfrac{c_{N}}{\varepsilon q_{N}^{2}\|h\|}\Big)+O\!\Big(\tfrac{\varepsilon}{\|h\|}\Big).$
-Choosing, e.g., $c_{N}=\tfrac{1}{2}\log\log N$ and any $\eta_{N}\downarrow 0$ with
-$N\,e^{c_{N}}\,\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}\to\infty$
-makes the RHS of (47) equal to $o_{p}(1)+O(\varepsilon)$ for fixed $\varepsilon>0$; then let $\varepsilon\downarrow 0$.
-With $N\to\infty$ first and fixed $\varepsilon>0$, (47) implies $1-v_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}0$.
-Then $\varepsilon\downarrow 0$ removes the $O(\varepsilon)$ residual, yielding $v_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}1$.
-Since $M_{N}-q_{N}\to 0$ in probability by Lemma 1, the unit vectors $v_{N}=(g+M_{N}h)/\|g+M_{N}h\|$ and $\tilde{v}_{N}:=(g+q_{N}h)/\|g+q_{N}h\|$ satisfy $\|v_{N}-\tilde{v}_{N}\|\to 0$ in probability.
-Hence $\widehat{u}_{N}\xrightarrow{p}\tilde{v}_{N}$, which is the claimed gradient direction of $A_{\beta_{N}}$ with $\beta_{N}=q_{N}$.
-By Lemma 3,
-$A_{\beta_{N}}(x+\Delta x^{(N)})=A_{\beta_{N}}(x)+\varepsilon\nabla A_{\beta_{N}}(x)^{\top}\widehat{u}_{N}+O(\varepsilon^{2}).$
-Write $\nabla A_{\beta_{N}}(x)=g+q_{N}h$.
-Using $\tilde{v}_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}1$ gives $\nabla A_{\beta_{N}}(x)^{\top}\widehat{u}_{N}=\|\nabla A_{\beta_{N}}(x)\|\,\tilde{v}_{N}^{\top}\widehat{u}_{N}=\|\nabla A_{\beta_{N}}(x)\|(1-o_{p}(1)).$ Thus $A_{\beta_{N}}(x+\Delta x^{(N)})\geq A_{\beta_{N}}(x)+\varepsilon\|\nabla A_{\beta_{N}}(x)\|-o_{p}(\varepsilon),$ as claimed.
-∎
-Remark 1 (Edge cases).
-(i) If $\|h\|=0$, then $\nabla A_{\beta_{N}}(x)=g$ and the direction reduces to $\nabla\mu(x)$.
-(ii) If $\sigma(x)=0$ but $\|h\|>0$, the result holds unchanged; only the linearization constants change.
-(iii) If $\sigma(x)=0$ and $\|h\|=0$, the model is locally deterministic and the problem is degenerate; selection aligns with $\nabla\mu(x)$ or is undefined if $g=0$.
-Appendix E GP-UCB and its evaluation efficiency
-Setup.
-Let $f$ be a unknown, black-box function that maps a compact and convex feasible region $\mathcal{D}\subset\mathbb{R}^{d}$ to $\mathbb{R}$.
-At each time $t$, a sampling algorithm $\mathcal{A}$ selects a point $x_{t}\in\mathcal{D}$ to evaluate.
-This selection is adaptive, based on the history of previously chosen points and their observed outcomes.
-The evaluation at $x_{t}$ yields a noisy observation $y_{t}=f\left(x_{t}\right)+\epsilon_{t}$, where $\epsilon_{t}$ is a random noise term, where $\left\{\epsilon_{t}:t=1,2,\ldots\right\}$ are independent, sub-Gaussian random variables.
-The unknown function $f$ is assumed to be an element of a Reproducing Kernel Hilbert Space (RKHS), $\mathcal{N}_{\Psi}(\mathcal{D})$, induced by a stationary kernel $\Psi$.
-Functions within this RKHS possess a specific smoothness property, the nature of which is determined by the kernel used.
-The space $\mathcal{N}_{\Psi}(\mathcal{D})$ is endowed with an inner product and is formally defined as the set of functions
-[…]
-where the norm is given by: $\|g\|_{\mathcal{N}_{\Psi}(\mathcal{D})}:=\left(\sum_{j,l=1}^{\infty}c_{j}c_{l}\Psi\left(\mathbf{x}_{j}-\mathbf{x}_{l}\right)\right)^{\frac{1}{2}}$.
-For any two functions $g_{1}(\mathbf{x})=\sum_{j=1}^{\infty}a_{j}\Psi\left(\mathbf{x}-\mathbf{x}_{j}\right)$ and $g_{2}(\mathbf{x})=\sum_{l=1}^{\infty}b_{l}\Psi\left(\mathbf{x}-\mathbf{x}_{l}\right)$ in the space, their inner product is $\left\langle g_{1},g_{2}\right\rangle_{\mathcal{N}_{\Psi}(\mathcal{D})}:=\sum_{j,l=1}^{\infty}a_{j}b_{l}\Psi\left(\mathbf{x}_{j}-\mathbf{x}_{l}\right)$.
-A key feature of RKHS space is the reproducing property, $\langle g,\Psi(\mathbf{x}-\cdot)\rangle_{\mathcal{N}_{\Psi}(\mathcal{D})}=g(\mathbf{x})$, which holds for all $\mathbf{x}\in\mathcal{D}$.
-The norm of the function $f$ in this space is assumed to be bounded such that $\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B$ for some constant $B>0$.
-The objectives.
-The performance of $\mathcal{A}$ is often evaluated through two types of regret over a time horizon $T$. One is cumulative regret, defined as
-$\mathcal{R}_{C}(T;f,\mathcal{A}):=\sum_{t=1}^{T}\left[\max_{\mathbf{x}\in\mathcal{D}}f(\mathbf{x})-f\left(\mathbf{x}_{t}\right)\right]$ (48)
-The other is simple regret, defined as
-$\mathcal{R}_{S}(T;f,\mathcal{A}):=\max_{\mathbf{x}\in\mathcal{D}}f(\mathbf{x})-f\left(\mathbf{x}^{(T)}\right)$ (49)
-where $\mathbf{x}^{(T)}$ is the output of the algorithm $\mathcal{A}$ after taking $T$ function evaluations.
-The lower bound.
-scarlett2017lower showed that a lower bound exists on the regret of any algorithm for $f\in\mathcal{N}_{\Psi}(\mathcal{D})$. Specifically, for any constant $B>0$,
-$\inf_{\mathcal{A}}\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}\mathbb{E}\left[\mathcal{R}_{C}(T;f,\mathcal{A})\right]=\begin{cases}\Omega(T^{\frac{\nu+d}{2\nu+d}}),&\text{ for Matérn kernels, }\\ \Omega(T^{\frac{1}{2}}\ln^{\frac{d}{2}}(T)),&\text{ for SE kernels.
-}\end{cases}$ (50)
-$\inf_{\mathcal{A}}\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}T\left(\epsilon,f,\mathcal{A}\right)=\begin{cases}\Omega\left(\left(\frac{1}{\epsilon}\right)^{2+d/\nu}\right),&\text{ for Matérn kernels, }\\ \Omega\left(\frac{1}{\epsilon^{2}}\left(\log\frac{1}{\epsilon}\right)^{d/2}\right),&\text{ for SE kernels.
-}\end{cases}$ (51)
-where Matérn kernels are defined as
-$\Psi_{\mathrm{M}}\left(\mathbf{x}-\mathbf{x}^{\prime}\right):=\frac{1}{\Gamma(\nu)2^{\nu-1}}\left(\frac{2\sqrt{\nu}\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}}{\ell}\right)^{\nu}K_{\nu}\left(\frac{2\sqrt{\nu}\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}}{\ell}\right),\quad\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$ (52)
-and SE kernels are defined as
-$\Psi_{\mathrm{SE}}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)=\exp\left(-\frac{\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}^{2}}{2\ell^{2}}\right),\quad\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$ (53)
-where $\nu>0$ is the smoothness parameter, $\ell>0$ is the length-scale parameter, $\Gamma(\cdot)$ is the gamma function, $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind of order $\nu$, and $\|\cdot\|_{2}$ denotes the Euclidean norm.
-GP-UCB algorithm.
-The Gaussian Process Upper Confidence Bound (GP-UCB) algorithm, introduced in the seminal work of srinivas2012information, operates from a Bayesian perspective by placing a Gaussian Process (GP) prior on the unknown objective function $f$.
-This GP is characterized by a zero-mean function and a symmetric, positive-definite covariance function, or kernel, $k:\mathcal{D}\times\mathcal{D}\mapsto\mathbb{R}$, where $k\left(\mathbf{x},\mathbf{x}^{\prime}\right)=\Psi\left(\mathbf{x}-\mathbf{x}^{\prime}\right)$ for all $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$.
-For any finite set of points $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{t}\}\subset\mathcal{D}$, the corresponding function values $(f(\mathbf{x}_{1}),\ldots,f(\mathbf{x}_{t}))^{\top}$ are assumed to follow a multivariate normal distribution with a zero mean vector and a covariance matrix with entries $\mathbf{K}_{t}]_{jl}=k(\mathbf{x}_{j},\mathbf{x}_{l})$.
-Under the assumption of independent, zero-mean Gaussian observation noise with variance $\sigma^{2}$, conditioning the GP prior on a set of observations $\mathbf{D}_{t}:=\{(\mathbf{x}_{j},y_{j}):j=1,\ldots,t\}$ yields a closed-form posterior distribution.
-The posterior mean and variance at any point $\mathbf{x}\in\mathcal{D}$ are given by:
-$\mathbb{E}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]=\mathbf{k}_{t}^{\top}(\mathbf{x})\left(\mathbf{K}_{t}+\sigma^{2}\mathbf{I}_{t}\right)^{-1}\mathbf{y}_{t},\operatorname{Var}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]=k(\mathbf{x},\mathbf{x})-\mathbf{k}_{t}(\mathbf{x})^{\top}\left(\mathbf{K}_{t}+\sigma^{2}\mathbf{I}_{t}\right)^{-1}\mathbf{k}_{t}(\mathbf{x}),$ (54)
-where $\mathbf{k}_{t}(\mathbf{x})$ is the vector of covariances between $\mathbf{x}$ and the observed points, $\mathbf{K}_{t}$ is the covariance matrix of the observed points, $\mathbf{y}_{t}$ is the vector of observed values, and $\mathbf{I}_{t}$ is the identity matrix.
-Inspired by upper confidence bound methods in the multi-armed bandit literature (Auer et al., 2002), GP-UCB employs an optimistic acquisition strategy to select subsequent evaluation points.
-At each step $t$, the next point $\mathbf{x}_{t+1}$ is chosen by maximizing an upper confidence bound on the function’s value $\mathbf{x}_{t+1}=\underset{\mathbf{x}\in\mathcal{D}}{\arg\max}\left(\mathbb{E}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]+\sqrt{\beta_{t}\operatorname{Var}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]}\right).$ This acquisition function naturally balances exploitation, driven by the posterior mean $\mathbb{E}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]$, with exploration, driven by the posterior standard deviation $\sqrt{\operatorname{Var}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]}$.
-The tunable parameter $\beta_{t}>0$ explicitly governs this trade-off, making its specification critical to the algorithm’s performance.
-The evaluation efficiency of GP-UCB.
-whitehouse2023sublinear, wang2023regret showed that GP-UCB srinivas2012information upper bound on the cumulative regret of any algorithm for $f\in\mathcal{N}_{\Psi}(\mathcal{D})$ that is almost nearly tight.
-Specifically, for any constant $B>0$, the GP-UCB algorithm $\mathcal{A}_{\text{GPUCB}}$ achieves
-$\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}\mathbb{E}\left[\mathcal{R}_{C}(T;f,\mathcal{A}_{\text{GPUCB}})\right]=\begin{cases}O(T^{\frac{\nu+d}{2\nu+d}}),&\text{ for Matérn kernels, }\\ O(T^{\frac{1}{2}}\ln^{\frac{d}{2}}(T))&\text{ for SE kernels.
-}\end{cases}$ (55)
-$\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}T(\epsilon,f,\mathcal{A}_{\mathrm{GPUCB}})=\begin{cases}O\left(\left(\frac{1}{\epsilon}\right)^{2+d/\nu}\right),&\text{for Matérn kernels,}\\ O\left(\frac{1}{\epsilon^{2}}\cdot\left(\log\frac{1}{\epsilon}\right)^{d+3}\right),&\text{for SE kernels.}\end{cases}$ (56)
-wang2023regret’s approach to proving the regret optimality of GP-UCB follows the two-step framework.
-The first critical component is to construct a high-probability uniform error bound that quantifies the difference between the true objective function $f(x)$ and its posterior mean estimate $\mu_{t}(x)$ at any given step $t$.
-This bound takes the form $|f(x)-\mu_{t}(x)|\leq\sqrt{\beta_{t}}\sigma_{t}(x)$ for all $x\in\mathcal{D}$, where $\sigma_{t}(x)$ is the posterior standard deviation and $\beta_{t}$ is a carefully chosen exploration parameter.
-The second component involves bounding the cumulative sum of the instantaneous regrets, which, under the uniform error bound, can be related to the sum of the posterior variances $\sum_{t=1}^{T}\sigma_{t-1}^{2}(x_{t})$.
-This sum is, in turn, bounded by the maximal information gain, $\gamma_{T}$.
-The final regret bound is thus a function of $T$, $\beta_{T}$, and an upper bound on $\gamma_{T}$.
-wang2023regret employs tools from empirical process theory and decomposes the estimation error, $f(x)-\mu_{t}(x)$, into a bias term and a random error term.
-By leveraging the connection between Gaussian process regression and kernel ridge regression, along with properties of the Fourier transform for stationary kernels, they show that the bias term is bounded by $f||_{\mathcal{N}_{\Psi}(\mathcal{D})}\sigma_{t}(x)$.
-The more challenging random error term is handled by viewing it as an empirical process indexed by a class of functions and bounding its supremum.
-They bound the $\epsilon$-entropy of this function class, which allows for a high-probability bound on the random error.
-With the new uniform error bound, we can select a much smaller, dimension-independent exploration parameter $\beta_{t}$ that grows only logarithmically with $T$.
-Combining this sharper $\beta_{t}$ with the tightest known bounds on the maximal information gain $\gamma_{T}$ for Matérn and SE kernels, we can achieve the tight cumulative regret of GP-UCB.
+GEPA [agrawal2025gepa] is an iterative, prompt-optimization-based, self-improving AI method.
+The key idea of GEPA is to “combine reflection with evolution”.
+Here, an evolution is the same concept as taking a textual gradient in textgrad.
+GEPA iterates over three steps in each optimization iteration: Step 1) choosing a prompt and evaluating it; Step 2) updating the meta-reflection; and Step 3) improving the prompt.
+For Step 1), it uses instance-level information to (i) maintain a Pareto archive of which prompts in the candidate prompt set currently win on which problem instances, (ii) stochastically sample a non-dominated candidate prompt (favoring candidates that win on more instances), and (iii) evaluate that sampled candidate by rolling it out on a minibatch of problem instances.
+In Step 2), GEPA updates a global meta-reflection $R_{t}$ that summarizes what has tended to work well or poorly so far.
+Concretely, after collecting rollouts and their scores, it applies a reflection operator to the accumulated history, $R_{t}:=\texttt{MetaReflect}(H_{t};M_{\text{critic}})$, where the critic is shown a context-limited subset of representative high-performing and low-performing rollouts (including the prompts and outcomes) and asked to (i) identify recurring success patterns and failure modes and (ii) distill them into a short list of actionable, prompt-level guidelines.
+This shared reflection is carried forward and provided as context to subsequent prompt-evolution (textual-gradient) steps.
+This step is fundamentally the same as TextBO’s meta-reflection steps.
+For Step 3), we take one step of evolution (or equivalently, one step of textual gradient) of the prompt we evaluated in Step 1) and accept it only if it has a better evaluation than the original prompt.
+When the offspring prompt is accepted, we append it to the candidate prompt set without removing the original prompt from the candidate set.
+TextBO-GEPA
+TextBO-GEPA is identical to GEPA except for the prompt-improvement step (Step 3): whenever GEPA would take a single evolution step (one textual-gradient update) to propose an offspring prompt, TextBO-GEPA instead performs multiple successive Best-of-N textual-gradient steps (as in §5) before evaluating and applying the same acceptance rule as GEPA.
+That is, at each inner step, it samples $N$ local textual edits, applies them to form $N$ candidate prompts, generates the corresponding candidate outcomes, and uses the critic (conditioned on the current reflection) to select the most promising candidate; repeating this for several steps yields a refined offspring prompt, which is then accepted and appended to the candidate prompt set only when it improves upon its parent under the evaluation procedure.